A Complete Bibliography of Publications in Mathematical and Computational Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 March 2024 Version 1.03

Title word cross-reference

(2 + 1) [BB16, MJK12, SK18, qS13]. (3 + 1) [LDL10]. (G′/G) [ME12]. (k + 1) [EG00]. (N + 1) [HL16]. (p, q) [BVCS18]. (v, k, 3) [LCL17]. 0− → 0+ [KSK05, KÜM08]. 120 ≤ A ≤ 150 [TM11]. 5(4) [DSI16]. < [BK10]. ^{(λ1,λ2)} [WT15]. ^{114−120} [MGT10]. ^{2} [ETW^{+}23]. ^{2+} [PD19]. ^{4+} [Mer09]. _{1−x−y} [Bül99]. _{2} [AMR^{+}23, AR23, DEL23, JKS23, KNRC22, NKS22a]. _{3} [AR23, DEL23, JKS23, PD19]. _{4} [AMR^{+}23]. ∞ [ASK21, JY23]. _{x} [Bül99]. _{y} [Bül99]. α [LC16, Sav16]. B_{1}(h, k) [Cet16]. β [GVC20, KSK05, KÜM08]. C^{2} [Asl18]. δ [Gun96]. E_{v}^{n+1} [EG00]. ℓ_{p} [SB02]. G [LZ17]. G′/G [Zhu10]. γ [AT05a]. Gϕ [ES06b]. Gϕn [ES06b, ES06a]. H_{0}^{2} [ÖK05]. L_{p} [GOA03, STB04]. L_{p}(x) [Kay15]. λ [CBM^{+}19]. m [ASA20]. M/G/1 [CT14b, KKC08, LZ17]. N [AAD11, AADS11, PKC22]. O(1/n) [RC16]. p [CT14b, DA19]. π [PS19, ASJQ24]. ψ [ENVPF19]. q [MAT20]. R = P (Y < X) [KN18]. σ [SS03]. S̃_{1}^{2} [Kaz05]. y′′ = f(x)y^{2} [SM99]. Z [AT05a].

-Adhesive [AR23]. -adic [DA19]. -Averaged [BVCS18]. -Catmull [LC16]. -Cells [GVC20]. -Decay [KSK05, KÜM08]. -Dimensional [BB16, CUA04, CO16, EG00, GE00, Li10, LDL10, SK18, qS13, HL16]. -Entering [CT14b]. -Expansion [ME12, Zhu10]. -Lidstone [MAT20]. -Modes [CBM^{+}19]. -Operator [Gun96].

-Order [AAD11]. -Polar [ASA20]. -Polygons [PKC22]. -Queue [LZ17]. -Statistical [SS03]. -th [AADS11]. -Zakharov [MJK12].

/M/1/WV [WT15]. /Titanium [AMR^{+}23]. /Water [JKS23, NKS22a].

0-1 [JLST04].

1 [İYK99, KMÇ^{+}05]. 1- [KSK01]. 100Ru [TU97]. 12 [Çat10]. 120 [TM11]. 122 [KMÇ^{+}05, TM11]. 122-128Te [KMÇ^{+}05]. 128Te [KMÇ^{+}05]. 130 [İTM12]. 132Te [İTM12]. 144 [İTİO08]. 144-154Nd [İTİO08]. 15 [SGZ11]. 154Nd [İTİO08]. 16 [KA16]. 162Dy [BUKM96]. 19 [CRA^{+}21, FSHGGB^{+}21, GM21, MPA^{+}22, SPCPMM21, SSGCPR23, TRD^{+}22].

2 [DK03, İTİO08, İTM12]. 2-Crossed [Mut00a]. 2-D [DB17a, DB17b]. 2-Domination [Tur16]. 2007 [Çat10]. 22 [HLX^{+}18]. 233U [Übe04]. 24 [KF19a]. 2=z2 [AA03]. 2D [AOL19, ÇÇÇ07, FMF20, FLYY21, HWA99, PPB^{+}20, VC23a, VVN23]. 2P [ENT05].

3 [Mat10b, OA09, YH05]. 3-D [OA09]. 3-Homogeneous [LCL17]. 3-Index [Man16]. 3-Machine [Cha05]. 3-Space [BY13, UÇ96, Yay00]. 300 [KA16]. 3D [AMR^{+}23, ACM19, AA13b, ÇÇÇ07, NSCTPW22, SSA22b, TUK19, ZJ10]. 3D-Printed [AMR^{+}23].

4-Dimensional [YH05]. 4-Point [PTLZ18]. 40 [HLX^{+}18, ŞİH13]. 40/50 [ŞİH13]. 4D [DJM18]. 4p [AE02].

5-Point [Asl18]. 50 [KD03, ŞİH13]. 56 [KF19a].

90Nb [BSK05].

= [IMWY18, Wan13].

AA [DUM04, MAE97]. Abilities [Kay16]. Ability [AG12]. above [KZAN98]. Abrasive [MWM16]. Absolute [ÇA04]. Absorber [ES06b, ES06a]. Abstract [Sha99]. Abutment [MAS^{+}11]. AC [GM03]. Accelerated [MCNM22]. Accelerating [Ese11, JL24, Mat10a, Öz00b]. Acceleration [DSU22]. Accepting [FSHRCV^{+}21]. Access [HSG^{+}20]. According [EBD11, Kız14, OFCP21, Özy11, KAY09]. Account [DB17a, DB17b]. Accounts [HMKS18]. Accuracy [Ara08, CC24, DL23, Jia17]. Accurate [EAEAA19, How22]. Acknowledgement [MCA17, MCA18, MCA19, MCA20]. Acknowledgment [MCA21, MCA22, MCA23]. Acoustic [ESNB19, PSW24, ZLZ17b]. Acoustics [DHK23, MCBM23]. Action [CGMCCA23, Shi06]. Activation [TKK^{+}22]. Active [AK09, ES06b, ES06a, EPDG21, Guc04, LL23, TGK96, TGG96]. Activity [AAR24, IÇÜ12]. Actual [SS07]. Actuator [HCdCTAAMEJ20]. Actuators [VJ23, VC23a]. Adaption [PS22]. Adaptive [AB03, AK11, CA19, CT14b, CKLZ21, DSMP19, FA22b, GC23a, KD19, KAM20, LOSCO^{+}20, PC23, SKB22, WZ17, ZLLX20]. Added [BIM^{+}20]. Added-Mass [BIM^{+}20]. Addendum [GÖA96]. Additive [Bad22, KAKB08]. Address [ARCRMM20, SPCPMM21]. Adherends [ABO^{+}16]. Adhesive [ABO^{+}16, AR23, GGT^{+}23, GA99a]. Adhesive-Butt [GA99a]. Adhesively [KT98, KTC99, ST11, YK96a]. Adhesively-Bonded [ST11]. ADI [IA04, IZS04]. adic [DA19]. Adjustable [Yan16]. Adjustment [TGK96]. Administration [Egb18]. Adomian [AJ20, Doğ12, HMDG11, RAA10]. Advanced [EGB99, KNHK20, LEVP21]. Advection [DRCSF22, GC23a, GZ20, SGZ10, SGZ11]. Advection-Aware [DRCSF22]. Advection-Diffusion [SGZ10, SGZ11]. Advection-Reaction-Diffusion [GZ20]. Adversely [Sha19]. Aerator [OA09]. Aero [AMR^{+}23]. Aero-Composite [AMR^{+}23]. Aerodynamic [MARL21]. Affect [AÇ11, AARW22]. Affecting [GLAA08, Sha19]. Affordability [HC22]. African [MMM23]. Ag [DEL23]. Ag/Al [DEL23]. Against [GNI19, KG13, BY17, YS17]. Age [MPA^{+}22]. Aged [DUM04]. Agent [MPA17, NMCM19]. Agent-Based [NMCM19]. Agents [AAGH21, GGMO21]. Agglomerated [dAPdSL^{+}22]. Aging [AMK96]. Agricultural [CÖGD11, SÇ10]. Ahead [Alm10]. AHP [ASA20, BK09]. AHP-Assisted [ASA20]. Aided [SBF96, TE05]. Air [BA99, BAUO09, GERV19, Kop02, SS07]. Air-Fuel [SS07]. Air-Jet [BA99]. Aircraft [BAS96]. Airdrop [GT22]. Airflow [BA99]. Airfoils [MARL21]. Airplane [HB99]. Airways [UÖKE14]. Al [Bül99, JKS23, AR23, DEL23, KA11b, TÇ06]. Al/Sicp [TÇ06]. Algebra [Can96, QN18]. Algebraic [CCWY10, Gha18b, KA96a, KS18, LXZ10, TB21]. Algebras [AHK18, Arv04, ENVPF19, FTGSNV21, KUI18, Mut00a]. Algorithm [AK09, AK18, AÖ09, Aks05a, Aks05b, AKHHM17, Alm10, BY16, BMV22, BD09, BGN13, CAQCMM22, CACRF^{+}21, CSSOB20, Cui07, ÇA06, DB03, GLA05, Gun96, GZ14, HJL10, KSTJ18, KB03, KPSS18, KD03, LP17, LCB18, mLkWbPxX17, LBY^{+}22, Mat10a, Mat10b, MFF20, PBY08, PKB09, PR21, PRFH23, PGB^{+}11, RC99, Sal02b, SAH98, SH18, SSGCPR23, SKKC18, SKK19, TDCT11, Tur16, TY11, UKÇ^{+}09, Ugu13, YRB16, ZLZ^{+}17a, ZWZ16]. Algorithmic [AFGV20, ETW^{+}23, FHRS18]. Algorithms [Aks08, AMMMM^{+}23, CBSTG13, CS22, CKLZ21, CLGGLB^{+}22, CS19, Çun08, HDAA20, Kay16, NSO^{+}21, PBY07, PS22, SAK01, Yen05, Yen14]. Aligned [AM05]. Alignment [Eli16, GLA05]. Allee [GK12]. Allocation [KKL18, LZC22]. Allowed [SBA^{+}04]. Alloy [AMK96, Bül99, DUM04]. Alloy/B [AMR^{+}23]. Alloys [KTS13, MAE97, TDC08]. Almost [CC24, MR19a, Sav12, SB02, Yao15, ZPL16]. along [DL24, ZZ17]. Alternate [CMCM21]. Alternative [PLW21]. Aluminium [AMK96]. Aluminum [ABO^{+}16, DUM04, KTS13, Oku11c]. Alzheimer [PM22]. Am [LGT^{+}18]. Amalgam [AT05b]. Ambartsumian [EAEAA19]. Ambient [DÇ06]. Ambulance [ACGC21]. American [FIJ21, SW10]. among [MWM22]. Amplification [DRKW21]. Amplification-Fitted [DRKW21]. Amplitude [AKT04, Kol11, KA16, ZZ10]. Analog [ÇT14a]. Analogue [BP99]. Analyses [CDA10, SS07, SY11]. Analysing [CPA20]. Analysis [AJ12, Ace19, AJ20, AO17, AVC23, AW19, AG03, ABO^{+}16, Alp02, AAR24, AB23, AK02, AÇ11, ATSA97, AT05b, AMM23, BT19, BG02, BG19, BdCCF19, BV19, BSSK12, BDC^{+}18, BÖ07, BO10, BBPL19b, CTL14, ÇŞS01, ÇKAK04, ÇDYS13, CPP19, Can14, Çat10, Cha97, CT14b, CÜ04, CK07, CG09, Dar11, DSMP19, DJY^{+}22, DCA10, Den16, DD11, DLL10, ESNB19, EBD11, Erp12, Ese01, ETL22, FLYY21, GVC20, GA10, GOOM18, GBİ14, GAAAS17, GCGZ10, GB00, GA99a, GM17, GC23b, GÇÖ12, HB99, HLX^{+}17, HLX^{+}18, Har21, HMKS18, HK23, HCdCTAAMEJ20, HXY15, JLZ04, JLST04, JLS^{+}10a, JLS^{+}10b, Ji13, JH13, JWP23a, KTK11, Kan00, KEA11, KC08, Kar10, Kat17, KTC99, KKC08, KR23, KWA^{+}10, KH22, KYG20, KPS04, KP12, KA11b, KIY^{+}23, KÖAG00, KE01, Küç01, KPSZ24, LLTC17, LTYG17, MK96]. Analysis [MC20b, Mit21a, MRG23, NCDM22, NSG^{+}23, NSCTPW22, NKI19, OMX18, Oku11a, OYCS11, OWO^{+}23, ÖG13, OD19, Özb02, Özc00, Öze02, Oze03, OC05, PT21, Pin22, Qad24, Rod20, SQA11, Sal06a, SÖS13, SS06, SSGCPR23, ST22, SHE05, Tan08, TÇ04, TE05, TA96, TGS98, TB21, ÜIG03, UAI21, UC19, VMRVF^{+}23, WFV21, WT15, Wu17, YD11, YY11, YQG22, YUKA03, YA14, ZBBZ18, ZdlÁGLTT22, ZJ10, ZCC15, ZA98]. Analytic [Bil02, Cet16, HL16]. Analytical [AYG14, AK19, ALFBPLE20, All17, ÇŞS01, DSG19, FKRG10, GGJ10, HL14, How22, Jab18, KLGP24, KG10, KAAK10, Kum07, MMF18, Mos18, OSAAA18, OFCP21, RSS^{+}21, RFD^{+}10, Sha19, SAM^{+}22, VRMJG19, YD11]. Analytically [ÇT14a]. Analyticity [BMT^{+}21]. Analyze [ACC19]. Analyzing [PRT^{+}19]. Anastomoses [AK11]. Anaysis [GB98, KT98]. Ancient [ZH10]. Aneurysm [SGC23]. ANFIS [JTM^{+}22, ŞE17]. ANFIS-Type [JTM^{+}22]. Angle [Akb09, EÜA13]. Angles [ST11]. Angular [Har21]. Anisotropic [Akt01, Boo17, CK07, Koi19, UY11]. ANN [KAK96, TDC08, AÇ11, RJ17]. Annealed [TB21]. Annealing [BŞ13b, FSHRCV^{+}21, SHFSGB^{+}21, Sar09, ÜC11, VMRVF^{+}23]. Annular [CG09]. Ant [BD09]. Antenna [KIY^{+}23]. Antennas [SGE98, UY03]. Anterior [GGT^{+}23]. Anti [CMB19, FLYY21, SD13, TEK11]. Anti-Periodic [SD13]. Anti-Plane [FLYY21]. Anti-Roll [TEK11]. Anti-Torpedo [CMB19]. Antibody [FA22a]. Antimagic [YBYL21]. Antiprotons [ENT05]. Aortic [DL23]. Aperture [Rou21]. Apostol [Saj18]. Apostol-Type [Saj18]. Apparatus [Haj05]. Appearance [DMN22]. Appearing [GLS18]. Appl [HLX^{+}18, KA16, KF19a]. Applicability [AMR^{+}23]. Application [AS09, AA16, AA13a, AHID18, ASAH20, AD11, BG02, BGA10, CT01, CA19, Çev10, DCC23, Erp12, GZY20, GFPF22, GBİ14, GHN11, GM03, GK11b, HC22, IMKEK18, IWH21, ICDM21, Koi19, KSS19, KTD12, LY15, LC18, LBY^{+}22, MSSZ20, MAE97, MVC23, ME12, Muc01, MCNM22, OSAAA18, PMF20, RC16, RMBTGH^{+}20, Rou21, SY11, SKM^{+}22, ÜIG03, WLW^{+}10, Xu10]. Applications [Alt09, ANK23, BSA^{+}23, Çat10, CW16, CKKA22, ÇÇ03, DW16, Fan20, Guy20, JTM^{+}22, JAT^{+}21a, JHA^{+}22, KHR^{+}22, KGCFK22, KPTM20, KMCI22, LB20, Loj22, LEVP21, MAS^{+}11, MAT20, MC20a, MKK97, MM10a, NVC22, NAC22, NAC23, NU19, Ole20, OV18, ÖG03, ÖE03, PRFP^{+}23, PSF22, RKX23, RSF23, Sağ98, SGZ11, SKB22, SJ19, TAS^{+}23, TGC22, YZHY17]. Applied [AVC23, BP99, CTH^{+}22, CP22, DK99, DSD^{+}21, FP21, FCGL19, HCdCTAAMEJ20, MSPC23, MF21, OFCP21]. Applying [AHVG23, HM23, QN18, WN21]. Approach [ARCRMM20, ADS19, AN19, ASA20, AAS11a, AA13b, BGG96, BŞ13b, BPC^{+}23, BCGS23, CMB19, Cha05, CCWY10, DSG19, DSMP19, DNS23, Egb18, EÇ99, EGIK97, GCCMD23, GWC98, GZ20, GPMRVM21, GU05, GGS11, GGMMDI20, HT19, HNB20, Kah11, KA96a, KHL06, KWA^{+}10, KYG20, KLRA10, KAK00, LVBV19, MPA^{+}22, MM17, MRIR20, MARL21, OM20, PB97, PAB11, Pak16, PRT^{+}19, PPB^{+}20, RFD^{+}10, Sal09, Sam99, SAZ18, ŞİH13, TDC08, TGG96, WCH23, Wu10, Xu10, Yur99, ZH10]. Approaches [CK12, DRH21, MPP^{+}24, SW22, TCA96]. Appropriate [Erp12]. Approximant [PKO11]. Approximate [BÖ07, CSSOB20, Çat10, EAEAA19, HLL15, KKS11, MR19a, NM10, OSAAA18, PÖ98, PY00, PŞ05, Raw14, Üna03, XWCF19, YÖS10, Yal13]. Approximating [NSCTPW22, QN18, SSA22b]. Approximation [AAS11a, CC24, Ese11, GHN11, Har21, Jia17, KAS19a, KB11, MR05, MV21, ORC19, Qua18, UAT19, ZZ17]. Approximations [CS22, How22, PALL18]. Aquifer [KA11a]. Aquila [TCC^{+}23]. Arbitrage [Hon20]. Arbitrarily [How22, KA02]. Arbitrary [BVCS18, Ese01, KE01, PB99]. Arc [RKX23]. Arch [DCC23, DJ10]. Archimedean [DJL^{+}05]. Archimedes [PS19]. Architecture [Sal03]. Architectures [WN21]. Architextured [TSC11]. Archiver [CS22]. Arctan [NAC23]. Area [Du 22, MWM22, RC16]. Areas [BBK20]. Areostationary [SGBR23]. Argument [ASJQ24]. Arid [JMB22]. Arising [KSS19, Saj18]. Arithmetic [Ara08, GU05, dlF21]. Arm [AAR24]. Arma [ÜIG03]. Arrestor [YYWX16]. Arrhythmia [AG12]. Arrival [KHL06, SHE05, Sul07, WT15]. Arterial [LVHL22, ÜIG03]. Arteries [SGC23]. Artifical [AMK96]. Artificial [BY11, CYAK03, ÇT12, ÇS10, DK03, DÇ06, DUM04, FPCA18, GGMO21, GÇŞT07, GNK19, KAK96, Kay16, KY13, MKK97, OOI^{+}22, Sağ98, Sar09, ŞÖK11, SGEP06, SÇGA12, SWJS22, TÇ06, TDC08, ÜDMA04, UKY96, YS02]. Aspect [KAAK10]. Aspects [ABH^{+}22, ÖG03, PT21, Sta23b]. Asphalt [Özg10]. Assembly [BD09, SKKC18]. Assessing [MCBM23, SGC23, Yur99]. Assessment [BP99, CPP19, ÇYAÖ05, ÇS10, Çiç07, FSH^{+}23, MLBR20]. Assigment [BAS96]. Assignment [AA16, BPC^{+}23, KSTJ18, Man16]. Assimilation [IWH21]. Assisted [ASA20]. Associated [AHK18, CBM^{+}19, DPC09]. Association [DT08, SGCRRL^{+}23, SSGCPR23]. Assumed [Dar11]. Assumption [KYG20]. Assurance [CKCC16, DJL^{+}05]. Astrocytes [PM22]. Asymmetric [BO10, HA10, IYKA20, LL24]. Asymmetric-Plan [BO10]. Asymptotic [ABKU10, Cai06, CCY08, CC09, Den16, KAGAM23, Mub23, UY03, UAI21]. Asymptotics [KSS19]. Asynchronous [Guy20]. Atmosphere [GS01, GS02, GS10, Shi06]. Atomization [JWP23a]. Atoms [CB04, ENT05]. Attached [Xu10]. Attachment [GGT^{+}23]. Attack [OV20]. Attainable [GOA03]. Attendance [ŞE17, TKK^{+}22]. Attention [EEJRLB^{+}22]. Attribute [Fu15, FbF16, Fu16, JZL10, Liu11]. Attributes [WOL15]. Attribution [Can14]. Augmentation [DCAR21]. Augmented [HT19, Mut00b, WZT20]. Austria [LSB^{+}20]. Authentication [BBSC18, KCMLTR18]. Authorship [Can14]. Autism [EEJRLB^{+}22]. Autocorrelated [Kar11]. Autoencoders [DRCSF22]. Automata [FPCA18, PSF22]. Automated [OPUÖ11]. Automatic [DCAR21, Li16, OC09, SWA^{+}22, WA01]. Automation [UC99]. AutoML [CTH^{+}22]. Automobile [ZWZ16]. Automotive [ACL^{+}22]. Autonomous [GLS18, NSG^{+}23, PSF22]. Auxetic [YQG22]. Auxiliary [HMDG11, RC16]. Average [AÇ11, JS18, Tur16]. Averaged [BVCS18]. Aware [DRCSF22, LP17]. AXB [Wan13]. Axial [IE10, WFV21]. Axial-Flux [WFV21]. Axially [Alt10, BIM^{+}20, Öz00b]. Axis [RKR^{+}22b, TK04].

B [Çat10, AMR^{+}23, CT01, DS04, GA10, mLkWbPxX17, OWO^{+}23]. B-Spline [CT01, DS04, GA10, mLkWbPxX17]. Back [DSMP19, VRMJG19]. Back-EMF [VRMJG19]. Back-Stepping [DSMP19]. Background [YS17]. Backorder [TCC11].

Backstepping [ZLLX20]. Backward [SW10]. Bacterial [SGCRRL^{+}23, SWA^{+}22]. Bad [UKY96]. Baffle [SJ22]. Balancing [AE03, BD09, SKKC18]. Ball [GB18, TW18, YA14]. Ballistical [Sha99]. Banach [KAM20]. Bandwith [SGE98]. Banking [BG02]. Bar [GAAAS17, TEK11]. Barcodes [AB96a]. Barley [DPC^{+}10]. Barriers [Unv07]. Bars [Özd96a, Özd97]. Base [Sal03, TMK21]. Based [Abi03, AE16, AKHHM17, AB19, AAS11a, AFGV20, AK11, AARW22, BT19, BhMsG16, BG99, BGRV21, BIM^{+}20, BG04, BBSC18, BD09, BPC^{+}23, BDC^{+}18, BS23, BS15, CGMCCA23, CA19, ÇGV02, ÇŞÖG13, ÇSB10, CWF^{+}22, CDDM23, Çiç07, CDA10, Cor19, DCA10, EE22, Eli16, Emi03, EBD11, FSH^{+}23, Fan20, Fu15, FbF16, Fu16, FLYY21, GT22, GGMO21, GDE12, GNK19, GAAAS17, GCPV18, Gun96, GC23b, HLX^{+}17, HLX^{+}18, HMKS18, HCGFOON20, HXY15, HSG^{+}20, HBS^{+}17, ID14a, JY23, JZL10, JSTGV^{+}19, KD19, KYG20, KPTM20, KA09, KA11b, LVHL22, LY15, LP17, LZC22, LL17, MECRMT^{+}21, MRIR20, MMM23, Mat10a, MPP^{+}24, NMCM19, ÖKK03, PYR22, PPRVC20, SHFSGB^{+}21, SC19, SAZ18, SÖS13, SÇGA12, SA19, ŞİH13, SGC23, ST22, TCC^{+}23, TB16, UKÇ^{+}09, UKY96, VMRVF^{+}23, VVN23, WJ23, WCH23, WW20, XWCF19]. Based [YL16, YW24, Yur99, ZJ10, ZGV10, Zha14, ZG17, ZWZ16, İBB98]. Bases [AFN^{+}19, BBH19]. Basic [Wu17]. Basis [ACM19, BMV22, BP21, EE22, GZ14, KYG20, KF19a, KF19b, UAT19]. Basket [Hon20]. Batch [KHL06]. Batch-Arrival [KHL06]. Bay [BÖ07]. Bayes [CDDM23]. Bayesian [ARCRMM20, CPA20, Erp12, Kar12, KAY09, Kay16, OM20, PLW21, Rou21, ÜC11]. BDDC [CKLZ21]. Be [BD22]. Beam [AJ20, Akt01, ANA20, BÖÖ11, BSSK12, ÇŞS01, ÇDYS13, DPC^{+}10, DCA10, Ese11, Ese01, Has03b, KEA11, KA02, Kir04, KÖAG00, KE01, KKA04, MRG23, Öz00a, Öz00b, ÖÖ05, Özk01, PÖ98, PB01, PB03, TZ19, TB16, YG02]. Beam-Mass [ÖÖ05]. Beam-to-Column [Kir04]. Beam/Columns [DPC^{+}10]. Beams [Alt10, ÇKAK04, CÜ04, GEKB13, KTK11, MRG23, ÖD06, TDT^{+}11, TK04]. Bearing [Yür02, YA14]. Bearings [BY06, ÜDMA04]. Bed [BG19, BGA10]. Bed-Nets [BG19]. Behavior [ALFBPLE20, EY13, GVC20, GGMO21, GEKB13, KAGAM23, KEA11, Kir04, LZL^{+}23, LGT^{+}18, MÖ96, MLBR20, OSAAA18, PTLZ18, TDC08, Tay01]. Behaviors [Guy20]. Behaviour [BCN22, RKX23, TM10]. Belief [NÇK14, PLW21]. Belt [YYWX16]. Benchmarking [VPJ22]. Bend [ZHN10]. Bending [CDA10]. Bends [TEK11]. Bernoulli [Alt10, LSARZ23, TZ19, TB16]. Bernstein [DI13, YG15]. Bertrand [BY13]. Bessel [Den16, TAS^{+}23]. Beta [DA19, HBFM22, SÇGA12]. Better [PKB09]. Between [GGJ10, YU07, BE99, CLY10, DMN22, DAD^{+}18, Du 22, ES06b, ES06a, Erg01, GB18, Has03b, IWH21, KBE01, PD19, Qad24, Sal09, ŞÖ11, SPCPMM21, TM11, UT96, Yür04]. Bézier [CT01, Kus19, PQ23, TT11, Yan16]. Bézier-Spline [PQ23]. Bi [ARCRMM20, Cha08, Sul07]. Bi-Level [Sul07]. Bi-Objective [ARCRMM20]. Bi-Tandem [Cha08]. Bias [ARCRMM20, CZC17]. Bias-Variance [ARCRMM20]. Biased [KN18, MC20a, MC20b]. Bicoherence [HCGFOON20]. Bicoherence-Based [HCGFOON20]. Bidispersive [RKR^{+}22a]. Bifurcation [Ars06, CTL14, Saj18, Sin10, ZCC15]. Bifurcations [PS20]. Biharmonic [Mat19a]. Bilevel [YKG20]. Billiard [BdCCF19]. Binary [ETL22]. Binocular [BVvdH^{+}23]. Binomial [GB23, SKM^{+}22].

Bio [LOSCO^{+}20]. Bio-Inspired [LOSCO^{+}20]. Biochemical [RAK22]. Biodiesel [SY11]. Biological [ABH^{+}22, BBK20, PGB^{+}11, SGCRRL^{+}23, UAT19]. Biomechanism [AFJ22]. Biomedical [NAC23, TGC22]. Biosciences [MQ18]. Biostructures [PRT^{+}19]. Bioventing [SZ23, SEZ24]. Bipolar [KUI18, SAZ18]. Biquaternionic [TD12]. Bishop [Kız14, Özy11]. Bisimulation [NKI19]. BisQ [HCGFOON20]. Bit [KS18, Kle18]. Bit-Streams [KS18, Kle18]. Bivariate [HC22, HBFM22, PMF20, YZHY17]. Black [PMS03, TSS17]. Blades [CA19, GÇÖ12, ZdlÁGLTT22]. Blasius [PKO11]. Blast [TMB96]. Blind [ATZES^{+}07]. Bloat [JSTGV^{+}19]. Block [AO16, CBM^{+}19, KO15, KA11c, LCL17, NU19]. Block-Transitive [LCL17]. Blood [DL23]. Blot [LRBMAM^{+}23]. Blowflies [Yao15]. Blowing [NKS22a, NKS22b]. Bluff [TCY06]. Bodies [NUM^{+}19, NU19]. Body [ÇÇ03, QM16, TCY06]. Boltzmann [PSW24]. Bonded [KBE01, KT98, KTC99, ST11, YK96a]. Bonding [TDC08]. Bone [Alt12, TKCC19]. Boolean [KB03]. Booms [OM20]. Boronizing [KST10]. Bot [NAA20]. Both [FA22b]. Bottle [Kaz04]. Bottom [KPS04]. Boundaries [LL24]. Boundary [ABKU10, BG18b, EL04b, EM08, FJ21, GB98, GB00, GKE05, Hak22, HM10, HPA13, JS22b, JKS23, Ken96, KG10, Mat19b, Mor22, MM10d, NNW14, OA13, PÖ98, PB01, PMR14, SWJS22, TY15, YRB16, ZLZ17b]. Bounded [CS22, Yao15]. Bounds [Dil02, Hon20, SB02, STB04]. Boussinesq [HL16, Li10, Öze00]. Bovines [DAD^{+}18]. Bow [GCPV18, LL17]. Bozdogan [Çat10]. Bracing [KG13]. Braided [ZJ10]. Brain [BNSDC20, FPCA18, HFATZR^{+}20]. Brainstem [HFATZR^{+}20]. Branch [TOM^{+}22]. Brand [CMB19]. Brazing [ŞT96]. Break [Cha05]. Breakdown [SHE05, TCC11]. Breakdowns [LZ17]. Breast [LRBMAM^{+}23, OMX18]. Breeding [Übe04]. Bridge [DCC23, LL23]. Bridging [ACC19]. Broadcast [BBSC18]. Broken [GAAAS17]. Brownian [SSV23]. Brushless [ÇA10, RCTR20]. Buckley [AGKT21]. Buckling [AG03, Dar11, EY13, LSARZ23, Özd96a, SUE13, TA96]. Budget [EÇ99]. Buffer [YYWX16]. Building [LZ21, UC99]. Buildings [Bak11, BO10]. Bulbous [GCPV18, LL17]. Bulk [Cha97, DZA19, SHE05, Sul07]. Bulk-Service [Cha97]. Bull’s [Fu15]. Bull’s-Eye [Fu15]. Bundles [Civ96, Muc01]. Burgers [BA11, CO16, ME12, Sar11]. Burgers-Type [ME12]. Burr [JAT^{+}21a]. Bus [TGK96]. Butt [GA99a]. BVPs [Faz21].

C [AMR^{+}23, KIY^{+}23, DD11, SC19, ŞİH13]. C/ZrO [AMR^{+}23]. C60 [KTK97]. Ca [PD19]. Cage [GAAAS17, MSSZ20]. Calcium [DL24]. Calculate [BBK20]. Calculating [CB04]. Calculation [AA01, ATK99, BK10, BhMsG16, BDH19, CTAA11, LL17, OD19, Öz00a, PCCT11]. Calculations [DE04, Dik06, İTİO08, KMÇ^{+}05, Man04, VRMJG19]. Calculus [AA13b]. Calibration [Haj05, HRM19, Sat23, SEZ24]. Call [TSS17]. Campaign [KY13]. Can [BD22, CMM19]. Canal [TY11]. Cancer [Amo21, LRBMAM^{+}23, OMX18, Rod20]. Candidate [GDE12]. Candidates [JLKK12]. Canine [Küç01]. Cantilever [AO17, Akt01, BSSK12, CDA10, KÖAG00, Özc00]. Capabilities [Bha24]. Capacitor [MFF20]. Capillary [Koi19]. Capital [Dan22, EÇ99]. Caputo [OWO^{+}23]. Carbide [Oku11a]. Carbon [AARW22, DCA10, MKK97]. Carbonization [GM17]. Cardiovascular [HGS^{+}22]. Career [CGM^{+}23]. Carleman [All12, AD04]. Carlo [AÖ05, CATCdlF20, ÇT14a, Man04]. Carotid [Ars06]. Carrying [Özk01, Özy09]. Cartesian [DB17a]. Cartilage [GZ20]. Cascades [PS20]. Case [AHVG23, AK18, DDPK21, EEJRLB^{+}22, GCPV18, HXY15, KSTJ18, LTYG17, LDCST21, YY11]. Cases [FSHGGB^{+}21, NSG^{+}23]. Casson [ETL22, KH15, RKS23]. Catalan [QSC19]. Categorized [JMS18]. Catmull [LC16]. Cattaneo [JS22b]. Cauchy [CP22, RAA10, SB02, STB04]. Causal [SSGCPR23]. Caused [YUKA03]. Cavitation [DHK23]. Cavity [HFAAM^{+}22, HWA99, SJ22]. Cd [Mer09]. Ce [İYK99]. Celine [Shr04, Shr05]. Cell [Amo21, FA22b, JL24, Mer09, Öze02, Oze03, SWA^{+}22]. Cells [GVC20, Mer09, Rod20]. Cellular [FPCA18, GFPF22, OWO^{+}23, UC99]. Cement [CYAK03]. Censored [IMKEK18, MCNM22, SK05, Sul08]. Central [HIVH21]. CeO2 [TSC11]. Cerebellar [Öze02]. Cerebral [SGC23]. Certain [Shr98, Shr05]. Certified [FdSPW20]. CFD [HTLY18, KR23]. Chain [KC08, KPTM20]. Chain-Based [KPTM20]. Challenges [CGM^{+}23, LCW^{+}22]. Chance [ABKU10, GBİ14, Kha99]. Change [HBFM22, Kar12, Kur98]. Channel [ADE99, HA10, IYKA20, IE04, Ken96, MAKH10, PD19, RRO16b, RRO16a, SJ22, Sul07]. Channel-Flow [IE04]. Channels [PD19]. Chaos [ALK19, BS23, HJL10, KLRA10, PPRVC20, Saj18, Sin10]. Chaotic [CLY10, CTL14, FSHRCV^{+}21, JFRZ10, KAK00]. Characterisation [TGG96]. Characteristic [RAA10]. Characteristics [GPEG10, JH13, Kop02, MMM23, TGS98, WSLZ17, YDD11]. Characterization [Alh21, GPACC^{+}19, Kız14, MECRMT^{+}21, PPRF22, VVV23]. Charged [KM10]. Chart [EB00]. Chebyshev [Dol04, GHN11, YÖS10]. Chemical [BP23, ETL22, JKS23, SSV23]. Chicken [KSTJ18]. Chikungunya [GPACC^{+}19]. Children [MWM22]. China [Bec17, HXY15, Ji13]. Chinese [LTYG17, WW20, ZH10]. Chopper [AÖ09]. Chopper-Controlled [AÖ09]. Chord [AFGV20]. Chord-Based [AFGV20]. Christov [JS22b]. Chromatic [YBYL21]. Chrome [LZL^{+}23]. CI [SY11]. Circuit [AMMMM^{+}23, BDH19, KIY^{+}23, PPB^{+}20]. Circuits [BÇ13, KAK00]. Circulant [SB02]. Circular [DJH18, EY13, HB01, IMWY18, KY96, KWA^{+}10, LC18, ÖD06, RFD^{+}10, TA96, Yür04]. CIS [MCB19]. City [SÇGA12]. Claim [EA22, MVC23]. Claims [SW10]. Clamped [BKOS11]. Class [All12, AD04, AHR05, DKÇ97, HNB20, Kar10, Kha03, KM10, Li16, MHC21, NK10, Raj10, RC99, Shr98, Shr05, ŞİH13, TYS13, VSAMMM23]. Classes [KKL18, SQA11]. Classical [Bro19, Özy11]. Classification [AG12, BG99, BY11, CACRF^{+}21, Emi03, KAY09, KBH06, KY13, MSPC23, MM10a, OC09, Pak02, SM99, SKAK05, XWCF19, ZMA23]. Classify [FTGSNV21]. Clearance [UÖKE14]. Climate [JMB22]. Climatic [Pin22]. Clinical [BPD08]. Clinker [CYAK03]. Close [Car21]. Closed [Koi19, NT04, PRFP^{+}23]. Closed-Form [PRFP^{+}23]. Cloud [CWF^{+}22, TUK19]. Cloudino [GERV19]. Cluster [AÇGJ99, DJY^{+}22, HXY15]. Clustering [AE16, LZC22, PGB^{+}11, UB10]. Clusters [GGJ99, LP17, PD19]. Clutch [ZWZ16]. CM [GM03]. CMOS [CATCdlF20]. CMRH [TW18]. CNC [SBF96]. CNN [FSHGGB^{+}21]. Co [HXY15]. Co-Evolution [HXY15]. Coal [HXY15, JB15]. Coal-Power [HXY15]. Coarse [BD22, CKLZ21]. Coated [Küç01, Mub23, TSC11]. Coating [Oku11c]. Coatings [OYCS11, PCCT11]. Code [EBD11, KCMLTR18]. Codes [JS18, Yil11]. Coding [HDAA20]. Coefficient [GÖA96, IMWY18, LNBM10, MJK12, PŞ06, ŞÖK11, ÜDMA04]. Coefficients [CKLZ21, GGS11, HL14, KKS11, KAAK10, KSS19, MO04, SRLEPVP20, SAM^{+}22, Yal13]. Cognitive [MECRMT^{+}21]. Coherent [Boo17]. Cold [JMS18]. Collaborative [LSB^{+}20]. Collection [RSF23]. Collective [EGIK97]. Collision [AB99, BDÖ^{+}99]. Collisions [ENT05]. Collocation [All12, BS15, DS04, DI13, Dol04, FLYY21, GY13, KYS13, TO13, TYS13, TY15, YÖS10, YG15, YW24]. Colombia [GPACC^{+}19]. Colonies [SWA^{+}22]. Colony [BD09]. Color [KBH06]. Coloring [Bad22]. Colouring [ÇŞÖG13]. Column [HCdCTAAMEJ20, Kir04, ŞÖ11, YG02]. Column-Beam [YG02]. Columns [DPC^{+}10, LSARZ23]. Combination [ESNB19, ZCC15]. Combinators [ÇGV02]. Combined [Doğ12]. Combining [BDH19]. Combustor [KK97]. Commitment [IGSB18, LTYG17]. Common [GM03, JLKK12]. Commonality [CKCC16, HT19]. Communication [CLY10, KIY^{+}23]. Community [CGM^{+}23]. Commutative [Kil03]. Commutators [Mut98]. Compact [PRFP^{+}23]. Company [BGG96, NYH^{+}17, TWL10]. Comparative [BPD08, EBD11, ETL22, Kay16, KY13, MSPC23, Oku11b, ŞE17, TCA96, Yen14, YK96b]. Comparing [KAK96]. Comparison [AE16, ABH^{+}22, AT05b, BY06, BP99, BDMS19, CCY08, ÇA06, ES06b, ES06a, Flu15, FHRS18, GN19, HCdCTAAMEJ20, İTİO08, İBB98, JAT21b, Kil03, KAK00, MDNN10a, PB99, SKGP22, SPCPMM21, VHMMG20, VC23a]. Comparisons [CC09, PBY07]. Compartment [Egb18]. Compatibility [TOM^{+}22]. Compensation [VVN23]. Competition [PAR23, SAZ18]. Competitive [Can14]. Complementarity [SA97, ZG17]. Complementary [AJ12, MAT20]. Complete [FLYY21, KA11c]. Complex [AN19, ASK21, BBPL19b, BB16, ÇA04, HAK03, ID14b, Ken96, LH10a, LC18, PGRL21, Qad24, RAK22, SQA11, TGS98]. Complexes [Mut00a, RG23]. Complexity [KB03, Nur03]. Compliant [DRS17]. Complicated [CML10, PSW24]. Component [BG99, Can14, FSHGGB^{+}21, HT19, MF21, WW20]. Components [ADJ18, Bay03]. Composite [AMR^{+}23, Akt01, AK02, AT05b, BKOS11, BSSK12, BE99, ÇŞS01, ÇKAK04, EBY06, EY13, Ese01, KÖAG00, KE01, Oku11b, Oku11a, Oku11c, Oku11d, ÖG13, Özc00, SS06, ST22, SMB96, TÇ06, TMB96]. Composites [BS11, KA11b, PPRF22, STK07, ZJ10]. Composition [AFGV20]. Compositional [AMR^{+}23]. Compound [MR05, MVC23, ME12]. Comprehensive [NSG^{+}23, PRFP^{+}23]. Compressible [AM10, Has01, LH10b]. Compression [AB03, CD04, CC22, KAK96]. Comput [HLX^{+}18, KA16, KF19a]. Computation [ACC19, CK12, CP22, DEL23, GB23, GH11, GU05, KWP11, LB20, Loj22, MK97, MR19a, NCDM22, PSW24, ST22]. Computational [AMAGA22, Ars06, Çat10, DCAR21, EARE18, FSV^{+}22, FdSPW20, Flu15, GGT^{+}23, GNK19, LEVP21, LL17, MCBM23, Ole20, PQ23, Qua18, RKX23, RSF23, SGZ11, SGC23, SHE05]. Computations [ETW^{+}23, SKGP22]. Compute [CBM^{+}19]. Computer [Can96, QN18, Sam99, Sat23, SBF96, TE05]. Computerized [CPA20, KY96]. Computing [ASJQ24, BNSDC20, Ert11, KNHK20, SGE98]. Conbustion [SS96]. Concentrated [Özk01, Özy09]. Concentration [TEK11]. Concentric [GB18, Yür04]. Concept [Akh04]. Concomitant [AM23b]. Concrete [Bağ10, GEKB13, Guy20, KEA11, Özg10, ŞİH13, TA96]. Condition [MM17, Pur21, SK03, WSLZ17].

Conditions [Akh04, AMK96, BG18b, BCN22, JKS23, Ken96, KGCFK22, MCBM23, Mut04, ÖÖ05, PB01, SBF96, SM21, YG15, Yan16]. Conduction [MMS19, PŞ05]. Conductivity [PŞ06]. Conductors [TSC11]. Confidence [CMG22, WOL15]. Configuration [SUE13]. Confirmed [FSHGGB^{+}21]. Conformable [CMCM21, CMG22]. Conical [MÖ96]. Conjectured [PKC22]. Conjugate [MMS19, Umu11]. Connected [TGK96]. Connection [ÖS05, ÜU05]. Connectionist [AM09]. Connections [DSD^{+}21, KÖ00, MF21, Mor22, Qad24, YG02]. Conoids [Özy98]. Consensus [MPA17]. Consequences [Yur99]. Conservation [AM10, KSM23, Kar10, NK10, NNM10, PMS03, ZMA23]. Conservative [CML10]. Conserved [FMM10, NNM10]. Consideration [Mub23]. Considering [AJ12, ALFBPLE20, CATCdlF20, JLS^{+}10a, KYG20, SGCRRL^{+}23, TZ19, Tsa09, WOL15, YLX^{+}13]. Constant [Akb09, EA22, PS19, YB03]. Constants [Man16]. Constitutive [ORC19, UY11]. Constrained [BGRV21, BMV22, BCS20, CAQCMM22, GBİ14, IGSB18, SBF96, VVV23, WZT20, WED^{+}23, YBTH17, Yen05, ZBBZ18]. Constraint [CCL06, LOSCO^{+}20, YZBZ16]. Constraints [DSD^{+}21, MF21, SGCRRL^{+}23]. Construct [KSS19]. Constructing [Erp12]. Construction [GÖD03, Kus19, LCB18, OL21, RG23]. Consumption [AARW22, KAY08]. Consumption-Based [AARW22]. Contact [BE99, EB02, GB00, KBE01, STA23a, Xia10]. Containing [SUE13]. Contaminant [MAH10]. Contaminant-Modified [MAH10]. Contingent [SW10]. Continuation [BV19]. Continuity [Yan16]. Continuous [Bha24, CCGHdCJ21, GAAAS17, Gun96, JZL10, Liu11, ÖB17, PBY97, PB97, YG11]. Continuum [BP23, Bro19, CDA10, Sta23b]. Contracting [GGJ10, RSS^{+}21, RRO16a]. Contrast [CKLZ21]. Contrecoup [TCALK20]. Contribution [DB17a, DB17b]. Contributions [MC20b]. Control [AAGH21, Abi03, AERE18, AB19, Ara08, AMM23, BDC^{+}18, BS23, CST20, DSMP19, DKÇ97, ES06b, ES06a, EARE18, EÜA13, FYuR^{+}23, Gha18a, Gha18c, Guc04, Gul03, GOA03, HBFM22, JY23, JSTGV^{+}19, Kar03b, KD19, Ken11, LL23, LESEVP^{+}20, LEVP21, Mak99, OMX18, PD18, Sam99, TT11, TCA96, TB21, Yag02, YE03, Yaz16, YK96b, YS02, ZLLX20]. Controllability [CK18, DVFA23]. Controllable [KKC08]. Controlled [AERE18, AÖ09, ÇA10, WT15, YUKA03]. Controller [ASK21, BP99, Büy16, LOSCO^{+}20, TGK96]. Controllers [NSG^{+}23, PS22]. Controlling [CTL14, HJL10]. Controls [DVFA23, GOA03, IMWY18]. Convection [HFAAM^{+}22, IYKA20, MMS19, RKR^{+}22a, RKR^{+}22b, SJ22, TKK^{+}22]. Convective [DJH18, JS22b, JKS23]. Conventional [BP99]. Convergence [AD11, ANK23, Bad22, Guy20, JL24, Kay15, PKB09, RC16, SS03, Sav14, Sav16, ZWZ17]. Convergent [Sav99, Sav12, Sav13]. Converters [HCdCTAAMEJ20]. Convex [QM16, RC16]. Convexity [SAI04b, VC23b]. Convexity-Preserving [VC23b]. Conveying [AO17, ETL22, ZCC15]. Conveyors [YYWX16]. Convolution [CK07, CG09, DCA10]. Convolutional [FSHGGB^{+}21, LRBMAM^{+}23, MRIR20, MSPC23, SSA22b, ZPW^{+}22]. Conway [MVC23]. Cooling [OYCS11]. Cooperation [PAR23]. Cooperatives [CÖGD11]. Coordinate [ACC20]. Coordinates [DB17a, DB17b]. Copper [RRO16b, RRO16a]. Copper-Water [RRO16b, RRO16a]. Coproducts [Mut04]. Core [ÇA04, KNHK20]. Cores [DK03, KD03]. Coriolis [RKR^{+}22a]. Cork [dAPdSL^{+}22]. Corn [ZPW^{+}22]. Corneal [AK19]. Corner [HWA99, TEK11]. Coronavirus [FYuR^{+}23]. Correction [HMDG11, KA96b, KA16, KF19a, MLBR20, YGS13]. Corrective [ZLZ17b]. Correlation [DAD^{+}18, Erp12, KLRA10, VVN23, YB97]. Correspondence [YU07]. Corrugated [AVC23, CC22]. Cortex [RM20]. Cosine [KUÖ05, Kaz05, ÖK05, UG00]. Cost [BDC^{+}18, CKCC16, KYG20, NISC18, SGBR23, TWL10, UK96a]. Cost-Effective [BDC^{+}18, UK96a]. Costs [ZS16]. Count [SKM^{+}22]. Counterpart [Gun96]. Coup [TCALK20]. Coupled [BUAM19, BÖ07, Çat10, MM10a, MK10, NM10, PB97, PM22, Raw14, STA23a, WSLZ17, ZH10]. Couples [TÇ06]. Coupling [AKT04, AGKT21, BHN10, GKE05, HMDG11, MDNN10a, NM10]. Couplings [AT05a]. Course [Öze02]. Court [Gia21]. CoV [GPMRVM21, UYAA22]. Coverings [Kaz04, Muc00]. COVID [CRA^{+}21, FSHGGB^{+}21, GM21, MPA^{+}22, SSGCPR23, TRD^{+}22, SPCPMM21]. COVID-19 [CRA^{+}21, FSHGGB^{+}21, GM21, MPA^{+}22, SSGCPR23, TRD^{+}22, SPCPMM21]. Cp [Kaz04]. Cp-Coverings [Kaz04]. CPPN’s [VC23a]. Cr [OYCS11]. Cr-Ni [OYCS11]. Crack [AUA03, ÖD06]. Cracked [LSARZ23]. Cracks [LVBV19]. Cramér [Dil02]. Cramér-Rao [Dil02]. Crane [LESEVP^{+}20, PR21]. Cranes [LL23]. Credit [LL24, Tsa09]. Credits [HH07]. Criteria [AA16, AE16, AD11, Bor11, CACRF^{+}21, DVFA23, DW16, YY11]. Critical [İTİO08, Oku11d, Pur21]. Crop [GNI19, KPSS18]. Cropped [ÇŞÖG13]. Crops [KP12]. Cross [BT19, BY11, ÇDYS13, EL04a, KAEÖ13, Pin22, TDT^{+}11, TY11]. Cross-Docking [KAEÖ13]. Cross-Ply [BT19]. Cross-Section [ÇDYS13, EL04a]. Cross-Sectional [TDT^{+}11]. Cross-Wavelet [Pin22]. Crossed [Arv04, Mut00a]. Crossing [MCNM22]. Crown [ATK99, GGT^{+}23]. Crowned [ATSA97]. Cryogenic [Kan00]. Cryptographic [OV18]. Cryptography [BBSC18]. Crystals [FLYY21, Kas00]. CSLS [AM09]. CT [FSHGGB^{+}21]. CTL [FA22b]. Cu [RRO16b, RRO16a, JKS23, KNRC22, MAE97]. Cu-H [KNRC22]. Cu-Water [RRO16b, RRO16a]. Cubic [DS04, Dal11, EESK^{+}02, LC16, Li16, VC23b]. Cuckoo [PS22]. Cuda [ÇŞÖG13]. Cultural [GZ14]. Culture [LTYG17]. Cumulant [KLRA10]. Currency [Hon20]. Current [Akb09, BCK^{+}19, BBD^{+}18, BDH19, BODH20, BBPL19a, OD19]. Currents [Öze02]. Curvature [AK19, Bağ10, GE00, Har21, PQ23, TT11]. Curvatures [EG00]. Curve [BBSC18, GA10, LLTC17, OL21, PQ23, Tsa09]. Curved [AJ20, KA02, ÖD06, ÖÖT04, Sin10]. Curves [BY13, CT01, Çal06, KÇ96, Kız14, Li16, MCNM22, OFCP21, Özy11, Özy14, UK96b, UT96, VC23b, Yan16]. Curvilinear [AS04, SAI04a]. CuSn7 [MAE97]. Customer [Ji13]. Customers [BGG96]. Cut [EY13]. Cut-Outs [EY13]. Cutaneous [BDC^{+}18]. Cutting [Cui07, Kah11, SK03, SBF96]. CuZn30 [MAE97]. CXT [Wan13]. Cyberinfrastructure [LSB^{+}20]. Cycle [SS07]. Cycles [KLGP24, Özg10]. Cyclical [WW20]. Cyclonic [BCN22]. Cylinder [HB01, HFAAM^{+}22, IMWY18, JS22a, NKS22a, NKS22b]. Cylinders [Yür04]. Cylindric [UY03]. Cylindrical [Du 22, GS01, GS02, GA99a, Shi06, SMB96, ZHN10]. Cytomatrix [UC99]. CyVerse [LSB^{+}20].

D [Çat10, DB17a, DB17b, HBS^{+}17, OA09, Wan13]. D2 [BDÖ^{+}99]. Dam [DJ10]. Damage [BS11, DLL10]. Damped [Bad22, Dil02, MOBA21, YG11, ZMA23]. Damping [AGANA22, ANA20, CJD24, UAI21, WSLZ17]. Darboux [ED11, KÇ96, Özy14, UK96b, UT96]. Darcy [AGKT21, ETL22, KH22]. Data [AJ12, ARCRMM20, Alp02, AA13b, AMM23, BG02, CPA20, CPP19, DCAR21, DJY^{+}22, ETW^{+}23, GBİ14, HRM19, IMKEK18, IWH21, JLZ04, JLST04, JLS^{+}10a, JLS^{+}10b, JLKK12, Kat17, Kay16, KYG20, LL23, LF19, LNBM10, MMM23, MGGHCR^{+}20, MC20b, MVC23, MCNM22, NAC22, NAC23, NISC18, NYH^{+}17, PGB^{+}11, PMF20, Sal03, SKM^{+}22, SSGCPR23, Sul08, TGC22, ÜÖ03, ÜC11, UKY96, YB97, YY11]. Data-Driven [ARCRMM20, AMM23, LL23, LF19]. Dc [AÖ09, ÇA10, Gul03, HCdCTAAMEJ20, MPA17, RCTR20]. DC-DC [HCdCTAAMEJ20]. DEA [DJL^{+}05, KJZL12]. Deb [CGM^{+}23]. Deblurring [AM23a]. Decay [KSK05, KÜM08]. Decision [AHID18, ADS19, AN19, ASA20, ASAH20, Aks05b, Bor11, CMS20, CWF^{+}22, CCC08, CS19, DW16, Fu15, FbF16, Fu16, GLAA08, GÖV96, JLS^{+}10b, JZL10, LY15, Liu11, LNBM10, MECRMT^{+}21, PVCRRV^{+}21, SAZ18, WOL15, YY11, ZGV10]. Decision-Maker [WOL15]. Decision-Making [AHID18, AN19, ASA20, ASAH20, CWF^{+}22, DW16, Fu15, FbF16, JZL10, Liu11, PVCRRV^{+}21]. Decoding [CD04]. Decomposition [AJ20, ARCRMM20, AFN^{+}19, CAQCMM22, Doğ12, EAEAA19, NNW14, Öze10, RAA10, VMRVF^{+}23]. Deconvolution [ATZES^{+}07]. Decoy [CMB19]. Dedicated [SS96]. Deep [CC24, CLGGLB^{+}22, ETW^{+}23, LC18, MKK97, TOM^{+}22, ZPW^{+}22]. Defect [Har21]. Defection [UKY96]. Defined [Eİ06b, Faz21, Sav13]. Definite [RS21].

Definition [All99, CMCM21]. Deflection [Akt01, RFD^{+}10]. Deformable [DCA10, MAKH10]. Deformation [LC18]. Deformed [SBA^{+}04]. Degree [ZHN10]. Degrees [Guc04, Üna03]. Del [CC09]. Delay [CK18, DMN22, EAEAA19, GK12, GGS11, JFRZ10, KAGAM23, KLGP24, LXZ10, MOBA21, OA13, SKGP22, TWL10, YY07, YGS13]. Delayed [EA22, MR05, ZPL16, Zhu17]. Delaying [Unv07]. Delays [DVFA23, SD13]. Delta [KS18, Kle18]. Delta-Sigma [KS18, Kle18]. Demands [BGG96]. Dendrite [Oze03]. Dengue [GOOM18]. Densities [BK10, LV19]. Density [Alh21, AFJ22, Flu15, LZC22, LS20, OC09]. Dental [MAS^{+}11, ZA98]. Dependence [Dik06, Dil02, KÜM08, YG11]. Dependent [DSG19, KC08, PŞ05, Yaz16]. Depending [KA11b]. Depends [Özy98]. Deployment [BGG96]. Deposition [BP23, ZHN10]. Deregulated [UC10]. Derivation [KO15, PBY07]. Derivative [ALFBPLE20, BP21, CMCM21, CMG22, Dal11, GNI19, RLY^{+}17, SKB23, SBB23]. Derivative-Free [BP21, SKB23]. Derivatives [CC11, NK10, PBY07, PBY08]. Derived [CC11]. Descent [BP21]. Describes [AK19]. Describing [Ji13, SGCRRL^{+}23]. Description [TD12]. Design [AMR^{+}23, BG04, Bay03, BS23, BP99, ÇA06, Çun08, DPC^{+}10, DB20, GFPF22, GZ14, KKC08, KA11c, KAEÖ13, LV23, LL17, Mak99, ÖKK03, PS22, PLLRCC20, RMBTGH^{+}20, SGBR23, SÇ10, ŞİH13, VJ23, VC23a, Yaz16]. Design-Optimization [Bay03]. Designed [IA04, YDD11]. Designing [Sar09]. Designs [BPD08, Bac13, LCL17, LCB18]. Desirability [ŞİH13]. Detached [MM10c]. Details [AHVG23, BBPL19b]. Detection [ARN96, DK99, NAA20, SKAK05, WA01, ZdlÁGLTT22, ZP11, ZPW^{+}22]. Detectors [OOI^{+}22]. Determination [Akt01, AB06, AMK96, CBSTG13, CGCC11, DUM04, GEKB13, GA99b, HAK03, KE98, KA02, KP12, NLMU19, ŞÖK11, Üna96, ÜDMA04]. Determine [Boo17, FKRG10, GLAA08, YDD11]. Determining [BAUO09, CCWY10, HB99, HBFM22, UC10, YB03]. Deterministic [GN19, MPP^{+}24]. Deterministic/Stochastic [GN19]. Developable [Kus19]. Developed [TE05]. Developing [SZ23]. Development [AARW22, CGM^{+}23, Özd97, YDD11]. Developments [LB20, Loj22]. Device [ŞÖ11]. Devices [AÖ05, TGG96]. Diabetes [GVC20]. Diagnose [MCBM23]. Diagnosing [DT08]. Diagnosis [BLEB^{+}18, CTH^{+}22, CLGGLB^{+}22, Erp12, GAAAS17, LRBMAM^{+}23, YA14, ZP11]. Diagnostics [EGB99]. Diagnostics-a [EGB99]. Diaphragm [ZWZ16]. Dichotic [VÇÖ11]. Dielectric [HAK03, KIY^{+}23, TA99, YB03]. Diesel [SGEP06]. Difference [CMCM21, CMG22, CJD24, Egb18, EAT16, FIJ21, FJ21, GH11, GGS11, JWP23b, KYS13, LY15, OA13, PMR14, PALL18, SGZ10, SGZ11, SGA18]. Difference-Index [LY15]. Different [Alt09, CLY10, DNS23, KAK96, LV19, Oku11b, OFCP21, ÖÖ05, ÖS05, PB99, RRO16b, ŞT96, STK07, UC19, YB03]. Differentiable [TAS^{+}23]. Differential [AV02, AS09, AAD11, AADS11, AO16, AK18, Akh04, AHR05, BhMsG16, BG18b, BS23, Cai06, Can96, ÇKÇK10, CK18, CCGHdCJ21, CÜ04, DI13, DSD^{+}21, DVFA23, Doğ12, Dol04, EAEAA19, Ert07, EM08, Ert11, FKRG10, Gha18b, GK11a, GH11, GLS18, GY13, GÖD03, GGMMDI20, HNB20, ID14a, ID14b, IÇÜ12, JR18, KTK11, KSTJ18, KA11a, KLGP24, KO10, KKS11, KPSS18, KAAK10, KSS19, KO15, KYS13, LH10a, LLL14, LXZ10, LS20, MR19a, MG11a, OA13, Özy11, PAB11, PKO11, Raw14, SSA22a, SQA11, Sar11, Sha99, SKGP22, SZTY23, SW10, SKKC18, SKK19, TYS13, TY11, VA12, WN21, YÖS10, YG15, YY07, YGS13, dlF21]. Differential-Algebraic [LXZ10]. Differential-Difference [GH11]. Differentialalgebraic [GHN11]. Diffraction [KY96, TA99, UY03, Umu11, YD11]. Diffrence [EB00]. Diffusion [BBK20, CBM^{+}19, DSG19, DZA19, EAT16, GC23a, GZ20, IZS04, MM10a, RAA10, SGZ10, SGZ11, SGA18, TO13, TZ19, TÇ06, TDC08]. Diffusion-Reaction [BBK20]. Diffusive [NM10, RKR^{+}22a, Zhu17]. Digamma [TAS^{+}23]. Digital [BBSC18, ÖKK03, RWL23, RG23, VVN23]. Dilatometer [GAS11]. Dimension [JS18, Mat23, PS23]. Dimensional [AK19, Alt16, ATSA97, ATK99, BBD^{+}18, BB16, CUA04, CO16, EG00, EBY06, GGJ10, GE00, HL16, IZS04, JS22b, Kah11, KT98, KTC99, KA11a, KAEÖ13, Li10, LDL10, MAKH10, PRFP^{+}23, RSS^{+}21, SK18, qS13, SZTY23, SEZ24, TÇ04, TSS17, VRMJG19, WN21, YH05, ÇT09]. Dipole [YB97]. Direct [AO16, AB19, Gha18a, KO15, ENT05]. Directed [KA02, UÇ96]. Direction [Mol18, ZZ17]. Directional [BG99]. Dirichlet [Mat19a]. Disc [SS06]. Discharges [JFRZ10]. Discipline [CT14b]. Discontinuity [ZZ10]. Discontinuous [mLkWbPxX17, RC99]. Discovery [VSAMMM23]. Discrete [ABKU10, BP99, CK07, CG09, DSG19, DCA10, EARE18, FS23, FHRS18, GA10, ICDM21, JY23, Kha99, SKM^{+}22, Sol19]. Discrete-Time [DSG19, EARE18, JY23]. Discretionary [AJ12, JLS^{+}10a]. Discretization [CMG22]. Discriminating [JLKK12]. Discrimination [AG12]. Discussion [Çat10]. Disease [FSH^{+}23, HBS^{+}17, PM22, Wu17]. Diseased [Ars06]. Disk [Ers10, ID14b, KH15].

Disorder [EEJRLB^{+}22]. Dispatch [RC16, UC10]. Dispersion [KVR23, SAM^{+}22]. Dispersive [BB16]. Displacement [FKRG10]. Display [BVvdH^{+}23]. Disposability [KYG20]. Dissipation [MM10d]. Dissipative [AMAGNA23, Sta23b]. Dissolution [RKR^{+}22b]. Dissolution-Driven [RKR^{+}22b]. Distance [AAS11b, BVCS18, SÖS13]. Distillation [HCdCTAAMEJ20]. Distinction [NÇK14]. Distributed [AE03, DVFA23, Ese01, KE01, MGGHCR^{+}20, MOBA21]. Distribution [ATK99, BY06, CZC17, CKKA22, CKCC16, CK12, GB23, GGMO21, HC22, HJL10, IMKEK18, ICDM21, JAT^{+}21a, KHR^{+}22, KMCI22, MKMB11, MC20a, MC20b, MFF20, NVC22, NAC22, NAC23, Oui23, PR21, PRFH23, ŞT96, SÇGA12, SKM^{+}22, ST11, TGC21, TGC22, Yür02, ZS04]. Distributional [Kan16]. Distributions [AE16, AM23b, BSA^{+}23, CTJJ22, JHA^{+}22, KN18, KAKB08, Kil03, Kil04, MHC21, Pek96]. Distributor [DM21]. Disturbance [LOSCO^{+}20]. Diuretics [DL24]. Divergence [Guy20]. Diversity [CS19]. Divided [EB00]. DLA [Hei23]. Do [AARW22]. Docking [KAEÖ13]. DOF [LESEVP^{+}20]. Domain [AG12, BCGS23, MMS19, SK18, SWJS22, Yao15]. Domains [Mat19b]. Dombi [ADS19]. Dominance [BVvdH^{+}23, DE23]. Dominant [CBM^{+}19]. Dominated [CSSOB20, ZWZ16]. Domination [Tur16]. Doppler [ÜIG03]. Double [Ara08, KSM23, RKR^{+}22a, Sam99, Sav00, Sav10, Sav16, TKK^{+}22]. Double-Diffusive [RKR^{+}22a]. Doubly [AB19]. Down [Cha05]. Downward [YYWX16]. Drag [AHM99, IMWY18, KP12]. Drainage [CMS20]. Drawing [MKK97]. Drawings [ÇÇÇ07]. Drift [BHN10]. Drift-Flux [BHN10]. Drilling [SBF96]. Drillstrings [YAAK96]. Drip [YDD11]. Drive [AÖ09, ÇD03, Gul03]. Driven [ARCRMM20, AMM23, IÇÜ12, LCW^{+}22, LL23, LF19, RKR^{+}22b, SBDN^{+}22, Sul08]. Drives [BA03, ÇA10, GM03]. Driving [LGT^{+}18, PSF22]. Drop [MM10c]. Drug [BG19]. Drugs [Egb18]. Dtm [LYYX15]. Dtm-Bf [LYYX15]. DtN [BD22]. Dual [Çal06, GK11b, KUÖ05, Kaz05, KKL18, KÇ11, UG00]. Duct [EL04a, EI05, EI06a]. Ductile [ABO^{+}16]. Due [BS13a, DMN22, EY13, Ers10, Ese11, RKR^{+}22b, ENT05]. Duffing [ALK19, CCY08, CTL14]. Dufour [JS22a, KH22, KNRC22, NKS22b, RKS23]. Duplication [LP17]. Duplication-Based [LP17]. Dura [HFATZR^{+}20]. Durable [Ji13]. During [JB15, BY17, OYCS11]. Dust [AB99]. Dynamic [AE03, Aks05a, ACC19, AK02, AA13b, BCN22, Con22, DJ10, EÇ99, Ese11, FBR17, GS10, GLA05, HFATZR^{+}20, KD19, MÖ96, OC05, PS22, SMB96]. Dynamical [Ara08, Bİ99, Sol19]. Dynamics [AÖ05, ADJ18, AÇGJ99, BIM^{+}20, DSG19, FJ21, FA22a, FPCA18, GCGZ10, GK12, KKA04, LL17, MCBM23, Ole20, OWO^{+}23, PM22, Saj18, SKB23, SGC23, ŚLO19, TRD^{+}22, UC99, ZBBZ18, Zhu17]. Dynamics-Based [SGC23].

E2 [BUKM96]. E2/M1 [BUKM96]. E3 [KYK02, Kız14]. Early [CLGGLB^{+}22, ŚLO19]. Earth [KODM11, VPJ22]. Earthquake [DLL10, EBD11]. Earthwork [GLAA08]. East [FYuR^{+}23]. Eastern [LDCST21]. Easy [CBSTG13]. Echo [Bha24]. Ecological [CMM19]. Ecology [WLW^{+}10]. Economic [RC16, UC10]. Economics [ACC19]. Ecosystem [HXY15]. Eddy [BCK^{+}19, BDH19, BODH20, OD19, TCY06]. Eddy-Current [BDH19, BODH20, OD19]. Edge [UY03]. Edged [Che18]. Edges [YD11]. Editing [NYH^{+}17]. Education [KA09]. EEEG [VÇÖ11]. EEG [EEJRLB^{+}22, SKAK05]. EFEF [AB06]. EFEF/VFVF [AB06]. Effect [ANJ20, Alt10, Alt12, BA03, DZA19, DL24, EL04a, GGT^{+}23, HTLY18, HK23, HFAAM^{+}22, JS22a, JKS23, Kop02, KSK01, Kur98, LZL^{+}23, MAS^{+}11, NKS22b, OSAAA18, PVCRRV^{+}21, RKR^{+}22a, SSV23, SJ22, Tay04, TCY06]. Effective [BDC^{+}18, RC99, UK96a, ZLZ^{+}17a]. Effectiveness [MV21]. Effects [BS11, EI05, EI06a, GK12, GÖA96, Has03b, IYKA20, JFRZ10, KH22, Kir04, KNRC22, KKA04, Oku11d, RRO16b, RKS23, SSV23, ST11, TWL10]. Efficacies [BG19]. Efficacy [FdSPW20]. Efficiency [AB99, CÖGD11, DL23, GBİ14, GPEG10, JLZ04, Kan00, KYG20]. Efficient [AK19, Ala11, BIM^{+}20, BBK20, CMS20, CJL18, EE22, HM10, JLS^{+}05, JLS^{+}10b, JLKK12, KF19a, KF19b, OV18, SKB22, SKB23, YQG22]. Eguation [BA11]. Eigenelements [Ert11]. Eigenvalue [NU19]. Eigenvalues [MR19a, NLMU19]. Eight [Guc04]. Elastic [AMAGA22, BPC^{+}23, BKOS11, BSSK12, BE99, BÇE02, ÇŞS01, ÇKAK04, ÇÇ03, CÜ04, FSH^{+}23, FLYY21, GFPF22, KTK11, KBE01, KÖAG00, KPSZ24, LSARZ23, MOBA21, Mub23, NSCTPW22, Özb02, Özc00, TB16, UC19, XS22, Xu10]. Elastic-Plastic [BKOS11, ÇŞS01, ÇKAK04, KÖAG00, Özb02, Özc00]. Elasticity [Akt01, ACC19, BA03, KJZL12, LVBV19, Mat19b, Raj10]. Elasto [ABO^{+}16, Ese01, KE01, Oku11a, ORC19, ÖG13, SS06, Sha19]. Elasto-Plastic [ABO^{+}16, Ese01, KE01, Oku11a, ÖG13, SS06]. Elasto-Viscoplasticity [ORC19]. Elasto-viscous [Sha19]. Elastodiffusive [TZ19]. Elastodynamics [DJM18]. Elastoplastic [GB98, GB00]. Electric [Amo21, Çun08, Has02, IGSB18, KVR23, Üna96]. Electrical [DB20, Ken11]. Electro [Kur98]. Electro-Optical [Kur98].

Electrode [Kir04]. Electrokinetic

[ZHN10]. Electromagnetic

[AT05a, CJL18, VRMJG19, YU07]. Electron [AÖ05, CBT21, Kas00, Yil11]. Electronic [SSA22b]. Electrostatic [GNK19]. Element [ATK99, BT19, BCK^{+}19, CTAA11, DCC23, Dar11, DD11, DJM18, Ese11, Fan20, FdSPW20, GC23a, GC23b, GKE05, JLW20, KEA11, KODM11, KTS13, Mol18, NSCTPW22, Öz00a, PCCT11, TA96, VVV23, WFV21, YAAK96, YRB16]. Element-Based [Fan20]. Elements [Hak22, KSK05, KÜM08, OF19, ŞT96, SGCRRL^{+}23]. Elevated [EAT05]. Elicitability [JAT21b]. Elicitation [PLW21]. Elimination [Egb18, UKY96]. Elliptic [BBSC18, CKLZ21, ÖG13, ÖT02, Wu10]. Elliptical [DJH18]. Embedded [CD04, ÇKÇK10, DSI16, LSARZ23]. Emden [Ert07, HM10, MK10]. Emenet [GB98]. Emergence [GS10]. Emergency [ACGC21, CWF^{+}22, Pur21]. Emerging [JAT21b]. EMF [VRMJG19]. EMG [AAR24]. Emission [NCM19, SGEP06]. Emissions [AARW22]. Emitter [YDD11]. EMO [CGM^{+}23, SW22]. Empirical [FHRS18, Kay16, VHMMG20]. Employee [LTYG17]. Enclosed [GM21]. Encoding [CD04, VC23a]. Encryption [AKHHM17, XWCF19]. End [KA02, ÖÖ05, PMR14]. Endoscope [HAMM10]. Ends [KE01]. Energies [AK05, GGJ99]. Energy [Ala11, Alh21, ALK19, AARW22, BA03, GS01, KAY08, Koi19, LP17, Man04, MS05, Pek96, SY11, ŞÖ11, SJ22, TKK^{+}22, VHMMG20]. Energy-Aware [LP17]. Energy-Momentum [ALK19]. Energy-Transport [MS05]. Engine [SGEP06, SS07, SY11]. Engineering [LEVP21, ÖE03]. Engineers [Kat17, PS19]. Engines [SS96]. Enhanced [ACM19, BWK23, HDAA20, TCC^{+}23].

Enhancement [DJH18]. Enhancing [DSU22, RWL23]. Enlarged [Şen05]. Enough [CPP19]. Ensemble [MRIR20, PGB^{+}11]. ENSO [DJY^{+}22, HIVH21, IWH21]. Entering [CT14b]. Enterprises [LTYG17]. Entropy [AMR^{+}23, Fu16, JKS23, KH22, VÇÖ11]. Envelopment [AJ12, Alp02, BG02, GBİ14, JLZ04, JLST04, JLS^{+}10a, JLS^{+}10b, Kat17, KYG20, YY11]. Enviromental [Yur99]. Environment [AN19, HAH^{+}20, SAZ18]. Environmental [RMBTGH^{+}20, TCM21]. Environments [JTM^{+}22]. Enzyme [BG18a]. Epidemic [AERE18, NMCM19, SPCPMM21, YA23]. Epileptic [SKAK05]. Epistemic [YQG22]. Epitaxial [Bül99]. EPQ [CBSTG13, CC11, CCL06, CCC08, CCWY10, CGCC11]. Epsilon [CSSOB20, DJL^{+}05]. Equation [AV02, AO16, AEB01, AA03, AK19, Akh04, AGANA22, AMAGNA23, AD11, AE02, ALK19, Cai06, CCY08, CBM^{+}19, CK18, CJD24, CRA^{+}21, DS04, Dal11, Den16, EAEAA19, EB02, GN19, GB00, HM10, HSG^{+}20, HL14, HLL15, HL16, ID14b, İB00, IA04, IZS04, JM03, KSM23, Li10, LDL10, LLL14, MDNN10a, MJK12, NSO^{+}21, ÖT02, PY00, Pak02, PŞ05, PŞ06, PT21, PKO11, PMS03, RAA10, SK18, SGZ10, Sar11, SGZ11, SGA18, qS13, SJ19, SW10, TO13, TD12, Wan13, WZ23, YG11, YW24, ZMA23]. Equations [AS09, AAD11, AADS11, AYG14, ACM19, AM05, All12, Alt16, AD04, AHR05, AWK^{+}20, BY16, BK99, BG18a, BG18b, BS15, BMT^{+}21, Can96, Can02, CCGHdCJ21, Civ96, DI13, DSD^{+}21, DB16, Doğ12, Dol04, ESNB19, EB06, EAT16, Ert07, FBR17, Gha18b, GHN11, GC23a, GK11a, GH11, GLS18, GY13, GGS11, HBN96, HPA13, HNB20, HM23, IÇÜ12, JR18, Kar03a, Kar10, KLGP24, KO10, KKS11, Kha03, KAAK10, KPS04, KSS19, KO15, KYS13, LH10a, LXZ10, LS20, MG11b, ME12, MDNN10b, OFCP21, ORC19, Öze10, PAB11, SSA22a, SQA11, Sal06b, ST18, Sha99, SKB22, SKGP22, SZTY23, TYS13, UY11, VA12, WN21, YÖS10, YAA10, YA11, Yal13, YG15, YY07, YGS13, Zhu10]. Equilibrium [Koi19]. Equities [JAT21b]. Equity [Sho22]. Equivalence [ÇA04, PPB^{+}20]. Equivalent [AMMMM^{+}23, BDH19, KIY^{+}23]. Erbium [YU07]. Ergodic [Aki11]. Erlangian [Sul07]. Erosion [ZdlÁGLTT22]. Erratum [SGZ11]. Error [CCY08, CA19, HB99, Har21, How22, OA13]. ESP [VJ23]. Essentiality [DDPK21]. Established [BDMS19]. Estimate [Ara08, GM03]. Estimates [OA13]. Estimating [AAR24, BM21, Kan16, OM20]. Estimation [AMMMM^{+}23, CMM19, CK12, DÇ06, Gha18b, HJL10, ICDM21, JLZ04, Kan16, KN18, KA11c, LVHL22, LS20, Mes96, MWM22, OC09, PYR22, PR21, PRFH23, SRLEPVP20, Sul08, TOM^{+}22, UÖKE14, UKY96, WCH23, Wu17]. Estimator [Kar12]. Estimators [CZC17]. ETEM [DB16]. Ethics [CMG22]. Ethiopia [MWM22]. Euclidean [BY13]. Euler [Alt10, CMCM21, GK11b, İBB98, LSARZ23, Özd97, RC99, TZ19, TB16, UYAA22]. Eulerian [Shr98, Shr05]. Europe [KNHK20]. European [TSS17]. Evaluation [Bak11, Car21, CLGGLB^{+}22, DL23, EBD11, FdSPW20, IGSB18, KYG20, NCM19, STK07, Shr98, Shr05, UY03, YE03]. Evaporation [JMB22]. Even [KMÇ^{+}05, MGT10]. Even-Even [KMÇ^{+}05, MGT10]. Event [JY23, MPA17]. Event-Triggered [JY23]. Events [Pin22]. Evidence [HBS^{+}17, WW20]. Evn [GE00]. Evolution [AK18, BS23, DLL10, FTGSNV21, GGMO21, GGMMDI20, HXY15, KSTJ18, Kar03a, KPSS18, MQ18, MG11b, NSG^{+}23, Sol19, SKKC18, SKK19, TY11, UC99, Zhu10, dlF21]. Evolutionary [AMMMM^{+}23, AMM23, CACRF^{+}21, CSSOB20, CS22, CS19, Gia21, HDAA20, QCdlFL^{+}23, QRdlFS22, SGCRRL^{+}23, SSGCPR23, VHMMG20]. Evospace [JSTGV^{+}19]. EW [DS04]. Exact [BBD^{+}18, CMG22, DB16, HAMM10, HL14, IÇÜ12, Li10, MQ18, MR19a, MJK12, NCM19, qS13, SUE13, VRMJG19]. Examination [CC24, ŞT96]. Examinations [GÖV96]. Examining [CY05]. Exascale [KNHK20]. Exceedances [CDDM23]. Exchanger [ZMCL24]. Excitable [BUAM19]. Excitation [BBD^{+}18, FKRG10, ENT05, ŚLO19, TMK21]. Excited [EESK^{+}02]. Execution [ZE16]. Exemplification [SM21]. Exergy [SY11]. Existence [BG18b, FBR17, Faz21, RAA10, SAI04b, WZT20]. Existing [ÇA06, TCA96]. Exp [MG11b]. Exp-Function [MG11b]. Expanding [GGJ10, RSS^{+}21, RRO16a]. Expanding/Contracting [RSS^{+}21]. Expansion [BIM^{+}20, Boo17, CCY08, GC23b, HH22, ME12, Zhu10]. Expansions [ABKU10, ACC20]. Expediting [CCC08]. Experiencing [ETL22]. Experimental [CC22, Oku11b, RFQCMMCR23, ST11, STA23a, ÜDMA04]. Experimetntal [ARN96]. Expert [BG04, SÇ10, Yur99]. Explicit [DSI16, DRKW21, NSO^{+}21, SK18, WZ23]. Exploration [ETL22]. Explore [HCGFOON20]. Exploring [KPTM20, PRFP^{+}23]. Explosion [WH10]. Explosion-Proof [WH10]. Exponent [AMAGNA23]. Exponential [ACC20, CKKA22, CTJJ22, FbF16, JHA^{+}22, KKY05, RRR^{+}22, Sol19, SJKCD20]. Exponential-Logarithmic [CTJJ22]. Exponentially [EAT16]. Exponentials [RS21]. Exponents [KAGAM23, ŚLO19]. Expressed [Pin22, RS21]. Expression [GK11b, Jab18, MKMB11, SC19]. Expressions [TB21]. EXT [Mat10b]. Extend [GLS18]. Extended [Civ96, KPTM20, SAZ18, TDCT11, WZ23, Zhu10].

Extending [OV20]. Extension [Hak22]. Extensions [Bİ99, DA19]. Exterior [Mat19b]. External [BBPL19a]. Extracting [ÇÇÇ07]. Extraction [OPUÖ11]. Extrahepatic [FA22b]. Extrapolation [AV04]. Extreme [AMR^{+}23, CDDM23, JLS^{+}10b, Pin22]. Eye [Fu15, Sat23]. Eyring [DSI^{+}18, HPA13].

F [GW97, KF19a]. F-Monotone [GW97]. Fabrication [CKCC16, CLC16]. Fabrication-Distribution [CKCC16]. Face [MSPC23]. Faced [CGM^{+}23]. Faces [Umu11]. Facility [Bor11, EKAÖ11]. Factor [AA01, CÖGD11, DE04, GLAA08, KA96b, LLTC17, MLBR20, WZ23]. Factored [CYAK03]. Factorial [KKC08, ŞİH13]. Factorisation [OV19]. Factorizable [JM03]. Factorize [OV20]. Factors [AJ12, AUA03, CCC^{+}18, JLS^{+}10a, MCB19, SK18]. Factory [AA13a, EKAÖ11]. Failure [CCC08, Oku11b, ST11, ST22]. Failures [All02]. Fall [BNSDC20]. Family [Aks08, Asl18, BSA^{+}23, CTJJ22, JHA^{+}22, KAM20, Özy98, Saj18, SKB22, Sol19, SJKCD20]. Fanno [AM10]. Far [KY96]. Far-Field [KY96]. Farm [GÇÖ12]. Fast [DMN22, JMS18, MK97, OV19, YRB16]. Fault [CTH^{+}22, KA09, LESEVP^{+}20, PS22, Sam99, YA14, ZP11]. Fault-Tolerance [Sam99]. Fault-Tolerant [LESEVP^{+}20, PS22]. Faulth [EGB99]. Faults [ARN96]. Faulty [All99]. FCC [ÇÇ03]. FDTD [YB03]. FE [ETW^{+}23]. Feasibility [TKGS20]. Feasible [ZWZ17]. Feature [BG04, CTH^{+}22, DCAR21, Eli16, HBS^{+}17, RSF23, ŞÜ13]. Feature-Based [Eli16]. Features [AG12, CTJJ22]. Featuring [KLGP24]. FeB [CTAA11]. Fed [AB19]. Feed [HTLY18]. Feedback [ES06b, ES06a, JY23, PD18]. Fem [MAS^{+}11, DNS23, SS06, UC19]. Fema [EBD11]. Fema-440 [EBD11]. Fermat [TY15]. Fermion [AKT04]. Fever [GOOM18]. FFT [ÜIG03]. Fiber [Oku11a, Oku11c]. Fibers [Oku11b]. Fibonacci [BY16, KYS13]. Fibre [ÇŞS01]. Fibrous [BP23]. Fictitious [BCGS23]. Field [AOL19, Amo21, AR23, Ars07, BY17, CGMCCA23, Has02, Has03a, HAMM10, HFAAM^{+}22, JLW20, KY96, KBH06, KVR23, KKA04, MAKH10, MMF18, MWM16, RKR^{+}22b, SWJS22, VRMJG19]. Fields [HBN96, JS18, Mol18]. Fifth [EM08, PBY08]. Fifth-Order [EM08]. Figure [ÇT09]. Fill [KODM11]. Fillet [GAÖ96]. Filling [BA99]. Film [MM10b]. Filter [ÖKK03, YL16]. Filtering [Ala11]. Filters [Haj04, WJ23]. Filtration [Pak02]. Fin [OSAAA18, PŞ06]. Financial [AARW22, CMG22, HC22, PC23, Sho22]. Find [AA03]. Finder [BBMA22, HK23]. Finding [PBY07, PBY08, PQ23, PKC22]. Fingerprint [BG99]. Finite [AS04, ATK99, BT19, BBK20, CO16, CMCM21, CMG22, CJD24, CTAA11, DCC23, Dar11, DD11, DJM18, Egb18, EAT16, Ese11, Fan20, FIJ21, FJ21, FdSPW20, GC23a, GC23b, GKE05, Hak22, IÇÜ12, JY23, JS18, JLW20, JWP23b, KEA11, KODM11, KTS13, KF19a, KF19b, Mol18, NCDM22, NMCM19, NSCTPW22, OF19, Öz00a, PCCT11, PS20, PMR14, PALL18, ŞT96, SGZ10, SGZ11, SGA18, TA96, VVV23, WFV21, YAAK96, ZWZ17]. Finite-Time [JY23]. Fire [OFCP21]. Fires [PGRL21]. Firing [Akb09]. Firms [MCB19]. First [Alt16, Den16, KSK05, KÜM08, MK10, NM10, PAB11, SPD18, TYS13, TY15, Üna03, YG15]. Fisher [BA11, RZ18, Sar11, WZ23]. Fisher/KPP [RZ18]. Fit [SK05]. Fitler [UKY96]. Fitted [DSI16, DRKW21, EAT16]. Five [AO16, MWM22, NUM^{+}19, TWS17]. Five-Layer [NUM^{+}19]. Five-Point [TWS17]. FIWARE [GERV19]. Fixed [AERE18, ANK23, Büy16, Che18, HB01, KAM20, Pop98, TK04, dlF21]. Fixed-Rigid [HB01]. Fixing [FIJ21]. Fizzle [CRA^{+}21].

Flat [FJ21, MMS19]. Flexibility [Kat17]. Flexible [MCNM22, SKK19]. Flierl [MDNN10a]. Fling [Xia10]. Floating [MV21]. Floating-Point [MV21]. Flow [AGKT21, AO17, AHM99, ADE99, AM05, AM10, Ars06, Ars07, BCN22, ÇSB10, Cha05, DJH18, DSI^{+}18, DE04, DL23, EL04a, Ers10, GGJ10, Has03a, HM23, IE04, JS22a, JS22b, JWP23b, JKS23, KA11a, KH15, KH22, KP12, Kop02, KNRC22, LYYX15, MCBM23, MAKH10, MWM16, MMS19, MAH10, MM10b, MM10c, Mos18, MM10d, NKS22a, NKS22b, NKI19, PRT^{+}19, RSS^{+}21, RKS23, SSV23, SJ22, TKK^{+}22, TGS98, YLX^{+}13, Yür04, ZS04]. Flow-Induced [MCBM23]. Flower [DPC09, PS22]. Flowing [SAI04a, SUE13]. Flows [BAUO09, BGA10, EL04b, ELT21, ETL22, FJ21, KG10, Kız14, TCY06]. Fluctuations [Hei23]. Fluent [BAUO09]. Fluid [ADE99, BIM^{+}20, DSU22, DSI^{+}18, Ers10, FJ21, HPA13, HWA99, IYKA20, KH15, LL17, MCBM23, MAH10, RRR^{+}22, Rod20, Sha19, SGC23, SUE13, SWJS22, STA23a, Xia10, Yür04, ZCC15]. Fluid-Structure [DSU22]. Fluids [AYG14, KVR23, LV19, Yür02]. Flux [BHN10, DEL23, GS02, JS22b, PFR19, TZ19, WFV21]. Flying [LH10b]. Focus [GVC20, ZMCL24]. Fokker [AB23]. Fold [BY11]. Food [KC08]. Foot [Wu17]. Forbidden [KSK05, KÜM08]. Force [AHM99, Dan22, EA22, KA02, Özc00]. Forced [MM10d, ÖÖT04, PKB11]. Forced-Free [MM10d]. Forces [KP12]. Forchheimer [AGKT21, ETL22, KH22]. Forecasting [KAY08, PC23, ŞE17, WW20]. Forecasts [GP21]. Forest [PGRL21]. Form [Den16, PRFP^{+}23, PKO11]. Formability [EAT05]. Formal [Özd96b]. Formation [AR23, ABH^{+}22, CGM^{+}23, GZ20]. Forming [EAT05]. Forms [All17, GGJ99]. Formula [HLX^{+}17, HLX^{+}18, Sho22]. Formulas [AA03, CSOPM24, KKY05].

Formulated [Sta23b]. Formulation [GB98, IZS04, KAS19a, UC19]. Forward [Tay04]. Foundation [BÇE02, KTK11, KPSZ24, Mub23, TB16]. Foundations [Bro19]. Four [Rob22, TWS17]. Four-Parameter [Rob22]. Four-Point [TWS17]. Fourier [BDH19, CC24, DEL23, GA10, JWP23a, PRFP^{+}23]. Fourth [KAAK10, PBY07, SKB23]. Fourth-Order [KAAK10, SKB23]. FP16 [dlF21]. Fractal [PTLZ18, SA16, VC23b, Yav22]. Fractals [PS23]. Fraction [Tan08]. Fractional [AM23a, ALFBPLE20, Alh21, ALK19, BG18b, CMCM21, CMG22, CJD24, Dal11, DVFA23, Ert11, ID14a, ID14b, KKS11, LH10a, MG11a, OWO^{+}23, PT21, RAK22, RRR^{+}22, RMBTGH^{+}20, SGA18, TCC^{+}23, TAS^{+}23, UYAA22, VA12]. Fractional- [Dal11]. Fractional-Order [AM23a, Alh21, TCC^{+}23, UYAA22]. Fracture [JLW20, Kir04, TKCC19]. Fractures [GGT^{+}23]. Fragmentation [AÇGJ99]. Frame [BO10, DD11, Kız14]. Frames [OC05, ÖS05, ÖÖT04, Özy09, ŞÖK11]. Framework [AMM23, CMS20, GCPV18, LV23, MGT10, Mit21a, Sta23b]. France [SPCPMM21]. Fredholm [AEB01, Alt16, BS15, EB02, HNB20, ST18, YAA10, YA11, Yal13, YGS13]. Free [AERE18, AJ20, AK05, BP21, BBPL19a, BÖ07, Çat10, CÜ04, CK07, CG09, CDA10, DCA10, JLW20, KTK11, KA02, KPSZ24, MMS19, MM10d, Mut00a, SKB23, YK96a, Yil11, Fan20]. Freedom [Guc04, Üna03]. Freeness [Mut04]. Freezing [Özg10]. Frenet [KÇ96, UK96b, Yay00]. Frequencies [DNS23, HAK03, Öz00a, ÖÖ05]. Frequency [ADJ18, BT19, BCK^{+}19, CLC16, GAAAS17, NCDM22, ZZ10]. Frequency-Amplitude [ZZ10]. Friction [DE04, KPS04, KTS13, ÜDMA04].

Friction-Factor [DE04]. Frictional [AGANA22, GB00]. Frictionless [BE99, KBE01]. Friendliness [UC99]. Front [CS22, FIJ21]. Front-Fixing [FIJ21]. Fronts [LV19]. FRP [NCDM22]. Fruits [DPC09]. FSI [BCGS23]. ft [SBA^{+}04]. Fuel [SS07, SY11]. Fueled [KK97]. Full [ŞİH13]. Function [ASJQ24, Akt01, BGG96, BP21, BB16, Eİ06b, ENVPF19, EA22, FA22a, Gun96, GZ14, HNB20, How22, IE04, Kar12, MKMB11, MG11b, ÖB17, Pin22, QM16, RS21, RZ18, Saj18, SPD18, Shr98, Shr01, ŞİH13, Sol23, UAT19, YBTH17, Yen05, YB97]. Functional [Akh04, DSD^{+}21, Flu15, Gha18b, Gha18c, GZ14, HMDG11, KAKB08, MF21, Mor22, PD19, ZPL16]. Functionally [AR23, BT19, ÇDYS13, KPSZ24, MLBR20, MRG23]. Functionals [ABKU10, FS23, Flu15]. Functions [ASK21, DA19, EE22, FLYY21, KB03, KRD^{+}22, Kay98, KA02, KPSZ24, LGT^{+}18, MMM23, Mus16, NÇK14, PRFP^{+}23, QN18, Qua18, RC99, RS21, TAS^{+}23, Wu10]. Functions-Based [MMM23]. Fundamental [UKÖ05]. Funnel [GÖD03]. Furnaces [PFR19]. Furniture [TOM^{+}22]. Fusion [HBS^{+}17]. Fuss [QSC19]. Futures [WW20, YS17]. Fuzzy [AV02, AV04, AAD11, AADS11, Abi03, AHID18, ADS19, AN19, ASA20, ASAH20, AAS11a, AAS11b, AK11, BK09, BG18a, Bor11, ÇGK96, Çun08, ÇA10, DW16, Emi03, EPLGGS^{+}21, FBR17, FbF16, GDE12, GLAA08, GU05, HCdCTAAMEJ20, JR18, KUI18, KHL06, Ken11, KA96b, KA11c, KTD12, LY15, Özg10, PS22, PPRVC20, SAZ18, Sav99, Sav00, Sav10, Sav12, Sav13, Sav14, Sav16, SA19, TGK96, YL16]. Fuzzy-Based [PPRVC20]. FW [ESNB19]. FW-H [ESNB19].

G [KA16]. Ga [Bül99]. Gain

[RMBTGH^{+}20]. Galerkin [KAS19a, ST18]. Galilean [ED11]. Galileo [KSK01]. Galois [KCMLTR18]. Game [GM17]. Games [ŞE17]. Gamma [DA19, HBFM22]. Gamow [BSK05, SBA^{+}04]. GAP [Mut98]. Gaps [RSS^{+}21]. Gas [GS10, KR23, SS96]. Gas-Solid [KR23]. Gasdynamic [GS01]. Gaseous [KK97]. Gasification [RKX23]. Gasket [Hei23]. Gate [BAS96]. Gauge [VVN23]. Gaussian [ATZES^{+}07, Bac13, Har21]. GCD [BS02, NT04, TB04]. GCD-Closed [NT04]. GCD-Reciprocal [NT04]. Ge [TM10]. Gear [ADJ18, GÖA96]. Gearbox [ARN96]. Gearboxes [CTH^{+}22]. Gene [AAGH21, DDPK21, TMK21]. GenEO [BD22]. General [AKT04, AS00, BCS20, HL14, HLL15, LZ17, LZ21, Mak99, NSG^{+}23, PBY97, Sal02a, Sal02b, SAI04b, Sal06a, UAI21]. Generalised [AFN^{+}19, KVR23]. Generalization [Özd96b, QSC19]. Generalized [ATZES^{+}07, Alh21, AD04, BKA08, BK10, BSA^{+}23, Bil02, Cai07, CC09, DE23, EBY06, GW97, HC22, HL16, Hus01, JHA^{+}22, JWP23b, KO10, KKS11, PB97, PS20, Saj18, SAK01, SI04, Sar11, Sav16, SGA18, Sha99, SA97, SAH98, WZT20, YH05]. Generate [LGT^{+}18]. Generated [EL04b, MHC21, TUK19]. Generating [CUA04, Cui07, GC23b, Saj18, Shr01]. Generation [ETW^{+}23, FS23, IE10, JKS23, LDCST21, Sal06b, SÇGA12, VC23a]. Generative [VJ23, VC23a]. Generator [AB19, HH22, TGK96]. Generators [Kar03a]. GENESIS [Yil11]. Genetic [BGN13, CAQCMM22, ÇA06, Çun08, JSTGV^{+}19, KAY08, KD03, LCB18, LGT^{+}18, MFF20, TMK21, UKÇ^{+}09, Yen05, ZWZ16]. Geo [ACGC21]. Geo-Indistinguishability [ACGC21]. Geodesic [AA13b, Mor22, ÖK05, UKÖ05]. Geodesics [KÖ00]. Geodesy [RM20]. Geometric [CY05, Çiç07, GT22, Mus16, Oku11d, QM16, SBF96, TUK19, Yan16]. Geometrical [DPC09]. Geometrically [BV19]. Geometry [BhMsG16, KÖ00, Koc00, Özy11, RM20, SA16, Şen05, ZJ10]. Geophysical [KGCFK22]. Gerber [EA22]. Germany [SPCPMM21]. Getting [CGM^{+}23]. Ginzburg [İB00, qS13]. GIS [AB03]. GLAD [Yil11]. Glial [PM22]. Global [BG19, BC97, Eli16, HB99, HDAA20, KNHK20, Nur03, OD19, SH18, YA23, ZG17]. Globally [BUAM19]. Glycocalyx [Ars07]. GM [MM17]. GMRES [YRB16]. GNPs [PPRF22]. Go [NKS22a]. Go-MoS [NKS22a]. Goal [GU05, KA11c, MR19b]. Gold [GP21]. Golden [OV18]. Goodness [SK05]. Goods [Ji13]. Gordon [ZMA23]. Goursat [JM03]. Grade [AYG14, ADE99, AM05, Ers10, Yür02, Yür04]. Graded [AR23, BT19, ÇDYS13, KPSZ24, MLBR20, MRG23]. Gradient [BWK23, FSV^{+}22, LVHL22, LVBV19, YKG20]. Gradient-Based [LVHL22]. Graft [AK11]. Graft-To-Vein [AK11]. Grafting [GC23b]. Graph [AHK18, ÇGV02, Hei23, MPA^{+}22, Üna96]. Graphical [Alt09]. Graphs [AHID18, ADS19, ASAH20, Che18, HAH^{+}20, SAZ18, Tur16, YBYL21]. Gravitating [GS01, GS10]. Gravitation [Shi06]. Gravity [CC24]. Great [PSW24]. Green [Akh04, IE04, RZ18]. Grey [BNSDC20, Fu15, Fu16, SH18, YL16]. Grid [FS23, NISC18, Sal06b, SPCPMM21]. Grids [LDCST21]. Gross [GN19]. Ground [GÇŞT07]. Group [ASA20, AB96b, BPD08, CWF^{+}22, DW16, MBA11, MM10a, Mut00b, Pak02, SM99, WOL15]. Grouping [CAQCMM22, RFQCMMCR23]. Groupoid [Muc01]. Groups [LCL17, MPA^{+}22, Mut04]. Grown [Bül99]. Growth [CÖGD11, FA22a, GÖV96, Ji13, Pek96, ZGV10, Zhu17]. GSVD [AKHHM17]. Guessing [Rob22]. GUI [KA09]. Guided [RWL23]. Gurson [ÖG03].

H [AS00, ESNB19, Shr98, ASK21, JY23, KNRC22, ENT05]. H-Function [Shr98]. H-Spaces [AS00]. Hadamard [SPD18, TAS^{+}23]. Hadamard-Type [TAS^{+}23]. Haghighi [NVC22]. Half [DZA19, HC22, Mub23, TKGS20]. Half-Logistic [HC22]. Half-Space [DZA19, Mub23, TKGS20]. Hall [MAE97]. Hall-Petch [MAE97]. Halo [OTŞ99]. Hamilton [Civ96]. Hamiltonian [All17, KWA^{+}10, Üna03, Xu10]. Hammerstein [EB06]. Hand [Sat23, Wu17]. Hand-Foot-Mouth [Wu17]. Hang [HLX^{+}18]. Hankel [SB02, STB04]. Hantavirus [TABT19]. Hard [ŞT96, Xia10]. Hard-Brazing [ŞT96]. Hardness [DUM04]. Hardware [PLLRCC20]. Harmonic [CÜ04, Hak22, SPD18, ŚLO19, TMK21]. Harmonics [MFF20, YUKA03]. Harnack [ÖT02]. Harvesting [PALL18, Yao15]. Hassell [ZPL16]. Hausdorff [BVCS18, CS22, Mat23]. Having [AS04, Dal11, ÖÖT04]. Hazard [Kar12, MCNM22]. Hazelnut [CÖGD11]. HBV [FA22b]. HDQ [CÜ04]. Head [BVvdH^{+}23, FP21, GPEG10]. Head-Mounted [BVvdH^{+}23]. Heading [BhMsG16]. Heads [PPB^{+}20]. Healing [TKCC19]. Health [DCC23]. Healthy [FSH^{+}23]. Healty [All02]. Heart [FSH^{+}23, HCGFOON20]. Heat [BS13a, DJH18, DSI^{+}18, DEL23, GS02, HA10, IZS04, Kop02, KNRC22, LYYX15, NSO^{+}21, OSAAA18, PŞ05, PŞ06, RRO16a, SAI04a, WFV21, ZMCL24]. Heated [HFAAM^{+}22, KZAN98]. Heating [HCdCTAAMEJ20, JKS23, NKS22a, RRR^{+}22]. Heats [AK05]. Heavy [Has03b, Özd96a, Özd97, TW18]. Heavy-Viscous [Has03b]. Hedging [GNI19, SW10, YS17]. Height [AÇ11]. Helical [ADJ18]. Helices [BY13, YH05]. Helicobacter [KAY09]. Helium [LZL^{+}23]. Helix [ZJ10]. Helmets [FdSPW20].

Helmholtz [BD22]. Help [CRA^{+}21]. Hemodynamics [AK11]. Hepatitis [OWO^{+}23]. Hermite [KK13, OL21, Shr01, Shr02, TAS^{+}23, YA11]. Hesitant [DW16]. Hestenes [AWK^{+}20]. Heterogeneous [BD22, Sul07, WN21]. Heuristic [Cha05, TDCT11, Ugu13]. Hidden [Ace19, AHVG23, AMM23]. Hiemenz [SSV23]. Hierarchical [DDPK21, GC23b, LV23]. High [AMR^{+}23, BCK^{+}19, CA19, CKLZ21, EAT05, FCGL19, GC23b, GGS11, Jia17, KRD^{+}22, KKS11, MO04, MMF18, PFR19, RMBTGH^{+}20, SGZ10, Sar11, SGZ11, SZTY23, TB21, ZZ17]. High-Dimensional [SZTY23]. High-Entropy-Alloy [AMR^{+}23]. High-Entropy-Alloy/B [AMR^{+}23]. High-Flux [PFR19]. High-Frequency [BCK^{+}19]. High-Gain [RMBTGH^{+}20]. High-Order [FCGL19, GGS11, KRD^{+}22, KKS11, SGZ10, Sar11, SGZ11]. High-Pressure [CA19]. High-Speed [MMF18]. High-Temperature [TB21]. Higher [BBK20, FSV^{+}22, JLW20, SBB23, SJ19, YÖS10]. Higher-Order [JLW20, SBB23]. Highway [EB00]. Highways [GLA05, GLAA08]. Hilbert [Pop98]. HIV [AAGH21, CST20, FA22a]. Holder [AD04]. Holditch [CT01]. Hole [AS04, Oku11b, Oku11a, Oku11d, ÖG13, SAI04a, TA96]. Holes [STK07]. Holographic [SA16]. Holomorphic [KSS19]. Homeomorphism [Bİ99]. Homogeneous [LCL17, MMF18]. Homogenisation [AVC23, BDMS19]. Homogenization [GFPF22, GNK19, KF19a, KF19b]. Homoglyph [MRIR20]. Homology [NAA20]. Homomorphic [XWCF19]. Homothety [KYK02]. Homotopic [HLL15]. Homotopy [Alt16, KG10, Mer09, OSAAA18, RFD^{+}10, TSS17]. Hopf [CTL14, KC08]. Hopping [ZBBZ18]. Hopscotch [NSO^{+}21]. Horticultural [TÇ04]. Hourly [DÇ06]. Household [HC22]. HTLV [Mer09]. HTLV-I [Mer09]. Hull [JS18, LL17]. Human [AK19, AAR24, ATSA97, CGMCCA23, CCC^{+}18, Mer09, Sar09, ZA98]. Hunter [KSM23]. Hybrid [AO16, AOL19, BODH20, Dar11, EE22, EY13, GZ14, HNB20, KN13, LOSCO^{+}20, NCDM22, NKS22b, PR21, PRFH23, RWL23, RKS23, SC19, SH18, SW22, WOL15]. Hydraulic [BLEB^{+}18, YYWX16]. Hydro [FJ21]. Hydrocyclone [BCN22, HK23]. Hydrodynamic [ESNB19, GS02]. Hydrodynamics [HTLY18]. Hydrogen [KR23, LZL^{+}23]. Hydrogenerator [BP99]. Hydromagnetic [KH22]. Hydroxyapatite [PCCT11]. Hygrothermal [MÖ96]. Hyper [BDMS19, FHRS18]. Hyper-Reduction [BDMS19, FHRS18]. Hyperbola [Özd97]. Hyperbolic [BB16, JM03, MQ18, ÖT02, ÖK05, PŞ05, UG00, UKÖ05, YG11]. Hyperchaos [CMG22]. Hyperchaotic [PPRVC20]. Hypercube [All99, All02, KA96a]. Hypergraphs [SA19]. Hyperlipidemia [DT08]. Hyperplanes [QM16]. Hyperspectral [ZPW^{+}22]. Hypersurfaces [Koi19]. Hypervolume [WED^{+}23]. Hz [KD03].

I-Convergent [Sav13]. IBA [MGT10, YU07]. IBA-I [YU07]. IBA-II [YU07]. IBM [İYK99, İTİO08, İTM12, KMÇ^{+}05, TU97, TM11]. IBM-1 [İYK99, KMÇ^{+}05]. IBM-2 [İTİO08, İTM12]. ICA [RJ17]. Ice [KZAN98]. Ideal [DMN22, PB01, PB03]. Identification [Bak11, JLS^{+}05, UKY96]. Identities [AKT04]. IFE [Übe04]. II [ZWZ16, ES06b, KUÖ05, LC18, MFF20, NGBLC22, YU07]. III [JAT^{+}21a]. Image [AB03, AKHHM17, AM23a, BGA10, Eli16, JWP23a, KAK96, Kar03b, KGCFK22, LPAD20, mLkWbPxX17, NYH^{+}17, SWA^{+}22, SÖS13, UB10, VVN23, YZBZ16].

Images [AHVG23, AB23, BG99, ÇGK96, ÜÖ03]. Imaging [KGCFK22, UK96a, ZPW^{+}22]. Immigrants [TRD^{+}22]. Immune [FA22b]. Impact [AK02, BS11, DSG19, DSI^{+}18, GPMRVM21, GEKB13, HIVH21, IGSB18, JH13, KEA11, NISC18, PAR23, TRD^{+}22]. Impacts [BG19, KVR23, NKS22b, RKS23]. Impedance [GM03]. Impeller [GPEG10]. Impersonation [KCMLTR18]. Impingement [BS13a, Kop02]. Impinging [ES13]. Implant [MAS^{+}11, ZA98]. Implanted [CB04]. Implementation [AA02, BGG96, ETW^{+}23, NISC18, PLLRCC20]. Implicit [FIJ21]. Importance [AA01, PYR22]. Important [CGM^{+}23]. Improper [AA02, Ara08]. Improve [Alm10]. Improved [BBSC18, ÇTY96, CBSTG13, DW16, HJL10, KPSS18, KB11, LBY^{+}22, SKK19]. Improvement [ŞÜ13]. Improvements [Mat10b]. Improving [AV04, LS20, MM17, VVV23]. Impulses [CK18, SD13]. Impulsive [DVFA23, EL04b, Yao15]. In-Plane [AJ20, CC22, Oku11d, ÖD06, Özb02, ÖÖT04]. Inactivation [BY17, Öze02]. Inadequate [CMM19]. INAR [ICDM21]. Incidence [Wu17]. Incident [OOI^{+}22, WA01]. Incinerator [ZMCL24]. Inclination [OSAAA18]. Inclined [BY06, RRR^{+}22, SJ22]. Including [YUKA03]. Inclusion [GÖD03]. Inclusions [Hus01, SAK01, SI04, SAH98]. Incompatible [GKE05]. Incompressible [ACM19, FJ21, Sal06b]. Increase [All02, CBT21]. Increased [AFJ22]. Independent [GB23, JM03, Ugu13]. Index [GZY20, HCGFOON20, Kur98, LY15, LLL14, Man16]. Indexes [Bec17]. Indicator [CA19, DRS17]. Indicator-Based [CA19]. Indistinguishability [ACGC21]. Individual [CW16]. Induced [Amo21, MCBM23, TCALK20]. Inducing [BVvdH^{+}23]. Inductance [VRMJG19]. Induction [Akb09, AB96a, Aks05a, AOL19, AB19, ÇA06, GAAAS17, Mat10b, MSSZ20]. Inductive [Aks05b]. Industries [HXY15]. Industry [ADS19, BG02, CCC^{+}18]. Inequalities [AS00, GW97, ÖT02, Sal02b, SAI04b, Sal04, Sal06a, TAS^{+}23]. Inequality [BCS20, KAM20, Sal02a, ZWZ17]. Inequivalence [SQA11]. Inexact [BGRV21]. Inextensible [Kız14]. Infant [LCW^{+}22]. Infection [FA22a, FA22b, Mer09, OWO^{+}23, TABT19, Zhu17]. Infective [TRD^{+}22]. Inference [Abi03, BSA^{+}23, CDDM23, OM20, ÜC11]. Infinite [AYG14, AS04, CK18, Faz21, KH15, KAM20, MOBA21, SAI04a, Sav13, TGK96]. Infinitesimal [FMM10]. Infinitesimals [JLZ04]. Infinity [Ers10, KSS19]. Inflammation [TKCC19]. Inflection [LLTC17]. Inflow [TCY06]. Influence [MLBR20, TABT19, YLX^{+}13]. Influences [HAMM10]. Information [AE16, CW16, Fu16, NKI19, SÖS13]. Informativeness [Haj04]. Informed [CMS20, DL23, HM23]. Infrared [WZ17, Yil11]. Initial [AV04, DSI16, MM17, OA13, PLW21, SM21]. Initial-Boundary [OA13]. Injection [BAUO09]. Injuries [TCALK20]. Injury [BNSDC20]. Innovation [MCB19]. Innovative [GFPF22]. InP [Bül99]. Input [Akb09, ZP11]. Insertion [BA99]. Inset [VRMJG19]. Insight [Faz21]. Inspection [RWL23]. Inspired [HBFM22, LOSCO^{+}20, Zha14]. Instabilities [LV19]. Instantaneous [CK18, KÇ96, UK96b, UT96]. Insulation [OYCS11]. Insurance [KTD12, MVC23]. Insurer [Kas19b]. Integer [BPC^{+}23, Kah11]. Integers [AA03]. Integral [AEB01, AD11, All12, Alt16, AD04, Boo17, BS15, DSD^{+}21, EB02, EB06, Eİ06b, GB00, GÖD03, ÖY01, RS21, ST18, YAA10, YA11, Yal13]. Integrals [AA02, Ara08, CP22, MK10, NM10, Shr98, Shr05, Üna03]. Integrated [Bor11, Kar11, LRL23, NCDM22]. Integrating [CMB19, SK18, WZ23]. Integration [CACRF^{+}21, FMR10, HLX^{+}17, HLX^{+}18, Mol18, Sar11]. Integro [AS09, AHR05, BG18b, DSD^{+}21, HNB20, YY07, YGS13]. Integro-Differential [AS09, AHR05, BG18b, DSD^{+}21, HNB20, YY07, YGS13]. Intelligence [OOI^{+}22, ÖE03]. Intelligent [NYH^{+}17]. Intensity [AA01, AUA03]. Inter [SÖS13]. Inter-Slice [SÖS13]. Interaction [BIM^{+}20, CMM19, DSU22, Has03b, MPP^{+}24, ŞÖ11]. Interactions [AT05a]. Interactive [ACC19, Con22, MECRMT^{+}21]. InterCriteria [AAR24]. Interdiction [ZS16]. Interdisciplinary [Ole20]. Interest [EA22]. Interface [Alt09]. Interfaces [GKE05, JWP23a]. Interference [ABKU10, Kha99]. Intermediate [PB03, VC23a]. Internal [BÖÖ11, Hei23, SUE13, ZCC15]. Interpolant [OL21]. Interpolating [Asl18, MAT20]. Interpolation [FHRS18, Li16, mLkWbPxX17, PTLZ18, PMF20, VC23b, YZHY17]. Interruption [WT15, YLX^{+}13]. Interruptions [Cha97]. Intersecting [QM16]. Intersection [Du 22]. Intersections [Mat23, Yav22]. Interval [AAS11a, DJL^{+}05, DW16, Fu16, MVC23, NAC22, ST22]. Interval-Based [ST22]. Interval-Valued [DW16]. Intervals [Faz21]. Interview [CGM^{+}23]. Intimate [DAKCPACC19]. Intra [PAR23]. Intra-Specific [PAR23]. Intraspecies [SAM^{+}22]. Intrauerine [GÖV96]. Intrauterine [ZGV10]. Introducing [HCGFOON20]. Intrusions [Aki11]. Intuitionistic [Bor11, LY15, Sav14, Sav16]. Invariance [KSK01]. Invariant [AM10, BK99, ENVPF19, Kha03, PMS03]. Invariants [ÖY01]. Invasive [LVHL22]. Inventory [HH07, Tsa09, Unv07]. Inverse [Bac13, GFPF22, KAS19a, MK97, VVV23]. Invesments [EÇ99]. Investigate [BGA10]. Investigating [Bha24]. Investigation [BSK05, BODH20, BBPL19b, CC22, ES13, HTLY18, İTM12, JLH21, MMM23, MGT10, MKK97, Oku11c, Oku11d, ÖS05, RRO16b, RFD^{+}10, SBA^{+}04, TSC11, TU97, YG02, YB97]. Investments [Kas19b]. Inviscid [STA23a]. Involved [TKCC19]. Involving [Cha05, RS21, Sal02a, Sal02b]. Ion [CB04, DB16]. Ionic [Öze02]. Ionization [Şen05]. IP [PD19]. IR1n [CUA04]. Iron [DB17a, DB17b, GBİ14, KR23]. Irradiation [MMM23]. Irrigation [SÇ10, YY11, YDD11]. Ishikawa [SAK01]. Iso [Koc00]. Iso-Taxicab [Koc00]. Isogeometric [Rod20]. Isomorphism [Bİ99]. Isoperimetric [AERE18, EARE18]. Isothermal [BS13a]. Isotonic [PMF20]. Isotopes [BSK05, İYK99, KMÇ^{+}05, MGT10, TM10, YU07]. Isotropic [AUA03, Oku11b]. Issue [Fan20]. Issues [CGM^{+}23]. Itô [FMM10]. Italy [SPCPMM21]. Item [Rob22, Tsa09]. Items [JMS18]. Iteractios [BM11]. Iterated [ST18, Sol23]. Iteration [AS09, AADS11, CML10, GHN11, LXZ10, MDNN10b, PAB11, YRB16]. Iterations [Guy20]. Iterative [ASJQ24, ANK23, BBMA22, BA11, Faz21, SAK01, SKB22, SBB23, Tan08]. IUGR [ZGV10]. IV [LCB18]. IV-Optimal [LCB18].

Jacob [FMF20]. Jacobi [BS15, Wu10]. Jacobi-Elliptic [Wu10]. Jaulent [SK18]. Jeffrey [JS22b]. Jet [BA99, Kop02, MWM16]. Jets [ES13, NNM10]. Job [Cha05, FSHRCV^{+}21, NISC18, SKK19]. Jobs [Cha05]. Jockeying [Cha08]. Join [Mat10a, Mut00b, YBYL21]. Join-Based [Mat10a]. Joining [KA11b]. Joint [AR23, CBSTG13, CGCC11, KTS13, KVR23, Sar09, TWL10]. Joints [GAÖ96, GA99a, KT98, KTC99, Kir04, OF19, ST11]. Joule [JKS23, NKS22a]. Journal [ÜDMA04]. Jump [Dan22, JY23]. Justify [CMCM21]. JWKB [Den16].

K-Fold [BY11]. Kadomtsev [LDL10]. Kalman [Haj04]. Kalyanmoy [CGM^{+}23]. Kantorovich [AD11]. KBM [Cai07]. KdV [HL14, KO10, ME12]. Keller [RZ18]. Kernel [AEB01, LS20]. Kernels [RRR^{+}22]. Keys [DJ10]. KF [PC23]. Khz [DK03]. Kidney [DL24]. Kind [Alt16, ST18]. Kinds [Alt09]. Kinematic [LCW^{+}22]. Kinetic [BG18a, Can02, LCW^{+}22]. Kinetic-Kinematic [LCW^{+}22]. Klein [Kaz04, ZMA23]. Knapsack [BGN13]. Knee [Sar09]. Knife [YD11]. Knock [TCALK20]. Knock-Out [TCALK20]. Knots [AB96b]. Knowledge [SBDN^{+}22, WJ23]. Knowledge-Driven [SBDN^{+}22]. Kolkiran [KA16]. Kolmogorov [WZ23]. Korteweg [BMT^{+}21]. KPP [RZ18]. Kr [ÇA04]. Kr-Core [ÇA04]. KSA [HC22]. Kunc [KF19a]. Kutta [DSI16, DRKW21, AAD11, BG18a, WZ23]. Kuznetsov [Kar10, MJK12]. KX [KG13].

L [AÖ09, KIY^{+}23, Übe04, VC23a].

L-Systems [VC23a]. L3 [YH05]. L4 [YH05]. Labor [Dan22]. Labyrinth [DE04]. Lacunary [SS03, Sav10, Sav14]. Ladder [FMF20]. Lagrange [WZT20]. Lagrangian [HT19]. Lags [AW19]. Laguerre [GGS11]. Laminar [HM23, NNM10, WFV21]. Laminated [ANA20, AK02, BS11, ÇKAK04, EY13, MÖ96, NCDM22, ÖG13, Özb02, STK07, TA96, TMB96]. Laminates [BKOS11]. Landau [İB00, qS13]. Lane [Ert07, HM10, MK10]. Langmuir [DB16]. Langrange [Civ96]. Lap [AR23, ST11]. Laplace [Doğ12, KWP11, KLGP24, MK97, SKGP22, TSS17]. Laplace-Stieljes [KWP11]. Laplacian [PS23]. Large [BUAM19, CAQCMM22, CWF^{+}22, GCGZ10, NGBLC22, SSA22a, TCY06]. Large-Scale [CAQCMM22, GCGZ10, NGBLC22]. Largest [PKC22]. Laser [Has03b, Yil11]. Late [DMN22]. Latency [AAGH21]. Lateral [DD11, STA23a]. Laterals [YDD11]. Lathe [BT03, SBF96]. Latitude [Pin22]. Lattice [PSW24]. Laurent [Özd96b]. Law [AYG14, AHM99, Aki11, JB15, MAH10, ZMA23]. Laws [AM10, KSM23, Kar10, NK10, NNM10, PMS03, ZMA23]. Layer [AUA03, BODH20, FJ21, HPA13, KBE01, KG10, MM10d, NUM^{+}19, Oku11c, PÖ98, PMR14, CTAA11]. Layered [BE99, EBY06]. Layers [Bül99, Hak22, VC23a]. Layout [EKAÖ11]. LCM [NT04]. Leadership [Kat17]. Leading [ADS19]. Leak [BLEB^{+}18]. Learning [ARCRMM20, Aks05b, AM09, CKLZ21, Con22, CLGGLB^{+}22, DDPK21, GCCMD23, HMKS18, JMB22, KAY09, LL23, NSG^{+}23, OOI^{+}22, PSF22, Rou21, Tsa09, ZE16]. Learning-Based [HMKS18]. Least [Kan16, KAS19a]. Least-Squares [Kan16]. Left [DAD^{+}18]. Legendre [GY13, YAA10, YGS13]. Leishmaniasis [AW19, BDC^{+}18]. Length [KN18, MC20a, MC20b, ST11, YDD11]. Length-Biased [KN18, MC20a]. Lerch [RS21]. Less [CPP19]. Level [BK10, BPC^{+}23, CCL06, MWM22, Sul07]. Levels [RLY^{+}17]. Levenberg [Kay16]. Leverett [AGKT21]. Levy [IÇÜ12, Dan22]. Lexicon [OPUÖ11]. LFP [Tan08]. Libraries [VPJ22]. Lidstone [MAT20]. lie [Öze00, Can96, FMM10, ZMA23]. Life [DM21]. Lifetime [MCNM22, NAC22]. Lifting [GA99b]. Lifts [Civ96]. Light [MR19b, Pin22]. Lightweight [AMR^{+}23]. Like [CUA04, FMR10, KÇ96, ÖY01, Özy14, Sal02a, UK96b, UT96, UÇ96]. Likelihood [CZC17]. Limit [KLGP24]. Limited [CMCM21]. Limiting [CC24]. Lindley [ICDM21, KHR^{+}22, SKM^{+}22, TGC21, TGC22]. Lindstedt [PKB09, PKB11, PS15]. Line [BD09, Bül99, mLkWbPxX17, MK97, SKKC18]. Linear [AYG14, Akh04, AOL19, BÖÖ11, BPC^{+}23, CK18, DSD^{+}21, Dol04, ES06b, ES06a, GC23a, GLS18, GGS11, JHA^{+}22, KLGP24, KKS11, KSS19, KYS13, MQ18, MRG23, NM10, Özk01, ÖÖT04, Özy09, RG23, SKB23, SKGP22, Tan08, TW18, TB21, Üna96, YÖS10, YA11, Yal13, YG15, Yao15, Yaz16, YGS13, ZG17, ZLLX20]. Linearities [EESK^{+}02]. Linearity [PY00, ZMA23]. Linearizable [SQA11]. Linearized [RZ18]. Linearly [JS22b, NSCTPW22]. Lines [LL17, MK96]. Linguistic [Ace19, CWF^{+}22]. Lining [LC18]. Link [All02]. Linked [HH07]. Liouville [Ert11]. Lipschitz [Sal02a, Sal02b]. Liquid [GB18, Has03a, MM10c]. Liquids [Has01, Has03b]. Listening [VÇÖ11]. LiTaO3 [Kur98]. Literature [ACL^{+}22]. Lithiated [Boo17]. Liver [FA22b]. Load [AE03, Akb09, BGA10, Ese01, KE01, Oku11d, RFD^{+}10]. Loaded [Alt10, BKOS11, KÖAG00, Özc00]. Loading [AO17, ÇŞS01, DM21, DD11, IE10, JB15, KG13, KAEÖ13, Özb02, SMB96, TA99, YG02]. Loadings [HFATZR^{+}20, Oku11d, TMB96]. Loads [AÖ09, CDDM23, MÖ96]. Local [Akh04, BG18b, JSTGV^{+}19, KAS19a, LSB^{+}20, MM10d, TB16, YBYL21]. Localization [HRM19, YK96b]. Localized [FLYY21, UAT19]. Location [AK18, Bor11, HB99]. log [SBA^{+}04]. Logarithmic [AGANA22, CTJJ22, RZ18]. Logic [ÇA10, GDE12, Ken11, KA96b, Kle18, Özg10, TGK96]. Logistic [HC22, Ji13, KY13, Zhu17]. Logistics [GM17]. Lognormal [KWP11]. Lomax [KN18, KHR^{+}22, MC20b, NVC22]. Long [BB16, KAY08, KPS04]. Longitude [Pin22]. Longitudinal [Çev10, ES06b]. Look [Alm10]. Look-Ahead [Alm10]. Loop [ASK21, BS23]. Lorentzian [Kaz05, YH05]. Lorenz [PS20, PPRVC20]. Lorenz-96 [PS20]. Lorenz-Type [PPRVC20]. Loss [AK09, BDH19, HAK03, SRLEPVP20]. Losses [BCK^{+}19, BODH20, OD19]. Lot [CBSTG13, CC11, CCWY10, CGCC11, CLC16, TCC11]. Lot-Size [CLC16]. Lotka [HXY15, SAM^{+}22]. Low [AK02, CYAK03, EAT05, GM17, İTİO08, MKK97, SGBR23]. Low-Clinker [CYAK03]. Low-Cost [SGBR23]. Low-Lying [İTİO08]. Low-Velocity [AK02]. Lower [Tur16]. LPV [WCH23]. LPV-Based [WCH23]. LQR [MPA17]. LSM [Haj05]. Lubricated [Yür02]. Lung [MCBM23]. Lyapunov [ŚLO19]. Lying [İTİO08]. Lymphotropic [Mer09].

M [WT15, YRB16, WT15]. M1 [BUKM96]. Machine [BhMsG16, Cha05, CLGGLB^{+}22, GCCMD23, HMKS18, JMB22, MMF18, OOI^{+}22, RFQCMMCR23, TE05, WFV21, ZE16]. Machines [BBD^{+}18, BODH20, DB20, Jab18, Ken11, KNHK20, MSSZ20, Mes96, VRMJG19, ZGV10]. Macroeconomic [PC23]. Macrophage [TKCC19]. Macrophage-Mediated [TKCC19]. Made [ALFBPLE20]. Magnesium [DL24, Oku11a]. Magnet [BODH20, Jab18, MMF18, OD19, VRMJG19, WFV21, ZLLX20]. Magnetic [AOL19, BDH19, DM99, DK03, EPDG21, Has03a, HAMM10, HFAAM^{+}22, KD03, KVR23, KKA04, MAKH10, MMF18, RKR^{+}22b, VRMJG19]. Magneto [BDMS19, FJ21, GS01, GS02, GS10, TKK^{+}22]. Magneto-Gas [GS10]. Magneto-Gasdynamic [GS01]. Magneto-Hydro [FJ21]. Magneto-Hydrodynamic [GS02]. Magneto-Mechanics [BDMS19].

Magnetoencephalography [CPP19]. Magnetohydrodynamic [RKX23, Sha19]. Magnetostatic [FCGL19]. Maintain [CS19]. Maintenance [ACL^{+}22]. Make [VSAMMM23]. Maker [PVCRRV^{+}21, WOL15]. Making [AHID18, ADS19, AN19, ASA20, ASAH20, Bor11, CWF^{+}22, CCC08, DW16, Fu15, FbF16, Fu16, JLS^{+}10b, JZL10, LY15, Liu11, LNBM10, MECRMT^{+}21, PVCRRV^{+}21, SAZ18, WOL15]. Malaria [BG19]. Male [DL24]. Malnutrition [MWM22]. Management [AK09, YY11]. Managing [CCC^{+}18]. Mandible [ZA98]. Maneuvers [SGBR23]. Manganese [KR23]. Manifolds [ÜU05]. Mann [SAK01]. Manual [GA99b]. Manufacture [GBİ14]. Manufactured [TDC08]. Manufacturing [AA13a, Kar03b, SBDN^{+}22, TCA96]. Manuscript [Ace19]. Many [ÇÇ03, GCCMD23]. Many- [ÇÇ03]. Many-Objectives [GCCMD23]. MAPLE [Qua18, PQ23]. Mapped [LZ21]. Mapping [UÇ96]. Mappings [KAM20, Sal02b, TAS^{+}23]. Maps [SJKCD20]. Marble [EKAÖ11]. Marine [DHK23, ESNB19, JTM^{+}22]. Marital [CCGHdCJ21]. Market [BM21, CMG22, GGMO21, Sho22]. Marketing [KY13, MCB19]. Markets [JAT21b, WW20]. Markov [Ace19, AFGV20, JY23, KPTM20]. Markovian [HLX^{+}18, HLX^{+}17]. Marquardt [Kay16]. Marshall [MC20a, MC20b]. Mask [MSPC23]. Mass [BK10, BIM^{+}20, DK99, Ese11, JH13, Öz00a, ÖÖ05, Özy09, Xu10]. Masses [Özk01]. Mat [LH10b]. Matching [Den16]. Matematical [PF10a]. Material [Alt10, AFJ22, DNS23, EBY06, FSH^{+}23, KPSZ24, NLMU19, Oku11d, dAPdSL^{+}22, TÇ06, VVV23]. Materially [BV19]. Materials [GGT^{+}23, GFPF22, HAK03, ŞT96, ST22, TSC11, UY11, UC19]. Math [HLX^{+}18, KA16, KF19a]. Mathematica [GK11a]. Mathematical [Amo21, Ars07, BÇ13, Bas99, BDC^{+}18, Çat10, CCC08, DSMP19, DAKCPACC19, DAD^{+}18, DB20, GPACC^{+}19, GPMRVM21, GLS18, Has03a, IMWY18, JLH21, KBH06, KAEÖ13, NU19, OMX18, OM20, OV20, PFR19, RSF23, SGZ11, SS07, SEZ24, SM21, TK04, UYAA22, Üna96, ZH10]. Mathematics [LEVP21]. MATLAB [CST20, KA09, ÇA10]. MATLAB/GUI [KA09]. Matlab/Simulink [ÇA10]. Matrices [BS02, ÇA04, CBM^{+}19, CJL18, NT04, NSCTPW22, NU19, SC19, SB02, STB04]. Matrix [BÇ13, Bil02, BS15, BO10, ÇKAK04, Çev10, Ese01, JLW20, KK13, KSK05, KÜM08, KE01, KKY05, Oku11a, Oku11c, Oku11d, Özb02, SSA22a, Sav13, SB02, TÇ06, TB04, UC19, Wan13]. Matrix-Free [JLW20]. Matter [BNSDC20, BM11, MO04, Man04, MKMB11]. Maxillary [GGT^{+}23]. Maximal [HAH^{+}20, dlFH19]. Maximization [Tsa09]. Maximum [CZC17, PQ23, TMK21, Ugu13]. Maxwell [MVC23, BDH19, MC20a]. MCDM [SW22]. McLaren [Şen05]. MCSA [GAAAS17]. Mean [Bac13, DÇ06, GU05, SH18]. Means [AMK96, DUM04, TÇ06, ÜDMA04, UÖKE14, YÖS10]. Measurable [BVCS18]. Measure [Özd96b]. Measurement [EEJRLB^{+}22, KY96, MAE97]. Measurements [DM99, LVHL22]. Measures [FS23]. Measuring [Haj05, Sal09]. Mechanical [Bay03, Çiç07, ÇÇ03, CTAA11, Kol11, KA16, LESEVP^{+}20, PCCT11, PS19, ST22]. Mechanics [BKA08, BDMS19, Bro19, CA19, Fan20, SM21, ZJ10]. Mechanism [HA10, TCALK20]. Media [ALFBPLE20, SWJS22]. Mediated [PD19, TKCC19]. Medical [DAD^{+}18, Erp12, KGCFK22, SÖS13].

Medium [LSARZ23, MAKH10, RKR^{+}22a, RKR^{+}22b, SSV23, TKK^{+}22]. MEG [CPP19]. Mellin [Kil04]. Melting [KZAN98]. Member [ÖÖT04]. Membrane [HTLY18]. Membranes [CG09]. Memory [AMAGA22, CJL18, MOBA21, SKB22, SKB23]. Memory-Type [AMAGA22]. Merit [ÇT09]. Mesh [AA01, ADJ18, Fan20]. Mesh-Free [Fan20]. Meshless [FSV^{+}22, KAS19a, TO13]. Metaheuristic [MSPC23]. Metal [EAT05, KE01, Oku11a, Oku11c, Özb02, TÇ06]. Metals [ÇÇ03]. Metalurgy [TDC08]. Metamaterial [GC23b]. Metamaterial-Based [GC23b]. Meteorological [DJY^{+}22]. Methane [KK97]. Method [AV04, AS09, AAD11, AADS11, AO16, AERE18, ASJQ24, AJ20, All02, AAS11b, AD11, All12, Alt16, AHR05, AMK96, AWK^{+}20, ATK99, BT19, BG18a, BCK^{+}19, Bay03, BY17, BDH19, BP21, Boo17, Bor11, BS15, BBK20, BÖ07, BO10, Cai07, CPP19, CB04, CBM^{+}19, CACRF^{+}21, Çat10, ÇÇÇ07, ÇKÇK10, Çev10, CO16, CML10, CWF^{+}22, CCC08, CBT21, CMCM21, CJD24, CTAA11, CJL18, DS04, DI13, DW16, Doğ12, Dol04, DAD^{+}18, DJM18, DUM04, EAEAA19, ESNB19, EB06, EARE18, EE22, Eli16, EB00, Ert07, Ert11, Faz21, FIJ21, FKRG10, FHRS18, FbF16, Fu16, GT22, GMV19, Gha18a, Gha18c, GHN11, GC23a, GCGZ10, GB00, GY13, HBS^{+}17, HH22, IZS04, JLST04, JLKK12, Jia17, JZL10, KTK11, KRD^{+}22, KSM23, KODM11, Ken96, KO10, KG10, KN13, KSS19, KB11]. Method [KO15, KPSZ24, KYS13, LY15, LYYX15, LXZ10, Liu11, MQ18, Man16, MGGHCR^{+}20, Mer09, MG11a, MG11b, ME12, MMS19, MFF20, MDNN10b, NNW14, OSAAA18, OV18, Öz00a, Öze10, PB99, PKB09, PKB11, PS15, PKO11, PMR14, PQ23, PMF20, PSW24, Qua18, Raw14, RFD^{+}10, RAA10, SK18, ŞT96, SAZ18, ST18, SGA18, SKB23, SKGP22, SZTY23, Sol23, TKGS20, TO13, Tan08, TÇ06, TDC08, TYS13, TY15, TGS98, TMK21, TSS17, UK96a, UAT19, UYAA22, ÜDMA04, UAI21, VVV23, VVN23, Wan13, WZ23, WED^{+}23, WOL15, YÖS10, YG15, YB03, YW24, YD11, YKG20, YQG22, YRB16, ZLZ17b, ZZ17, ZH10, Zhu10]. Methodology [GDE12, KBH06, LRL23, LZ21, SHFSGB^{+}21]. Methods [AVC23, Bad22, BK09, BGRV21, BA11, BBK20, BDMS19, ÇT14a, DRH21, DSI16, Fan20, FCGL19, GN19, GK11a, GKE05, İBB98, JTM^{+}22, KAM20, KKL18, LS20, MV21, Ole20, Özg10, PCCT11, SSA22a, SI04, Sal04, SJ19, TW18, UB10, YE03, Yen05, YK96b, ZG17]. Metrics [CW16, Che18]. Mexico [LDCST21]. MHD [AM05, Ers10, JS22a, KNRC22, MMS19, SSV23]. Microanalysis [CBT21]. Microarray [UB10]. Microelectromechanical [GCGZ10]. Microflow [ZHN10]. Microhardness [MAE97, TÇ06]. Microparticle [ZHN10]. Microplastic [SAM^{+}22]. Micropolar [NU19]. Microscobic [Kas00]. Microscopy [AB23, SWA^{+}22]. Microstrip [SGE98]. Microstructure [LF19]. Microtubules [CDA10]. Microwave [HAK03, KY96, PPRF22, TGG96]. Middle [FYuR^{+}23]. Migration [Amo21, LL24]. Mildly [AS00, SAI04b, Sal06a]. Mine [EÇ99]. Mineral [BY11, CYAK03]. Minimal [Çal06]. Minimax [DKÇ97]. Minimizaion [Ken96]. Minimization [CKCC16, GC23a]. Minimizing [Kas19b, KÖ00]. Minimum [KHR^{+}22, Mes96]. Mining [MGGHCR^{+}20, SSGCPR23]. Minkowski [CUA04, UÇ96, Yay00]. Minor [SRLEPVP20]. Miodek [SK18]. Mission [SGBR23]. Mixed [AS00, Bül99, Hus01, Mat19b, Mat19a, NK10, PRFP^{+}23, Sal02a, SAI04b, Sal06a, SJ22, TKK^{+}22, YG15]. Mixing [BUKM96]. Mixture [AE16, DDPK21, LCB18]. MLE [SKAK05]. Mobile [GÇŞT07, LOSCO^{+}20]. Modal [NAC23]. Mode [Gul03, GM03, PPRVC20, YK96b]. Model [AAGH21, AERE18, AJ12, ATZES^{+}07, AGKT21, AA16, Ace19, AW19, AHM99, AE03, AM23a, ALFBPLE20, AOL19, Amo21, AM10, ABH^{+}22, Bağ10, BÇ13, BG19, BIM^{+}20, Bas99, BUKM96, BY17, BODH20, BDC^{+}18, BBH19, CBSTG13, CA19, ÇT14a, CTH^{+}22, ÇSB10, CO16, CCGHdCJ21, CC11, CW16, CCL06, CCC08, CCWY10, CGCC11, CDA10, DCC23, Dan22, DCAR21, DSG19, DAKCPACC19, DE04, Dik06, DEL23, Erp12, ELT21, FSH^{+}23, FIJ21, FA22b, Fu15, GVC20, GS02, GOOM18, GGMO21, GZ20, GDE12, Guc04, GM03, HGS^{+}22, HT19, Has03a, HIVH21, HXY15, IE10, JLS^{+}10a, JMB22, Ji13, JZL10, JLH21, JAT21b, KD19, KPTM20, Kum07, KVR23, KAEÖ13, LCB18, LLL14, LL24, MM17, MSSZ20, Mer09, MVC23, MPP^{+}24, MCNM22, OMX18, OM20, ÖG03, Oze03, OA09, PB99, PT21, PCCT11, dAPdSL^{+}22, PF10a]. Model [PC23, PALL18, RAK22, RKS23, Rob22, RMBTGH^{+}20, SA19, SPCPMM21, SGC23, SAM^{+}22, SZ23, SEZ24, Sta23b, TSS17, UAT19, UC99, UYAA22, Üna96, Unv07, UÖKE14, VVV23, VSAMMM23, WW20, WA01, Wu17, YA23, YLX^{+}13, Yao15, YU07, ZBBZ18, ZJ10, Zha14, ZPL16, Zhu17]. Model-Based [FSH^{+}23, MPP^{+}24]. Model-Robust [LCB18]. Modeling [ACM19, Aks05a, AA01, Ars06, AK11, BhMsG16, BP23, BCK^{+}19, BAUO09, BLEB^{+}18, Bha24, BWK23, BBK20, ÇTY96, CTJJ22, DRH21, DHK23, DL23, DRCSF22, EB00, EPLGGS^{+}21, EPDG21, EÜA13, FYuR^{+}23, FA22b, GM21, Gha18b, GPACC^{+}19, GPMRVM21, GLS18, IMWY18, Kas00, KTS13, KIY^{+}23, LRL23, LCW^{+}22, LL17, Mec22, NAC22, NUM^{+}19, NU19, OF19, PMF20, RWL23, RJ17, RRR^{+}22, SS07, TÇ04, TUK19, TK04, TKCC19, Ull17, YLX^{+}13, YAAK96, ZLZ17b]. Modelization [SGBR23]. Modelling [BDMS19, ÇD03, FP21, HSG^{+}20, KODM11, KR23, MC20b, NCDM22, Özg10, PGRL21, PPB^{+}20, PPRF22, RKX23, Rou21, ŞÖ11, SS96, TÇ06, TCY06, VPJ22]. Models [AJ12, BHN10, BP21, CPA20, ÇÇÇ07, CJL18, DJL^{+}05, DDPK21, DB20, Egb18, FSV^{+}22, FdSPW20, HCdCTAAMEJ20, HSG^{+}20, HRM19, JLS^{+}10a, JL24, KC08, KBH06, MQ18, MS05, NMCM19, ÖS05, PAR23, PD18, Pek96, PS20, PFR19, Raj10, Rob22, RG23, SKB23, SM21, Tsa09, ZS16]. Modes [Bül99, CBM^{+}19, SHE05, Yag02]. Modification [BCN22, CMCM21, GÖA96, RG23]. Modified [AJ20, AK18, AHR05, AFGV20, Car21, CKKA22, Haj05, JAT^{+}21a, Ji13, MG11a, MDNN10b, MAH10, SK05, TAS^{+}23, TGC21, TGC22, Übe04, UYAA22]. Modular [ASK21]. Modulated [TABT19]. Modulation [BPC^{+}23]. Modules [Arv04, HTLY18]. Modulus [ZG17]. Modulus-Based [ZG17]. Molecular [AÇGJ99, BP23, KTK97]. Moment [Bağ10]. Moment-Curvature [Bağ10]. Moments [ABKU10, IMKEK18, NCM19, Oui23]. Momentum [ALK19]. Monitoring [BT03, DCC23, GERV19, RMBTGH^{+}20, SK03]. Monochromatic [Shi06]. Monodromy [Muc00, Muc01]. Monotone [AWK^{+}20, GW97, Sal02b]. Monte [AÖ05, CATCdlF20, ÇT14a, Man04]. Moody [DE04]. Mooney [VVV23]. MOR [GCGZ10]. MOR-PIM [GCGZ10]. Morlet [Pin22]. MoS [NKS22a]. Mosaicing [Eli16]. Motion [EL04b, Has01, KY05, KÇ11, ÖÖ05, PSF22, SSV23, Yay00]. Motions [VVN23]. Motor [Akb09, ÇA06, ÇA10, KA96b, RCTR20, ZLLX20]. Motorcycle [FdSPW20]. Motors [Akb09, AOL19, Alt09, Çun08, GAAAS17, MPA17]. Mountain [Kum07]. Mounted [BVvdH^{+}23, DMN22, Jab18]. Mouth [Wu17]. Moveable [RRR^{+}22]. Moving [BIM^{+}20, BS13a, ES13, Ese11, GB18, KAS19a, PÖ98, RKS23]. MPPT [TCC^{+}23]. MPPT-Based [TCC^{+}23]. Mucus [UÖKE14]. Multi [AA16, AA13a, ADJ18, BMV22, BBSC18, BCS20, BODH20, BS23, Bor11, BÖ07, CMS20, CACRF^{+}21, CS22, CKCC16, CS19, CJL18, DE23, DW16, EPLGGS^{+}21, EPDG21, FSHRCV^{+}21, Fu16, GFPF22, HBS^{+}17, KIY^{+}23, LL24, LBY^{+}22, Mak99, MPA17, MMS19, MFF20, NGBLC22, NKI19, RC16, SSGCPR23, SBDN^{+}22, SHE05, Sul07, TOM^{+}22, TMK21, Tsa09, VHMMG20, VMRVF^{+}23, WED^{+}23, WJ23, XS22, YA23, YY11, ZG17, ZWZ16]. Multi-Agent [MPA17]. Multi-Area [RC16]. Multi-Attribute [Fu16]. Multi-Bay [BÖ07]. Multi-Branch [TOM^{+}22]. Multi-Channel [Sul07]. Multi-Credit-Rating [LL24]. Multi-Criteria [AA16, CACRF^{+}21, DW16]. Multi-Domain [MMS19]. Multi-Feature [HBS^{+}17]. Multi-Frequency [ADJ18]. Multi-Gene [TMK21]. Multi-item [Tsa09]. Multi-Layer [BODH20]. Multi-Objective [AA16, BMV22, BCS20, BS23, CMS20, CACRF^{+}21, CS22, CS19, DE23, EPLGGS^{+}21, FSHRCV^{+}21, Mak99, MFF20, NGBLC22, SSGCPR23, SBDN^{+}22, VHMMG20, VMRVF^{+}23, WED^{+}23, WJ23, XS22, ZWZ16]. Multi-Parameters [ZG17]. Multi-Phase [BODH20]. Multi-Physics [EPDG21, GFPF22]. Multi-Product [CKCC16]. Multi-Region [CJL18]. Multi-Server [SHE05]. Multi-Stacked [KIY^{+}23]. Multi-Strain [YA23]. Multi-Strategy [LBY^{+}22]. Multi-Threaded [NKI19]. Multi-User [BBSC18]. Multiatomic [Kas00]. Multiattribute [LY15]. Multibay [Çat10]. Multibody [ZBBZ18]. Multichannel [ATZES^{+}07]. Multicomponent [DZA19]. Multidimensional [BGN13, Sal03].

Multigrid [NSCTPW22]. Multigroup [TO13]. Multiindex [Shr01, Shr02]. Multilevel [KSTJ18, LTYG17, Sal03]. Multinomial [Oui23]. Multiobjective [BGRV21, BP21, Çun08, DRH21, PD18]. Multiphase [AGKT21, JWP23a, JWP23b, MSSZ20]. Multiphysics [BY17]. Multiple [AW19, BÖÖ11, BY11, BBMA22, BPC^{+}23, CT14b, Dal11, Fu15, FbF16, JLST04, JZL10, KHL06, KUE13, Liu11, ORC19, PB99, PKB09, PKB11, PS15, Sav14, TKGS20, WOL15, ZS16, Bay03]. Multiple-Set [TKGS20]. Multiplicative [SBB23]. Multipliers [WZT20]. Multipole [BUKM96, YRB16]. Multiprocessor [All99]. Multirate [BP99]. Multispectral [KBH06]. Multisplitting [ZG17]. Multistability [PS20]. Multivalued [AS00, KAM20, SAI04b, Sal06a]. Multivariable [DJY^{+}22, Shr98, Shr02, Shr05]. Multivariate [AE16]. Municipalities [MPA^{+}22]. Müntz [YGS13]. Müntz-Legendre [YGS13]. Muscles [AAR24]. Musculoskeletal [LCW^{+}22]. Mutation [RFQCMMCR23]. Mutations [ŚLO19].

N [Cha05, KB03]. N-Job [Cha05]. n=12 [GGJ99]. n=12-14 [GGJ99]. n=7 [BDÖ^{+}99]. Nadarajah [NVC22]. Nano [BP23, KG10, Kol11, KA16]. Nano-Mechanical [Kol11, KA16]. Nano-Scale [BP23]. Nanobeam [Alh21, TB16]. Nanofluid [DJH18, DEL23, KNRC22, NKS22b, RRO16b, RKS23, TKK^{+}22]. Nanofluids [ES13, LYYX15, MMS19, RRO16a]. Nanoliquid [JS22b]. Nanoparticles [KNRC22, RRO16b, SSV23]. Nanotubes [DCA10, SUE13]. Nanowires [Boo17, UC19]. National [KAY09]. Natural [Alt12, BT19, HFAAM^{+}22, Öz00a, ÖÖ05, SKM^{+}22]. Naturel [SS96]. Navier [ACM19, HM23, Sal06b]. Navigation [NSG^{+}23, SPD18]. Near [BdCCF19, HWA99, SWJS22, GS10, KH15]. Near-Field [SWJS22]. Nearest [KPTM20]. Nearly [HB01]. Necessary [PD19]. Needed [AERE18]. Needle [RKS23]. Negacyclic [JS18]. Negative [ES06b, ES06a, Fu15, GB23, Oui23]. Neighbor [KPTM20]. Neighborhood [All99]. Nematics [AK05]. Nets [BG19]. Network [Abi03, ARCRMM20, ABH^{+}22, Bağ10, BY11, BLEB^{+}18, CYAK03, ÇSB10, Cha08, DK03, DL23, Erp12, FSHGGB^{+}21, GZY20, GNK19, GZ14, JLS^{+}05, Kar11, Kar03b, KKL18, KD03, KAEÖ13, MRIR20, MFF20, RCTR20, SK03, ŞÖK11, SGEP06, SZTY23, SKAK05, TOM^{+}22, TDC08, TGG96, UKY96, WA01, Zha14, ZS16, ZPW^{+}22]. Networks [AB03, Ala11, Alm10, AMK96, BBSC18, BPC^{+}23, Bha24, BBPL19a, BBPL19b, Can14, ÇT12, ÇS10, DÇ06, DUM04, ETW^{+}23, EG11, GÇŞT07, GAS11, GM03, GÖV96, HM23, KAK96, KE98, Kay16, KY13, LRBMAM^{+}23, MSPC23, MKK97, Mes96, NAA20, OC09, PGRL21, PLW21, Sağ98, ŞE17, SWA^{+}22, SD13, SSA22b, TÇ06, ÜDMA04, WCH23, YS02]. Neural [Abi03, AB03, AMK96, Bağ10, BY11, BBPL19a, Can14, CYAK03, ÇT12, ÇSB10, ÇS10, DK03, DÇ06, DL23, DUM04, ETW^{+}23, EG11, FSHGGB^{+}21, GZY20, GÇŞT07, GAS11, GNK19, GM03, GÖV96, HM23, Kar11, KAK96, KE98, Kar03b, Kay16, KY13, KD03, LRBMAM^{+}23, MRIR20, MSPC23, MKK97, Mes96, OC09, PM22, RCTR20, Sağ98, SGE98, SK03, ŞE17, SWA^{+}22, ŞÖK11, SGEP06, SD13, SZTY23, SSA22b, SKAK05, TÇ06, TDC08, TGG96, ÜDMA04, UKY96, WCH23, WA01, YS02, Zha14, ZPW^{+}22]. Neuro [AK11]. Neuro-Fuzzy [AK11]. Neuroevolution [LRBMAM^{+}23]. Neuron [UC99]. Neuronal [JFRZ10]. Neutrix [Kil03]. Neutron [CBM^{+}19, TO13]. Newhouse [Yav22]. Newly [DK99]. Newsvendor [HT19]. Newton [AHR05, CBM^{+}19, DSU22, EE22, EB00, SJ19, WED^{+}23]. Newtonian [Bro19, ETL22]. Ni [OYCS11, TDC08, TSC11]. Ni-Ti [TDC08]. Ni19 [AÇGJ99]. Nicholson [Yao15]. Nin [BDÖ^{+}99, GGJ99]. NMR [YB97]. No [SGZ11]. Node [All02]. Noether [MK10]. Noise [DSG19, DHK23, Kol11, KA16]. Noisy [ÜC11]. Nominal [OFCP21]. Non [AYG14, AJ12, Aki11, AÇGJ99, BÖÖ11, BG18b, BBH19, CSSOB20, ÇT14a, CK18, DJL^{+}05, DEL23, DMN22, EESK^{+}02, ES06b, ES06a, ETL22, FJ21, GS01, GS02, GS10, Gha18c, GLS18, İB00, JS22b, JLS^{+}10a, JLS^{+}10b, LVHL22, LC18, LVBV19, MQ18, MMF18, MM10d, PY00, PB01, PB03, PMF20, RRR^{+}22, SKB23, Shi06, TB16, YG15, Yaz16, ZMA23, ZWZ16]. Non-Analog [ÇT14a]. Non-Archimedean [DJL^{+}05]. Non-Autonomous [GLS18]. Non-Circular [LC18]. Non-Dominated [ZWZ16]. Non-Epsilon [CSSOB20]. Non-Ergodic [Aki11]. Non-Extreme [JLS^{+}10b]. Non-Fourier [DEL23]. Non-homogeneous [MMF18]. Non-Ideal [DMN22, PB01, PB03]. Non-Instantaneous [CK18]. Non-Invasive [LVHL22]. Non-Isotonic [PMF20]. Non-Linear [AYG14, BÖÖ11, ES06b, ES06a, MQ18, SKB23, YG15, Yaz16]. Non-Linearities [EESK^{+}02]. Non-Linearity [PY00, ZMA23]. Non-Linearly [JS22b]. Non-Local [BG18b, TB16]. Non-Newtonian [ETL22]. Non-Orthonormal [BBH19]. Non-Perturbative [İB00]. Non-Poisson [Aki11]. Non-Programmer [Gha18c]. Non-Rotating [AÇGJ99]. Non-Similarity [MM10d]. Non-Singular [LVBV19, RRR^{+}22]. Non-Standard [FJ21].

Non-Uniform [GS01, GS02, GS10, Shi06]. Noncoaxial [Ers10]. Nondominated [JLS^{+}05]. Nonexistence [BC97]. Nonexpansive [KAM20]. Nonextensive [AK05, KST10]. Nonholonomic [GMV19]. Nonhomogeneous [PLLRCC20]. Nonlinear [AS09, AO17, AS00, AGANA22, AD11, All12, AD04, AHR05, ANA20, ADJ18, Asl18, ALK19, AWK^{+}20, BBMA22, BV19, Cai06, Cai07, CA19, CW16, DI13, DVFA23, DD11, DJ10, DLL10, Egb18, EE22, FKRG10, FCGL19, FHRS18, GVC20, GW97, GC23a, GH11, GCGZ10, GY13, HBN96, HCGFOON20, HLL15, KAGAM23, KEA11, Kar03a, KWA^{+}10, LLL14, MG11a, MG11b, MOBA21, NK10, Ole20, Öz00b, Öze10, PB97, Pak02, PKB11, PS15, PT21, PLLRCC20, RAK22, Raw14, RMBTGH^{+}20, SSV23, SAK01, Sal02a, SAI04b, SI04, Sal06a, Sha99, SKB22, SAH98, SBB23, SJ19, TE05, TYS13, TY15, TB16, TMB96, UAT19, UAI21, YG15, YG11, YBTH17, YW24, ZZ10, Zhu10]. Nonlinearities [CML10, Dal11, MG11a, PBY97, PB99]. Nonlinearity [AMAGNA23]. Nonlinearly [LYYX15, UY11]. Nonlocal [Akh04, CDA10, FMR10, GZ20, SM21, UC19]. Nonnegative [ÖT02]. Nonstandard [Egb18]. Nonsymmetric [SSA22a]. Nonsymmetrical [Alt10]. Nonuniform [ANJ20]. Nonylphenol [Bas99]. Normal [AE16, All17, Den16, SK05, ZZ17]. Normalized [Mus16]. Normed [Sav14, Sav16]. Norms [BS02, SB02, STB04, TB04]. Northeastern [KPSS18]. Note [EA22, ID14a, Kil04, RRO16a, TB04]. Notes [GLS18, KWP11, Sol23]. Novel [AN19, CBT21, DA19, GT22, Gha18a, GERV19, KG13, LP17, Liu11, MFF20, PMF20, ZZ17]. Nozzle [MWM16]. NSGA [ZWZ16, MFF20, NGBLC22]. NSGA-II [ZWZ16, MFF20, NGBLC22]. Nuclear [EGIK97, MO04, Man04, MKMB11, OTŞ99].

Nuclei [BK10, KSK05, KSK01, OTŞ99, SBA^{+}04, TU97, TM11]. Nucleus [BKA08, BUKM96]. Null [YH05]. Number [CB04, CBSTG13, CC11, CGCC11, EI05, EI06a, Fu16, GCCMD23, Tur16, Wu17]. Numbers [AA03, AAS11a, AAS11b, Erg01, FMF20, FbF16, LY15, QSC19, SN14, Sav99, Sav00, Sav10, Sav12, Sav13, YBYL21, dlF21]. Numerical [AV02, AA02, AAD11, AB99, AVC23, AGANA22, Amo21, BY16, BG18a, BS13a, BAUO09, BA11, BSSK12, BDMS19, Cai06, CCY08, CST20, CC22, DSG19, DSI16, DEL23, DHK23, EE22, ES13, FP21, GZ20, GGS11, HK23, HFAAM^{+}22, İBB98, KZAN98, KK97, LB20, Loj22, MV21, MWM16, MM10b, OA09, PT21, PKC22, PF10b, QCdlFL^{+}23, QRdlFS22, RRO16b, SSA22a, Sal06b, SGA18, SJ22, ST11, Sta23b, TO13, TSC11, VA12, YW24, ZS04]. Numerics [AMAGNA23]. Nusselt [EI05, EI06a]. Nutritional [KE98, KVR23]. Nyström [DSI16, DRKW21].

O [AR23, DEL23, JKS23, KB03, KF19a, KNRC22]. Object [ÇŞÖG13, Cor19, GWC98, KUE13, WZ17, YL16]. Objective [AA16, ARCRMM20, BMV22, Bay03, BCS20, BS23, CMS20, CATCdlF20, CACRF^{+}21, CS22, CS19, DE23, EPLGGS^{+}21, FSHRCV^{+}21, JLST04, Mak99, MFF20, NGBLC22, PD18, SSGCPR23, SBDN^{+}22, VHMMG20, VMRVF^{+}23, WED^{+}23, WJ23, XS22, ZWZ16]. Objectives [GCCMD23]. Objector [Xia10]. Objects [ÇÇÇ07, NLMU19, TDCT11]. Observable [BUAM19]. Observations [MR05]. Observer [PPRVC20, RMBTGH^{+}20, TJCGAPA18, WCH23, ZP11]. Observer-Based [WCH23]. Observers [HCdCTAAMEJ20]. Obtain [Gun96]. Obtained [SQA11]. Occurring [TCC11]. Octave [CST20]. Octave/MATLAB [CST20]. Odd [CTJJ22, JHA^{+}22, MHC21, PBY97, PB99].

ODE [ÇKÇK10, İBB98, MM10b]. ODEs [DSD^{+}21, FMR10, FMM10]. Off [EB00]. Off-Highway [EB00]. Offset [ZZ17]. Oil [ALFBPLE20, Sar09]. Olkin [MC20a, MC20b]. Omega [Alm10]. One [AK19, ÇT09, DKÇ97, Egb18, EBY06, KA11a, PRFP^{+}23, PMR14, YA14]. One-Compartment [Egb18]. One-Dimensional [AK19, KA11a, PRFP^{+}23, ÇT09]. One-End [PMR14]. One-Way [YA14]. onto [BS13a]. OP [LZ21]. OP-Mapped [LZ21]. OPC [Yil11]. Open [EÇ99, GPEG10, GCPV18, IE04, ÖD06, TY11, TA99, VSAMMM23]. Open-Set [VSAMMM23]. Open-Source [GCPV18]. Operated [Akb09]. Operating [BK09]. Operational [BS15, Mit21a, Mit21b]. Operations [KS18, SBF96]. Operator [FTGSNV21, Gun96, ID14a, SB02, YW24]. Operators [FMR10, GW97, GH11, Kle18, MK10, RFQCMMCR23, Sal02a, YZHY17]. Optical [BPC^{+}23, Kur98]. Optimal [AAGH21, AERE18, Bac13, Bak11, BBMA22, BP99, CST20, CCWY10, DKÇ97, EARE18, Gha18a, Gha18c, GK18, GA99b, HH07, Kha03, KKL18, KCMLTR18, LCB18, Mit21b, MFF20, OMX18, Pak16, PS22, PD18, PALL18, SRLEPVP20, SKB23, TCC11, Yaz16, YS17, ZMA23]. Optimality [AFN^{+}19]. Optimisation [SC19, SBF96, TEK11]. Optimization [AA13a, AMR^{+}23, AB19, BGRV21, BMV22, Bay03, BCS20, BS23, CAQCMM22, CATCdlF20, ÇA06, Çun08, DRH21, DE23, DPC09, EE22, EÇ99, EPLGGS^{+}21, GCCMD23, GLA05, GLAA08, GCPV18, GGMMDI20, HDAA20, Jab18, LVHL22, LDCST21, LL17, MM17, Mak99, Man16, MARL21, MSPC23, MFF20, NGBLC22, Nur03, PYR22, Pek96, PFR19, PKC22, QCdlFL^{+}23, QRdlFS22, RLY^{+}17, ŞİH13, SH18, SBDN^{+}22, TY11, VVV23, VHMMG20, WZT20, WED^{+}23, WJ23, XS22, YBTH17, Yen05, Yen14, Yil11, YKG20, YYWX16, ZWZ16]. Optimizer [SH18, TCC^{+}23]. Optimizing [EPLGGS^{+}21, GZ14, dlFH19]. Optimum [ÇYAÖ05, Sar09, YDD11]. Option [LLL14, TSS17]. Options [FIJ21, Hon20, Mol18, YS17]. Orbitals [KTK97]. Order [AAD11, AADS11, AO16, Akh04, ACM19, AM23a, Alh21, AE02, BDMS19, BBH19, Büy16, Cai06, CA19, CK18, Dal11, Den16, DRCSF22, Ers10, EM08, Ert11, FSV^{+}22, FCGL19, GLS18, GGS11, Jia17, JLW20, KRD^{+}22, KKS11, KAAK10, KO15, KKY05, MSSZ20, MM10b, NNW14, OWO^{+}23, PBY07, PBY08, PAB11, PT21, SQA11, SPD18, SGZ10, Sar11, SGZ11, Sav16, SKB23, SBB23, SJ19, TCC^{+}23, TYS13, TY15, UYAA22, YÖS10, YG15, dlFH19]. Order/Grade [Ers10]. Ordinal [CACRF^{+}21]. Ordinary [AO16, Akh04, DSD^{+}21, Doğ12, GK11a, KO15, SQA11]. Organisational [CCC^{+}18]. Organization [AA16]. Organizational [LTYG17]. Orientation [Ese01, KE01]. Oriented [DCC23, GWC98]. Originating [QM16]. Orthogonal [YQG22]. Orthonormal [BBH19]. Orthorhombic [Mub23]. Orthotropic [AB06, BT19, KÖAG00, TZ19, YK96a]. Oscillating [DRKW21, ŞÖ11]. Oscillation [Xu10]. Oscillations [FKRG10]. Oscillator [CC09, Kol11, KA16, ZZ10, ZH10]. Oscillators [Cai07, CML10, MG11a, NM10]. Oscillatory [CKLZ21, DSI^{+}18, DSI16]. Out-of [ÖÖT04]. Out-of-Plane [VVN23]. Outer [VRMJG19]. Outliers [KN18]. Output [CW16, JY23, WCH23]. Output-Feedback [JY23]. Output-Predicting [WCH23]. Outputs [KYG20]. Outs [EY13]. Outsourcing [CLC16]. Overall [DJL^{+}05]. Overland [ELT21]. Overlap [ST11]. Overlapping [MMS19]. Overview [CGMCCA23, Ken11].

Owa [ID14a]. Owned [LTYG17]. Oxides [KR23]. Öztürk [Çat10].

PA6 [PPRF22]. PA6/GNPs [PPRF22]. Package [Kas00, Mut98]. Packages [PF10a, PF10b]. Packed [ZS04]. Packet [AK09]. Padé [ÇKÇK10, PKO11]. Pair [DSI16, DRKW21]. Pairing [KSK01]. Pairs [QM16, ST18]. Pallet [DM21]. Pancreatic [GVC20]. Panel [DJY^{+}22, ESNB19]. Panels [BIM^{+}20, CC22, TMB96]. Paper [RSF23, YS02]. Paperboard [AVC23, CC22]. Parabolic [BY06, KAAK10, OA13]. Paraboloidal [UY03]. Paradigm [Gha18b, Gha18c]. Parafoil [GT22]. Parallel [BCGS23, Haj04, IE04, JLW20, LP17, PSW24, RFQCMMCR23, TA99]. Parallel-Machine [RFQCMMCR23]. Parallelization [CSOPM24]. Parameter [AWK^{+}20, Bak11, BMV22, CC09, CK12, ENVPF19, Fu16, Gha18b, HMDG11, KY05, KN18, PS22, PC23, Rob22, Saj18, Sul08, TABT19, Yaz16]. Parameter-Dependent [Yaz16]. Parameters [AMMMM^{+}23, Cai07, CZC17, EPLGGS^{+}21, FSH^{+}23, GM03, GK11b, JL24, Kan16, KHL06, KÜM08, KA11c, Li16, OM20, Sha19, TWS17, YZBZ16, Yil11, YYWX16, ZG17]. Parametric [BG04, FKRG10, KHL06, KN13, LLTC17, LL17, MCNM22, ORC19, ŚLO19, TEK11, ZCC15]. Parametrically [EESK^{+}02]. Parametrization [Bil04]. Pareto [BCS20, CS22, DRS17]. Parking [IGSB18]. Part [ÇTY96, ES06b, ES06a]. Partial [BBK20, Gha18b, KAAK10, MK10, NK10, Özd96b, SZTY23, VA12, WN21]. Partially [ZS04]. Particle [Ala11, HM23, Man04, WJ23, YL16, ZLZ17b]. Particles [AKT04, AB99, ETL22, SAM^{+}22]. Parties [GDE12]. Partner [DAKCPACC19]. Parts [Çiç07, DB17a, DB17b, HLX^{+}17, HLX^{+}18, KÜM08]. Pascoletti [BMV22]. Pascoletti-Serafini [BMV22]. Passive [ES06b, ES06a, IMWY18]. Past [FJ21, KNRC22, MM10c]. Pasternak [TB16]. Patches [CY05, Kus19, WZ17]. Path [Boo17, MBA11]. Paths [JLS^{+}05]. Patients [FSH^{+}23]. Pattern [GZ20]. Patterns [Cui07, DNS23, GC23b, PRT^{+}19, Yav22]. Payment [HH07]. Payments [TWL10]. PDE [ABH^{+}22, BGRV21, BMV22]. PDE-Constrained [BGRV21, BMV22]. PDEs [NK10, Raw14]. Peaceman [IZS04]. Peak [CDDM23, LZC22]. Peak-Over-Threshold [CDDM23]. Pediatric [KE98]. Peiffer [Mut98]. Penalty [YBTH17, Yen05]. Pendulum [ES06b, ES06a, ŚLO19]. Pentanacci [SN14]. Peptides [SHFSGB^{+}21]. Percent [EG11]. Perceptual [BVvdH^{+}23]. Perforated [Mec22]. Performance [AÖ09, Alm10, BA03, ÇD03, ÇT12, ESNB19, EB00, EBD11, HK23, KRD^{+}22, KD03, PGB^{+}11, ŞÜ13, SGEP06, SÖS13, VMRVF^{+}23, ZLLX20]. Period [TCC11]. Periodic [GNK19, KKA04, LDL10, SD13, Yao15, ZPL16]. Peristalsis [IYKA20]. Peristaltic [ADE99, HAMM10, HA10]. Permanent [BODH20, GGT^{+}23, Jab18, MMF18, OD19, VRMJG19, WFV21, ZLLX20]. Permanent-Magnet [BODH20, MMF18, OD19]. Permeability [GGJ10, MMF18]. Permeable [RSS^{+}21]. Permeance [Mes96]. Permissible [TWL10]. Permittivity [HAK03]. Persistent [NAA20]. Perspective [GNI19]. Perturbated [SAH98]. Perturbation [KG10, Mer09, OSAAA18, PKB09, PAB11, Pak16, PMR14, RFD^{+}10, TSS17, TB21, UAI21, dlFH19]. Perturbation-Iteration [PAB11]. Perturbations [DZA19]. Perturbative [ACC20, İB00, PBY07]. Perturbed [HLL15, Kar03a, SAK01]. Petch [MAE97]. Petrov [KAS19a, ST18]. Petviashivili [MDNN10a]. Petviashvili [LDL10]. Pharmacokinetic [Egb18]. Phase [AGKT21, AO17, BY17, BODH20, BBPL19a, DRKW21, Dil02, JLW20, Mos18, Oze03, Umu11, WLW^{+}10]. Phase-Field [JLW20]. Phase-Fitted [DRKW21]. Phase-Plane [Oze03]. Phenomenon [DJY^{+}22]. Phonon [Bül99]. Phononic [FLYY21]. Photocatalytic [RJ17]. Photograph [ZLZ^{+}17a]. Physics [Bro19, DCAR21, DL23, EPDG21, GFPF22, HM23, MQ18]. Physics-Informed [DL23, HM23]. PID [TCC^{+}23]. Piecewise [ÇKÇK10, Qua18, ZBBZ18]. Piecewise-Smooth [ZBBZ18]. Piezoelectric [Mec22, UY11]. Pigs [HGS^{+}22]. PIM [GCGZ10]. PINN [LRL23]. Pipe [TGS98, ZCC15]. Pipeline [SRLEPVP20]. Pipes [AO17, Has01, Sin10]. Pit [EÇ99]. Pitaevskii [GN19]. Pitch [EÜA13]. Pitchfork [PS20]. Placement [Bak11, MFF20]. Plan [BO10]. Planar [ADE99, Man16, TDCT11]. Planck [AB23]. Plane [AJ20, CC22, Du 22, FLYY21, LVBV19, Oku11d, ÖD06, Özb02, Oze03, ÖÖT04, TA99, VVN23]. Planes [IE04]. Planning [ÇTY96, GT22, KPSS18, TCA96]. Plans [ZE16]. Plant [GÇÖ12, HBS^{+}17, Kar03b]. Plasma [BY17, RKX23]. Plastic [ABO^{+}16, BKOS11, BSSK12, ÇŞS01, ÇKAK04, Ese01, KÖAG00, KE01, Oku11b, Oku11a, ÖG13, Özb02, Özc00, SS06]. Plate [AYG14, AS04, AGANA22, AK02, BS13a, DEL23, DMN22, ES13, FSV^{+}22, FJ21, Kop02, KNRC22, RRR^{+}22, RFD^{+}10, SAI04a, TA99]. Plates [AG03, AB06, BT19, CK07, EY13, KPSZ24, Mec22, MMS19, MLBR20, Oku11d, ÖG13, Özb02, TA96, TDT^{+}11, YK96a]. Platform [KNHK20]. Ply [BT19]. Poincaré [PKB09, PKB11, PS15]. Point [AERE18, ANK23, Asl18, HBFM22, İTİO08, KC08, Kar12, KAM20, MV21, PB03, PTLZ18, Pop98, TWS17, TUK19, dlFH19, dlF21].

Points [DM99, LLTC17, ŞÜ13]. Poisson [Aki11, AM23b, IA04, MVC23]. Pol [CTL14, CC09, NM10]. Polar [ASA20, DB17b]. Poles [AS04]. Policy [CLC16, HH07, TWL10]. Political [GDE12]. Pollination [PS22]. Pollution [RJ17]. Polygons [PKC22]. Polyline [TT11]. Polymer [Ese01]. Polymers [Can02]. Polynomial [Pak16, Qua18, Sol19, SJKCD20, YQG22, YGS13]. Polynomials [BY16, GGS11, HMDG11, KK13, KKS11, MAT20, MDNN10a, Saj18, Shr01, Shr02, Shr04, Shr05]. Pool [JSTGV^{+}19]. Pool-Based [JSTGV^{+}19]. Poppy [DPC09]. Population [AERE18, GPMRVM21, GK12, PALL18, SÇGA12, SK05, Sul08, UAT19]. Populations [AW19]. Porous [AYG14, AMAGA22, ALFBPLE20, DSI^{+}18, Ers10, JKS23, KPSZ24, LH10b, MAKH10, MLBR20, MOBA21, OSAAA18, RKR^{+}22a, RKR^{+}22b, SSV23, SWJS22, TKK^{+}22]. Porous-Elastic [AMAGA22]. Portfolio [EPLGGS^{+}21]. Portuguese [MCB19]. Posedness [MOBA21]. Positioning [HB99]. Positions [CGM^{+}23]. Positive [Fu15, Yao15, ZPL16]. Possibilistic [Alp02]. Possibility [Has02]. Possible [ÖE03]. Postponement [CKCC16]. Potential [AEB01, ÇÇ03, HBFM22, Kar03a, KÜM08, PRFP^{+}23, PSW24]. Powder [TDC08]. Powel [HPA13]. Powel-Eyring [HPA13]. Powell [DSI^{+}18]. Power [AYG14, AHM99, Aki11, AB19, GPEG10, GÇÖ12, HXY15, KA96b, KA09, LDCST21, MFF20, MAH10, NAC23, RC16, TGK96, TÇ04, Üna96, UC10, YUKA03, ZMA23]. Power-Law [AYG14]. Powers [RS21]. Practical [AÖ09, CLY10, Gha18c, Has02, LC18, SC19]. Pre [DUM04]. Pre-Aged [DUM04]. Precariat [Dan22]. Precipitation [ELT21, IWH21]. Precision [MM17, MV21, ZZ17]. Preconditioning [AM23a, CBM^{+}19]. Predator [DSMP19, MPP^{+}24, PAR23, SAM^{+}22, ZPL16]. Predict [ACGC21]. Predicting [HIVH21, WCH23]. Prediction [Alt12, CYAK03, ÇT12, ÇSB10, DK03, EG11, GZY20, GPEG10, JMB22, KD03, MM17, MSSZ20, MVC23, Mit21a, SHFSGB^{+}21, SGEP06, TDC08, TMK21, TB21, YL16, Zha14, ZPW^{+}22]. Prediction-Based [YL16]. Predictive [ACL^{+}22, FYuR^{+}23, Kay16]. Preestimation [UKY96]. Prefabricated [TDT^{+}11]. Preface [LB20, Loj22]. Preferences [CACRF^{+}21]. Preinvex [TAS^{+}23]. Preliminary [SGBR23, SM99]. Preparing [KNHK20]. Preprocessing [SKAK05]. Prescribed [GÖD03, ZLLX20]. Presence [HFAAM^{+}22, KN18, MFF20]. Presentation [Kay98]. Presentations [Kaz04]. Preserved [XWCF19]. Preserving [ACGC21, Sar11, TWS17, VC23b, dlFH19]. Pressure [ALFBPLE20, BY06, CA19, EG11, GA99a, LVHL22, TMK21, Yür02]. Pressurized [AO17]. Prevalence [BDC^{+}18]. Prevalent [SGCRRL^{+}23]. Previous [AJ12, JLS^{+}10a]. Prey [DSMP19, MPP^{+}24, PAR23, SAM^{+}22, ZPL16]. Price [ACC19, GZY20, UC10, YS17]. Pricing [Sho22]. Primary [JWP23a, RM20]. Prime [Erg01, FMF20, OV20]. Primes [OV19, OV20]. Primitive [Sal06b]. Primitives [PSF22]. Principal [BG99, Can14, Muc01]. Principle [Muc00, RC16]. Printed [AMR^{+}23]. Printing [TUK19]. Priori [KGCFK22]. Prismatic [NUM^{+}19]. Privacy [XWCF19]. Privacy-Preserved [XWCF19]. Probabilities [KCMLTR18]. Probability [Kas19b, PLW21, QM16]. Probe [CBT21]. Problem [AS04, AERE18, AA16, AK18, AM23a, AA13b, BŞ13b, BGN13, BÇE02, Car21, ÇT14a, Cha05, CKCC16, CLC16, DPC09, DM21, EKAÖ11, EB02, EBY06, EARE18, EE22, EPLGGS^{+}21, FSHRCV^{+}21, FHRS18, HM10, HH22, IGSB18, IE10, JLST04, JM03, KSTJ18, Kah11, KBE01, KAS19a, KN13, Koi19, KSS19, KB11, KAEÖ13, Nur03, OA13, Öze00, PÖ98, PR21, PRFH23, PALL18, RFQCMMCR23, RC16, Sal02a, Sha99, SA97, SWJS22, TKGS20, Tan08, Ugu13, VSAMMM23]. Problems [AV04, Akh04, BBMA22, BV19, BCS20, BCGS23, BD22, CAQCMM22, CKLZ21, DSI16, DJM18, EM08, Ert11, FCGL19, GFPF22, Gha18a, Gha18c, GNK19, GB00, JLW20, KRD^{+}22, KAM20, LS20, Man16, Mat19b, Mat19a, MR19a, MMS19, Mor22, NU19, NNW14, PS15, PMR14, PSW24, SBB23, SKK19, TY15, WZT20, WED^{+}23, YKG20, ZG17, ZWZ17]. Procedural [SÇGA12]. Procedure [BV19, Mos18, PBY97]. Procedures [BAS96, EBD11]. Process [Abi03, AB96a, ABKU10, AFGV20, Bas99, ÇTY96, Cha08, CDDM23, CCWY10, GHN11, GDE12, HSG^{+}20, HBFM22, ICDM21, JB15, KST10, Kha99, KAKB08, OYCS11, PVCRRV^{+}21, TKCC19, ZP11]. Processed [TÇ06]. Processes [IÇÜ12, Kar11]. Processing [AWK^{+}20, BGA10, Emi03, EGB99, MBA11, PRFP^{+}23]. Procudere [All99]. Produced [Kur98]. Product [CKCC16, HAH^{+}20]. Production [CYAK03, CGCC11, HIVH21, VHMMG20]. Productivity [CÖGD11]. Products [BM21, Kil03]. Profile [Bül99, PVCRRV^{+}21]. Profiles [OFCP21]. Profit [RLY^{+}17, Tsa09]. Program [BG04]. Programmer [Gha18c]. Programming [Alt09, BPC^{+}23, GWC98, GU05, JLST04, JSTGV^{+}19, Kah11, KAY08, KHL06, KA11c, LGT^{+}18, MR19b, RC16, SBF96, Tan08, TMK21]. Programs [Mol18, NKI19]. Progressive [LZC22]. Projection [AWK^{+}20, JZL10, KAM20, Sal02b, Sal04, TKGS20]. Projects [Yur99]. Prolate [DAD^{+}18]. PROMETHEE [ASA20]. Prometheus [Übe04]. Prometheus-L [Übe04]. Prony [HH22]. Proof [WH10]. Propagation [FLYY21, GS01, KRD^{+}22, PD19, Shi06, ZLZ17b]. Propellers [DHK23, ESNB19]. Proper [AFN^{+}19]. Properties [AT05a, AA03, Alt10, Alt12, BM11, CY05, CUA04, Cet16, ÇÇ03, CTAA11, DK03, DNS23, Eİ06b, GOA03, HC22, ICDM21, JAT^{+}21a, KYK02, KHR^{+}22, KLRA10, Mus16, NAC22, NAC23, PŞ05, PKB09, PCCT11, QM16, SKM^{+}22, ŞİH13, ST22, TU97]. Property [LF19]. Proposal [AA16, NÇK14]. Proposals [NSCTPW22]. Proposed [GDE12]. Proposing [AJ12, JLS^{+}10a]. Protocol [BBSC18]. Protocols [BBPL19a]. Protons [ENT05]. Prototype [EÜA13]. Proximal [SI04, YKG20]. Pruning [Aks05b, HRM19]. Pseudo [AM23b, ED11, ICDM21, OA13, Yur99]. Pseudo-Galilean [ED11]. Pseudo-Parabolic [OA13]. Pseudo-Poisson [AM23b]. Pso [Zha14]. Psoriasis [GK18]. Public [HSG^{+}20]. Published [CGM^{+}23]. Pulmonary [LVHL22, UÖKE14]. Pulsating [ZCC15]. Pulsations [MSSZ20]. Pulsed [BBPL19a]. Pumps [SÇ10]. Punch [TCALK20]. Punched [Cui07]. Punkinje [Öze02]. Pure [Kaz05]. Purkinje [Oze03]. Put [FIJ21]. PV [JL24, TCC^{+}23]. PVT [CATCdlF20]. Pylori [KAY09]. Pythagorean [AHID18, ADS19, AN19, SA19].

Q [Ull17, TAS^{+}23]. q-Digamma [TAS^{+}23]. Q-Sequence [Ull17]. qLPV [LESEVP^{+}20]. Quad [ÇŞÖG13]. Quad-Tree [ÇŞÖG13]. Quadratic [CT01, Dal11, EESK^{+}02, HDAA20, Ken96, PY00, PS15]. Quadrature [KA11a, Sar11]. Quadraure [CÜ04]. Quadrilaterals [FS23]. Qualitative [GOOM18]. Quality [BGG96, CKCC16, DRS17, FS23, GERV19, KA11b]. Quantify [JWP23a]. Quantile [MMM23, NAC22, NAC23]. Quantities [FMM10, NNM10]. Quantization [KSS19]. Quantum [Boo17, ÇT09, SM21, Zha14]. Quasi [ALFBPLE20, BV19, DSU22, EE22, Oku11b, SI04, Sal04, SAH98, YZHY17]. Quasi-Analytical [ALFBPLE20]. Quasi-Interpolation [YZHY17]. Quasi-Isotropic [Oku11b]. Quasi-Newton [DSU22, EE22]. Quasi-Static [BV19]. Quasi-Variational [SI04, Sal04, SAH98]. Quasivariational [Hus01]. Quaternions [AHVG23]. Quay [PR21]. Queries [Mat10a]. Query [Emi03]. Questionnaire [BM21]. Queue [AK09, KHL06, LZ17, WT15]. Queueing [Cha97, Cha08, CT14b, Sul07]. Queueing-System [Cha97]. Quintic [YZHY17].

R [AÖ09, KIY^{+}23, KPSZ24, PD19, CPA20, DD11]. R-Functions [KPSZ24]. R-L [AÖ09]. R-L-C [KIY^{+}23]. R-Mediated [PD19]. R/C [DD11]. R13 [UÇ96]. Rachford [IZS04]. Radar [Rou21]. Radial [BP21, EE22, KÜM08, MKMB11, NNM10, PRFH23, UAT19]. Radiated [JS22b]. Radiation [DSI^{+}18, ETL22, GS02, GS10, NKS22a, NKS22b, RKS23, SSV23, Sha19, Shi06]. Radius [OTŞ99]. Rails [RWL23]. Random [CPP19, CLC16, CRA^{+}21, GB23, GLS18, JZL10, KBH06, Liu11, LS20, TCC11, VC23a]. Randomized [KA11c]. Ranked [JLKK12]. Ranking [AAS11a, AAS11b, JLS^{+}10b, LY15, LNBM10]. Rao [Dil02]. Rape [HBS^{+}17]. Rapid [Gha18b, Gha18c]. Rat [DL24]. Rate [AK09, CS19, FA22a, RC16, ŞE17, WT15, Wu17]. Rates [BAUO09, CMM19, DHK23, HFATZR^{+}20]. Rating [LL24]. Ratio [CYAK03, ÇYAÖ05, DSG19, HK23, KC08, OV18]. Ratio-Dependent [DSG19, KC08]. Rational [PTLZ18, RS21, VC23b, YÖS10, YZBZ16]. Ratios [BUKM96, HBFM22]. Raw [JB15].

Rayleigh [Mub23]. RBF [TO13]. RBS [CB04]. RC [TDT^{+}11]. Re [GGT^{+}23, IMWY18, KSS19]. Re-Attachment [GGT^{+}23]. Re-Quantization [KSS19]. Reacting [TDCT11]. Reaction [BBK20, EAT16, ETL22, GC23a, GZ20, JKS23, LV19, RAK22, RAA10]. Reaction-Diffusion [EAT16, RAA10]. Reactions [KR23]. Reactive [TGK96]. Reactor [Übe04, YUKA03]. Ready [NGBLC22]. Real [DM21, Erp12, KUE13, SQA11, Saj18]. Real-Life [DM21]. Real-Time [KUE13]. Reality [BVvdH^{+}23]. Realization [ÇA10]. Realized [GP21]. Recharge [ELT21]. Reciprocal [NT04]. Recognition [CGMCCA23, HBS^{+}17, Kar03b, MRIR20, VSAMMM23]. Recognize [AHVG23, AB96a]. Recognizing [Kar11]. Recommendation [DCAR21, MECRMT^{+}21, ZE16]. Recommender [JMS18]. Reconfigurable [PLLRCC20, SBDN^{+}22]. Reconstruction [CBT21, KY96, WLW^{+}10]. Recordings [VÇÖ11]. Recovering [ÜÖ03, ÜC11]. Recovery [BA03, Öze02]. Rectangular [AB06, CK07, DMN22, EL04a, IE04, SGE98, TA96, TDT^{+}11, TMK21]. Rectifier [Akb09]. Recuperative [ZMCL24]. Recurrence [BdCCF19, BBPL19b]. Recurrent [AA03, WCH23]. Recursion [GH11]. Recursive [Cui07]. Recursively [TUK19]. Reddy [MRG23]. Redefinition [NK10]. Reduce [BDC^{+}18, CCC^{+}18, CRA^{+}21, Oku11c]. Reduced [ACM19, BMV22, BDMS19, CA19, DRCSF22, KO10, KF19a, KF19b, MSSZ20, Raw14]. Reduced-Order [BDMS19, MSSZ20]. Reducing [AK09, GCCMD23, ŞÖK11]. Reduction [ASJQ24, AM05, BDMS19, BBH19, ÇGV02, FHRS18, IMWY18, KSM23, KR23, Kol11, KA16, MS05, MM10b, PD18, TEK11].

Reductions [KSM23, MAH10, MJK12]. Redundant [Sam99]. Reference [DM99]. Reflectance [ZPW^{+}22]. Reflected [SZTY23]. Reflector [UY03]. Refractive [Kur98]. Refrigeration [EPDG21]. Refuge [PAR23]. Regenerative [EPDG21]. Regime [Aki11, ÇT09]. Region [BK10, BMV22, BP21, CJL18, KY96, KPSS18, Şen05, TM11]. Regression [Erp12, KY13, KTD12, MVC23, NAC22, Pak16, PMF20, UÖKE14]. Regressions [NAC23]. Regridding [VPJ22]. Regular [Kaz04, Kus19, Özy14, ST18]. Regularization [Alt16, Kay16, YD11]. Regularization-Homotopy [Alt16]. Regulate [MPA17]. Rehabilitation [CMS20]. Reinforced [AG03, Bağ10, BSSK12, ÇŞS01, GEKB13, Mub23, Oku11a, Oku11c, OF19, TA96]. Reinsurance [Kas19b]. Rejection [LOSCO^{+}20]. Related [GK12, Kar03a, Rob22, Sal09, Tur16]. Relatedness [Sal09]. Relation [Erg01, JLH21, MAE97, UT96]. Relations [CCGHdCJ21, KK13, LF19, OPUÖ11, Shr01]. Relationship [Bağ10, IWH21, ZZ10]. Relationships [CY05]. Relative [LZC22]. Relatively [KAM20]. Relativistic [ALK19, CO16, KM10, ZH10]. Relaxation [TKGS20, TZ19, YB97]. Relaxed [KAM20, Sal02a, Sal02b]. Reliability [BS23, CMM19, CT14b, JHA^{+}22, JLH21, PYR22]. Reliability-Based [BS23]. Reliable [ZS16]. Reluctance [BBD^{+}18, Mes96]. Remeshing [DJM18]. Removal [RJ17]. Reneging [Cha08]. Renewable [AARW22]. Renewal [Aki11, ABKU10, KAKB08]. Renovations [MK96]. Repair [CCC08]. Repairable [CT14b]. Replenishing [TCC11]. Replenishment [CBSTG13, CC11, TWL10]. Replication [NISC18]. Reporting [EA22]. Representation [ZLZ^{+}17a]. Representations [Car21]. Reproduction [Wu17]. Requests [BPC^{+}23]. Requirements [FdSPW20, KE98, TÇ04]. Research [CW16]. Reservoir [ALFBPLE20]. Residual [BKOS11, GC23a, Özb02, TSC11, YGS13]. Resin [AT05b]. Resistance [KD19]. Resolution [BBK20, CBT21, EARE18]. Resolving [Hak22]. Resonance [YUKA03]. Resonances [BÖÖ11, ZCC15]. Resonant [ÇT09, SGE98]. Resonator [KIY^{+}23, Yil11]. Resources [AERE18, KKL18]. Respiratory [FYuR^{+}23]. Response [AA13a, AB06, ACGC21, EESK^{+}02, Ese11, FA22b, Rob22, SMB96, TMB96, Yen14, ZPL16]. Responses [DJ10]. Restarted [TW18]. Resting [BE99, BÇE02]. Restoration [AB03]. Restored [AT05b]. Restricted [AERE18, Bac13, Tay01]. Restriction [ZGV10]. Result [Kay15, ŞT96]. Results [AMAGA22, AMAGNA23, DRH21, Kha99, KY13, MOBA21, TM11]. Retailer [HH07, HH07]. Retaining [MMF18]. Retardation [GÖV96]. Retarder [EB00]. Retraction [HLX^{+}18]. Retrial [LZ17]. Retrieval [ZLZ^{+}17a]. Returns [GP21, KJZL12]. Reversals [MPP^{+}24]. Reverse [Mit21b]. Reversing [AAGH21]. Review [AJ12, Aks08, ACL^{+}22, EGB99, EPDG21, JLS^{+}10a, OOI^{+}22, SW22, TCM21]. Reviewers [MCA17, MCA18, MCA19, MCA20, MCA21, MCA22, MCA23]. Revision [OD19]. Revolution [RM20]. Reward [ABKU10, KAKB08]. Rework [CBSTG13, CCL06, CCC08, CCWY10, CGCC11]. Rheumatic [FSH^{+}23]. Rhythm [HCGFOON20]. Ribs [AG03]. Ricci [GE00]. Rice [HIVH21]. Ricochet [Xia10]. Riga [DEL23]. Rigid [BE99, GB18, HB01, KBE01, KWA^{+}10, OC05, ÖS05, ŞÖK11, ŚLO19]. Rigorous [YD11]. Ring [AUA03]. Ring-Shaped [AUA03]. Rings [KCMLTR18]. Riser [BC97]. Risk [CMS20, CCC^{+}18, CMG22, JZL10, JAT21b, Liu11, Mit21a, Mit21b, SGC23, WOL15, YS17, ZGV10]. Risk-Informed [CMS20]. Risks [CRA^{+}21]. Rivalry [BVvdH^{+}23]. Rivers [RMBTGH^{+}20]. Rivets [KTC99]. Rivlin [VVV23]. RLC [BÇ13]. RNA [KPTM20]. Road [OOI^{+}22]. Robot [LOSCO^{+}20, VJ23, VC23a, Yag02, YS02]. Robotic [Sağ98]. Robots [LV23]. Robust [ASK21, BD22, GGMMDI20, KTD12, LCB18, LESEVP^{+}20, MR19b, SZ23, WZ17, Yag02]. Robustness [JLH21]. Rocket [ANJ20]. Rod [KWA^{+}10, XS22]. Rodrigues [GK11b]. Rods [Çev10]. Role [KNRC22, MPP^{+}24, PM22]. Roll [TEK11, UAI21]. Roll-Damping [UAI21]. Rolling [Tay04]. ROM [BGRV21, LC16]. ROM-Based [BGRV21]. Roofs [TMK21]. Root [BBMA22, CSOPM24, GGT^{+}23, PBY07, PBY08]. Root-Finding [PBY07, PBY08]. Rope [IE10]. Rotating [AJ20, AÇGJ99, GS01, GS02, GS10, KH15, Shi06, TK04, YK96a]. Rotation [BUKM96, ÇS10, Ers10, KÇ96, RKR^{+}22b, UK96b, UT96]. Rothko [AHVG23]. Rotor [GÇÖ12, JH13, MSSZ20, STA23a, VRMJG19]. Rotor-to-Stator [STA23a]. Roughness [SRLEPVP20, Ull17]. Routes [Egb18]. Routing [AK18, BPC^{+}23, PRFH23]. RSA [OV20]. Rub [JH13]. Rub- [JH13]. Ruin [Kas19b]. Rule [Den16, Emi03, EGIK97, KUÖ05, SSGCPR23, WN21]. Rule-Based [Emi03]. Ruled [CUA04, EG00, ED11, GE00, KYK02, KÇ11, Özy98, ÖY01, Yay00]. Rules [Aks08, AG12, DT08, Kaz05, ÖK05, SGCRRL^{+}23, UG00, Mat10b]. RULES-3 [Mat10b]. Rules3 [Aks05a, Aks05b, Mat10b]. RULES3-EXT [Mat10b]. Runge [AAD11, BG18a, DSI16, DRKW21, WZ23]. Rupture [SGC23].

S [HLX^{+}18, KA16, Rou21, HBS^{+}17]. SAFE [KRD^{+}22]. Safeguarding [BY17]. Safety [CDDM23]. Sagitally [GA99b]. Saint [ELT21]. Saint-Venant [ELT21]. Sales [CÖGD11, Tsa09]. Salesman [BŞ13b, KN13]. Sample [PYR22, SK05]. Sample-Based [PYR22]. Samples [CB04, Sul08]. Sampling [CMM19, CPP19, KPTM20, KF19a, KF19b, MGGHCR^{+}20, PYR22]. Sandwich [ÇDYS13]. SARS [GPMRVM21, UYAA22]. SARS-CoV-2 [GPMRVM21, UYAA22]. Satisfaction [Ken96, LTYG17]. Satisfying [AA03, BGG96]. Saturated [FA22a, SWJS22]. Saxon [KÜM08]. Saxton [KSM23]. Scale [AA13b, BP23, BBPL19a, CAQCMM22, GCGZ10, KJZL12, NGBLC22]. Scale-Free [BBPL19a]. Scales [Dal11, FBR17, PB99, PKB09, PKB11, PS15]. Scanning [mLkWbPxX17]. Scarce [PYR22]. Scattering [CJL18, HB01]. Scenario [HBFM22]. Scene [ACGC21]. Scheduling [Cha05, FSHRCV^{+}21, LP17, NISC18, PR21, RFQCMMCR23, SKK19]. Scheme [CD04, EE22, FJ21, GNK19, JY23, JWP23b, MPA17, PTLZ18, SBB23, TWS17, YW24]. Schemes [ANK23, Asl18, EAT16, LZ21, OA13, SGZ10, SGZ11]. Scholes [PMS03, TSS17]. Schrodinger [MDNN10b, HLL15, TD12, YW24]. Schur [Büy16]. Schwarz [Bad22]. Science [MM10a]. Sciences [TCM21]. Scientific [DB17a, DB17b]. Scimitar [AJ20]. Scimitar-Type [AJ20]. Scission [Can02]. Scoring [LGT^{+}18]. scppband [Kas00]. Scrap [CCC08, CLC16, TCC11]. Scrap-or-Rework [CCC08]. Screw [GK11b]. Seals [DE04]. Search [Bay03, HDAA20, JSTGV^{+}19, LBY^{+}22, PS22, PVCRRV^{+}21, TM11]. Searching [MBA11]. Second [AYG14, AO16, Akh04, AM05, CK18, Ers10, GLS18, KO15, SQA11, ST18, Yür02]. Second-Order [AO16]. Secondary [KPTM20]. Section [ÇDYS13, EL04a]. Sectional [TDT^{+}11]. Sections [Bağ10, TY11]. Sector [ACL^{+}22, GBİ14, WW20]. Secure [CLY10, NKI19, OV18]. Security [AA16, CD04, IGSB18]. Security-Constrained [IGSB18]. Seed [ZPW^{+}22]. Seen [PS19]. Seepage [JB15]. Segel [RZ18]. Segment [DB03]. Segmentation [ÇGK96, GT22, KGCFK22, SÖS13, UB10]. Segmented [ZLLX20]. SEIR [YA23]. Seismic [KG13, YG02]. Seizure [SKAK05]. Selecting [CMB19, SÇ10]. Selection [ADS19, BK09, Bor11, CTH^{+}22, DCAR21, GDE12, SGCRRL^{+}23]. Self [AT05a, DMN22, KAM20]. Self-Adaptive [KAM20]. Self-Couplings [AT05a]. Self-Synchronization [DMN22]. Semantic [OPUÖ11, Sal09]. Semi [AK19, CK18, DDPK21, EG00, Faz21, GPEG10, GE00, KYG20, KKY05, OV19, OV20, OC05, ÖS05, RSS^{+}21, ŞÖK11, ÜU05]. Semi-Analytical [AK19, RSS^{+}21]. Semi-Disposability [KYG20]. Semi-Infinite [Faz21]. Semi-Linear [CK18]. Semi-Open [GPEG10]. Semi-Prime [OV20]. Semi-Primes [OV19]. Semi-Rigid [OC05, ÖS05, ŞÖK11]. Semi-Ruled [EG00]. Semi-Supervised [DDPK21]. Semi-Symmetric [ÜU05]. Semiconductor [AÖ05]. Semiconductors [MS05]. Semicontinuity [GOA03]. Semidefinite [GWC98]. Sensing [HDAA20]. Sensitivity [BG19, HLX^{+}17, HLX^{+}18, OTŞ99, Sal06a, Tan08]. Sensor [Ala11, Bak11, BBSC18, ZP11]. Sensorless [RCTR20]. Sentiment [HMKS18]. Separable [RC16]. Separating [LV19]. Separation [HK23, Has03a]. Sequence [Bil02, PKC22, SS03, Sav10, Sav12, Sav13, Ull17, ZWZ17]. Sequences [AA03, ÇA04, SWA^{+}22, Sav99, Sav00, Sav14]. Sequential [BPD08, MARL21]. Serafini [BMV22]. Serial [BPC^{+}23]. Series [CC24, Özd96b, PRFP^{+}23, SÖS13, YAA10, YA11]. Server [SHE05]. Servers [KHL06, Sul07]. Service [Cha97, CCL06, GM17, Sul07]. Services [Pur21]. Set [All02, CSSOB20, Che18, SGCRRL^{+}23, TKGS20, Ugu13, VSAMMM23]. Sets [AA13b, BPC^{+}23, BVCS18, ÇGK96, GOA03, NT04, PS22, PGB^{+}11, dlFH19]. Setting [RLY^{+}17]. Settle [TDCT11]. Several [KKL18, Shr04]. Severity [BNSDC20, CTH^{+}22]. SH [HB01]. Shadowed [PS22]. Shaft [BA03, YA14]. Shaft- [YA14]. Shallow [Tay01]. Shape [AK19, Bül99, EE22, EI05, EI06a, GCPV18, Jab18, KRD^{+}22, LLTC17, MARL21, TWS17]. Shape-Preserving [TWS17]. Shaped [AUA03, DCC23, SKKC18]. Shapes [MAS^{+}11, RRO16b]. Shaping [ASK21]. Share [BM21]. Shear [BÖ07, Çat10, DCA10, DJ10, Kum07, MLBR20, Rod20, TDC08]. Sheet [JKS23, LYYX15, SSV23]. Sheets [EAT05, MKK97]. Shell [Dik06, MÖ96]. Shells [KUE13, SMB96]. Shift [AD11, All12, AD04]. Shifted [BS15, NSO^{+}21]. Shifted-Hopscotch [NSO^{+}21]. Shifting [BS23]. Ship [Tay04, UAI21]. Shipments [CBSTG13, CC11, CGCC11]. Shipping [CLC16]. Ships [Tay01]. Shiu [EA22]. Shock [EBY06, GS01, GS02, GS10, Shi06, SMB96]. Shoes [BGG96]. Shop [Cha05, FSHRCV^{+}21, SKK19]. Short [AERE18]. Short-Term [AERE18]. Shortest [MBA11]. Shrink [GLAA08]. Shrinking [NKS22a]. SI [SS07]. Siamese [TOM^{+}22]. Sicp [TÇ06]. Side [EL04a]. Sierpinski [Hei23, PS23]. Sigma [KS18, Kle18]. Signal [AWK^{+}20, EGB99, PRFP^{+}23]. Signals [EEJRLB^{+}22, RWL23, SKAK05, ÜIG03]. Signature [BBSC18]. Silicon [Boo17, Oku11a]. Similarity [CWF^{+}22, Çiç07, HPA13, KPS04, MS05, MM10d, Pak02, PŞ06, SÖS13]. Simple [AA03, BD09, BE99, CB04, ES06b, ES06a, Erp12, TJCGAPA18, TB21, Yan16]. Simplicial [Mut00a, Mut00b]. Simplification [KB03]. Simplified [ÇT14a, Oze03]. Simply [Özk01, PB03]. Simulated [BŞ13b, FSHRCV^{+}21, LVHL22, SHFSGB^{+}21, Sar09, ÜC11, VMRVF^{+}23]. Simulation [AÖ05, AÖ09, Amo21, AK11, BT03, BLEB^{+}18, EKAÖ11, FP21, GGMO21, GÇŞT07, GAS11, HGS^{+}22, JTM^{+}22, KZAN98, KA09, KK97, MWM16, MCNM22, ÖE03, PFR19, TCY06, YB03, YLX^{+}13, ZMCL24]. Simulations [CATCdlF20, PF10b, ZS04]. Simulink [ÇA10]. Simultaneous [SRLEPVP20]. Sine [BSA^{+}23, Kaz05, ÖK05, TGC21, TGC22, UG00]. Single [AO16, AR23, ADJ18, BS23, CATCdlF20, Che18, KA02, Özc00, ST11]. Single-Edged [Che18]. Single-Lap [ST11]. Single-Loop [BS23]. Single-Mesh [ADJ18]. Single-Objective [CATCdlF20]. Single-Step [AO16]. Singular [AD11, All12, AD04, AHR05, BdCCF19, CK07, CG09, CP22, DCA10, JY23, LLTC17, LVBV19, PMR14, RRR^{+}22]. Singularities [HWA99]. Singularity [Qad24]. Sink [DSI^{+}18, DEL23, KNRC22, OSAAA18]. Sink/Source [DEL23]. Sintering [Sta23b]. Sinusoidal [BBD^{+}18]. Sinusoids [Dil02, ÜC11]. SIR [AERE18]. SIRS [EARE18]. Sisko [IYKA20]. SISO [ASK21]. Sister [Shr04, Shr05]. Sitter [CO16]. Situations [Guy20]. Six [KO15]. Six-Step [KO15]. Size [All02, Alt12, CBSTG13, CC11, CCWY10, CGCC11, CLC16, KAM20, PALL18, SJ22]. Size-Structured [PALL18]. Sizes [SGC23]. Sizing [TCC11]. Skew [KKY05]. Skew-Symmetric [KKY05]. Skin [BY17, GC23b]. Skull [HFATZR^{+}20]. Skyrme [BM11]. Sleeve [MMF18]. Slender [ANJ20]. Slice [Alt12, SÖS13]. Slider [BY06, Yür02]. Sliding [Gul03, PPRVC20, Yag02]. Sliding-Mode [PPRVC20]. Slightly [Sin10]. Slip [BA03, JS22a, JKS23, RRO16b]. Slowly [Cai07, CC09, GGJ10, RRO16a]. Small [MWM22, PKC22]. SMoG [Rou21]. Smoke [SGEP06]. Smooth [PRFP^{+}23, ZBBZ18]. Smoothed [ZLZ17b]. Smoothing [YBTH17]. Smoothness [Asl18]. Sobolev [EB06]. Soccer [ŞE17]. Social [Kay16, NAA20]. Societal [OM20]. Soft [LV23, LH10b, ÖB17, SA19, VJ23, VC23a]. Software [IA04, Kas00, TE05, YDD11]. Soil [BK99, KAGAM23, Kha03, MM10a, YB03]. Solar [MMM23, PFR19]. Solid [AHM99, ÇŞÖG13, ÇÇÇ07, KR23]. Solitary [LDL10, Wu10]. Solitary-Wave [LDL10]. Solution [AV02, AA02, AV04, AAD11, AO16, AK19, All17, AHR05, Ars07, BÇ13, BY16, BG18a, BA11, CCY08, ÇKÇK10, Çev10, Dal11, DSI16, Doğ12, DL23, EAEAA19, FKRG10, GGJ10, Gha18c, GNK19, HM10, HAMM10, ID14b, İB00, JR18, KA11a, KKS11, KNHK20, KO15, LC18, MMF18, MG11a, MM10b, Öze00, PBY97, PÖ98, PŞ06, PS15, PSW24, PALL18, RAK22, RSS^{+}21, RAA10, Sal06b, SGA18, SUE13, SEZ24, TO13, UKÇ^{+}09, VA12, YÖS10, Yao15, ZWZ17]. Solutions [AYG14, AE02, AM10, BK99, BC97, BG18b, BMT^{+}21, BB16, Cai06, CC09, CSSOB20, DB16, FBR17, HPA13, HL14, HLL15, HL16, KAGAM23, KLGP24, KG10, Kha03, KSS19, Li10, LDL10, MQ18, Mit21b, MJK12, MM10d, ÖT02, Pak02, PVCRRV^{+}21, PMS03, Raw14, RRO16a, SK18, Sar11, SD13, qS13, SAM^{+}22, Wu10, YAA10, YA11, Yal13, YGS13, ZPL16, Zhu10]. Solvability [Akh04, IÇÜ12]. Solvable [ÇT14a]. Solve [BGN13, HM23, Mor22, Mos18, SZTY23, SBB23, Ugu13]. Solver [BCGS23, FLYY21]. Solvers [İBB98, JLW20]. Solving [AADS11, AK18, BA11, BS15, CCL06, CLC16, CJD24, DI13, DM21, Dol04, Ert07, EM08, GMV19, GHN11, GK11a, GY13, HNB20, IGSB18, IA04, JLST04, KAM20, KYS13, Mer09, MG11b, NNW14, Öze10, SSA22a, SGZ10, SGZ11, ST18, SKB22, SKGP22, SJ19, TKGS20, Tan08, TYS13, WN21, Wan13, YG15, YKG20]. Some [AE02, ANK23, Bil02, ÇÇ03, GLS18, İYK99, Kar03a, KK13, KYK02, Kha99, MK96, MC20b, NU19, Özd96b, Pak02, PT21, Pop98, QM16, RRO16a, SSA22a, SS03, Sav10, Sav13, SBA^{+}04, Shr01, SJ19, SB02, TU97, YBYL21, YU07, Bİ99]. Sonic [BHN10]. Soret [JS22a, KH22, KNRC22, NKS22b, RKS23]. Sorting [Mat10a, ZWZ16]. Sound [DB16, ZLZ17b]. Source [ATZES^{+}07, AGANA22, DSI^{+}18, DEL23, GCPV18, KNRC22]. Source/Sink [DSI^{+}18, KNRC22]. Sources [DMN22]. Southern [MMM23]. Sovereign [KNHK20]. Space [BY13, CPA20, CUA04, CS19, DZA19, DJM18, ED11, GC23a, Kat17, KÇ96, MK96, Mub23, TKGS20, UÇ96, WLW^{+}10, YH05, Yay00]. Space-Like [KÇ96, UÇ96]. Space-Time [DJM18]. Spacelike [Yay00]. Spacer [HTLY18]. Spaces [AS00, AD04, Bil02, BD22, CKLZ21, GM21, Kay15, KAM20, Mat19a, ÖB17, Pop98, Sav00, SS03, Sav10, Sav12, Sav13, Sav14, Sav16, YZHY17]. Spacetime [CO16]. Spacing [Kop02]. Spanish [HSG^{+}20]. Sparrow [LBY^{+}22]. Sparse [HH22, SC19, ZLZ^{+}17a]. Spatial [BMT^{+}21, CBT21, Yag02]. Spatio [BWK23]. Spatio-Temporal [BWK23]. Spatiotemporal [Zhu17]. Special [AK18, Eİ06b, Fan20, KY05, MHC21, PMR14, ZMCL24]. Specially [AB06, DCC23]. Specific [AK05, AHID18, LCW^{+}22, PAR23].

Spectral [AWK^{+}20, BCK^{+}19, CMG22, MV21, MMS19, NSCTPW22, STB04, TJCGAPA18, ÜIG03]. Spectrum [BPC^{+}23, EEJRLB^{+}22, PS23]. Speed [AÇ11, ÇYAÖ05, EÜA13, Gul03, MPA17, MMF18, RCTR20, Tay04]. Speeds [BHN10]. Sphere [AHM99, BBK20, Du 22, KUÖ05, Kaz05, ÖK05, SPD18, UKÇ^{+}09]. Spheres [GB18, KM10]. Spherical [ASAH20, GS10, Kay98, KUÖ05, KSK05, KSK01, KÇ11, MM10c, UG00]. Spheroid [DAD^{+}18]. Spikes [KP12]. Spillovers [GP21]. Spillway [OA09]. Spin [AKT04]. Spin-1 [AKT04]. Spinning [KH15]. Spiral [HTLY18]. Spline [CT01, DS04, GA10, LC16, mLkWbPxX17, PQ23, YZBZ16, YZHY17]. Split [BM21, HK23, TKGS20]. Splitting [GN19, YW24]. Spot [KD19, UC10]. Spread [GM21, GPACC^{+}19, SPCPMM21]. Spreadsheet [Gha18a, Gha18b, Gha18c]. Spring [ZWZ16]. Spur [CTH^{+}22, GÖA96]. Square [CSOPM24, EY13, HFAAM^{+}22, Oku11a, Oku11d, STK07]. Square-Root [CSOPM24]. Square/Circular [EY13]. Squares [Kan16, KAS19a, OV19, OV20]. Squirrel [GAAAS17]. Srivastava [ID14a]. Srivastava-Owa [ID14a]. ssMousetrack [CPA20]. Stability [AGANA22, AMAGNA23, Büy16, ÇÇ03, DB03, GK12, Mak99, MOBA21, OWO^{+}23, Öz00b, Özg10, PAR23, Sar11, SJKCD20, Tay04, UAI21, WH10, YA23, YY07, ZCC15]. Stabilized [GC23a, ŞÜ13]. Stable [ACM19, NSO^{+}21]. Stacked [KIY^{+}23]. Stage [BPD08, BPC^{+}23, KJZL12, ŚLO19]. Staged [Cui07]. Stamp [BE99, KBE01]. Standard [Den16, FJ21, FdSPW20, Mol18, Sta23b, Wu17]. Star [GS10]. Starting [DK99]. State [AOL19, AB06, Bha24, CPA20, LOSCO^{+}20, LRL23, LTYG17, MK96, PPB^{+}20, Pin22, SSA22b, UKY96, WCH23]. State-Owned [LTYG17]. State-Space [CPA20, MK96].

States [BSK05, Boo17, İTİO08, KSK01]. Static [BV19, JY23, MLBR20, MRG23]. Stations [DJY^{+}22]. Statistical [BKA08, BSA^{+}23, KAK00, NCM19, Özg10, PMF20, SS03, Sav14, Sav16, TCM21]. Statistically [Sav99]. Stator [STA23a]. Steady [AO17, AOL19, AB06, LRL23, MM10c, PPB^{+}20, SSA22b]. Steady-State [AOL19, LRL23, SSA22b]. Steel [BSSK12, ÇŞS01, DCC23, DLL10, GBİ14, Kir04, MKK97, Oku11c, OFCP21, YG02]. Stefan [KAS19a, NKS22a, NKS22b]. Stein [SSA22a]. Steiner [WH10]. Steklov [Mat19a]. Steklov-Type [Mat19a]. Stellar [Shi06]. Stems [DPC^{+}10]. Step [AO16, Alt12, Çiç07, Eli16, KAM20, KO15, SKB22]. Stepper [Alt09]. Stepping [DSMP19]. Stiefel [AWK^{+}20]. Stieljes [KWP11]. Stiff [AG03]. Stiffened [MÖ96]. Stir [KTS13]. Stochastic [Ara08, CCGHdCJ21, CKLZ21, Dan22, DSMP19, FMM10, GN19, IÇÜ12, Kha99, LNBM10, MR05, MPP^{+}24, SWA^{+}22, SGA18, SW10]. Stochastically [Boo17]. Stock [GZY20]. Stocks [WW20]. Stokes [ACM19, HM23, Sal06b]. Storage [SC19]. Straight [DE04]. Straight-Through [DE04]. Strain [Alh21, BG19, FSV^{+}22, HRM19, HFATZR^{+}20, KF19a, KF19b, VVN23, YA23]. Strain-Gauge-Based [VVN23]. Strains [KAY09, YG02]. Strategies [AMM23, BDC^{+}18, BWK23, FYuR^{+}23, GK18]. Strategy [FSV^{+}22, GT22, LZC22, LBY^{+}22, SH18, YA23]. Stratification [JS22a]. Stratified [TKK^{+}22]. Stratigraphic [PMF20]. Streamlines [MM10c]. Streams [KS18, Kle18]. Strenght [AMK96]. Strength [DPC^{+}10, EAT05, Oku11b, TDC08]. Strength/Low [EAT05]. Strengthening [TDT^{+}11]. Stres [GB00, ZA98]. Stress [ABO^{+}16, AA01, AR23, AUA03, ATSA97, AT05b, BSSK12, ÇŞS01, ÇKAK04, Dar11, Ese01, GB98, GA99a, KT98, KTC99, KA02, KA11b, KÖAG00, KE01, Küç01, LC18, Mit21a, Mit21b, Oku11a, OYCS11, ÖG13, Özb02, Özc00, PRT^{+}19, Rod20, SS06, ST11, TEK11]. Stresses [BKOS11, BS11, GÖA96, GAÖ96, MAS^{+}11, Oku11c, Özb02, STK07, TSC11, YG02]. Stretched [PB01, Xu10]. Stretching [DSI^{+}18, JS22a, JKS23, LYYX15, NKS22b, RRO16b, SSV23]. Stretchy [JS22b]. Strip [DK03, NCDM22]. Strips [BÇE02, Cui07, LRBMAM^{+}23]. Strong [ANK23, GS10, HBN96, Sar11]. Strongly [Cai07, GW97, PKB11, Sav12]. Structural [ANJ20, ANA20, BV19, CA19, CDDM23, DCC23, Fan20, GGJ99, HSG^{+}20, JLH21, MRG23, SJKCD20, TMB96, YE03]. Structural-Health-Monitoring-Oriented [DCC23]. Structure [BIM^{+}20, DSU22, ETW^{+}23, Kar11, KPTM20, SHFSGB^{+}21, WSLZ17, WH10, YQG22]. Structured [PALL18]. Structures [DE23, DLL10, LV19, NCDM22, NSCTPW22, TB21]. Studies [Ken11]. Study [AB99, AHVG23, AGANA22, AÇGJ99, BG19, BS13a, DDPK21, DA19, EEJRLB^{+}22, GGT^{+}23, GLS18, HXY15, HFAAM^{+}22, İYK99, JB15, KSTJ18, KST10, Kay16, KÜM08, KY13, KTD12, LTYG17, LDCST21, MSPC23, MWM16, MPP^{+}24, OSAAA18, Oku11b, OFCP21, RFQCMMCR23, RKX23, ŞE17, Sha19, SJ22, ST11, STA23a, TCA96, UÇ96, WH10, Yen14, YK96b]. Sturm [Ert11]. Style [TOM^{+}22]. Sub [BPC^{+}23, MGGHCR^{+}20]. Sub-Sampling [MGGHCR^{+}20]. Sub-Sets [BPC^{+}23]. Subcubes [All02]. Subdivision [Asl18, BGRV21, PTLZ18, TWS17]. Subdomain [Bad22, BBD^{+}18, DB17a, DB17b]. Subgradient [Man16]. Subject [Cha97, DSD^{+}21, KH22, LCW^{+}22, RRR^{+}22]. Subject-Specific [LCW^{+}22]. Subjected [AR23, AK02, ÇŞS01, Ese01, KA02, TMK21, TMB96, YG02]. Subjects [FSH^{+}23].

Submerged [MWM16]. Subordination [ID14a]. Substitution [CYAK03, KCMLTR18]. Substrate [PCCT11]. Substrates [BP23, OYCS11]. Subsystem [Kas00]. Successive [DK99]. Sufficiently [BD22]. Suggestion [Erp12]. Sulphomethylation [Bas99]. Sum [Cet16, EGIK97, GB23, OV19, OV20]. Summability [ÇA04]. Sumudu [JR18]. Superradiant [NCM19]. Superresolution [AB23]. Supervised [DDPK21]. Supplier [HH07]. Supply [ACC19]. Support [CMS20, Con22, GLAA08, KBE01, PRFP^{+}23, PB03, Sam99, ZGV10]. Supported [AB06, KE01, Özk01, PB03]. Supports [BÖÖ11, BE99]. Suppression [BBPL19a]. Supreme [Gia21]. Surface [AAR24, BBK20, CY05, DZA19, DSI^{+}18, GS10, Jab18, JS22b, KWA^{+}10, KÇ96, Koi19, KÇ11, OL21, Özy98, ÖY01, Özy14, RM20, SK03, TDCT11, UK96b, UT96, Ull17, VRMJG19, Xia10, Yen14]. Surface-Mounted [Jab18]. Surfaces [CUA04, EG00, ED11, GE00, KYK02, Li16, VC23b, Yay00]. Surging [BCN22]. Surrogate [BP21, Bha24, BWK23, DRH21, GCPV18, LRL23, PD18]. Surrogate-Based [GCPV18]. Surveillance [CRA^{+}21, OC09]. Survey [BM21, Mat23, PD18, Sol23, Yav22, YS02]. Survival [GGMMDI20, MCNM22]. Suspension [Guc04, Yaz16]. Swarm [ÖE03]. Swell [EG11, GLAA08]. Swell/Shrink [GLAA08]. Swelling [AMAGA22, KAGAM23]. Swept [WN21]. Swirl [NNM10]. Switched [BBD^{+}18]. Symbolic [AM09, ACC19, CP22, GH11, LB20, Loj22, SKGP22]. Symmetric [BM11, GA99b, KKY05, ÜU05]. Symmetries [Can96, HBN96, Kar10, NMCM19, Öze00, PY00, PŞ05]. Symmetry [AM05, FMR10, İTİO08, KSM23, Kar03a, KTK97, MAH10, MJK12, MM10b, Qad24, SQA11, ZMA23]. Symmetry-Like [FMR10]. Symplectic [BBH19]. Synchronization [BBPL19a, BBPL19b, CLY10, DMN22, PPRVC20]. Synchronizer [PLLRCC20]. Synchronous [BODH20, KA96b, TGK96, ZG17, ZLLX20]. Syndrome [FYuR^{+}23]. Synovial [KVR23]. Synthesis [ASK21]. Synthetic [Rou21]. System [AS09, Abi03, AE03, Aki11, Aks05a, AM09, AMAGA22, All02, Ars07, BK09, BG04, BY11, Bİ99, BUAM19, BP99, BDÖ^{+}99, Bül99, BB16, CTL14, ÇTY96, ÇGV02, Cha97, CT14b, CMG22, DSMP19, DB16, Doğ12, EESK^{+}02, EARE18, GT22, GLAA08, HB99, HSG^{+}20, JMS18, JH13, KAGAM23, KA96a, KKC08, KY96, KA09, MECRMT^{+}21, Mak99, Mat19b, MM10a, MPP^{+}24, MOBA21, MK10, NM10, NYH^{+}17, NNW14, Öz00a, Öze10, PPRVC20, RC99, RZ18, SÇ10, Sul07, TGK96, TY15, TGS98, Üna96, Üna03, VPJ22, YE03, Yal13, Yaz16, YYWX16, Yur99, YA14, ZE16, ZMA23]. Systems [AB19, All17, ASK21, BS15, BBK20, CLY10, CMS20, ÇT09, DVFA23, DRKW21, DKÇ97, FMR10, GMV19, GCGZ10, Gul03, GM17, GOA03, Has02, JTM^{+}22, JY23, KLGP24, Kha03, KIY^{+}23, LRL23, LEVP21, OOI^{+}22, ÖÖ05, OC09, PBY97, PB97, PKB11, Pek96, QN18, Raw14, RC16, SQA11, Sam99, SBDN^{+}22, Sol19, Sol23, SSA22b, TCC^{+}23, TW18, UC10, VC23a, YS02, YUKA03].

T

[BDÖ^{+}99, FLYY21, GVC20, OF19, Mer09]. T-Cell [Mer09]. T-Cells [GVC20, Mer09]. T-Complete [FLYY21]. T-Joints [OF19]. Tables [PLW21, RG23]. Tablet [Sat23]. Tabu [Bay03]. Tachyon [SA16]. Taiex [LLL14]. Tails [Aki11]. Take [RLY^{+}17]. Take-Profit [RLY^{+}17]. Taking [DB17a, DB17b]. Talks [CGM^{+}23]. Taming [CMG22]. Tandem [Cha08]. Tangent [ASJQ24, MQ18, Özy98]. Tanks [TMK21]. Taper [ST11]. Target [Ala11, Fu15]. Task [LP17, NSG^{+}23]. Tasks [GA99b, LP17, MECRMT^{+}21]. Tau [GMV19]. Taxicab [KÖ00, Koc00]. Taxonomy [DRH21]. Taylor [BÇ13, Çev10, KKS11, TYS13]. Te [TM11]. Teaching [Con22]. Techmique [Ala11]. Technical [CÖGD11, ÇÇÇ07]. Technique [AB96a, AM23a, BBMA22, BBD^{+}18, DCA10, DB17a, DB17b, EM08, JLST04, KAK96, OSAAA18, ZLZ17b]. Techniques [Bak11, EGB99, KAK96, LVHL22, MK96, Sal06b, TCM21]. Technological [Abi03]. Technologies [GERV19]. Teeth [ATSA97]. Teichholz [DAD^{+}18]. Teissier [KMCI22]. Telecommunication [KKL18]. Telegraph [BY16, HBN96]. Teller [BSK05, SBA^{+}04]. Temperature [Alh21, AR23, ATK99, BS11, DNS23, OFCP21, PŞ05, ŞT96, SS06, Sha19, SSA22b, TB21]. Temperature-Time [OFCP21]. Temperatures [BY17, DÇ06, EAT05, MO04]. Template [ÇGV02, SC19]. Template-Based [SC19]. Temporal [BWK23]. Tensioned [AO17, BÖÖ11]. Tensor [GE00, NU19, ORC19]. Tensors [Bro19, Mut04]. Term [AERE18, AGANA22, KAY08, Yao15]. Termination [XS22]. Terms [GGS11, KKS11, MOBA21, RS21, Sal09]. Ternary [Asl18, PTLZ18]. Tertiary [SHFSGB^{+}21]. Testing [CRA^{+}21, Mit21b]. Tests [GAS11, SK05]. Tetragonal [CY05]. Text [CD04]. Textile [ADS19]. Textiles [WH10]. Texture [SK03]. Textures [KBH06]. th [AADS11]. Thailand [DJY^{+}22, HIVH21, KPSS18]. Thawing [Özg10]. Their [ANK23, BG19, CY05, GK11b, Guy20, KRD^{+}22, PS23, Rou21, SKB22, SKB23, YZHY17]. Theorem [CT01, Kay98, SAI04b]. Theorems [Pop98, UKÖ05]. Theoretical [AGANA22, AMAGA22]. Theory [AGKT21, Akt01, AMAGNA23, Alh21, BS23, DSD^{+}21, DCA10, Flu15, HBS^{+}17, JHA^{+}22, LVBV19, MPA^{+}22, MC20b, MF21, Mor22, MRG23, Sha99, TB16]. Therapy [AAGH21]. Thermal [AO17, AR23, BIM^{+}20, BS11, ÇŞS01, DSI^{+}18, EBY06, EY13, ETL22, FHRS18, Küç01, MRG23, NKS22a, NKS22b, OSAAA18, Oku11c, OYCS11, PŞ05, PŞ06, PPB^{+}20, RKS23, SS06, STK07, ZMCL24]. Thermally [JS22b]. Thermo [GFPF22]. Thermo-Elastic [GFPF22]. Thermodynamic [Has03b, KPTM20]. Thermoelastic [AS04, DZA19, EBY06]. Thermoelasticity [Alh21]. Thermofluid [LRL23]. Thermoplastic [BKOS11, BSSK12, ÇŞS01, ÇKAK04, Oku11d, ÖG13, Özc00, SS06, STK07]. Thermoplastic-Matrix [Oku11d]. Thermostatistics [AK05, KST10]. These [AA03]. Thick [AUA03]. Thickness [KPSZ24, Sar09, Yav22]. Thin [Alt10, MM10b, NUM^{+}19, NU19, RKS23]. Thin-Walled [Alt10]. Third [ADE99, MM10b, NNW14, Yür02, Yür04]. Third-Grade [ADE99]. Third-Order [MM10b, NNW14]. Threaded [NKI19]. Three [ALFBPLE20, ATSA97, ATK99, Fu16, IZS04, JS22b, Kah11, KT98, KTC99, Li10, SEZ24, Sul08]. Three-Dimensional [ATK99, IZS04, JS22b, SEZ24]. Three-Parameter [Fu16, Sul08]. Three-Wave [Li10]. Threshold [CDDM23, FSHRCV^{+}21]. Thrombosis [CLGGLB^{+}22]. Thrust [ANJ20]. Thyristor [YUKA03]. Ti [TDC08]. Ti6Al4V [PCCT11]. Tibial [ÇS10]. Time [AERE18, AW19, ACM19, ALFBPLE20, ACGC21, AG12, AA13b, ALK19, BdCCF19, Bha24, CUA04, Cha05, Dal11, DSG19, Dik06, DMN22, DJM18, EARE18, FBR17, Gun96, GGMMDI20, HH07, JY23, JFRZ10, KAGAM23, KUE13, OFCP21, Öze02, ÖY01, Özy14, Pin22, Pur21, Sar11, SAM^{+}22, SWJS22, Sul07, UK96b, UT96, UÇ96]. Time-Course [Öze02]. Time-Domain [AG12, SWJS22]. Time-Like [CUA04, ÖY01, Özy14, UK96b, UT96, UÇ96]. Time-Varying [SAM^{+}22]. Times [LZ17]. Timoshenko [ANA20]. Tiny [ETL22]. Tip [ÇYAÖ05]. Tire [ÇT12]. Tires [TÇ04]. Tissue [Alt12, FA22b]. Titanium [AMR^{+}23]. TL [IMKEK18]. TL-Moments [IMKEK18]. TM [YD11]. Toeplitz [SB02, STB04]. Tolerance [Sam99]. Tolerant [LESEVP^{+}20, PS22]. Tomographic [HRM19]. Tomography [KY96]. Tool [ACC19, KA09, SK03, TE05, YY11]. Tools [Con22]. Tooth [ATK99, AT05b, GGT^{+}23, GÖA96]. Top [AO17]. Top-Tensioned [AO17]. Topological [Çiç07, Muc01, RG23]. Topology [Cor19, MARL21, ÖB17]. TOPSIS [BK09, SAZ18]. TOR [ZG17]. Toroidal [DK03, KD03]. Torpedo [CMB19]. Torque [Akb09, MSSZ20, TÇ04, VRMJG19]. Torsion [Bil04]. Torsional [GAÖ96, STA23a]. Total [AM23a, CÖGD11, CK18]. Tracer [BCS20]. Tracking [Ala11, CPA20, Cor19, KUE13, LESEVP^{+}20, NYH^{+}17, RCTR20, SWA^{+}22, WZ17, YL16]. Traction [TÇ04]. Tractive [ÇD03, ÇT12]. Tractors [TÇ04]. Trade [Tsa09]. Trading [RLY^{+}17]. Traffic [ÇSB10, OC09, WA01, YLX^{+}13, Zha14]. Traffic-Flow [ÇSB10]. Trailer [PRFH23]. Train [ADJ18, ORC19]. Trajectory [GT22, GB18, GA99b, KÇ11]. Transcendence [Özd96b]. Transdisciplinary [PRT^{+}19]. Transfer [BS13a, BO10, DJH18, Gun96, HA10, Kop02, LYYX15, PŞ06, RRO16a, SJ22, WJ23, WFV21]. Transform [ÇKÇK10, Doğ12, EM08, Ert11, GA10, GAAAS17, ID14b, KTK11, KWP11, KLGP24, KO10, Kil04, LH10a, LPAD20, MK97, Raw14, SÖS13, SKGP22, TSS17].

Transformation [Ert07, Faz21, FKRG10, FSHGGB^{+}21, GK11b, MG11a, PKO11]. Transformational [AFGV20]. Transformations [ACC20, Bil02, KA96a]. Transformers [AMMMM^{+}23, BCK^{+}19]. Transforms [JR18]. Transient [AÖ09, ALFBPLE20, AK02, GCGZ10, MK96]. Transients [MK97]. Transition [Aki11, BBPL19b]. Transitions [SBA^{+}04, YU07]. Transitive [LCL17]. Transmission [CST20, ÇT09, FYuR^{+}23, FA22b, MPA^{+}22, MK96, MK97]. Transport [BGA10, DL24, HAMM10, KVR23, MO04, MS05, OOI^{+}22, ZHN10]. Transportation [Cha05, Has02, KSTJ18, KB11]. Transvalvular [DL23]. Transversality [Sol23]. Transversally [ES06a]. Transverse [BÖÖ11, Ese01, FLYY21, GAÖ96, KE01, Öz00b, ÖÖ05, ÖD06, Özk01]. Transversely [AUA03, BKOS11]. Trapezoidal [EPLGGS^{+}21, LY15]. Trauma [FP21]. Traumatic [BNSDC20]. Traveling [KN13]. Travelling [BŞ13b, Zhu10]. Traveltime [ÜÖ03]. Treatment [BG19, BY17, GK18, ZLZ17b]. Tree [ÇŞÖG13, WH10]. Trees [Aks05b]. Trending [Kar11]. Trends [PD18]. Tri [MPP^{+}24]. Tri-Trophic [MPP^{+}24]. Trials [BPD08]. Triangle [AKT04, KUÖ05]. Triangles [Kaz05, ÖK05, UKÖ05]. Triggered [JY23, MPA17]. Trigonometric [Li16, OL21, QN18, Qua18]. Trigonometrically [DSI16]. Trigonometrically-Fitted [DSI16]. Trigonometry [Koc00, UG00]. Trihedra [Özy11]. Trihedrons [Özy14]. Trihybrid [AMR^{+}23]. Trilinear [FA22a]. Triple [JS22a]. Trophic [MPP^{+}24]. Truck [EB00, PRFH23]. Truncated [DAD^{+}18]. Trust [BMV22, BP21]. Trustworthiness [OF19]. TSA [DNS23]. TSP [MBA11, UKÇ^{+}09]. Tube [DJH18]. Tubes [Mos18]. Tubular [GAÖ96]. Tuned [ES06b, ES06a]. Tungsten [LZL^{+}23].

Tuning [CTH^{+}22]. Tunnel [LC18]. Turbine [AB19, ADJ18, CA19, EÜA13, Kan00, ZdlÁGLTT22]. Turbines [ÇYAÖ05]. Turbulence [TCY06]. Turbulent [ACM19, AM10, DJH18]. Turkey [CÖGD11]. Turkish [BG02, EBD11, GBİ14, Ken11, OPUÖ11]. Tutorial [PRFP^{+}23, SW22, ZBBZ18]. Twin [Erg01, RWL23]. Twinjets [BS13a]. Twist [AB96b]. Twitter [HMKS18]. Two [AO17, Alh21, Alt16, ABH^{+}22, BPD08, BG19, BBD^{+}18, BÇE02, CLY10, Cui07, Eli16, ENVPF19, GGJ10, GB18, HH22, IMWY18, JM03, KY05, KJZL12, KE01, KAEÖ13, MAKH10, Mos18, NM10, RSS^{+}21, ŞT96, Saj18, SKB22, SHE05, TWS17, TSS17, Üna03, Unv07, VRMJG19, WN21, WW20]. Two-Component [WW20]. Two-Degrees [Üna03]. Two-Dimensional [Alt16, BBD^{+}18, GGJ10, MAKH10, RSS^{+}21, VRMJG19, WN21]. Two-Generator [HH22]. Two-Parameter [ENVPF19, Saj18]. Two-Phase [AO17]. Two-Stage [BPD08, KJZL12]. Two-Staged [Cui07]. Two-Step [Eli16, SKB22]. Two-Strain [BG19]. Two-Temperature [Alh21]. Type [AJ20, AMAGA22, BMT^{+}21, CP22, Ert07, GVC20, JTM^{+}22, Kız14, KG13, Mat19a, ME12, Özy11, PS22, PPRVC20, Saj18, SAK01, Şen05, SK05, TAS^{+}23, ZPL16, dlFH19, BD09, IMKEK18]. Type-1 [GVC20]. Type-2 [Kız14, Özy11, PS22]. Type-I [IMKEK18]. Types [AHID18, GZ14].

U [BD09, SKKC18]. U-Shaped [SKKC18]. U-type [BD09]. U.S [Gia21]. U1 [Bil04]. Uclear [BM11]. Ultimate [Kas19b]. Ultra [ÖT02]. Ultra-Hyperbolic [ÖT02]. Ultrasonic [Can02, ÇGK96]. Ultrasonographic [GÖV96]. Ultrasound [RWL23, UK96a]. Umbral [LPAD20]. UMF [VRMJG19]. Unbalance [JH13].

Unbounded [SK18]. Unbranched [Kaz04]. Uncertain [CWF^{+}22, DKÇ97, RAA10, Sho22]. Uncertainly [Kan00]. Uncertainties [CLY10, GNI19, YQG22]. Uncertainty [CMB19, GP21, ST22]. Uncovering [AMM23]. Undermine [CMM19]. Undesirable [KYG20]. Unequal [BHN10]. Unidimensional [Bec17]. Unidirectional [ST22]. Unified [QSC19]. Uniform [AR23, CT01, GS01, GS02, GS10, RFD^{+}10, SS06, Shi06]. Uniformly [Ese01, KÖAG00, KE01]. Unions [CÖGD11]. Uniqueness [FBR17, Faz21, Koi19, RAA10]. Unit [GERV19, ID14b, IGSB18, KUÖ05, Kaz05, KMCI22, NAC22, ÖK05, ZS16]. Units [Bil04, BUAM19, JLS^{+}10b, LNBM10]. Univariate [MF21]. Universal [Muc00]. Universities [Ken11]. University [HSG^{+}20]. Unknown [Liu11, ZP11]. Unpacked [ZS04]. Unreinforced [OF19]. Unrelated [RFQCMMCR23]. Unrestored [AT05b]. Unsteady [AM05, BGA10, DZA19, DSI^{+}18, EL04a, EL04b, GC23a, TGS98]. Unwinding [PF10a, PF10b]. UP-Algebras [AHK18, KUI18]. upon [CDDM23]. Upper [AAR24, Küç01, STB04]. Upscaling [YZBZ16]. Urban [GM17, LGT^{+}18]. Use [GÇŞT07, KRD^{+}22, Kat17, MCBM23, YY11]. Used [Alt09, Flu15, SÇ10, VPJ22]. User [Alt09, BBSC18, EEJRLB^{+}22, JMS18]. Users [KKL18]. Using [AO16, ATZES^{+}07, AJ20, AB96a, Aks05a, Akt01, AB03, AAS11b, AAR24, Ara08, AMMMM^{+}23, ATK99, BY16, BG99, BK09, BY11, Bay03, BD09, BAUO09, BA11, BP21, BS23, BBPL19a, BO10, BBPL19b, CLY10, Can14, ÇT12, CAQCMM22, CACRF^{+}21, ÇÇÇ07, ÇGK96, ÇS10, CÜ04, ÇA10, CJL18, DDPK21, DK03, DE04, DT08, DJM18, DRCSF22, Emi03, EG11, Ese11, EEJRLB^{+}22, FTGSNV21, GA10, GERV19, GAAAS17, GCPV18, GK11b, GGMMDI20, Has02, HIVH21, ID14b, IMWY18, JR18, JLST04, JMB22, JMS18, JSTGV^{+}19, KTK11, KAY08, KUE13, KY96, KG10, KBH06, KA96b, KD03, KF19a, KF19b, KPSZ24, LVHL22, LCB18, LLL14, LRBMAM^{+}23, LGT^{+}18, LNBM10, MQ18, MK97, MAS^{+}11, MG11a, Mes96, MFF20, Mor22, MPP^{+}24, Mub23, MRG23, Mut98, NAA20, OSAAA18, OL21, OV19]. Using [Öz00a, OC09, PRFP^{+}23, PYR22, PCCT11, Pek96, PGRL21, PSW24, PC23, RCTR20, Raw14, SK18, SK03, ŞÜ13, Sal06b, Sat23, ŞÖK11, SGEP06, SY11, SS06, SÇGA12, SKGP22, SAM^{+}22, SBF96, SKKC18, SKAK05, TCC^{+}23, TUK19, TEK11, TMK21, UAT19, Ull17, Üna96, UB10, UAI21, ÜC11, VJ23, Wu10, Yag02, YB97, YA14, ZE16, ZP11, ZLZ17b, ZPW^{+}22, Aks05b, KAY09, Öze00, TÇ04]. Utilizing [CRA^{+}21].

V6.2 [BAUO09]. Vacation [WT15]. Vacations [CT14b]. Vaccination [YA23]. Vaginosis [SGCRRL^{+}23]. Vague [HAH^{+}20]. Valid [AA02]. Validation [BY11, HGS^{+}22, SEZ24]. Validity [CMCM21]. Value [AV04, AAS11a, DSI16, EM08, HM10, Mat19b, Mor22, NNW14, OA13, TY15]. Valued [DW16]. Values [ŞÖK11, SBA^{+}04, TÇ06]. Vanadium [LZL^{+}23]. Vapor [BP23]. VAR [PC23]. VAR-KF [PC23]. Variable [AMAGNA23, ÇDYS13, CAQCMM22, EÜA13, GGS11, HL14, JZL10, KTK11, KAGAM23, KKS11, KAAK10, KPS04, KPSZ24, LC18, Liu11, MAKH10, Mes96, MJK12, PŞ06, PT21, WT15, Yal13]. Variable-Exponent [AMAGNA23]. Variables [AM23b, GB23, JM03, Pin22, PC23, Sal06b, Shr04]. Variance [ARCRMM20, GP21, HBFM22, YA14]. Variances [Kan16]. Variant [GPMRVM21, Man16, SKB23]. Variation [Akb09, AM23a, CS19, LNBM10, SPD18]. Variational [AS09, AADS11, AS00, CML10, GW97, GHN11, KAM20, LXZ10, Man04, MDNN10b, NK10, Rou21, SAK01, Sal02a, Sal02b, SAI04b, SI04, Sal04, Sal06a, SAH98, Wu10, ZWZ17, ZH10]. Variational-Like [Sal02a]. Various [Egb18, Özg10, PSW24]. Varley [ZPL16]. Varying [Cai07, CC09, GS01, SGC23, SAM^{+}22]. Vascular [Ars07]. VCO [CATCdlF20]. Vecor [Civ96]. Vector [BS23, ED11, ZGV10]. Vectors [UK96b, UT96, Yay00]. Vehicle [AA16, Guc04, IGSB18, MVC23, OC09, VSAMMM23, Yaz16]. Vehicles [GÇŞT07, LH10b]. Vein [AK11]. Velocities [NLMU19]. Velocity [AK02, JS22a]. Vélu’s [CSOPM24]. Venant [ELT21]. Venous [CLGGLB^{+}22]. Ventricular [DAD^{+}18]. Venturi [BAUO09]. Verification [Cai06, CCY08, LC18, OA09]. Version [MHC21]. Versions [PB99]. versus [Guy20]. Vertical [GLA05, KNRC22, RKR^{+}22b, STA23a, ZBBZ18]. Vertices [AMR^{+}23, Che18]. Vessels [GA99a, ZS04]. VFVF [AB06]. via [Aki11, CPA20, Den16, EAEAA19, EKAÖ11, Ert11, FMR10, GC23a, GZ14, KSM23, KAS19a, KWA^{+}10, mLkWbPxX17, PPRVC20, SK03, Sal09, UYAA22, WZ17]. Viability [ZPW^{+}22]. Vibrating [KKA04]. Vibration [AJ20, Alh21, Alt12, BUKM96, BÖ07, BO10, ÇDYS13, Çat10, Çev10, CÜ04, CK07, CG09, CDA10, DCA10, ES06b, ES06a, EGIK97, GÇÖ12, JH13, KTK11, KWA^{+}10, KPSZ24, STA23a, TB16, UC19, Yaz16, YK96b]. Vibrations [ANJ20, AO17, Alt10, BÖÖ11, BC97, Öz00b, ÖD06, Özk01, ÖÖT04, Özy09, PB97, PB01, PB03, PKB11, TE05, YK96a]. Video [WZ17]. View [Gia21]. Violence [DAKCPACC19]. Viral [OWO^{+}23, Zhu17]. Virtual [BVvdH^{+}23, KUE13]. Virus [FA22b, Mer09]. Viscoelastic [AGANA22, AMAGNA23, HWA99, WSLZ17]. Viscoplasticity [ORC19]. Viscosity [MAH10]. Viscous [GGJ10, GB18, Has03a, Has03b, MM10d, RSS^{+}21, RRR^{+}22, XS22, Sha19]. Vision [CGMCCA23]. Vision-Based [CGMCCA23]. Visual [Flu15, RM20]. Vivo [HGS^{+}22]. VOCs [ZMCL24]. Vogel [KB11]. Voigt [LPAD20]. Vol [SGZ11]. Volatility [WW20]. Voltage [Akb09]. Volterra [AS09, AEB01, BS15, EB02, HXY15, HNB20, SAM^{+}22]. Volume [Boo17, BBK20, CO16, DAD^{+}18]. Volumetric [UK96a]. Volumic [BODH20]. Vortex [HK23]. Voting [JLKK12]. Voynich [Ace19]. Vries [BMT^{+}21]. VRP [YRB16]. VRP-GMRES [YRB16]. Vslam [ŞÜ13].

Wakes [MM10c]. Wall [RRO16b]. Walled [Alt10]. Walls [BÖ07, Çat10, EL04a, GGJ10, Ken96, RSS^{+}21, RRO16a]. Ward [AKT04]. Warrants [Sho22]. Warships [CMB19]. Wastewater [RJ17]. Water [BK99, BB16, CC24, HK23, JKS23, JB15, Kha03, MWM16, NKS22a, RRO16b, RRO16a, ŞÖ11, TMK21, Yur99]. Waters [Tay01]. Wave [BB16, CJD24, FLYY21, GS01, KRD^{+}22, Kay98, KPS04, Kum07, Li10, LDL10, NLMU19, PY00, RWL23, ŞÖ11, SWJS22, TA99, Zhu10]. Waveguide [TA99]. Wavelet [GAAAS17, Pin22, WA01, Zha14]. Wavelet-Neural [WA01]. Waves [CC24, DB16, GS02, HB01, Mub23, PD19, ŞÖ11, Shi06, Umu11, YD11]. Wavy [BA99]. Waxy [ZPW^{+}22]. Way [YA14]. Weak [GGJ10]. Weakly [AMAGNA23, DRS17, Kay15]. Wealth [GGMO21]. Weather [GNI19]. Wedge [Umu11]. Weibull [CZC17, CK12, IMKEK18, KWP11, MHC21, Sul08]. Weibull-Generated [MHC21]. Weighing [DK99]. Weight [Liu11]. Weighted [AAS11a, AAS11b, CKKA22, Kan16, KN18, Mat19a, MHC21, SJ19, WZ17]. Weights [Cha05, JLKK12]. Weld [GAÖ96, Kir04]. Welded [Kir04, RWL23]. Welding [KD19, KTS13]. Well [MOBA21]. Well-Posedness [MOBA21]. WENO [LZ21]. Western [LRBMAM^{+}23]. Weyl [ÜU05]. Whale [SH18]. Wheat [DPC^{+}10]. Wheel [ÇD03]. Wheels [Tur16]. Which [Özy98]. while [dlFH19]. White [BNSDC20]. Wiley [Şen05]. Wiley-McLaren [Şen05]. Williamson [ETL22, TKK^{+}22]. Wind [AB19, ADJ18, AÇ11, ÇYAÖ05, CDDM23, EÜA13, GÇÖ12, Kum07, ZdlÁGLTT22]. Wind-Power [GÇÖ12]. Winding [PPB^{+}20, ZLLX20]. Windowing [MGGHCR^{+}20]. Windows [ÖKK03]. Windows-Based [ÖKK03]. Winkler [KTK11, LSARZ23, Mub23]. Wire [IE10, Xu10]. Wireless [Ala11, BBSC18, KIY^{+}23]. within [AK05, AFJ22, CS19, KST10, WFV21]. Without [CC11, SAI04b]. WOFOST [HIVH21]. Wolf [SH18]. Woods [KÜM08]. Word [OPUÖ11]. Words [Sal09]. Working [LZ17]. Works [CGM^{+}23]. World [SPD18]. Worm [LOSCO^{+}20]. Wound [DK03, HTLY18, KD03]. Woven [BSSK12]. Wright [Mus16]. Wrist [KÇ11]. WSNs [BBSC18]. WV [WT15].

X [HLX^{+}18, İTİO08, AA03]. x2 [AA03]. Xe [MGT10, TM11].

Yarn [BA99, PF10a, PF10b]. YBCO [TSC11]. YBCO/CeO2/YSZ/CeO2/Ni [TSC11]. Yield [AMK96, CB04]. Yields [GNI19]. YSZ [TSC11].

Zakharov [Kar10, MJK12]. Zipper [VC23b]. Zone [MWM22]. Zone-Level [MWM22]. ZrO [AMR^{+}23]. ZrO2 [OYCS11]. ZS3 [Bil04].

References

Aktas:2001:IMM

[AA01] Alaattin Aktaş and Mehmet Aktaş. The importance of mesh modeling in the calculation of stress intensity factor. Mathematical and Computational Applications, 6(2):97–101, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/97.

Abbasbandy:2002:NSI

[AA02] Saeid Abbasbandy and Mohammad Ali Fariborzi Araghi. Numerical solution of improper integrals with valid implementation. Mathematical and Computational Applications, 7 (1):83–91, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/83.

Abiyev:2003:SRF

[AA03] Asker Ali Abiyev and Azer Abiyev. The simple recurrent formulas to find a sequences of numbers satisfying equation x2+(x+1)2=z2 and the properties of these integers. Mathematical and Computational Applications, 8(2):173– 180, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/173.

Akcay:2013:MRO

[AA13a] Hakan Akçay and A. Sermet Anagün. Multi response optimization application on a manufacturing factory. Mathematical and Computational Applications, 18(3):531–538, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/531.

Atmaca:2013:TSC

[AA13b] Sibel Paşalı Atmaca and Ömer Akgüller. The time scale calculus approach to the geodesic problem in 3D dynamic data sets. Mathematical and Computational Applications, 18(3):421– 427, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/421.

Acar:2016:MPM

[AA16] Engin Acar and Hakan Soner Aplak. A model proposal for a multi-objective and multi-criteria vehicle assignment problem: an application for a security organization. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/39.

Abbasbandy:2011:NSO

[AAD11] S. Abbasbandy, T. Allahviranloo, and P. Darabi. Numerical solution of N-order fuzzy differential equations by Runge– Kutta method. Mathematical and Computational Applications, 16(4):935–946, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/935.

Abbasbandy:2011:VIM

[AADS11] S. Abbasbandy, T. Allahviranloo, P. Darabi, and O. Sedaghatfar. Variational iteration method for solving n-th order fuzzy differential equations. Mathematical and Computational Applications, 16(4):819–829, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/819.

Abernathy:2021:OCH

[AAGH21] Zachary Abernathy, Kristen Abernathy, Andrew Grant, and Paul Hazelton. Optimal control of an HIV model with gene therapy and latency reversing agents. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/77.

Angelova:2024:ESE

[AAR24] Silvija Angelova, Maria Angelova, and Rositsa Raikova. Estimating surface EMG activity of human upper arm muscles using InterCriteria analysis. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/8.

Awosusi:2022:HDF

[AARW22] Abraham Ayobamiji Awosusi, Tomiwa Sunday Adebayo, Husam Rjoub, and Wing-Keung Wong. How do financial development and renewable energy affect consumption-based carbon emissions? Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/73.

Allahviranloo:2011:AAR

[AAS11a] T. Allahviranloo, S. Abbasbandy, and R. Saneifard. An approximation approach for ranking fuzzy numbers based on weighted interval - value. Mathematical and Computational Applications, 16(3):588–597, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/588.

Allahviranloo:2011:MRF

[AAS11b] T. Allahviranloo, S. Abbasbandy, and R. Saneifard. A method for ranking of fuzzy numbers using new weighted distance. Mathematical and Computational Applications, 16(2):359– 369, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/359.

Aksoy:1996:NTP

[AB96a] M. S. Aksoy and M. Bayram.

A new technique to process and recognize barcodes using induction. Mathematical and Computational Applications, 1(2):1–6, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/1.

Allin:1996:GTK

[AB96b] Yılmaz Allin and Mehmet Emin Bozhüyük. The group of twist knots. Mathematical and Computational Applications, 1(2): 7–15, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/7.

Adanur:1999:NSC

[AB99] Sabit Adanur and Sayavur I. Bakhtiyarov. Numerical study of collision efficiency of dust particles. Mathematical and Computational Applications, 4(3):297–303, December 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/297.

Al-Bastaki:2003:AGI

[AB03] Yousif A. Latif Al-Bastaki. Adaptive GIS image compression and restoration using neural networks. Mathematical and Computational Applications, 8(2): 191–200, August 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/191.

Altintas:2006:DSS

[AB06] Gökhan Altıntaş and Muhiddin Bağcı. Determination of the steady state response of EFEF/ VFVF supported rectangular specially orthotropic plates. Mathematical and Computational Applications, 11 (1):51–63, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/51.

Alhato:2019:DPC

[AB19] Mohammed Mazen Alhato and Soufiene Bouallègue. Direct power control optimization for doubly fed induction generator based wind turbine systems. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/77.

Annunziato:2023:FPA

[AB23] Mario Annunziato and Alfio Borzì. Fokker– Planck analysis of superresolution microscopy images. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/6/113.

Astuto:2022:CTA

[ABH^{+}22] Clarissa Astuto, Daniele Boffi, Jan Haskovec, Peter Markowich, and Giovanni Russo. Comparison of two aspects of a PDE model for biological network formation. Mathematical and Computational Applications, 27(5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/5/87.

Abiyev:2003:FIS

[Abi03] Rahib Abiyev. Fuzzy inference system based on neural network for technological process control. Mathematical and Computational Applications, 8(2):245–252, August 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/245.

Aliyev:2010:AEM

[ABKU10] Rovshan Aliyev, Nurgul Okur Bekar, Tahir Khaniyev, and Ihsan Unver. Asymptotic expansions for the moments of the boundary functionals of the renewal reward process with a discrete interference of chance. Mathematical and Computational Applications, 15 (1):117–126, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/117.

Akdag:2016:EPS

[ABO^{+}16] Mustafa Akdag, Turan Batar, Fuat Okumus, Onur Sayman, and Volkan Arikan. Elastoplastic stress analysis in a ductile adhesive & aluminum adherends. Mathematical and Computational Applications, 21 (1):??, March 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/1/5.

Ata:2011:AHA

[AÇ11] Raşit Ata and Numan S.

Çetin. Analysis of height affect on average wind speed by ANN. Mathematical and Computational Applications, 16(3):556–564, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/556.

Andraz:2019:BSC

[ACC19] Jorge M. Andraz, Renato Candeias, and Ana C. Conceição. Bridging symbolic computation and economics: a dynamic and interactive tool to analyze the price elasticity of supply. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/87.

Arnal:2020:EPE

[ACC20] Ana Arnal, Fernando Casas, and Cristina Chiralt. Exponential perturbative expansions and coordinate transformations. Mathematical and Computational Applications, 25(3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/50.

Acedo:2019:HMM

[Ace19] Luis Acedo. A hidden Markov model for the linguistic analysis of the Voynich manuscript. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/14.

Arcolezi:2021:PGI

[ACGC21] Héber H. Arcolezi, Selene Cerna, Christophe Guyeux, and Jean-François Couchot. Preserving geo-indistinguishability of the emergency scene to predict ambulance response time. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/56.

Avci:1999:FNR

[AÇGJ99] Halil Avcı, Mehmet Çivi, Ziya B. Güvenç, and Julius Jellinek. Fragmentation of a non-rotating ni19 cluster: a molecular dynamics study. Mathematical and Computational Applications, 4(2): 99–106, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/99.

Arena:2022:PMA

[ACL^{+}22] Fabio Arena, Mario Collotta, Liliana Luca, Marianna Ruggieri, and Francesco Gaetano Termine. Predictive maintenance in the automotive sector: a literature review. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/2.

Akkari:2019:TSR

[ACM19] Nissrine Akkari, Fabien Casenave, and Vincent Moureau. Time stable reduced order modeling by an enhanced reduced order basis of the turbulent and incompressible 3D Navier–Stokes equations. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/45.

Amer:2004:CNS

[AD04] S. M. Amer and S. Dardery. On a class of nonlinear singular integral equations with Carleman shift in generalized Holder spaces. Mathematical and Computational Applications, 9(2): 225–236, August 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/2/225.

Allan:2011:ANC

[AD11] Mohamed M. Allan and Smah M. Dardery. On the application of new convergence criteria for Kantorovich method to nonlinear singular integral equation with shift. Mathematical and Computational Applications, 16(4):913–922, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/913.

Akyildiz:1999:PFT

[ADE99] F. T. Akyıldız, H. Demir, and V. S. Ertürk. Peristaltic flow of a thirdgrade fluid in a planar channel. Mathematical and Computational Applications, 4(2): 113–120, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/113.

Arreyndip:2018:NMF

[ADJ18] Nkongho Ayuketang Arreyndip, Alain Moise Dikande, and Ebobenow Joseph. Nonlinear multifrequency dynamics of wind turbine components with a single-mesh helical Gear train. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/12.

Akram:2019:DMA

[ADS19] Muhammad Akram, Jawaria Mohsan Dar, and Sundas Shahzadi. Decision making approach under Pythagorean dombi fuzzy graphs for selection of leading textile industry. Mathematical and Computational Applications, 24 (4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/102.

Altin:2002:SSE

[AE02] Abdullah Altın and Ayşegül Erençin. Some solutions for an equation of order of 4p. Mathematical and Computational Applications, 7 (1):33–40, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/33.

Akay:2003:DLB

[AE03] Oğuz Akay and Kayhan Erciyeş. A dynamic load balancing model for a distributed system. Mathematical and Computational Applications, 8(3):353–360, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/353.

Akogul:2016:CIC

[AE16] Serkan Akogul and Murat Erisoglu. A comparison of information criteria in clustering based on mixture of multivariate normal distributions. Mathematical and Computational Applications, 21(3):??, September 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/34.

Abdou:2001:FVI

[AEB01] M. A. Abdou and A. A.

El-Bary. Fredholm– Volterra integral equation with potential kernel. Mathematical and Computational Applications, 6(2):113–122, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/113.

Abouelkheir:2018:TNC

[AERE18] Imane Abouelkheir, Fadwa El Kihal, Mostafa Rachik, and Ilias Elmouki. Time needed to control an epidemic with restricted resources in SIR model with short-term controlled population: a fixed point method for a free isoperimetric optimal control problem. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/64.

Amram:2020:TMM

[AFGV20] Meirav Amram, Etan Fisher, Shai Gul, and Uzi Vishne. A transformational modified Markov process for chord-based algorithmic composition. Mathematical and Computational Applications, 25(3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/3/43.

Andreucci:2022:IMD

[AFJ22] Carlos Aurelio Andreucci, Elza M. M. Fonseca, and Renato N. Jorge. Increased material density within a new biomechanism. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/90.

Alameddin:2019:TOP

[AFN^{+}19] Shadi Alameddin, Amélie Fau, David Néron, Pierre Ladevèze, and Udo Nackenhorst. Toward optimality of proper generalised decomposition bases. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/30.

Akbulut:2003:BAP

[AG03] Harnit Akbulut and Ömer Gündoğdu. On the buckling analysis of plates reinforced with stiff ribs. Mathematical and Computational Applications, 8(1):103– 110, April 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/103.

Arikan:2012:DAT

[AG12] Umut Arıkan and Fikret Gürgen. Discrimination ability of timedomain features and rules for arrhythmia classification. Mathematical and Computational Applications, 17(2):111– 120, August 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/2/111.

Al-Gharabli:2022:NTS

[AGANA22] Mohammad M. Al-Gharabli, Adel M. Almahdi, Maher Noor, and Johnson D. Audu. Numerical and theoretical stability study of a viscoelastic plate equation with nonlinear frictional damping term and a logarithmic source term. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/1/10.

Abushaikha:2021:BLT

[AGKT21] Ahmad Abushaikha, Dominique Guérillot, Mostafa Kadiri, and Saber Trabelsi. Buckley–Leverett theory for a Forchheimer– Darcy multiphase flow model with phase coupling. Mathematical and Computational Applications, 26(3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/3/60.

Akram:2018:STP

[AHID18] Muhammad Akram, Amna Habib, Farwa Ilyas, and Jawaria Mohsan Dar. Specific types of Pythagorean fuzzy graphs and application to decision-making. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/3/42.

Ansari:2018:GAA

[AHK18] Moin A. Ansari, Azeem Haidar, and Ali N. A.

Koam. On a graph associated to UP-algebras. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/61.

Ahbasov:1999:DFS

[AHM99] T. Ahbasov, S. Herdem, and A. Memnedov. On the drag force of a solid sphere in power law flow model. Mathematical and Computational Applications, 4(2): 161–167, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/161.

Amer:2005:MNM

[AHR05] S. M. Amer, A. A. M.

Hassan, and A. H. Radwan. On the modified Newton’s method for the solution of a class of nonlinear singular integro-differential equations. Mathematical and Computational Applications, 10(2):269– 281, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/269.

Aharony:2023:AQR

[AHVG23] Adam Aharony, Ron Hindi, Maor Valdman, and Shai Gul. Applying quaternions to recognize hidden details in images: Rothko as a case study. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/66.

Abri:2012:CPN

[AJ12] A. Gholam Abri and M. Fallah Jelodar. A complementary on proposing a new model on data envelopment analysis by considering non discretionary factors and a review on previous models. Mathematical and Computational Applications, 17(3):244–246, December 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/244.

Adair:2020:PFV

[AJ20] Desmond Adair and Martin Jaeger. In-plane free vibration analysis of a scimitar-type rotating curved beam using the Adomian modified decomposition method. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/68.

Aslan:2002:TDA

[AK02] Züleyha Aslan and Ramazan Karakuzu. Transient dynamic analysis of laminated composite plate subjected to lowvelocity impact. Mathematical and Computational Applications, 7 (1):73–82, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/73.

Afsar:2005:FES

[AK05] O. Afsar and O. Kayacan. Free energies and specific heats of the nematics within nonextensive thermostatistics. Mathematical and Computational Applications, 10(3):307–312, December 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/307.

Abbasov:2009:AQM

[AK09] Babek Abbasov and Serdar Korukoğlu. An active queue management algorithm for reducing packet loss rate. Mathematical and Computational Applications, 14 (1):65–72, April 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/1/65.

Arslan:2011:SHG

[AK11] Nurullah Arslan and Ferhat Karaca. Simulation of hemodynamics in a graft-to-vein anastomoses by adaptive neuro-fuzzy based modeling. Mathematical and Computational Applications, 16(3):702– 711, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/702.

Akararungruangkul:2018:MDE

[AK18] Raknoi Akararungruangkul and Sasitorn Kaewman. Modified differential evolution algorithm solving the special case of location routing problem. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/3/34.

Abukhaled:2019:ESA

[AK19] Marwan Abukhaled and Suheil Khuri. An efficient semi-analytical solution of a one-dimensional curvature equation that describes the human corneal shape. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/8.

Akbaba:2009:MIV

[Akb09] Mehmet Akbaba. Motor input voltage and rectifier firing angle variation with load torque in constant current operated induction motors. Mathematical and Computational Applications, 14(1):73–84, April 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/73.

Akhiev:2004:SCG

[Akh04] Seyidali S. Akhiev. Solvability conditions and green functional concept for local and nonlocal linear problems for a second order ordinary differential equation. Mathematical and Computational Applications, 9(3): 349–358, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/349.

Al-Kufi:2017:ABG

[AKHHM17] Mohammed Abdul Hameed Jassim Al-Kufi, Hayder Raheem Hashim, Ameer Mohammed Hussein, and Hind Rustum Mohammed. An algorithm based on GSVD for image encryption. Mathematical and Computational Applications, 22(2):??, June 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/2/28.

Akin:2011:NEE

[Aki11] Osman Çağlar Akin. Non-ergodic to ergodic regime transition via intrusions in a non-Poisson renewal system with power law tails. Mathematical and Computational Applications, 16 (1):248–257, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/248.

Aksoy:2005:DSM

[Aks05a] M. Sabih Aksoy. Dynamic system modeling using rules3 induction algorithm. Mathematical and Computational Applications, 10 (1):121–132, April 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/121.

Aksoy:2005:PDT

[Aks05b] Mehmet Sabih Aksoy. Pruning decision trees using rules3 inductive learning algorithm. Mathematical and Computational Applications, 10 (1):113–120, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/113.

Aksoy:2008:RRF

[Aks08] Mehmet Sabih Aksoy. A review of rules family of algorithms. Mathematical and Computational Applications, 13 (1):51–60, April 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/1/51.

Aktas:2001:DDF

[Akt01] Alaattin Aktaş. Determination of the deflection function of a composite cantilever beam using theory of anisotropic elasticity. Mathematical and Computational Applications, 6(1):67– 74, April 2001. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/67.

Abdelhafiz:2004:GFT

[AKT04] M. I. Abdelhafiz, F. M.

Khaliel, and Ramadan S. Tantawi. General fermion triangle amplitude in the coupling of spin-1 particles and Ward identities. Mathematical and Computational Applications, 9(2):285–298, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/285.

Alaybeyoglu:2011:EET

[Ala11] Aysegul Alaybeyoglu. Energy efficient target tracking with particle filtering techmique in wireless sensor networks. Mathematical and Computational Applications, 16(4):809–818, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/809.

Alcantara-Lopez:2020:QAM

[ALFBPLE20] Fernando Alcántara-López, Carlos Fuentes, Fernando Brambila-Paz, and Jesús López-Estrada. Quasianalytical model of the transient behavior pressure in an oil reservoir made up of three porous media considering the fractional time derivative. Mathematical and Computational Applications, 25(4):??, December 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/74.

Alharthi:2021:CVS

[Alh21] Hamzah Abdulrahman Alharthi. Characterization of the vibration and strain energy density of a nanobeam under two-temperature generalized thermoelasticity with fractional-order strain theory. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/78.

Attia:2019:CRE

[ALK19] Raghda A. M. Attia, Dianchen Lu, and Mostafa M. A. Khater. Chaos and relativistic energymomentum of the nonlinear time fractional Duffing equation. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/10.

Allahverdi:1999:PDN

[All99] Novruz Allahverdi. Procudere for definition of neighborhood in faulty hypercube multiprocessor. Mathematical and Computational Applications, 4(1):17–23, April 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/17.

Allahverdi:2002:MIS

[All02] Novruz Allahverdi. A method to increase the size of healty subcubes set in a hypercube system with node or link failures. Mathematical and Computational Applications, 7(3):189–203, December 2002. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/3/189.

Allan:2012:CMC

[All12] Mohamed M. Allan. A collocation method for a class of nonlinear singular integral equations with a Carleman shift. Mathematical and Computational Applications, 17(2):121–131, August 2012. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/2/121.

Allahem:2017:ASN

[All17]  Ali Allahem.  Ana-

lytical solution to normal forms of Hamiltonian systems. Mathematical and Computational Applications, 22 (3):??, September 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/3/37.

Almazyad:2010:NLA

[Alm10] Abdulaziz S. Almazyad.

A new look-ahead algorithm to improve the performance of Omega networks. Mathematical and Computational Applications, 15(2):156– 165, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/156.

Alp:2002:PDE

[Alp02] Ihsan Alp. Possibilistic data envelopment analysis. Mathematical and Computational Applications, 7(1):5–14, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/5.

Altintas:2009:GUI

[Alt09]  Ahmet Altintas.  A

graphical user interface for programming stepper motors used at different kinds of applications. Mathematical and Computational Applications, 14(2):139– 146, August 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/2/139.

Altintas:2010:EMP

[Alt10] Gökhan Altıntas. Effect of material properties on vibrations of nonsymmetrical axially loaded thin-walled Euler–Bernoulli beams. Mathematical and Computational Applications, 15(1):96–107, April 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/1/96.

Altintas:2012:ESS

[Alt12] Gokhan Altintas. Effect of slice step size on prediction of natural vibration properties of bone tissue. Mathematical and Computational Applications, 17(3):235– 243, December 2012. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/3/235.

Alturk:2016:RHM

[Alt16]  Ahmet Altürk.  The

regularization-homotopy method for the twodimensional Fredholm integral equations of the first kind. Mathematical and Computational Applications, 21(2):??, June 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/9.

Ali:2005:SRU

[AM05] Asif Ali and Ahmer Mehmood. Symmetry reduction of unsteady MHD aligned second grade flow equations. Mathematical and Computational Applications, 10(3):395–402, December 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/395.

Aksoy:2009:CCS

[AM09] Mehmet Sabih Aksoy and Hassan Mathkou. CSLS: Connectionist symbolic learning system. Mathematical and Computational Applications, 14(3):177–186, December 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/3/177.

Anthonyrajah:2010:CLI

[AM10] M. Anthonyrajah and D. P. Mason. Conservation laws and invariant solutions in the Fanno model for turbulent compressible flow. Mathematical and Computational Applications, 15(4):529–542, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/529.

Al-Mahdi:2023:PTI

[AM23a] Adel M. Al-Mahdi. Preconditioning technique for an image deblurring problem with the total fractional-order variation model. Mathematical and Computational Applications, 28 (5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/97.

Arnold:2023:PPD

[AM23b] Barry C. Arnold and Bangalore G. Manjunath. Pseudo-Poisson distributions with concomitant variables. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/11.

Al-Mahdi:2022:TCR

[AMAGA22] Adel M. Al-Mahdi, Mohammad M. Al-Gharabli, and Mohamed Alahyane. Theoretical and computational results of a memory-type swelling porous-elastic system. Mathematical and Computational Applications, 27(2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/27.

Al-Mahdi:2023:SRW

[AMAGNA23] Adel M. Al-Mahdi, Mohammad M. Al-Gharabli, Maher Noor, and Johnson D. Audu. Stability results for a weakly dissipative viscoelastic equation with variableexponent nonlinearity: Theory and numerics. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/5.

AtiK:1996:DYS

[AMK96] Enver AtiK, Cevdet Meriç, and Bekir Karlik. Determination of yield strenght of 2014 aluminium alloy under aging conditions by means of artifical neural networks method. Mathematical and Computational Applications, 1(2): 16–20, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/16.

Azzaoui:2023:DDF

[AMM23] Nourddine Azzaoui, Tomoko Matsui, and Daisuke Murakami. Data-driven framework for uncovering hidden control strategies in evolutionary analysis. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/103.

Ascencion-Mestiza:2023:EEC

[AMMMM^{+}23] Hector Ascencion-Mestiza, Serguei Maximov, Efrén Mezura-Montes, Juan Carlos Olivares-Galvan, Rodrigo Ocon-Valdez, and Rafael Escarela-Perez. Estimation of the equivalent circuit parameters in transformers using evolutionary algorithms. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/36.

Amoddeo:2021:MMN

[Amo21] Antonino Amoddeo. Mathematical model and numerical simulation for electric field induced cancer cell migration. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/4.

Akinwande:2023:AEV

[AMR^{+}23] Abayomi Adewale Akinwande, Dimitry Moskovskikh, Elena Romanovskaia, Oluwatosin Abiodun Balogun, J. Pradeep Kumar, and Valentin Romanovski. Applicability of extreme vertices design in the compositional optimization of 3d-printed lightweight high-entropy-alloy/B_{4}C/ ZrO_{2}/titanium trihybrid aero-composite. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/54.

Akram:2019:NDM

[AN19] Muhammad Akram and Sumera Naz. A novel decision-making approach under complex Pythagorean fuzzy environment. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/73.

Apalara:2020:LTB

[ANA20] Tijani A. Apalara, Aminu M.

Nass, and Hamdan Al Sulaimani. On a laminated Timoshenko beam with nonlinear structural damping. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/35.

Adair:2020:ETS

[ANJ20] Desmond Adair, Aigul Nagimova, and Martin Jaeger. Effect of thrust on the structural vibrations of a nonuniform slender rocket. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/29.

Anku:2023:SFP

[ANK23] Anku, Mona Narang, and Vinay Kanwar. On some fixed point iterative schemes with strong convergence and their applications. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/45.

Akarsu:2005:MCS

[AÖ05] Mustafa Akarsu and Ömer Özbaş. Monte Carlo simulation for electron dynamics in semiconductor devices. Mathematical and Computational Applications, 10 (1):19–26, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/19.

Akbaba:2009:PAS

[AÖ09] Mehmet Akbaba and Orhan Özhan. A practical algorithm for simulation of transient performance of choppercontrolled r-l and dc drive loads. Mathematical and Computational Applications, 14(3):229– 239, December 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/3/229.

Abdelrahim:2016:DSS

[AO16] Ra’ft Abdelrahim and Zurni Omar. Direct solution of second-order ordinary differential equation using a single-step hybrid block method of order five. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/12.

Adegoke:2017:ANV

[AO17] Adeshina S. Adegoke and Ayo A. Oyediran. The analysis of nonlinear vibrations of top-tensioned cantilever pipes conveying pressurized steady two-phase flow under thermal loading. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/4/44.

Aleksandrov:2019:HSS

[AOL19] Samuil R. Aleksandrov, Timo T. Overboom, and Elena A. Lomonova. 2D hybrid steady-state magnetic field model for linear induction motors. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/74.

Apalak:2023:TSF

[AR23] Mustafa Kemal Apalak and Junuthula N. Reddy. Thermal stress formation in a functionally graded Al_{2}O_{3}-adhesive single lap joint subjected to a uniform temperature field. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/82.

Araghi:2008:DCA

[Ara08] Mohammad Ali Fariborzi Araghi. Dynamical control of accuracy using the stochastic arithmetic to estimate double and improper integrals. Mathematical and Computational Applications, 13 (2):91–100, August 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/2/91.

Aguilera-Rueda:2020:DDB

[ARCRMM20] Vicente-Josué Aguilera-Rueda, Nicandro Cruz-Ramírez, and Efrén Mezura-Montes. Datadriven Bayesian network learning: a Biobjective approach to address the bias-variance decomposition. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/37.

A:1996:EDF

[ARN96] Ekrekli A., Güçlü R., and Tahralı N. Experimetntal detection of the faults in a gearbox. Mathematical and Computational Applications, 1(2): 35–41, December 1996.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/35.

Arslan:2006:CMF

[Ars06] Nurullah Arslan. Computational modeling of flow inside a diseased carotid bifurcation. Mathematical and Computational Applications, 11(3):173–180, December 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/3/173.

Arslan:2007:MSF

[Ars07] Nurullah Arslan. Mathematical solution of the flow field over glycocalyx inside vascular system. Mathematical and Computational Applications, 12(3):173–179, December 2007. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/3/173.

Arvasi:2004:CMA

[Arv04] Zekeriya Arvasi. Crossed modules of algebras. Mathematical and Computational Applications, 9(2):173–182, August 2004. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/9/2/173.

Ahmad:2000:GMM

[AS00] Rais Ahmad and Hanif Salahuddin. General mixed multivalued mildly nonlinear variational inequalities in h-spaces. Mathematical and Computational Applications, 5(2):91–94, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/91.

Abdou:2004:TPI

[AS04] M. A. Abdou and F. A.

Salama. Thermoelastic problem for an infinite plate with a curvilinear hole having finite poles. Mathematical and Computational Applications, 9(2):237– 248, August 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/237.

Abbasbandy:2009:AVI

[AS09] Saeid Abbasbandy and Elyas Shivanian. Application of the variational iteration method for system of nonlinear Volterra’s integrodifferential equations. Mathematical and Computational Applications, 14(2):147–158, August 2009. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/14/2/147.

Akram:2020:PFP

[ASA20] Muhammad Akram, Shumaiza, and José Carlos R. Alcantud. An mpolar fuzzy PROMETHEE approach for AHP-assisted group decision-making. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/26.

Akram:2020:SFG

[ASAH20] Muhammad Akram, Danish Saleem, and Talal Al-Hawary. Spherical fuzzy graphs with application to decision-making. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/8.

Abrarov:2024:IMC

[ASJQ24] Sanjar M. Abrarov, Rehan Siddiqui, Rajinder Kumar Jagpal, and Brendan M. Quine. An iterative method for computing π by argument reduction of the tangent function. Mathematical and Computational Applications, 29(2):17:1– 17:23, April 2024. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/2/17.

Azar:2021:RHL

[ASK21] Ahmad Taher Azar, Fernando E. Serrano, and Nashwa Ahmad Kamal. Robust H_{∞} loop shaping controller synthesis for SISO systems by complex modular functions. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/21.

Aslam:2018:FPN

[Asl18] Muhammad Aslam. A family of 5-point nonlinear ternary interpolating subdivision schemes with C^{2} smoothness. Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/18.

Abdelhafiz:2005:EPS

[AT05a] M. I. Abdelhafiz and R. S. Tantawi. Electromagnetic properties of Z and γ self-couplings interactions. Mathematical and Computational Applications, 10(2):261– 267, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/2/261.

Aykul:2005:CSA

[AT05b] Halil Aykul and Mustafa Toparlı. A comparison of the stress analysis of an unrestored and restored tooth with amalgam and composite resin. Mathematical and Computational Applications, 10(1):89–98, April 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/89.

Aykul:1999:CTD

[ATK99] H. Aykul, M. Toparlı, and N. S. Köksal. Calculation of temperature distribution on a crown tooth by using threedimensional finite element method. Mathematical and Computational Applications, 4 (1):31–37, April 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/31.

Aykul:1997:TDS

[ATSA97] Halil Aykul, Mustafa Toparh, Onur Sayman, and Celal Artunç. Three dimensional stress analysis in the crowned human teeth. Mathematical and Computational Applications, 2 (1):13–20, April 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/1/13.

Abu-Taleb:2007:MBD

[ATZES^{+}07] A. S. Abu-Taleb, E. M. E.

Zayed, W. M. El-Sayed, A. M. Badawy, and O. A. Mohammed. Multichannel blind deconvolution using a generalized Gaussian source model. Mathematical and Computational Applications, 12 (1):1–9, April 2007. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/1.

Ataberk:2003:SIF

[AUA03] Necati Ataberk, Mesut Uyaner, and Ahmet Avci. Stress intensity factors for a ring-shaped crack in a transversely isotropic thick layer. Mathematical and Computational Applications, 8 (1):79–88, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/79.

Abbasbandy:2002:NSF

[AV02] S. Abbasbandy and T. Allah Viranloo. Numerical solution of fuzzy differential equation. Mathematical and Computational Applications, 7 (1):41–52, April 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/1/41.

Abbasbandy:2004:EMI

[AV04] S. Abbasbandy and T. Allah Viranloo. Extrapolation method for improving the solution of fuzzy initial value problems. Mathematical and Computational Applications, 9(2):205– 214, August 2004. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/2/205.

Aduke:2023:ANH

[AVC23] Rhoda Ngira Aduke, Martin P. Venter, and Corné J. Coetzee. An analysis of numerical homogenisation methods applied on corrugated paperboard. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/46.

Agyingi:2019:AML

[AW19] Ephraim Agyingi and Tamas Wiandt. Analysis of a model of leishmaniasis with multiple time lags in all populations.

Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/63.

Awwal:2020:PHS

[AWK^{+}20] Aliyu Muhammed Awwal, Lin Wang, Poom Kumam, Hassan Mohammad, and Wiboonsak Watthayu. A projection Hestenes–Stiefel method with spectral parameter for nonlinear monotone equations and signal processing. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/27.

Abbasbandy:2014:ASN

[AYG14] Saeid Abbasbandy, Muhammet Yürüsoy, and Hüseyin Güllüce. Analytical solutions of non-linear equations of power-law fluids of second grade over an infinite porous plate. Mathematical and Computational Applications, 19(2):124–133, August 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/124.

Bakhtiyarov:1999:AWY

[BA99] Sayavur Bakhtiyarov and Sabit Adanur. Airflow over wavy yarn in air-jet filling insertion. Mathematical and Computational Applications, 4(1): 1–7, April 1999. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/1.

Basurrah:2003:ESE

[BA03] Salva Saleh Basurrah and Mehmet Akbaba. Effect of shaft elasticity on the performance of slip energy recovery drives. Mathematical and Computational Applications, 8(1):1–11, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/1.

Behzadi:2011:NSS

[BA11] Sh. Sadigh Behzadi and M. A. Fariborzi Araghi. Numerical solution for solving Burgers–Fisher eguation by using iterative methods. Mathematical and Computational Applications, 16(2):443– 455, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/443.

Bacanli:2013:ORD

[Bac13]  Sevil Bacanlı.  Opti-

mal restricted designs for the inverse Gaussian mean. Mathematical and Computational Applications, 18(2):94– 102, August 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/2/94.

Badea:2022:CDA

[Bad22] Lori Badea. On the convergence of the damped additive Schwarz methods and the subdomain coloring. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/59.

Bagci:2010:NNM

[Bağ10] Muhiddin Bağcı. Neural network model for moment-curvature relationship of reinforced concrete sections. Mathematical and Computational Applications, 15 (1):66–78, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/66.

Bakir:2011:EOS

[Bak11] Pelin Gündeş Bakir.

Evaluation of optimal sensor placement techniques for parameter identification in buildings. Mathematical and Computational Applications, 16(2):456–466, August 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/456.

Bolat:1996:PAG

[BAS96] Ahmet Bolat and Khaled As-Saifan. Procedures for aircraft gate assigment. Mathematical and Computational Applications, 1(1):9–14, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/9.

Bashirov:1999:MMS

[Bas99] Novruz Bashirov. Mathematical model of sulphomethylation process of nonylphenol. Mathematical and Computational Applications, 4(1): 9–15, April 1999. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/9.

Baylar:2009:NMV

[BAUO09] Ahmet Baylar, M. Cihan Aydin, Mehmet Unsal, and Fahri Ozkan. Numerical modeling of venturi flows for determining air injection rates using fluent V6.2. Mathematical and Computational Applications, 14 (2):97–108, August 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/2/97.

Baykasoglu:2003:DOM

[Bay03] Adil Baykasoğlu. Designoptimization of mechanical components using the “Multiple objective tabu search” method. Mathematical and Computational Applications, 8(1):35–47, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/35.

Bulut:2016:NCH

[BB16] Hasan Bulut and Haci Mehmet Baskonus. New complex hyperbolic function solutions for the (2 + 1)-dimensional dispersive long water–wave system. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/6.

BenYahia:2018:TDE

[BBD^{+}18] Mohammed Ben Yahia, Kamel Boughrara, Frédéric Dubas, Lazhar Roubache, and Rachid Ibtiouen. Two-dimensional exact subdomain technique of switched reluctance machines with sinusoidal current excitation. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/59.

Buchfink:2019:SMO

[BBH19] Patrick Buchfink, Ashish Bhatt, and Bernard Haasdonk. Symplectic model order reduction with non-orthonormal bases. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/43.

Bowers:2020:EMC

[BBK20] Abigail Bowers, Jared Bunn, and Myles Kim. Efficient methods to calculate partial sphere surface areas for a higher resolution finite volume method for diffusionreaction systems in biological modeling. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/2.

Behl:2022:OIT

[BBMA22] Ramandeep Behl, Sonia Bhalla, Fouad Mallawi, and Majed Aali Alsulami. An optimal iterative technique for multiple root finder of nonlinear problems. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/74.

Boaretto:2019:SPS

[BBPL19a] Bruno Rafael Reichert Boaretto, Roberto C. Budzinski, Thiago L. Prado, and Sergio Roberto Lopes. Suppression of phase synchronization in scale-free neural networks using external pulsed current protocols. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/46.

Budzinski:2019:IDT

[BBPL19b] Roberto C. Budzinski, Bruno Rafael Reichert Boaretto, Thiago L. Prado, and Sergio Roberto Lopes. Investigation of details in the transition to synchronization in complex networks by using recurrence analysis. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/42.

Bashirpour:2018:IDS

[BBSC18] Hamed Bashirpour, Saman Bashirpour, Shahaboddin Shamshirband, and Anthony T. Chronopoulos. An improved digital signature protocol to multi-user broadcast authentication based on elliptic curve cryptography in wireless sensor networks (WSNs). Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/17.

Bayrak:1997:GNS

[BC97] Vural Bayrak and Mehmet Can. Global nonexistence of solutions of the vibrations of a riser. Mathematical and Computational Applications, 2(1):45–52, April 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/1/45.

Bahsi:2013:TMS

[BÇ13] M. Mustafa Bahşı and Mehmet Çevik. Taylor matrix solution of the mathematical model of the RLC circuits. Mathematical and Computational Applications, 18(3):467–475, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/467.

Birinci:2002:TSP

[BÇE02] Ahmet Birinci, Mehmet Çakıroğlu, and Ragıp Erdöl. Two strips problem resting on an elastic foundation. Mathematical and Computational Applications, 7(3): 275–286, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/275.

Boffi:2023:PSF

[BCGS23] Daniele Boffi, Fabio Credali, Lucia Gastaldi, and Simone Scacchi. A parallel solver for FSI problems with fictitious domain approach. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/59.

Bastiaens:2019:SEM

[BCK^{+}19] Koen Bastiaens, Mitrofan Curti, Dave C. J. Krop, Sultan Jumayev, and Elena A. Lomonova. Spectral element method modeling of eddy current losses in high-frequency transformers. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/28.

Bhamjee:2022:MDB

[BCN22] Muaaz Bhamjee, Simon H. Connell, and André Leon Nel. The modification of the dynamic behaviour of the cyclonic flow in a hydrocyclone under surging conditions. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/88.

Beltran:2020:PTG

[BCS20] Fernanda Beltrán, Oliver Cuate, and Oliver Schütze. The Pareto tracer for general inequality constrained multi-objective optimization problems. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/80.

Baykasoglu:2009:SUT

[BD09] Adil Baykasoğlu and Türkay Dereli. Simple and u-type assembly line balancing by using an ant colony based algorithm. Mathematical and Computational Applications, 14(1):1–12, April 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/1.

Bootland:2022:CDG

[BD22] Niall Bootland and Victorita Dolean. Can DtN and GenEO coarse spaces be sufficiently robust for heterogeneous Helmholtz problems? Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/35.

Biswas:2018:CEA

[BDC^{+}18] Dibyendu Biswas, Suman Dolai, Jahangir Chowdhury, Priti K. Roy, and Ellina V. Grigorieva. Cost-effective analysis of control strategies to reduce the prevalence of cutaneous leishmaniasis, based on a mathematical model. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/38.

Baroni:2019:TRA

[BdCCF19] Rodrigo Simile Baroni, Ricardo Egydio de Carvalho, Bruno Castaldi, and Bruno Furlanetto. Time recurrence analysis of a near singular billiard. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/50.

Benmessaoud:2019:CME

[BDH19] Youcef Benmessaoud, Frédéric Dubas, and Mickael Hilairet. Combining the magnetic equivalent circuit and Maxwell–Fourier method for eddy-current loss calculation. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/60.

Brands:2019:ROM

[BDMS19] Benjamin Brands, Denis Davydov, Julia Mergheim, and Paul Steinmann. Reduced-order modelling and homogenisation in magneto-mechanics: a numerical comparison of established hyperreduction methods. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/20.

Boyukata:1999:DNN

[BDÖ^{+}99] Mustafa Böyükata, Perihan Durmuş, Süleyman Özçelik, Ziya B. Güvenç, and Julius Jellinek. D2 +Nin(T), n=7 and 9, collision system. Mathematical and Computational Applications, 4 (1):61–67, April 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/61.

Birinci:1999:FCB

[BE99] Ahmet Birinci and Ragıp Erdöl. Frictionless contact between a rigid stamp and an elastic layered composite resting on simple supports. Mathematical and Computational Applications, 4(3): 261–272, December 1999.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/261.

Bechtel:2017:NUI

[Bec17] Gordon Bechtel. New unidimensional indexes for China. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/1/13.

Ballan:1999:PCB

[BG99] Meltem Ballan and Fikret Gurgen. On the principal component based fingerprint classification using directional images. Mathematical and Computational Applications, 4(2):91–97, August 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/2/91.

Bal:2002:DEA

[BG02] Hasan Bal and Ayhan Gölcükcü. Data envelopment analysis: an application to Turkish banking industry. Mathematical and Computational Applications, 7 (1):65–72, April 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/1/65.

Basak:2004:FBP

[BG04] Hüdayim Başak and Mahmut Gülesin. A feature based parametric design program and expert system for design. Mathematical and Computational Applications, 9(3):359–370, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/359.

Barazandeh:2018:NSF

[BG18a] Yousef Barazandeh and Bahman Ghazanfari. Numerical solution for fuzzy enzyme kinetic equations by the Runge– Kutta method. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/16.

Bazgir:2018:ESF

[BG18b] Hamed Bazgir and Bahman Ghazanfari. Existence of solutions for fractional integro-differential equations with nonlocal boundary conditions. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/3/36.

Bala:2019:GSA

[BG19] Saminu Bala and Bello Gimba. Global sensitivity analysis to study the impacts of bed-nets, drug treatment, and their efficacies on a two-strain malaria model. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/32.

Bombar:2010:AIP

[BGA10] Gökçen Bombar, M. Sükrü Güney, and Mustafa S. Altınakar. Application of image processing in unsteady flows to investigate bed load transport. Mathematical and Computational Applications, 15(3):420–427, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/420.

Baracli:1996:QFD

[BGG96] Hayri Baraçlı, Bahadir Gülsün, and Ali Fuat Güneri. Quality function deployment approach to satisfying customers demands and an implementation at shoes company.

Mathematical and Computational Applications, 1(1):1–8, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/1.

Berberler:2013:GAS

[BGN13] Murat Ersen Berberler, Asli Guler, and Urfat G. Nurıyev. A genetic algorithm to solve the multidimensional knapsack problem. Mathematical and Computational Applications, 18(3):486– 494, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/486.

Banholzer:2021:RBI

[BGRV21] Stefan Banholzer, Bennet Gebken, Lena Reichle, and Stefan Volkwein. ROM-based inexact subdivision methods for PDE-constrained multiobjective optimization. Mathematical and Computational Applications, 26(2):??, June 2021. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/26/2/32.

Bhatnagar:2024:ISM

[Bha24] Saakaar Bhatnagar. Investigating the surrogate modeling capabilities of continuous time echo state networks. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/29/1/9.

Bai:2016:MCH

[BhMsG16] Long Bai, Lu han Ma, and Xin sheng Ge. The modeling and calculation of the heading machine based on differential geometry. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/4/40.

Banda:2010:CDF

[BHN10] Mapundi K. Banda, Michael Herty, and Jean-Medard T. Ngnotchouye. Coupling drift-flux models with unequal sonic speeds. Mathematical and Computational Applications, 15(4):574– 584, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/574.

Bildik:1999:SED

[Bİ99] Necdet Bildik and Mustafa İnç. On some extensions of dynamical homeomorphism and dynamical isomorphism in dynamical system. Mathematical and Computational Applications, 4(2): 189–194, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/189.

Bilgin:2002:MTS

[Bil02] Tunay Bilgin. Matrix transformations of some generalized analytic sequence spaces. Mathematical and Computational Applications, 7(2): 165–170, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/165.

Bilgin:2004:PTU

[Bil04] Tevfik Bilgin. Parametrization of torsion units in U1(ZS3). Mathematical and Computational Applications, 9(1):73– 77, April 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/73.

Banichuk:2020:AMB

[BIM^{+}20] Nikolay Banichuk, Svetlana Ivanova, Evgeny Makeev, Juha Jeronen, and Tero Tuovinen. Added-mass based efficient fluid–structure interaction model for dynamics of axially moving panels with thermal expansion. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/9.

Baikov:1999:ISS

[BK99] V. A. Baikov and C. M. Khalique. Invariant solutions for soil water equations. Mathematical and Computational Applications, 4(3):217–223, December 1999. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/217.

Balli:2009:OSS

[BK09] Serkan Ballı and Serdar Korukoğlu. Operating system selection using fuzzy AHP and TOP-SIS methods. Mathematical and Computational Applications, 14(2):119– 130, August 2009. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/2/119.

Babacan:2010:CGL

[BK10] H. Babacan and O. Kayacan. Calculation of generalized level densities for nuclei in mass region 20 < a < 50. Mathematical and Computational Applications, 15(2):230– 239, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/230.

Babacan:2008:GSM

[BKA08] Halil Babacan, Ozhan Kayacan, and Ramazan Atici. Generalized statistical mechanics for nucleus. Mathematical and Computational Applications, 13(2):81– 90, August 2008. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/2/81.

Benli:2011:EPR

[BKOS11] Semih Benli, Mustafa Karamolla, Fuat Okumus, and Onur Sayman. Elastic-plastic and residual stresses in clamped thermoplastic composite laminates loaded transversely. Mathematical and Computational Applications, 16(4):849– 857, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/849.

Bermudez:2018:MSH

[BLEB^{+}18] José-Roberto Bermúdez, Francisco-Ronay López-Estrada, Gildas Besançon, Guillermo Valencia-Palomo, Lizeth Torres, and Héctor-Ricardo Hernández. Modeling and simulation of a hydraulic network for leak diagnosis. Mathematical and Computational Applications, 23(4):??, December 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/70.

Bicer:2011:PSU

[BM11] Ahmet Biçer and Kaan Manisa. Properties of symmetric uclear matter with Skyrme iteractios. Mathematical and Computational Applications, 16(4):900–912, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/900.

Balmelli:2021:EMS

[BM21] Simone Balmelli and Francesco Moresino. Estimating the market share for new products with a split questionnaire survey. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/7.

Bouhali:2021:SAS

[BMT^{+}21] Keltoum Bouhali, Abdelkader Moumen, Khadiga W. Tajer, Khdija O. Taha, and Yousif Altayeb. Spatial analyticity of solutions to Korteweg–de Vries type equations. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/75.

Banholzer:2022:TRR

[BMV22] Stefan Banholzer, Luca Mechelli, and Stefan Volkwein. A trust region reduced basis pascolettiserafini algorithm for multi-objective PDEconstrained parameter optimization. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/39.

Bastien:2020:CBW

[BNSDC20] Christophe Bastien, Clive Neal-Sturgess, Huw Davies, and Xiang Cheng. Computing brain white and grey matter injury severity in a traumatic fall. Mathematical and Computational Applications, 25(3):??, September 2020.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/3/61.

Bozdogan:2007:AMF

[BÖ07] Kanat Burak Bozdogan and Duygu Öztürk. An approximate method for free vibration analysis of multi-bay coupled shear walls. Mathematical and Computational Applications, 12 (1):41–50, April 2007. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/1/41. See dis-

cussion [Çat10].

Bozdogan:2010:VAA

[BO10] Kanat Burak Bozdogan and Duygu Ozturk. Vibration analysis of asymmetric-plan frame buildings using transfer matrix method. Mathematical and Computational Applications, 15(2):279–288, August 2010. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/15/2/279.

Benmessaoud:2020:IVP

[BODH20] Youcef Benmessaoud, Daoud Ouamara, Frédéric Dubas, and Mickael Hilairet. Investigation of volumic permanentmagnet eddy-current losses in multi-phase synchronous machines from hybrid multi-layer model. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/14.

Bagdatli:2011:NLT

[BÖÖ11] S. Murat Bağdatlı, H. Rıdvanz Öz, and Erdoğan Özkaya. Non-linear transverse vibrations and 3:1 internal resonances of a tensioned beam on multiple supports. Mathematical and Computational Applications, 16 (1):203–215, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/203.

Boone:2017:QCS

[Boo17] Donald C. Boone. Quantum coherent states and path integral method to stochastically determine the anisotropic volume expansion in lithiated silicon nanowires. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/4/41.

Boran:2011:IIF

[Bor11] Fatih Emre Boran. An integrated intuitionistic fuzzy multi criteria decision making method for facility location selection. Mathematical and Computational Applications, 16(2):487– 496, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/487.

Boglou:1999:ACD

[BP99] Anastasios K. Boglou and Demetrios P. Papadopoulos. Assessment and comparison of discrete optimal multirate and conventional analogue controller design applied to a hydrogenerator system. Mathematical and Computational Applications, 4(3): 209–215, December 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/209.

Berkemeier:2021:DFM

[BP21] Manuel Berkemeier and Sebastian Peitz. Derivativefree multiobjective trust region descent method using radial basis function surrogate models. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/31.

Barua:2023:CMM

[BP23] Himel Barua and Alex Povitsky. Continuum and molecular modeling of chemical vapor deposition at nano-scale fibrous substrates. Mathematical and Computational Applications, 28 (6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/112.

Benitez:2023:RML

[BPC^{+}23] Luis Víctor Maidana Benítez, Melisa María Rosa Villamayor Paredes, José Colbes, César F. Bogado-Martínez, Benjamin Barán, and Diego P. Pinto-Roa. Routing, modulation level, and spectrum assignment in elastic optical networks — a serial stage approach with multiple sub-sets of requests based on integer linear programming. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/67.

Bacanli:2008:TSG

[BPD08] Sevil Bacanlı, Özgür K. Peker, and Yaprak Parlak Demirhan. Twostage group sequential designs for comparative clinical trials. Mathematical and Computational Applications, 13 (1):31–40, April 2008. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/1/31.

Brovko:2019:TNP

[Bro19] George L. Brovko. Tensors in Newtonian physics and the foundations of classical continuum mechanics. Mathematical and Computational Applications, 24(3):??, September 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/3/79.

Bozkurt:2002:NGM

[BS02] Durmuş Bozkurt and Süleyman Solak. On the norms of GCD matrices. Mathematical and Computational Applications, 7(3):205–210, December 2002. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/3/205.

Benli:2011:ETT

[BS11] Semih Benli and Onur Sayman. The effects of temperature and thermal stresses on impact damage in laminated composites. Mathematical and Computational Applications, 16(2):392– 403, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/392.

Basaran:2013:NSH

[BS13a] Anıl Başaran and Fatih Selimefendigil. Numerical study of heat transfer due to twinjets impingement onto an isothermal moving plate. Mathematical and Computational Applications, 18(3):340– 350, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/340.

Bayram:2013:NSA

[BŞ13b] Hüsamettin Bayram and Ramazan Şahin. A new simulated annealing approach for travelling salesman problem. Mathematical and Computational Applications, 18(3):313–322, December 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/313.

Borhanifar:2015:SJC

[BS15] Abdollah Borhanifar and Khadijeh Sadri. Shifted Jacobi collocation method based on operational matrix for solving the systems of Fredholm and Volterra integral equations. Mathematical and Computational Applications, 20(2):76–93, August 2015. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/2/76.

Biswas:2023:SLM

[BS23] Raktim Biswas and Deepak Sharma. Singleloop multi-objective reliabilitybased design optimization using chaos control theory and shifting vector with differential evolution. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/26.

Benchiha:2023:NSF

[BSA^{+}23]  SidAhmed  Benchiha,

Laxmi Prasad Sapkota, Aned Al Mutairi, Vijay Kumar, Rana H. Khashab, Ahmed M.

Gemeay, Mohammed Elgarhy, and Said G. Nassr. A new sine family of generalized distributions: Statistical inference with applications. Mathematical and Computational Applications, 28 (4):??, August 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/83.

Babacan:2005:IGT

[BSK05] Tahsin Babacan, Djavad I.

Salamov, and Atalay Küçükbursa. An investigation of the Gamow– Teller 1+ states in 90Nb isotopes. Mathematical and Computational Applications, 10(3):359– 368, December 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/359.

Benli:2012:NEP

[BSSK12] Semih Benli, Onur Sayman, Metin Sayer, and Mustafa Karamolla. Numerical elastic plastic stress analysis in a woven steel reinforced composite thermoplastic cantilever beam. Mathematical and Computational Applications, 17(2):100– 110, August 2012. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/2/100.

Balan:2003:MSL

[BT03] George C. Balan and Calm Tarau. The monitoring simulation of a lathe. Mathematical and Computational Applications, 8(3):279–286, December 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/279.

Bacciocchi:2019:NFA

[BT19] Michele Bacciocchi and Angelo Marcello Tarantino. Natural frequency analysis of functionally graded orthotropic cross-ply plates based on the finite element method. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/52.

Boari:2019:OLS

[BUAM19] Santiago Boari, Gonzalo Uribarri, Ana Amador, and Gabriel B. Mindlin. Observable for a large system of globally coupled excitable units. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/37.

Baylan:1996:EMM

[BUKM96] Mehmet Baylan, İhsan Uluer, Atalay Küçükbursa, and İsmail Maraş. E2/M1 multipole mixing ratios of 162Dy nucleus in rotation vibration model. Mathematical and Computational Applications, 1(1):154–158, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/154.

Bulbul:1999:LSP

[Bül99] M. M. Bülbül. Line shape and profile of the LO phonon modes in mixed alloy system In_{1−x−y}Ga_{x}Al_{y}As epitaxial layers grown on InP. Mathematical and Computational Applications, 4(1):83– 89, April 1999. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/83.

Buyukkoroglu:2016:FOC

[Büy16]  Taner  Büyükköroğlu.

Fixed order controller for Schur stability. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/25.

Bellora:2019:CPQ

[BV19] Davide Bellora and Riccardo Vescovini. A continuation procedure for the quasi-static analysis of materially and geometrically nonlinear structural problems. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/94.

Bogoya:2018:AHD

[BVCS18] Johan M. Bogoya, Andrés Vargas, Oliver Cuate, and Oliver Schütze. A (p, q)-averaged Hausdorff distance for arbitrary measurable sets. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/51.

Blignaut:2023:IPD

[BVvdH^{+}23] Julianne Blignaut, Martin Venter, David van den Heever, Mark Solms, and Ivan Crockart. Inducing perceptual dominance with binocular rivalry in a virtual reality headmounted display. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/77.

Bouwer:2023:STG

[BWK23] Johann M. Bouwer, Daniel N. Wilke, and Schalk Kok. Spatiotemporal gradient enhanced surrogate modeling strategies. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/57.

Bayrakceken:2006:CPD

[BY06] H. Bayrakçeken and M. Yürüsoy. Comparison of pressure distribution in inclined and parabolic slider bearings. Mathematical and Computational Applications, 11(1):65–73, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/65.

Baykan:2011:MCS

[BY11] Nurdan Akhan Baykan and Nihat Yılmaz. A mineral classification system with multiple artificial neural network using k-fold cross validation.

Mathematical and Computational Applications, 16(1):22–30, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/22.

Babaarslan:2013:HBC

[BY13] Murat Babaarslan and Yusuf Yayli. On helices and Bertrand curves in Euclidean 3-space. Mathematical and Computational Applications, 18 (1):1–11, April 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/1/1.

Bahsi:2016:NAN

[BY16] Ayşe Kurt Bahşı and Salih Yalçınbaş. A new algorithm for the numerical solution of telegraph equations by using Fibonacci polynomials. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/15.

Beni:2017:SAI

[BY17] Mehrdad Shahmohammadi Beni and Kwan Ngok Yu. Safeguarding against inactivation temperatures during plasma treatment of skin: Multiphysics model and phase field method. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/24.

Cakan:2004:KCC

[ÇA04] Celal Çakan and Cafer Aydin. On the Kr-core of complex sequences and the absolute equivalence of summability matrices. Mathematical and Computational Applications, 9(3):409–416, December 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/409.

Cunkas:2006:DOI

[ÇA06] Mehmet Çunkaş and Ramazan Akkaya. Design optimization of induction motor by genetic algorithm and comparison with existing motor. Mathematical and Computational Applications, 11(3):193–203, December 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/193.

Cunkas:2010:RFL

[ÇA10] Mehmet Çunkas and Omer Aydoğdu. Realization of fuzzy logic controlled brushless DC motor drives using Matlab/ Simulink. Mathematical and Computational Applications, 15(2):218– 229, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/218.

Casenave:2019:EIB

[CA19] Fabien Casenave and Nissrine Akkari. An error indicator-based adaptive reduced order model for nonlinear structural mechanics — application to high-pressure turbine blades. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/41.

Castellanos-Alvarez:2021:MIP

[CACRF^{+}21] Alejandro Castellanos-Alvarez, Laura Cruz-Reyes, Eduardo Fernandez, Nelson Rangel-Valdez, Claudia Gómez-Santillán, Hector Fraire, and José Alfredo Brambila-Hernández. A method for integration of preferences to a multi-objective evolutionary algorithm using ordinal multi-criteria classification. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/27.

Cai:2006:NVO

[Cai06] Jianping Cai. Numerical verification of the order of the asymptotic solutions of a nonlinear differential equation. Mathematical and Computational Applications, 11 (1):85–90, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/85.

Cai:2007:GKM

[Cai07] Jianping Cai. A generalized KBM method for strongly nonlinear oscillators with slowly varying parameters. Mathematical and Computational Applications, 12 (1):21–30, April 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/21.

Caliskan:2006:DMC

[Çal06] Ali Çalıskan. Dual minimal curves. Mathematical and Computational Applications, 11 (1):19–23, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/19.

Can:1996:LSD

[Can96] Mehmet Can. Lie symmetries of differential equations by computer algebra. Mathematical and Computational Applications, 1(1):15– 20, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/15.

Can:2002:KEU

[Can02] M. Can. Kinetic equations for ultrasonic scission of polymers. Mathematical and Computational Applications, 7(3): 241–252, December 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/3/241.

Can:2014:AAU

[Can14]  Mehmet Can.  Au-

thorship attribution using principal component analysis and competitive neural networks. Mathematical and Computational Applications, 19 (1):21–36, April 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/21.

Carmona-Arroyo:2022:VDL

[CAQCMM22] Guadalupe Carmona-Arroyo, Marcela Quiroz-Castellanos, and Efrén Mezura-Montes. Variable decomposition for large-scale constrained optimization problems using a grouping genetic algorithm. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/23.

Carvalho:2021:MRC

[Car21] Camille Carvalho. Modified representations for the close evaluation problem. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/69.

Catal:2010:DAM

[Çat10] Hikmet Hüseyin Çatal.

Discussion of “An Approximate Method for Free Vibration Analysis of Multibay Coupled Shear Walls”, K. B. Bozdoğan and D. Öztürk, Mathematical and Computational Applications 12, 41–50, 2007”. Mathematical and Computational Applications, 15(3):503–505, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/503.  See

[BÖ07].

Castaneda-Avina:2020:SOO

[CATCdlF20] Perla Rubi Castañeda-Aviña, Esteban Tlelo-Cuautle, and Luis Gerardo de la Fraga. Singleobjective optimization of a CMOS VCO considering PVT and Monte Carlo simulations. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/76.

Can:2004:SMC

[CB04] N. Can and H. G. Belik. A simple method for calculating the number of atoms in ion implanted samples from the RBS yield. Mathematical and Computational Applications, 9 (1):91–99, April 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/1/91.

Carreno:2019:BPM

[CBM^{+}19] Amanda Carreño, Luca Bergamaschi, Angeles Martinez, Antoni Vidal-Ferrándiz, Damian Ginestar, and Gumersindo Verdú. Block preconditioning matrices for the Newton method to compute the dominant λmodes associated with the neutron diffusion equation. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/9.

Cardenas-Barron:2013:EIA

[CBSTG13] Leopoldo Eduardo Cárdenas-Barrón, Biswajit Sarkar, and Gerardo Treviño-Garza. Easy and improved algorithms to joint determination of the replenishment lot size and number of shipments for an EPQ model with rework. Mathematical and Computational Applications, 18(2):132– 138, August 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/2/132.

Claus:2021:NRM

[CBT21] Tamme Claus, Jonas Bünger, and Manuel Torrilhon. A novel reconstruction method to increase spatial resolution in electron probe microanalysis. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/3/51.

Ciftci:2003:MSE

[ÇÇ03] Yasemin Öztekin Çiftci and Kemal ÇolakoğIu. Mechanical stability and elastic properties of some FCC metals: a manybody potential applications. Mathematical and Computational Applications, 8(3):295–302, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/295.

Cai:2009:ASC

[CC09] Ping Cai and Jianping Cai. Asymptotic solutions and comparisons of a generalized Van del Pol oscillator with slowly varying parameter. Mathematical and Computational Applications, 14(2):159–168, August 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/2/159.

Chen:2011:RLS

[CC11] Kuang-Ku Chen and Singa Wang Chiu. Replenishment lot size and number of shipments for EPQ model derived without derivatives. Mathematical and Computational Applications, 16(3):753–760, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/753.

Cillie:2022:ENI

[CC22] Johan Cillie and Corné Coetzee. Experimental and numerical investigation of the in-plane compression of corrugated paperboard panels. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/108.

Chen:2024:AEF

[CC24] Yang-Yih Chen and Hsien-Kuo Chang. Accuracy examination of the Fourier series approximation for almost limiting gravity waves on deep water. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/5.

Cayiroglu:2007:NME

[ÇÇÇ07] İbrahim Çayıroglu, Abdullah Çavusoglu, and Veli Çelik. A new method for extracting 3D solid models of objects using 2D technical drawings. Mathematical and Computational Applications, 12(1):31–39, April 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/31.

Chiu:2008:MME

[CCC08] Singa Wang Chiu, Kuang-Ku Chen, and Huei-Hsin Chang. Mathematical method for expediting scrap-or-rework decision making in EPQ model with failure in repair. Mathematical and Computational Applications, 13(3):137–145, December 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/3/137.

Carpitella:2018:MHF

[CCC^{+}18] Silvia Carpitella, Fortunato Carpitella, Antonella Certa, Julio Benítez, and Joaquín Izquierdo. Managing human factors to reduce organisational risk in industry. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/67.

Chen-Charpentier:2021:CMM

[CCGHdCJ21] Benito Chen-Charpentier, Clara Eugenia Garza-Hume, and María del Carmen Jorge. A continuous model of marital relations with stochastic differential equations. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/3.

Chiu:2006:SEM

[CCL06] Yuan-Shyi Peter Chiu, Singa Wang Chiu, and Hong-Dar Lin. Solving an EPQ model with rework and service level constraint. Mathematical and Computational Applications, 11(1):75– 84, April 2006. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/75.

Chiu:2010:AAD

[CCWY10] Singa Wang Chiu, Chi-Bin Cheng, Mei-Fang Wu, and Jyh-Chau Yang. An algebraic approach for determining the optimal lot size for EPQ model with rework process. Mathematical and Computational Applications, 15(3):364–370, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/364.

Cai:2008:NVC

[CCY08] Jianping Cai, Shuhui Chen, and Cuihong Yang. Numerical verification and comparison of error of asymptotic expansion solution of the Duffing equation. Mathematical and Computational Applications, 13(1):23–29, April 2008. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/1/23.

Carman:2003:MTP

[ÇD03] Kazim Çarman and Fariz Dünyamalioğlu. Modelling tractive performance for drive wheel. Mathematical and Computational Applications, 8(1):49–54, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/49.

Celikel:2004:NED

[CD04] Ebru Celikel and Mebmet E. Dalkilic. A new encoding decoding scheme for text compression with embedded security. Mathematical and Computational Applications, 9(3):475–484, December 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/475.

Civalek:2010:FVB

[CDA10] Ömer Civalek, Çiğdem Demir, and Bekir Akgöz. Free vibration and bending analyses of cantilever microtubules based on nonlocal continuum model. Mathematical and Computational Applications, 15(2):289–298, August 2010. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/15/2/289.

Chiodo:2023:BIS

[CDDM23] Elio Chiodo, Fabio De Angelis, Bassel Diban, and Giovanni Mazzanti. Bayes inference of structural safety under extreme wind loads based upon a peak-over-threshold process of exceedances. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/111.

Callioglu:2013:VAF

[ÇDYS13] Hasan Çallıoğlu, Ersin Demir, Yasin Yılmaz, and Metin Sayer. Vibration analysis of functionally graded sandwich beam with variable cross-section. Mathematical and Computational Applications, 18(3):351– 360, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/351.

Cetin:2016:APS

[Cet16]  Elif Cetin.  Analytic

properties of the sum B_{1}(h, k). Mathematical and Computational Applications, 21(3):??, September 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/31.

Cevik:2010:ATM

[Çev10] Mehmet Çevik. Application of Taylor matrix method to the solution of longitudinal vibration of rods. Mathematical and Computational Applications, 15(3):334– 343, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/334.

Civalek:2009:DSC

[CG09] Ömer Civalek and Murat Gürses. Discrete singular convolution for free vibration analysis annular membranes. Mathematical and Computational Applications, 14(2):131– 138, August 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/2/131.

Chiu:2011:JDP

[CGCC11] Singa Wang Chiu, Dah-Chuan Gong, Chun-Lin Chiu, and Chia-Ling Chung. Joint determination of the production lot size and number of shipments for EPQ model with rework. Mathematical and Computational Applications, 16(2):317– 328, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/317.

Celebi:1996:SUI

[ÇGK96] H. Çelebi, M. Gündüzalp, and M. Karaman. Segmentation of ultrasonic images using fuzzy sets. Mathematical and Computational Applications, 1(2):29–34, December 1996. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/1/2/29.

Coello:2023:IKD

[CGM^{+}23] Carlos Coello, Erik Goodman, Kaisa Miettinen, Dhish Saxena, Oliver Schütze, and Lothar Thiele. Interview: Kalyanmoy Deb talks about formation, development and challenges of the EMO community, important positions in his career, and issues faced getting his works published. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/34.

Camarena:2023:OVB

[CGMCCA23]  Fernando  Camarena,

Miguel Gonzalez-Mendoza, Leonardo Chang, and Ricardo Cuevas-Ascencio. An overview of the vision-based human action recognition field. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/61.

Cavusoglu:2002:TBG

[ÇGV02]  Abdullah  Çavuşoğlu,

H. Haldun Göktaş, and Necla Vardal. A template based graph reduction system based on combinators. Mathematical and Computational Applications, 7(3): 253–262, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/253.

Chandramouli:1997:ABS

[Cha97] A. B. Chandramouli.

Analysis of a bulk-service queueing-system subject to interruptions. Mathematical and Computational Applications, 2 (2):53–59, August 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/2/53.

Chandramouli:2005:HAJ

[Cha05] A. B. Chandramouli.

Heuristic approach for n-job, 3-machine flow shop scheduling problem involving transportation time, break down time and weights of jobs. Mathematical and Computational Applications, 10(2):301–305, August 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/301.

Chandramouli:2008:NTB

[Cha08] A. B. Chandramouli.

Network of tandem and Bi-tandem queueing process with reneging and jockeying. Mathematical and Computational Applications, 13(3):175– 182, December 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/3/175.

Chen:2018:MSE

[Che18] Ray-Ming Chen. Metrics for single-edged graphs over a fixed set of vertices. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/66.

Cicek:2007:SBG

[Çiç07]  Adem Çiçek.  Step

based geometric and topological similarity assessment of mechanical parts. Mathematical and Computational Applications, 12(3):141– 150, December 2007. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/3/141.

Civalek:1996:LLH

[Civ96]  Şevket Civalek. The lifts

of langrange and Hamilton equations to the extended vecor bundles. Mathematical and Computational Applications, 1(1):21–28, June 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/21.

Cui:2024:FDM

[CJD24] Manruo Cui, Cui-Cui Ji, and Weizhong Dai. A finite difference method for solving the wave equation with fractional damping. Mathematical and Computational Applications, 29(1):??, February 2024. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/2.

Custers:2018:MEM

[CJL18] C. H. H. M. Custers, J. W. Jansen, and E. A. Lomonova. Memory efficient method for electromagnetic multi-region models using scattering matrices. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/71.

Civalek:2007:DSC

[CK07] Ömer Civalek and Okyay Kiracioglu. Discrete singular convolution for free vibration analysis of anisotropic rectangular plates. Mathematical and Computational Applications, 12(3):151– 160, December 2007. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/3/151.

Chu:2012:CAP

[CK12] Yunn-Kuang Chu and Jau-Chuan Ke. Computation approaches for parameter estimation of Weibull distribution. Mathematical and Computational Applications, 17(1):39–47, April 2012. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/1/39.

Chalishajar:2018:TCS

[CK18] Dimplekumar N. Chalishajar and Avadhesh Kumar. Total controllability of the second order semi-linear differential equation with infinite delay and noninstantaneous impulses. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/3/32.

Callioglu:2004:EPS

[ÇKAK04] Hasan Çallioğlu, N. Sman Köksal, Necati Ataberk, and İdris Kaynak. Elasticplastic stress analysis of thermoplastic matrix laminated composite beams. Mathematical and Computational Applications, 9(1):49– 64, April 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/49.

Chiu:2016:CMM

[CKCC16] Singa Wang Chiu, Jyun-Sian Kuo, Victoria Chiu, and Yuan-Shyi Peter Chiu. Cost minimization for a multi-product fabrication-distribution problem with commonality, postponement, and quality assurance. Mathematical and Computational Applications, 21 (3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/38.

Cenesiz:2010:PEP

[ÇKÇK10] Y. Çenesiz, A. B. Koç, B. Çitil, and A. Kurnaz. Padé embedded piecewise differential transform method for the solution of ODE’s. Mathematical and Computational Applications, 15(2):199– 207, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/199.

Chesneau:2022:MWE

[CKKA22] Christophe Chesneau, Vijay Kumar, Mukti Khetan, and Mohd Arshad. On a modified weighted exponential distribution with applications. Mathematical and Computational Applications, 27(1):??, February 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/17.

Chung:2021:LAC

[CKLZ21] Eric Chung, Hyea-Hyun Kim, Ming-Fai Lam, and Lina Zhao. Learning adaptive coarse spaces of BDDC algorithms for stochastic elliptic problems with oscillatory and high contrast coefficients. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/44.

Chiu:2016:SFL

[CLC16] Yuan-Shyi Peter Chiu, Gang-Ming Liang, and Singa Wang Chiu. Solving a fabrication lotsize and shipping frequency problem with an outsourcing policy and random scrap. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/45.

Contreras-Lujan:2022:EML

[CLGGLB^{+}22] Eduardo Enrique Contreras-Luján, Enrique Efrén García-Guerrero, Oscar Roberto López-Bonilla, Esteban Tlelo-Cuautle, Didier López-Mancilla, and Everardo Inzunza-González. Evaluation of machine learning algorithms for early diagnosis of deep venous thrombosis. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/24.

Cai:2010:SCU

[CLY10] Jianping Cai, Meili Lin, and Zhengzhong Yuan. Secure communication using practical synchronization between two different chaotic systems with uncertainties. Mathematical and Computational Applications, 15(2):166–175, August 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/166.

Carreno:2019:AIU

[CMB19] Rafael M. Carreño, Javier Martínez, and José Benito Bouza. An approach for integrating uncertainty when selecting an anti-torpedo decoy in brand new warships. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/5.

Clemence-Mkhope:2021:LVC

[CMCM21] Dominic P. Clemence-Mkhope and Belinda G. B. Clemence-Mkhope. The limited validity of the conformable Euler finite difference method and an alternate definition of the conformable fractional derivative to justify modification of the method. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/66.

Clemence-Mkhope:2022:THE

[CMG22] Dominic P. Clemence-Mkhope and Gregory A. Gibson. Taming hyperchaos with exact spectral derivative discretization finite difference discretization of a conformable fractional derivative financial system with market confidence and ethics risk. Mathematical and Computational Applications, 27(1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/4.

Chen:2010:VIM

[CML10] Y. M. Chen, G. Meng, and J. K. Liu. Variational iteration method for conservative oscillators with complicated nonlinearities. Mathematical and Computational Applications, 15(5):802–809, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/802.

Cabella:2019:ISR

[CMM19] Brenno Cabella, Fernando Meloni, and Alexandre S. Martinez. Inadequate sampling rates can undermine the reliability of ecological interaction estimation. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/48.

Cai:2020:EFM

[CMS20] Xiatong Cai, Abdolmajid Mohammadian, and Hamidreza Shirkhani. An efficient framework for multi-objective riskinformed decision support systems for drainage rehabilitation. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/73.

Ceylan:2016:FVM

[CO16] Tuba Ceylan and Baver Okutmustur. Finite volume method for the relativistic Burgers model on a (1+1)-dimensional de Sitter spacetime. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/16.

Candemir:2011:TET

[CÖGD11] Mehmet Candemir, Mümin Özcan, Mustafa Güneş, and Ertuğrul Deliktaş. Technical efficiency and total factor productivity growth in the hazelnut agricultural sales cooperatives unions in Turkey. Mathematical and Computational Applications, 16(1):66–76, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/66.

Conceicao:2022:DIT

[Con22] Ana C. Conceição. Dynamic and interactive tools to support teaching and learning. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/1.

Corcoran:2019:TBO

[Cor19] Padraig Corcoran. Topology based object tracking. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/84.

Conceicao:2022:SCA

[CP22] Ana C. Conceição and Jéssica C. Pires. Symbolic computation applied to Cauchy type singular integrals. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/3.

Calcagni:2020:SAC

[CPA20] Antonio Calcagnì, Massimiliano Pastore, and Gianmarco Altoé. ssMousetrack — analysing computerized tracking data via Bayesian state-space models in R. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/41.

Campi:2019:LEA

[CPP19] Cristina Campi, Annalisa Pascarella, and Francesca Pitolli. Less is enough: Assessment of the random sampling method for the analysis of magnetoencephalography (MEG) data. Mathematical and Computational Applications, 24 (4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/98.

Cullenbine:2021:FTE

[CRA^{+}21] Christopher A. Cullenbine, Joseph W. Rohrer, Erin A. Almand, J. Jordan Steel, Matthew T. Davis, Christopher M. Carson, Steven C. M. Hasstedt, John C. Sitko, and Douglas P. Wickert. Fizzle testing: an equation utilizing random surveillance to help reduce COVID-19 risks. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/16.

Cetiner:2010:TRA

[ÇS10] B. Gültekin Çetiner and Murat Sarı. Tibial rotation assessment using artificial neural networks. Mathematical and Computational Applications, 15(1):34–44, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/34.

Cuate:2019:VRM

[CS19] Oliver Cuate and Oliver Schütze. Variation rate to maintain diversity in decision space within multi-objective evolutionary algorithms. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/82.

Castellanos:2022:BAH

[CS22] Carlos Ignacio Hernández Castellanos and Oliver Schütze. A bounded archiver for Hausdorff approximations of the Pareto front for multiobjective evolutionary algorithms. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/48.

Cetiner:2010:NNB

[ÇSB10] B. Gültekin Çetiner, Murat Sari, and Oğuz Borat. A neural network based traffic-flow prediction model. Mathematical and Computational Applications, 15(2):269– 278, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/269.

Cavusoglu:2013:CQT

[ÇŞÖG13]  Abdullah  Çavuşoğlu,

Baha Şen, Caner Özcan, and Salih Görgünoğlu. Cropped quad-tree based solid object colouring with Cuda. Mathematical and Computational Applications, 18(3):301– 312, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/301.

Chavez-Saab:2024:PSR

[CSOPM24] Jorge Chávez-Saab, Odalis Ortega, and Amalia Pizarro-Madariaga. On the parallelization of square-root Vélu’s formulas. Mathematical and Computational Applications, 29(1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/14.

Callioglu:2001:AEP

[ÇŞS01] Hasan Çallıoğlu, Mehmet Şenel, and Mehmet Savran. An analytical elastic-plastic stress analysis for a steel fibre reinforced thermoplastic composite beam subjected to thermal loading. Mathematical and Computational Applications, 6(2):123–136, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/123.

Castellanos:2020:NED

[CSSOB20] Carlos Ignacio Hernández Castellanos, Oliver Schütze, Jian-Qiao Sun, and Sina Ober-Blöbaum. Nonepsilon dominated evolutionary algorithm for the set of approximate solutions. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/3.

Campos:2020:NOC

[CST20] Carlos Campos, Cristiana J. Silva, and Delfim F. M. Torres. Numerical optimal control of HIV transmission in Octave/MATLAB. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/1.

Caliskan:2001:AHT

[CT01] A. Caliskan and B. Tantay. Application of Holditch theorem to Bézier and quadratic uniform B-spline curves. Mathematical and Computational Applications, 6(3):169–176, December 2001. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/6/3/169.

Cipiloglu:2009:FMO

[ÇT09] Mustafa Ali Ali Çipiloğlu and Sadi Turgut. Figure of merit of onedimensional resonant transmission systems in the quantum regime. Mathematical and Computational Applications, 14(1):13–19, April 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/13.

Carman:2012:PTT

[ÇT12] Kazım Çarman and Alper Taner. Prediction of tire tractive performance by using artificial neural networks. Mathematical and Computational Applications, 17(3):182–192, December 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/182.

Cecen:2014:SAS

[ÇT14a] Yiğit Çeçen and Mehmet Tombakoğlu. Simplified, analytically solvable model problem for non-analog Monte Carlo methods. Mathematical and Computational Applications, 19(2):134– 143, August 2014. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/134.

Cheng:2014:RAR

[CT14b] Jiang Cheng and Yinghui Tang. Reliability analysis of M/G/1 repairable queueing system with multiple adaptive vacations and p-entering discipline. Mathematical and Computational Applications, 19(2):105– 114, August 2014. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/105.

Culha:2011:CMP

[CTAA11] Osman Culha, Mustafa Toparli, Tevfik Aksoy, and Mustafa Akdag. Calculation of the mechanical properties of FeB Layer’s by finite element method. Mathematical and Computational Applications, 16 (1):113–124, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/113.

Cerrada:2022:AFS

[CTH^{+}22] Mariela Cerrada, Leonardo Trujillo, Daniel E. Hernández, Horacio A. Correa Zevallos, Jean Carlo Macancela, Diego Cabrera, and René Vinicio Sánchez. AutoML for feature selection and model tuning applied to fault severity diagnosis in spur gearboxes. Mathematical and Computational Applications, 27(1):??, February 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/6.

Chesneau:2022:OEL

[CTJJ22] Christophe Chesneau, Lishamol Tomy, Meenu Jose, and Kuttappan Vallikkattil Jayamol. Odd exponential-logarithmic family of distributions: Features and modeling. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/68.

Cai:2014:ACH

[CTL14] Ping Cai, Jia-Shi Tang, and Zhen-Bo Li. Analysis and controlling of Hopf bifurcation for chaotic van der Pol–Duffing system. Mathematical and Computational Applications, 19(3):184–193, December 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/3/184.

Caliskan:1996:IPM

[ÇTY96] Ali Çalışkan, Bahadlr Tantay, and Filiz Yağcı. Improved part modeling in a process planning system. Mathematical and Computational Applications, 1(2): 21–28, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/21.

Civalek:2004:FVA

[CÜ04] Ömer Civalek and Mehmet Ülker. Free vibration analysis of elastic beams using harmonic differential quadraure (HDQ). Mathematical and Computational Applications, 9(2):257– 264, August 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/257.

Can:2004:PKD

[CUA04] S. Can, B. Uyar, and İ. Aydemir. On the properties of (k+1)- dimensional time-like ruled surfaces with the space like generating space in the Minkowski space IR1n. Mathematical and Computational Applications, 9(3): 393–398, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/393.

Cui:2007:RAG

[Cui07] Yaodong Cui. Recursive algorithm for generating two-staged cutting patterns of punched strips. Mathematical and Computational Applications, 12(2):107–115, August 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/2/107.

Cunkas:2008:DOE

[Çun08] Mehmet Çunkaş. Design optimization of electric motors by multiobjective fuzzy genetic algorithms. Mathematical and Computational Applications, 13(3):153– 163, December 2008. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/3/153.

Chen:2016:NNM

[CW16] Ming Chen and Zhong Wan. New nonlinear metrics model for information of individual research output and its applications. Mathematical and Computational Applications, 21 (3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/26.

Chen:2022:LGE

[CWF^{+}22] Gang Chen, Lihua Wei, Jiangyue Fu, Chengjiang Li, and Gang Zhao. A large group emergency decision-making method based on uncertain linguistic cloud similarity method. Mathematical and Computational Applications, 27(6):??, December 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/101.

Caliskan:2005:ETS

[CY05] A. Caliskan and Y. Yuzcelik. Examining of tetragonal surface patches with their geometric properties and relationships. Mathematical and Computational Applications, 10(1):155– 163, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/155.

Canpolat:2003:PRM

[CYAK03] Fethullah Canpolat, Kemalettin Yilmaz, Raşit Ata, and M. Metin Köse. Prediction of ratio of mineral substitution in the production of lowclinker factored cement by artificial neural network. Mathematical and Computational Applications, 8(2):209–216, August 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/209.

Cetin:2005:AOT

[ÇYAÖ05] N. S. Çetin, M. A. Yurdusev, R. Ata, and A. Özdamar. Assessment of optimum tip speed ratio of wind turbines. Mathematical and Computational Applications, 10(1):147– 154, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/147.

Chen:2017:BML

[CZC17] Man Chen, Zheng Zhang, and Cen Cui. On the bias of the maximum likelihood estimators of parameters of the Weibull distribution. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/19.

Duran:2019:SNE

[DA19] Ugur Duran and Mehmet Acikgoz. A study on novel extensions for the p-adic gamma and p-adic beta functions. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/53.

Domingues:2018:LVV

[DAD^{+}18] José Sérgio Domingues, Ravel Alvarenga Coelho, Francielly Dos Santos Bento, Brenda Cristina Almeida, and Ayanne Oliveira De Castro. Left ventricular volume in bovines: The correlation between Teichholz’s medical mathematical method and the volume of the truncated prolate spheroid. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/50.

Delgadillo-Aleman:2019:MMI

[DAKCPACC19] Sandra Delgadillo-Aleman, Roberto Ku-Carrillo, Brenda Perez-Amezcua, and Benito Chen-Charpentier. A mathematical model for intimate partner violence. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/29.

Dal:2011:MTS

[Dal11] Fadime Dal. Multiple time scales solution of an equation with quadratic and cubic nonlinearities having fractional- order derivative. Mathematical and Computational Applications, 16 (1):301–308, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/301.

Danane:2022:SCL

[Dan22] Jaouad Danane. Stochastic capital–labor Lévy jump model with the precariat labor force. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/93.

Pereira:2022:NMM

[dAPdSL^{+}22] Gabriel Thomaz de Aquino Pereira, Ricardo J. Alves de Sousa, I-Shih Liu, Marcello Goulart Teixeira, and Fábio A. O. Fernandes. A new material model for agglomerated cork. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/92.

Darilmaz:2011:ASH

[Dar11] Kutlu Darılmaz. An assumed stress hybrid finite element for buckling analysis. Mathematical and Computational Applications, 16(3):690– 701, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/690.

Dzhafarov:2003:ASS

[DB03] V. Dzhafarov and T. Büyükköroğlu. An algorithm for segment stability. Mathematical and Computational Applications, 8 (1):111–118, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/111.

Demiray:2016:NES

[DB16] Seyma Tuluce Demiray and Hasan Bulut. New exact solutions of the system of equations for the ion sound and Langmuir waves by ETEM. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/11.

Dubas:2017:NSCa

[DB17a] Frédéric Dubas and Kamel Boughrara. New scientific contribution on the 2-d subdomain technique in Cartesian coordinates: Taking into account of iron parts. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/17.

Dubas:2017:NSCb

[DB17b] Frédéric Dubas and Kamel Boughrara. New scientific contribution on the 2-d subdomain technique in polar coordinates: Taking into account of iron parts. Mathematical and Computational Applications, 22(4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/4/42.

Dubas:2020:MMD

[DB20] Frédéric Dubas and Kamel Boughrara. Mathematical models for the design of electrical machines. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/77.

Dombayci:2006:EHM

[DÇ06] Ömer Altan Dombaycı and Önder Çivril. Estimation of hourly mean ambient temperatures with artificial neural networks. Mathematical and Computational Applications, 11(3):215–224, December 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/3/215.

Demir:2010:FVA

[DCA10]  Çiğdem Demir, Ömer

Civalek, and Bekir Akgöz. Free vibration analysis of carbon nanotubes based on shear deformable beam theory by discrete singular convolution technique. Mathematical and Computational Applications, 15 (1):57–65, April 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/1/57.

Daniel:2021:DAF

[DCAR21] Thomas Daniel, Fabien Casenave, Nissrine Akkari, and David Ryckelynck. Data augmentation and feature selection for automatic model recommendation in computational physics. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/17.

Dai:2023:SHM

[DCC23] Li Dai, Mi-Da Cui, and Xiao-Xiang Cheng. Structural-health-monitoring-oriented finite element model for a specially shaped steel arch bridge and its application. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/33.

Dere:2011:NFE

[DD11] Yunus Dere and Fatma Tuba Dede. Nonlinear finite element analysis of an R/C frame under lateral loading. Mathematical and Computational Applications, 16(4):947– 958, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/947.

Daniels:2021:SSL

[DDPK21] Michael W. Daniels, Daniel Dvorkin, Rani K. Powers, and Katerina Kechris. Semi-supervised learning using hierarchical mixture models: Gene essentiality case study. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/40.

Dereli:2004:FCS

[DE04] Yilmaz Dereli and Dursun Eser. Flow calculations in straight-through labyrinth seals by using Moody’s friction-factor model. Mathematical and Computational Applications, 9(3):435–442, December 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/435.

Deb:2023:GDS

[DE23] Kalyanmoy Deb and Matthias Ehrgott. On generalized dominance structures for multiobjective optimization. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/100.

Divya:2023:NCA

[DEL23] S. Divya, S. Eswaramoorthi, and Karuppusamy Loganathan. Numerical computation of Ag/ Al_{2}O_{3} nanofluid over a Riga plate with heat sink/source and non-Fourier heat flux model. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/20.

Deniz:2016:NSF

[Den16] Coşkun Deniz. A normal and standard form analysis of the JWKB asymptotic matching rule via the first order Bessel’s equation. Mathematical and Computational Applications, 21(4):??, December 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/41.

Dlamini:2023:NMC

[DHK23] Kwanda Mercury Dlamini, Vuyo Terrence Hashe, and Thokozani Justin Kunene. Numerical modeling of cavitation rates and noise acoustics of marine propellers. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/42.

Dascioglu:2013:BCM

[DI13] Aysegul Akyuz Dascioglu and Nese Isler. Bernstein collocation method for solving nonlinear differential equations. Mathematical and Computational Applications, 18(3):293–300, December 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/293.

Dikmen:2006:TDS

[Dik06] E. Dikmen. Time dependence of Shell model calculations. Mathematical and Computational Applications, 11 (1):41–49, April 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/1/41.

Dilaveroglu:2002:PDC

[Dil02]  Erdoğan  Dilaveroğlu.

On the phase dependence of the Cramér-Rao bounds for damped sinusoids. Mathematical and Computational Applications, 7(2):93–102, August 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/2/93.

Du:2010:NDR

[DJ10] Chengbin Du and Shouyan Jiang. Nonlinear dynamic responses of arch dam with shear keys. Mathematical and Computational Applications, 15(5):828–833, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/828.

Danook:2018:NCH

[DJH18] Suad Hassan Danook, Qusay Kamel Jasim, and Adnan Mohammed Hussein. Nanofluid convective heat transfer enhancement elliptical tube inside circular tube under turbulent flow. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/78.

Daneshian:2005:OAI

[DJL^{+}05] B. Daneshian, G. R.

Jahanshahloo, F. Hosseimzadeh Lotfi, T. Allahviranloo, and S. Mehrabian. The overall assurance interval for the non-Archimedean epsilon in DEA models. Mathematical and Computational Applications, 10(3):387– 393, December 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/387.

Dumont:2018:RUS

[DJM18] Serge Dumont, Franck Jourdan, and Tarik Madani. 4D remeshing using a space-time finite element method for elastodynamics problems. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/29.

Dechpichai:2022:MPD

[DJY^{+}22] Porntip Dechpichai, Nuttawadee Jinapang, Pariyakorn Yamphli, Sakulrat Polamnuay, Sittisak Injan, and Usa Humphries. Multivariable panel data cluster analysis of meteorological stations in Thailand for ENSO phenomenon. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/37.

Danaci:1999:DSN

[DK99] Mustafa Danacı and Adem Kalınlı. Detection of the starting of the newly applied mass for successive weighing. Mathematical and Computational Applications, 4(1):53–59, April 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/53.

Derebasi:2003:PMP

[DK03] Naim Derebasi and Ilker Kucuk. Prediction of magnetic properties of strip wound toroidal cores up to 2 Khz using artificial neural network. Mathematical and Computational Applications, 8(2):217–223, August 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/217.

Dzhafarov:1997:MOC

[DKÇ97] V. Dzhafarov, A. Karamancıoğlu, and S. Çetintaş. Minimax optimal control for one class of uncertain systems. Mathematical and Computational Applications, 2 (1):21–29, April 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/1/21.

DuToit:2023:EPI

[DL23] Jacques François Du Toit and Ryno Laubscher. Evaluation of physicsinformed neural network solution accuracy and efficiency for modeling aortic transvalvular blood flow. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/62.

Dutta:2024:MCT

[DL24] Pritha Dutta and Anita T.

Layton. Magnesium and calcium transport along the male rat kidney: Effect of diuretics. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/29/1/13.

delaFraga:2021:DEU

[dlF21] Luis Gerardo de la Fraga. Differential evolution under fixed point arithmetic and FP16 numbers. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/13.

delaFraga:2019:OMP

[dlFH19] Luis Gerardo de la Fraga and Heriberto Cruz Hernández. Optimizing the maximal perturbation in point sets while preserving the order type. Mathematical and Computational Applications, 24(4):??, December 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/4/97.

Duan:2010:NAD

[DLL10] Hongxia Duan, Shouju Li, and Yingxi Liu. Nonlinear analysis of damage evolution for steel structures under earthquake. Mathematical and Computational Applications, 15(5):889–894, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/889.

Demirtas:1999:RPM

[DM99] Ömer Can Demirtaş and Rauf Mirzababayev. Reference points in magnetic measurements. Mathematical and Computational Applications, 4 (1):25–30, April 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/25.

DellAmico:2021:SRL

[DM21] Mauro Dell’Amico and Matteo Magnani. Solving a real-life Distributor’s pallet loading problem. Mathematical and Computational Applications, 26(3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/53.

Djanan:2022:AFL

[DMN22] Armand Anthelme Nanha Djanan, Steffen Marburg, and Blaise Roméo Nana Nbendjo. On appearance of fast or late self-synchronization between non-ideal sources mounted on a rectangular plate due to time delay. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/20.

Duarte:2023:TPT

[DNS23] Guilherme Duarte, Ana Neves, and António Ramos Silva. Temperature patterns in TSA for different frequencies and material properties: a FEM approach. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/8.

Dogan:2012:SSO

[Doğ12] Nurettin Doğan. Solution of the system of ordinary differential equations by combined Laplace transform– Adomian decomposition method. Mathematical and Computational Applications, 17(3):203– 211, December 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/203.

Dolapci:2004:CCM

[Dol04] İhsan Timuçin Dolapçi.

Chebyshev collocation method for solving linear differential equations. Mathematical and Computational Applications, 9(1):107–115, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/107.

Deger:2009:GOP

[DPC09] Gözde Değer, Mehmet Pakdemirli, and Feyza Candan. A geometrical optimization problem associated with fruits of poppy flower. Mathematical and Computational Applications, 14(3):169– 175, December 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/3/169.

Deger:2010:SWB

[DPC^{+}10] Gözde Değer, Mehmet Pakdemirli, Feyza Candan, Selda Akgün, and Hakan Boyacı. Strength of wheat and barley stems and design of new beam/columns. Mathematical and Computational Applications, 15 (1):1–13, April 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/1.

Dutta:2022:ROM

[DRCSF22] Sourav Dutta, Peter Rivera-Casillas, Brent Styles, and Matthew W. Farthing. Reduced order modeling using advection-aware autoencoders. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/34.

Deb:2021:SMA

[DRH21] Kalyanmoy Deb, Proteek Chandan Roy, and Rayan Hussein. Surrogate modeling approaches for multiobjective optimization: Methods, taxonomy, and results. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/5.

Demba:2021:PFA

[DRKW21] Musa Ahmed Demba, Higinio Ramos, Poom Kumam, and Wiboonsak Watthayu. A phasefitted and amplificationfitted explicit Runge– Kutta–Nyström pair for oscillating systems. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/59.

Dilettoso:2017:WPC

[DRS17] Emanuele Dilettoso, Santi Agatino Rizzo, and Nunzio Salerno. A weakly Pareto compliant quality indicator. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/25.

Dag:2004:CBS

[DS04] İdris Dağ and Bülient Saka. A cubic B-spline collocation method for the EW equation. Mathematical and Computational Applications, 9(3): 381–392, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/381.

DeFlorio:2021:TFC

[DSD^{+}21] Mario De Florio, Enrico Schiassi, Andrea D’Ambrosio, Daniele Mortari, and Roberto Furfaro. Theory of functional connections applied to linear ODEs subject to integral constraints and linear ordinary integro-differential equations. Mathematical and Computational Applications, 26(3):??, September 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/65.

Das:2019:DDI

[DSG19] Kalyan Das, M. N. Srinivas, and Nurul Huda Gazi. Diffusion dynamics and impact of noise on a discrete-time ratiodependent model: an analytical and numerical approach. Mathematical and Computational Applications, 24 (4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/103.

Demba:2016:EPE

[DSI16] Musa A. Demba, Norazak Senu, and Fudziah Ismail. A 5(4) embedded pair of explicit trigonometrically-fitted Runge–Kutta–Nyström methods for the numerical solution of oscillatory initial value problems. Mathematical and Computational Applications, 21(4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/46.

Dawar:2018:ITR

[DSI^{+}18] Abdullah Dawar, Zahir Shah, Muhammad Idrees, Waris Khan, Saeed Islam, and Taza Gul. Impact of thermal radiation and heat source/sink on Eyring– Powell fluid flow over an unsteady oscillatory porous stretching surface. Mathematical and Computational Applications, 23(2):??, June 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/20.

Das:2019:MAP

[DSMP19] Kalyan Das, M. N. Srinivas, V. Madhusudanan, and Sandra Pinelas. Mathematical analysis of a prey–predator system: an adaptive backstepping control and stochastic approach. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/22.

Davis:2022:EQN

[DSU22] Kyle Davis, Miriam Schulte, and Benjamin Uekermann. Enhancing quasi-Newton acceleration for fluid-structure interaction. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/40.

Dogan:2008:DHU

[DT08] Sengul Dogan and Ibrahim Turkoglu. Diagnosing hyperlipidemia using association rules. Mathematical and Computational Applications, 13(3):193– 202, December 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/3/193.

DuToit:2022:AIB

[Du 22] Charl Gabriël Du Toit. Area of the intersection between a sphere and a cylindrical plane. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/5/79.

Durmus:2004:DHP

[DUM04] Hülya (Kaçar) Durmuş, Bekir Sadik Unlü, and Cevdet Meriç. Determination of hardness of pre-aged AA 6063 aluminum alloy by means of artificial neural networks method. Mathematical and Computational Applications, 9(2): 249–256, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/249.

Debbouche:2023:CCN

[DVFA23] Amar Debbouche, Bhaskar Sundara Vadivoo, Vladimir E. Fedorov, and Valery Antonov. Controllability criteria for nonlinear impulsive fractional differential systems with distributed delays in controls. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/13.

Ding:2016:IIV

[DW16] Zhenhua Ding and Yingyu Wu. An improved interval-valued hesitant fuzzy multi-criteria group decision-making method and applications. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/22.

Davydov:2019:TDM

[DZA19] Sergey A. Davydov, Andrei V. Zemskov, and Elena R. Akhmetova. Thermoelastic diffusion multicomponent halfspace under the effect of surface and bulk unsteady perturbations. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/26.

Essiomle:2022:NGS

[EA22] Kokou Essiomle and Franck Adekambi. A note on Gerber–Shiu function with delayed claim reporting under constant force of interest. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/51.

Ebaid:2019:AAS

[EAEAA19] Abdelhalim Ebaid, Asmaa Al-Enazi, Bassam Z. Albalawi, and Mona D. Aljoufi. Accurate approximate solution of Ambartsumian delay differential equation via decomposition method. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/7.

ElKihal:2018:OCC

[EARE18] Fadwa El Kihal, Imane Abouelkheir, Mostafa Rachik, and Ilias Elmouki. Optimal control and computational method for the resolution of isoperimetric problem in a discrete-time SIRS system. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/52.

Erdin:2005:FHS

[EAT05] M. Emin Erdin, Halil Aykul, and M. Şafak Tunalıoğlu. Forming of high strength/low formability metal sheets at elevated temperatures. Mathematical and Computational Applications, 10(3):331–340, December 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/331.

Erdogan:2016:EFF

[EAT16] Utku Erdoğan, Kenan Akarbulut, and Neşet Özkan Tan. Exponentially fitted finite difference schemes for reaction-diffusion equations. Mathematical and Computational Applications, 21(3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/32.

Erarslan:2000:MPR

[EB00] Kaan Erarslan and Sunay Beyhan. Modeling of performance and retarder chart of off-highway truck by Newton divided diffrence method. Mathematical and Computational Applications, 5 (1):57–65, April 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/1/57.

El-Bary:2002:FVI

[EB02] A. A. El-Bary. Fredholm– Volterra integral equation in the contact problem. Mathematical and Computational Applications, 7(2):157–164, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/157.

El-Bary:2006:SMH

[EB06] A. A. El-Bary. Sobolev’s method for Hammerstein integral equations. Mathematical and Computational Applications, 11 (2):91–94, August 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/2/91.

Erdem:2011:CEP

[EBD11] R. Tuğrul Erdem, Muhiddin Bağcı, and Ali Demir. A comparative evaluation of performance based analysis procedures according to 2007 Turkish earthquake code and fema-440. Mathematical and Computational Applications, 16(3):605–616, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/605.

El-Bary:2006:TSP

[EBY06] A. A. El-Bary and H. M. Youssef. Thermal shock problem for one dimensional generalized thermoelastic layered composite material. Mathematical and Computational Applications, 11(2):103–110, August 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/2/103.

Erarslan:1999:DAO

[EÇ99] Kaan Erarslan and Neş’e Çelebi. A dynamic approach for optimization of open pit mine invesments with capital budget. Mathematical and Computational Applications, 4 (1):45–51, April 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/45.

Ekici:2011:DVR

[ED11] Cumali Ekici and Mustafa Dede. On the Darboux vector of ruled surfaces in pseudo-Galilean space. Mathematical and Computational Applications, 16(4):830–838, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/4/830.

ElYazidi:2022:ENS

[EE22] Youness El Yazidi and Abdellatif Ellabib. An efficient numerical scheme based on radial basis functions and a hybrid quasi-Newton method for a nonlinear shape optimization problem. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/67.

Esqueda-Elizondo:2022:AMA

[EEJRLB^{+}22] José Jaime Esqueda-Elizondo, Reyes Juárez-Ramírez, Oscar Roberto López-Bonilla, Enrique Efrén García-Guerrero, Gilberto Manuel Galindo-Aldana, Laura Jiménez-Beristáin, Alejandra Serrano-Trujillo, Esteban Tlelo-Cuautle, and Everardo Inzunza-González. Attention measurement of an autism spectrum disorder user using EEG signals: a case study. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/21.

Eissa:2002:RPE

[EESK^{+}02] M. Eissa, S. A. El-Serafi, M. M. Kamel, M. Z. Rageb, and Y. A. Amer. Response of a parametrically excited system with quadratic and cubic non-linearities. Mathematical and Computational Applications, 7(3): 263–274, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/263.

Ekici:2000:CDS

[EG00] Cumali Ekici and Ali Görgülü. On the curvatures of (k + 1)- dimensional semi-ruled surfaces in E_{v}^{n+1}. Mathematical and Computational Applications, 5(3): 139–148, December 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/3/139.

Erzin:2011:PSP

[EG11] Yusuf Erzin and Nazmi Güneş. The prediction of swell percent and swell pressure by using neural networks. Mathematical and Computational Applications, 16(2):425– 436, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/425.

Engin:1999:ASP

[EGB99]  Şeref Naci Engin, Kay-

han Gülez, and M. N. M. Badi. Advanced signal processing techniques for faulth diagnosticsa review. Mathematical and Computational Applications, 4(2):121– 136, August 1999. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/2/121.

Egbelowo:2018:NED

[Egb18] Oluwaseun Egbelowo.

Nonlinear elimination of drugs in one-compartment pharmacokinetic models: Nonstandard finite difference approach for various routes of administration. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/27.

Erbil:1997:SRA

[EGIK97] H. Erbil, M. Gerçeklioglu, M. Ilhan, and A. Kuliev. Sum rule approach to nuclear collective vibration. Mathematical and Computational Applications, 2(1):1–11, April 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/1/1.

Erdogan:2005:EDS

[EI05] M. Emin Erdoğan and C. Erdem Imrak. The effects of duct shape on the Nusselt number. Mathematical and Computational Applications, 10 (1):79–88, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/79.

Erdogan:2006:EDS

[EI06a] M. Emin Erdoğan and C. Erdem Imrak. The effects of duct shape on the Nusselt number. Mathematical and Computational Applications, 11(3):235–??, December 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/235.

Erdogan:2006:PSF

[Eİ06b] M. Emin Erdoğan and C. Erdem İmrak. On the properties of a special function defined by an integral. Mathematical and Computational Applications, 11 (1):11–18, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/11.

Edis:2011:FLP

[EKAÖ11] Rahime Sancar Edis, Bayram Kahraman, Özlem Uzun Araz, and M. Kemal Özfırat. A facility layout problem in a marble factory via simulation. Mathematical and Computational Applications, 16(1):97–104, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/97.

Erdogan:2004:ESW

[EL04a] M. Emin Erdogan and C. Erdem Lmrak. Effect of the side walls on the unsteady flow in a duct of rectangular cross-section. Mathematical and Computational Applications, 9(3): 443–448, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/443.

Erdogan:2004:UFG

[EL04b] M. Emin Erdogan and C. Erdem Lmrak. On the unsteady flows generated by impulsive motion of a boundary. Mathematical and Computational Applications, 9(2): 265–274, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/265.

Elibol:2016:TSG

[Eli16] Armagan Elibol. A twostep global alignment method for feature-based image mosaicing. Mathematical and Computational Applications, 21 (3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/30.

Ersoy:2021:SVM

[ELT21] Mehmet Ersoy, Omar Lakkis, and Philip Townsend. A Saint-Venant model for Overland flows with precipitation and recharge. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/1.

Erturk:2008:DTT

[EM08] Vedat Suat Ertürk and Shaher Momani. Differential transform technique for solving fifthorder boundary value problems. Mathematical and Computational Applications, 13(2):113– 121, August 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/2/113.

Eminov:2003:RBF

[Emi03] Mübariz Eminov. Rulebased fuzzy classification using query processing. Mathematical and Computational Applications, 8(2):253– 262, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/253.

Naqeeb:2005:DEH

[ENT05] Thanaa M. El-Naqeeb and Reda S. Tantawi. Direct excitation of H(2P) atoms due to collisions with protons and antiprotons. Mathematical and Computational Applications, 10(2):291– 299, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/291.

Escobar:2019:ITP

[ENVPF19] José María Escobar, Juan Núñez-Valdés, and Pedro Pérez-Fernández. The invariant two-parameter function of algebras ψ. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/89.

Eustache:2021:RMP

[EPDG21] Julien Eustache, Antony Plait, Frédéric Dubas, and Raynal Glises. Review of multi-physics modeling on the active magnetic regenerative refrigeration. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/47.

Estrada-Padilla:2021:MOM

[EPLGGS^{+}21] Alejandro Estrada-Padilla, Daniela Lopez-Garcia, Claudia Gómez-Santillán, Héctor Joaquín Fraire-Huacuja, Laura Cruz-Reyes, Nelson Rangel-Valdez, and María Lucila Morales-Rodríguez. Modeling and optimizing the multi-objective portfolio optimization problem with trapezoidal fuzzy parameters. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/36.

Ergin:2001:RBP

[Erg01] Abdullah Aziz Ergin. A relation between prime numbers and twin prime numbers. Mathematical and Computational Applications, 6(2):153– 154, August 2001. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/153.

Erpolat:2012:SCB

[Erp12] Semra Erpolat. A suggestion for constructing a Bayesian network model with simple correlation and an appropriate regression analysis: a real medical diagnosis application. Mathematical and Computational Applications, 17(3):212– 222, December 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/212.

Ersoy:2010:MFS

[Ers10] H. Volkan Ersoy. MHD flow of a second order/grade fluid due to noncoaxial rotation of a porous disk and the fluid at infinity. Mathematical and Computational Applications, 15(3):354– 363, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/354.

Erturk:2007:DTM

[Ert07] Vedat Suat Ertürk. Differential transformation method for solving differential equations of Lane– Emden type. Mathematical and Computational Applications, 12(3):135– 139, December 2007. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/3/135.

Erturk:2011:CES

[Ert11] Vedat Suat Ertürk. Computing eigenelements of Sturm–Liouville problems of fractional order via fractional differential transform method. Mathematical and Computational Applications, 16(3):712–720, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/712.

Eissa:2006:CBAb

[ES06a] M. Eissa and M. Sayed. A comparison between active and passive vibration control of non-linear simple pendulum. part i: Transversally tuned absorber and negative Gϕn feedback. Mathematical and Computational Applications, 11(2):137– 149, August 2006. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/2/137.

Eissa:2006:CBAa

[ES06b] M. Eissa and M. Sayed. A comparison between active and passive vibration control of nonlinear simple pendulum. part II: Longitudinal tuned absorber and negative Gϕ and Gϕn feedback. Mathematical and Computational Applications, 11(2):151– 162, August 2006. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/2/151.

Ersayin:2013:NII

[ES13] Erdem Ersayın and Fatih Selimefendigil. Numerical investigation of impinging jets with nanofluids on a moving plate. Mathematical and Computational Applications, 18(3):428–437, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/428.

Esendemir:2001:EPS

[Ese01] Ümran Esendemir. An elasto-plastic stress analysis in a polymer matrix composite beam of arbitrary orientation subjected to transverse uniformly distributed load. Mathematical and Computational Applications, 6(1):53–65, April 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/53.

Esen:2011:DRB

[Ese11] İsmail Esen. Dynamic response of a beam due to an accelerating moving mass using moving finite element approximation. Mathematical and Computational Applications, 16 (1):171–182, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/171.

Ebrahimi:2019:HAP

[ESNB19] Abouzar Ebrahimi, Mohammad Saeed Seif, and Ali Nouri-Borujerdi. Hydrodynamic and acoustic performance analysis of marine propellers by combination of panel method and FW-h equations. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/81.

Eswaramoorthi:2022:EDF

[ETL22] Sheniyappan Eswaramoorthi, S. Thamaraiselvi, and Karuppusamy Loganathan. Exploration of Darcy–Forchheimer flows of non-Newtonian Casson and Williamson conveying tiny particles experiencing binary chemical reaction and thermal radiation: Comparative analysis. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/52.

Eivazi:2023:FCD

[ETW^{+}23] Hamidreza Eivazi, Jendrik-Alexander Tröger, Stefan Wittek, Stefan Hartmann, and Andreas Rausch. FE^{2} computations with deep neural networks: Algorithmic structure, data generation, and implementation. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/91.

Evren:2013:MCV

[EÜA13] Sanem Evren, Mustafa Ünel, and Mahmut F. Akşit. Modeling and control of a variable speed variable pitch angle prototype wind turbine. Mathematical and Computational Applications, 18(3):408–420, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/408.

Erklig:2013:TBB

[EY13] Ahmet Erkliğ and Eyüp Yeter. On the thermal buckling behavior of laminated hybrid composite plates due to square/circular cut-outs. Mathematical and Computational Applications, 18(3):548–557, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/548.

Fikri:2022:HDT

[FA22a] Fatima Ezzahra Fikri and Karam Allali. HIV dynamics with a trilinear antibody growth function and saturated infection rate. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/85.

Fikri:2022:MAI

[FA22b] Fatima Ezzahra Fikri and Karam Allali. Modeling the adaptive immune response in an HBV infection model with virus to cell transmission in both liver with CTL immune response and the extrahepatic tissue. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/65.

Fantuzzi:2020:SIM

[Fan20] Nicholas Fantuzzi. Special issue “Mesh-Free and finite element-based methods for structural mechanics applications”. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/75.

Fazio:2021:EUB

[Faz21] Riccardo Fazio. Existence and uniqueness of BVPs defined on semiinfinite intervals: Insight from the iterative transformation method. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/18.

Fu:2016:MAD

[FbF16] Sha Fu and Guo bing Fan. A multiple attribute decision-making method based on exponential fuzzy numbers. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/19.

Fard:2017:EUS

[FBR17] Omid Solaymani Fard, Tayebeh Aliabdoli Bidgoli, and Azim Rivaz. On existence and uniqueness of solutions to the fuzzy dynamic equations on time scales. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/16.

Friedrich:2019:HOM

[FCGL19] Léo A. J. Friedrich, Mitrofan Curti, Bart L. J. Gysen, and Elena A. Lomonova. High-order methods applied to nonlinear magnetostatic problems. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/24/1/19.

Fernandes:2020:CMH

[FdSPW20] Fábio A. O. Fernandes, Ricardo J. Alves de Sousa, Mariusz Ptak, and Johannes Wilhelm. Certified motorcycle helmets: Computational evaluation of the efficacy of standard requirements with finite element models. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/12.

Fritzen:2018:ACH

[FHRS18] Felix Fritzen, Bernard Haasdonk, David Ryckelynck, and Sebastian Schöps. An algorithmic comparison of the hyperreduction and the discrete empirical interpolation method for a nonlinear thermal problem. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/8.

Fazio:2021:FFI

[FIJ21] Riccardo Fazio, Alessandra Insana, and Alessandra Jannelli. A frontfixing implicit finite difference method for the American put options model. Mathematical and Computational Applications, 26(2):??, June 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/30.

Fazio:2021:NSF

[FJ21] Riccardo Fazio and Alessandra Jannelli. A nonstandard finite difference scheme for magnetohydro dynamics boundary layer flows of an incompressible fluid past a flat plate. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/22.

Fereidoon:2010:ASD

[FKRG10] A. Fereidoon, N. Kordani, Y. Rostamiyan, and D. D. Ganji. Analytical solution to determine displacement of nonlinear oscillations with parametric excitation by differential transformation method. Mathematical and Computational Applications, 15(5):810– 815, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/810.

Flurchick:2015:VCC

[Flu15] K. M. Flurchick. Visual and computational comparison of functionals used in density functional theory. Mathematical and Computational Applications, 20(2):111– 120, August 2015. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/2/111.

Fu:2021:LCS

[FLYY21] Zhuo-Jia Fu, Lu-Feng Li, De-Shun Yin, and Li-Li Yuan. A localized collocation solver based on t-complete functions for anti-plane transverse elastic wave propagation analysis in 2D phononic crystals. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/2.

Fraile:2020:JLP

[FMF20] Alberto Fraile, Roberto Martínez, and Daniel Fernández. Jacob’s ladder: Prime numbers in 2D. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/5.

Fredericks:2010:LIC

[FMM10] E. Fredericks, F. M. Mahomed, and K. Masike. Lie infinitesimal conserved quantities for Itô stochastic ODEs. Mathematical and Computational Applications, 15(4):601–612, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/601.

Farooq:2010:ISO

[FMR10] M. U. Farooq, F. M. Mahomed, and M. A. Rashid. Integration of systems of ODEs via nonlocal symmetry-like operators. Mathematical and Computational Applications, 15(4):585– 600, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/585.

Fernandes:2021:NMS

[FP21] Fábio A. O. Fernandes and Mariusz Ptak. Numerical modelling and simulation applied to head trauma. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/50.

Fraile:2018:CAA

[FPCA18] Alberto Fraile, Emmanouil Panagiotakis, Nicholas Christakis, and Luis Acedo. Cellular automata and artificial brain dynamics. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/75.

Flores:2023:NQM

[FS23] Guilmer Ferdinand González Flores and Pablo Barrera Sánchez. New quality measures for quadrilaterals and new discrete functionals for grid generation. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/95.

Familusi:2023:MBA

[FSH^{+}23] Mary A. Familusi, Sebastian Skatulla, Jagir R. Hussan, Olukayode O. Aremu, Daniel Mutithu, Evelyn N. Lumngwena, Freedom N. Gumedze, and Ntobeko A. B. Ntusi. Model-based assessment of elastic material parameters in rheumatic heart disease patients and healthy subjects. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/106.

Frausto-Solis:2021:CNN

[FSHGGB^{+}21] Juan Frausto-Solís, Lucía J. Hernández-González, Juan J. González-Barbosa, Juan Paulo Sánchez-Hernández, and Edgar Román-Rangel. Convolutional neural network–component transformation (CNN–CT) for confirmed COVID-19 cases. Mathematical and Computational Applications, 26(2):??, June 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/2/29.

Frausto-Solis:2021:CMO

[FSHRCV^{+}21] Juan Frausto-Solis, Leonor Hernández-Ramírez, Guadalupe Castilla-Valdez, Juan J. González-Barbosa, and Juan P. Sánchez-Hernández. Chaotic multi-objective simulated annealing and threshold accepting for job shop scheduling problem. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/8.

Fabbrocino:2022:MCS

[FSV^{+}22] Francesco Fabbrocino, Serena Saitta, Riccardo Vescovini, Nicholas Fantuzzi, and Raimondo Luciano. Meshless computational strategy for higher order strain gradient plate models. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/19.

Fernandez-Ternero:2021:UEO

[FTGSNV21] Desamparados Fernández-Ternero, Víctor M. Gómez-Sousa, and Juan Núñez-Valdés. Using the evolution operator to classify evolution algebras. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/57.

Fu:2015:MAD

[Fu15] Sha Fu. Multiple attribute decision-making model of grey target based on positive and negative bull’s-eye. Mathematical and Computational Applications, 20(3):189–199, December 2015. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/3/189.

Fu:2016:TPI

[Fu16] Sha Fu. Three-parameter interval grey number multi-attribute decision making method based on information entropy. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/17.

Fatima:2023:PMC

[FYuR^{+}23] Bibi Fatima, Mehmet Yavuz, Mati ur Rahman, Ali Althobaiti, and Saad Althobaiti. Predictive modeling and control strategies for the transmission of middle East respiratory syndrome coronavirus. Mathematical and Computational Applications, 28 (5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/98.

Gunay:1999:SAA

[GA99a] Durmuş Günay and Alpay Aydemir. Stress analysis in adhesive-butt joints of cylindrical pressure vessels. Mathematical and Computational Applications, 4(3): 233–239, December 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/233.

Gundogdu:1999:OTD

[GA99b] Omer Gundogdu and Kurt S. Anderson. Optimal trajectory determination for sagitally symmetric manual lifting tasks. Mathematical and Computational Applications, 4(2):169– 174, August 1999. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/169.

Ganguly:2010:ABS

[GA10] Ashok Ganguly and Pranjali Arondekar. Analysis of B-spline curve using discrete Fourier transform. Mathematical and Computational Applications, 15(1):127– 136, April 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/127.

Granda:2017:BBD

[GAAAS17] Danilo Granda, Wilbert G.

Aguilar, Diego Arcos-Aviles, and Danny Sotomayor. Broken bar diagnosis for squirrel cage induction motors using frequency analysis based on MCSA and continuous wavelet transform. Mathematical and Computational Applications, 22(2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/2/30.

Gunay:1996:TST

[GAÖ96] Durmuş Günay, Alpay Aydemir, and Halil Özer. Torsional stresses in the transverse fillet weld tubular joints. Mathematical and Computational Applications, 1(1): 29–35, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/29.

Goktepe:2011:SDT

[GAS11] A. Burak Göktepe, Selim Altun, and Alper Sezer. Simulation of dilatometer tests by neural networks. Mathematical and Computational Applications, 16(2):535– 545, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/535.

Gun:1998:BEF

[GB98] Halit Gun and A. Adib Becker. Boundary emenet formulation in elastoplastic stress anaysis. Mathematical and Computational Applications, 3(3):139–151, December 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/3/139.

Gun:2000:BIE

[GB00] Halit Gun and A. Adib Becker. The boundary integral equation method for elastoplastic stres analysis of frictional contact problems. Mathematical and Computational Applications, 5(1):23–33, April 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/1/23.

Gladkov:2018:MTB

[GB18] Sergey Gladkov and Sophie Bogdanova. On the moving trajectory of a ball in a viscous liquid between two concentric rigid spheres. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/77.

Girondot:2023:CDS

[GB23] Marc Girondot and Jon Barry. Computation of the distribution of the sum of independent negative binomial random variables. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/63.

Gedik:2014:CCD

[GBİ14] Hakan Gedik, Hasan Bal, and Mahmut İzciler. Chance constrained data envelopment analysis for efficiency analysis: an application to Turkish manufacture of iron and steel sector. Mathematical and Computational Applications, 19 (1):1–11, April 2014. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/1.

Giraldo:2023:ASS

[GC23a] Juan F. Giraldo and Victor M. Calo. An adaptive in space, stabilized finite element method via residual minimization for linear and nonlinear unsteady advection– diffusion–reaction equations. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/7.

Gupta:2023:FEA

[GC23b] Vivek Gupta and Arnab Chanda. Finite element analysis of hierarchical metamaterialbased patterns for generating high expansion in skin grafting. Mathematical and Computational Applications, 28 (4):??, August 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/89.

Gaspar-Cunha:2023:MOO

[GCCMD23] António Gaspar-Cunha, Paulo Costa, Francisco Monaco, and Alexandre Delbem. Manyobjectives optimization: a machine learning approach for reducing the number of objectives. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/17.

Guan:2010:NTA

[GCGZ10] Le Guan, Jinkui Chu, Jiali Gao, and Ran Zhang. Nonlinear transient analysis for largescale dynamics of microelectromechanical systems with the MOR-PIM method. Mathematical and Computational Applications, 15(5):859– 865, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/859.

Gursel:2012:VAR

[GÇÖ12] K. Turgut Gürsel, Tufan Çoban, and Aydogan Özdamar. Vibration analysis of rotor blades of a farm windpower plant. Mathematical and Computational Applications, 17(2):164– 175, August 2012. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/2/164.

Guerrero:2018:SBO

[GCPV18] Joel Guerrero, Alberto Cominetti, Jan Pralits, and Diego Villa. Surrogate-based optimization using an opensource framework: The bulbous bow shape optimization case. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/60.

Goktas:2007:UAN

[GÇŞT07] Haldun Göktas, Abdullah Çavusoğlu, Baha Şen, and İhsan Toktaş. The use of artificial neural networks in simulation of mobile ground vehicles. Mathematical and Computational Applications, 12(2):87–96, August 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/2/87.

Goksen:2012:PMC

[GDE12] Yılmaz Gökşen, Onur Doğan, and Mete Eminağaoğlu. A proposed model for candidate selection process in political parties based on fuzzy logic methodology. Mathematical and Computational Applications, 17(2):152–163, August 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/2/152.

Gorgulu:2000:RCT

[GE00] Ali Görgülü and Cumali Ekici. On the Ricci curvature tensor of (k+1)- dimensional semi ruled surfaces evn+1. Mathematical and Computational Applications, 5(3): 149–155, December 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/3/149.

Gucuyen:2013:DIB

[GEKB13] Engin Gücüyen, R. Tuğrul Erdem, Erkan Kantar, and Muhiddin Bağcı. Determination of the impact behavior of concrete and reinforced concrete beams. Mathematical and Computational Applications, 18(3):502– 510, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/502.

Gomez:2019:NAQ

[GERV19] Yolanda Raquel Baca Gómez, Hugo Estrada Esquivel, Alicia Martínez Rebollar, and Daniel Villanueva Vásquez. A novel air quality monitoring unit using cloudino and FIWARE technologies. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/15.

Gavazzoni:2022:MPI

[GFPF22] Matteo Gavazzoni, Nicola Ferro, Simona Perotto, and Stefano Foletti. Multi-physics inverse homogenization for the design of innovative cellular materials: Application to thermo-elastic problems. Mathematical and Computational Applications, 27(1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/15.

Guvenc:1999:SFE

[GGJ99] Ziya B. Güvenç, Dilek Güvenç, and Julius Jellinek. Structural forms and energies of nin, n=12-14, clusters. Mathematical and Computational Applications, 4(1):75–81, April 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/75.

Ganji:2010:AST

[GGJ10] Z. Z. Ganji, D. D. Ganji, and A. Janalizadeh. Analytical solution of twodimensional viscous flow between slowly expanding or contracting walls with weak permeability. Mathematical and Computational Applications, 15(5):957–961, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/957.

Guzman-Gaspar:2020:DER

[GGMMDI20] José-Yaír Guzmán-Gaspar, Efrén Mezura-Montes, and Saúl Domínguez-Isidro. Differential evolution in robust optimization over time using a survival time approach. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/72.

Giunta:2021:MBE

[GGMO21] Andrea Giunta, Gaetano Giunta, Domenico Marino, and Francesco Oliveri. Market behavior and evolution of wealth distribution: a simulation model based on artificial agents. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/12.

Gurbuz:2011:NAH

[GGS11] Burcu Gürbüz, Mustafa Gülsu, and Mehmet Sezer. Numerical approach of high-order linear delay difference equations with variable coefficients in terms of Laguerre polynomials. Mathematical and Computational Applications, 16(1):267–278, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/267.

Garg:2023:EAM

[GGT^{+}23] Anshika Garg, Shubham Gupta, Nitesh Tewari, Sukeshana Srivastav, and Arnab Chanda. Effect of adhesive materials in re-attachment of crown and crown–root fractures of permanent maxillary anterior tooth: a computational study. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/41.

Goktas:2011:SCR

[GH11] Ünal Göktaş and Willy Hereman. Symbolic computation of recursion operators for nonlinear differential-difference equations. Mathematical and Computational Applications, 16(1):1– 12, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/1.

Ghaddar:2018:NSD

[Gha18a] Chahid Kamel Ghaddar. Novel spreadsheet direct method for optimal control problems. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/6.

Ghaddar:2018:RMP

[Gha18b] Chahid Kamel Ghaddar. Rapid modeling and parameter estimation of partial differential algebraic equations by a functional spreadsheet paradigm. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/39.

Ghaddar:2018:RSO

[Gha18c] Chahid Kamel Ghaddar. Rapid solution of optimal control problems by a functional spreadsheet paradigm: a practical method for the nonprogrammer. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/54.

Ghovatmand:2011:ACA

[GHN11] M. Ghovatmand, M. M. Hosseini, and M. Nilli. Application of Chebyshev approximation in the process of variational iteration method for solving differentialalgebraic equations. Mathematical and Computational Applications, 16(4):969– 978, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/969.

Giansiracusa:2021:EVU

[Gia21] Noah Giansiracusa. An evolutionary view of the U.S. supreme court. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/37.

Goktas:2011:MMS

[GK11a] Ünal Göktaş and Devendra Kapadia. Methods in Mathematica for solving ordinary differential equations. Mathematical and Computational Applications, 16(4):784– 796, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/784.

Gursoy:2011:EDE

[GK11b] Ayşın Erkan Gürsoy and İlhan Karakılıç. Expression of dual Euler parameters using the dual Rodrigues parameters and their application to the screw transformation. Mathematical and Computational Applications, 16(3):680– 689, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/680.

Gumus:2012:SDP

[GK12] O. A. Gumus and H. Kose. On the stability of delay population dynamics related with Allee effects. Mathematical and Computational Applications, 17(1):56–67, April 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/1/56.

Grigorieva:2018:OSP

[GK18] Ellina Grigorieva and Evgenii Khailov. Optimal strategies for psoriasis treatment. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/45.

Guzelbey:2005:CFB

[GKE05] Ibrahim H. Guzelbey, Bahattin Kanber, and Ahmet Erklig. Coupling of finite and boundary element methods with incompatible interfaces. Mathematical and Computational Applications, 10(3):321– 330, December 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/321.

Goktepe:2005:DOA

[GLA05] A. Burak Goktepe, A. Hilmi Lav, and Selim Altun. Dynamic optimization algorithm for vertical alignment of highways. Mathematical and Computational Applications, 10(3):341–350, December 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/341.

Goktepe:2008:FDS

[GLAA08] A. Burak Göktepe, A. Hilmi Lav, Selim Altun, and Gökhan Altıntaş. Fuzzy decision support system to determine swell/ shrink factor affecting earthwork optimization of highways. Mathematical and Computational Applications, 13 (1):61–70, April 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/1/61.

Gregori:2018:SNE

[GLS18] Julia Calatayud Gregori, Juan Carlos Cortés López, and Marc Jornet Sanz. Some notes to extend the study on random non-autonomous second order linear differential equations appearing in mathematical modeling. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/76.

Gulez:2003:ANN

[GM03] Kayhan Gulez and Nobuyoshi Mutoh. Application of neural networks to estimate common mode (CM) model impedance parameters in AC drives. Mathematical and Computational Applications, 8(2):225–233, August 2003. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/8/2/225.

Guo:2017:GAL

[GM17] Jidong Guo and Shugang Ma. Game analysis of low carbonization for urban logistics service systems. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/1/12.

Gaddis:2021:MSC

[GM21] Matthew David Gaddis and Valipuram S. Manoranjan. Modeling the spread of COVID-19 in enclosed spaces. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/79.

Gavina:2019:SNS

[GMV19] Alexandra Gavina, José M. A. Matos, and Paulo B. Vasconcelos. Solving nonholonomic systems with the Tau method. Mathematical and Computational Applications, 24(4):??, December 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/91.

Geiser:2019:CSM

[GN19] Jürgen Geiser and Amirbahador Nasari. Comparison of splitting methods for deterministic/ stochastic Gross–Pitaevskii equation. Mathematical and Computational Applications, 24(3):??, September 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/76.

Gyamerah:2019:HCY

[GNI19] Samuel Asante Gyamerah, Philip Ngare, and Dennis Ikpe. Hedging crop yields against weather uncertainties — a weather derivative perspective. Mathematical and Computational Applications, 24(3):??, September 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/3/71.

Gokuzum:2019:ANN

[GNK19] Felix Selim Göküzüm, Lu Trong Khiem Nguyen, and Marc-André Keip. An artificial neural network based solution scheme for periodic computational homogenization of electrostatic problems. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/40.

Gunay:1996:EAM

[GÖA96] Durmuş Günay, Halil Özer, and Alpay Aydemir. The effects of addendum modification coefficient on tooth stresses of spur Gear. Mathematical and Computational Applications, 1(1): 36–43, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/36.

Guseinov:2003:SPA

[GOA03] Kh. G. Guseinov, Orhan Ozer, and Emrah Akyar.

On the semicontinuity properties of the attainable sets of control systems with controls in L_{p}. Mathematical and Computational Applications, 8(1):127–134, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/127.

Guseinov:2003:CDI

[GÖD03] Kh. G. Guseinov, O. Özer, and S. A. Düzce. On the construction of differential inclusion with prescribed integral funnel. Mathematical and Computational Applications, 8(1):119–126, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/119.

Gbadamosi:2018:QAD

[GOOM18] Babatunde Gbadamosi, Mayowa Micheal Ojo, Segun Isaac Oke, and Maba Boniface Matadi. Qualitative analysis of a dengue fever model. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/33.

Gurgen:1996:DIG

[GÖV96] Fikret Gürgen, Emrah Önal, and Füsun G. Varol. The decision of intrauerine growth retardation from ultrasonographic examinations with neural networks. Mathematical and Computational Applications, 1(1):44–49, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/44.

Gupta:2021:USF

[GP21] Rangan Gupta and Christian Pierdzioch. Uncertainty, spillovers, and forecasts of the realized variance of gold returns. Mathematical and Computational Applications, 26(3):??, September 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/49.

Gonzalez-Parra:2019:MMC

[GPACC^{+}19] Gilberto C. González-Parra, Diego F. Aranda, Benito Chen-Charpentier, Miguel Díaz-Rodríguez, and Jaime E. Castellanos. Mathematical modeling and characterization of the spread of Chikungunya in Colombia. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/24/1/6.

Golcu:2010:PHE

[GPEG10] Mustafa Gölcü, Yasar Pancar, H. Sevil Ergür, and Esrah Ö. Göral. Prediction of head, efficiency, and power characteristics in a semiopen impeller. Mathematical and Computational Applications, 15 (1):137–147, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/137.

Gonzalez-Parra:2021:INS

[GPMRVM21] Gilberto Gonzalez-Parra, David Martínez-Rodríguez, and Rafael J. Villanueva-Micó. Impact of a new SARS-CoV-2 variant on the population: a mathematical modeling approach. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/25.

Ganguly:2001:PCM

[GS01] Ashok Ganguly and Shilpa Shinde. Propagation of cylindrical magneto-gasdynamic shock wave with varying energy in a rotating, gravitating, non-uniform atmosphere. Mathematical and Computational Applications, 6(3):177–188, December 2001. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/3/177.

Ganguly:2002:CMH

[GS02] Ashok Ganguly and Shilpa Shinde. Cylindrical magneto-hydrodynamic model of shock waves with radiation heat flux in non-uniform rotating atmosphere. Mathematical and Computational Applications, 7(3): 287–296, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/287.

Ganguly:2010:ESM

[GS10] Ashok Ganguly and Pankaj Sharma. The emergence of spherical magneto-gas dynamic strong shock with radiation near the surface of a star with a rotating, gravitating, non-uniform atmosphere. Mathematical and Computational Applications, 15(3):461– 471, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/461.

Gao:2022:NTP

[GT22] Haitao Gao and Jin Tao.

A novel trajectory planning method for parafoil airdrop system based on geometric segmentation strategy. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/75.

Gunes:2005:FGP

[GU05] Mustafa Güneş and Nurullah Umarosman. Fuzzy goal programming approach on computation of the fuzzy arithmetic mean. Mathematical and Computational Applications, 10(2):211– 220, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/211.

Guclu:2004:ASC

[Guc04] Rahmi Guclu. Active suspension control of eight degrees of freedom vehicle model. Mathematical and Computational Applications, 9(1): 1–10, April 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/1.

Guldemir:2003:SMS

[Gul03] Hanifi Guldemir. Sliding mode speed control for DC drive systems. Mathematical and Computational Applications, 8(3): 377–384, December 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/377.

Gundogdu:1996:NAO

[Gun96] Omer Gundogdu. A new algorithm to obtain δoperator based transfer function from its continuous time counterpart. Mathematical and Computational Applications, 1(2):42–46, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/42.

Guyeux:2020:CVD

[Guy20] Christophe Guyeux. Convergence versus divergence behaviors of asynchronous iterations, and their applications in concrete situations. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/69.

Gamboa:2020:NAT

[GVC20] Diana Gamboa, Carlos E. Vázquez, and Paul J. Campos. Nonlinear analysis for a type-1 diabetes model with focus on t-cells and pancreatic β-cells behavior. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/23.

Ganguly:1997:FMO

[GW97] Ashok Ganguly and Kamal Wadhwa. On fmonotone operators and generalized strongly nonlinear variational inequalities. Mathematical and Computational Applications, 2(3):155–162, December 1997. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/3/155.

Ge:1998:OOA

[GWC98] Yuzhen Ge, Layne T.

Watson, and Emmanuel G. Collins, Jr. An object oriented approach to semidefinite programming. Mathematical and Computational Applications, 3(1):1–10, April 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/1/1.

Guner:2013:LCM

[GY13] Aysun Güner and Salih Yalçınbaş. Legendre collocation method for solving nonlinear differential equations. Mathematical and Computational Applications, 18(3):521– 530, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/521.

Guo:2014:OHB

[GZ14] Peigang Guo and Yongquan Zhou. Optimizing hybrid basis function types in functional network design via cultural algorithm. Mathematical and Computational Applications, 19(1):50– 68, April 2014. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/50.

Glimm:2020:NAN

[GZ20] Tilmann Glimm and Jianying Zhang. Numerical approach to a nonlocal advection-reaction-diffusion model of cartilage pattern formation. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/36.

Gao:2020:ASI

[GZY20] Penglei Gao, Rui Zhang, and Xi Yang. The application of stock index price prediction with neural network. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/53.

Hayat:2010:PMA

[HA10] T. Hayat and F. M. Abbasi. Peristaltic mechanism in an asymmetric channel with heat transfer. Mathematical and Computational Applications, 15(4):621– 637, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/621.

Hoseini:2020:MPG

[HAH^{+}20] Behnaz Sheikh Hoseini, Muhammad Akram, Mehrnaz Sheikh Hosseini, Hossein Rashmanlou, and Rajab Ali Borzooei. Maximal product of graphs under vague environment. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/10.

Hajiyev:2004:IPK

[Haj04] Chingiz Hajiyev. Informativeness of parallel Kalman filters. Mathematical and Computational Applications, 9(3): 457–462, December 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/3/457.

Hajiyev:2005:MLC

[Haj05] Chingiz Hajiyev. Modified LSM for calibration of measuring apparatus. Mathematical and Computational Applications, 10(1):71–77, April 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/71.

Hasar:2003:DCD

[HAK03] U. C. Hasar, M. F. Akay, and S. N. Kharkovsky. Determination of complex dielectric permittivity of loss materials at microwave frequencies. Mathematical and Computational Applications, 8(3):319–326, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/319.

Hakula:2022:RBL

[Hak22] Harri Hakula. Resolving boundary layers with harmonic extension finite elements. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/57.

Hayat:2010:MFE

[HAMM10] T. Hayat, S. Abelman, E. Momoniat, and F. M. Mahomed. Magnetic field and endoscope influences on peristaltic transport: an exact solution. Mathematical and Computational Applications, 15(4):638–657, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/638.

Hartig:2021:AGC

[Har21]  Marie-Sophie  Hartig.

Approximation of Gaussian curvature by the angular defect: an error analysis. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/15.

Hassanov:2001:MCL

[Has01] Hikmet Hassanov. Motion of compressible liquids in pipes. Mathematical and Computational Applications, 6 (1):75–83, April 2001. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/6/1/75.

Hassanov:2002:PPU

[Has02] Hikmet G. Hassanov. Possibility of practical using of electric field in transportation systems. Mathematical and Computational Applications, 7(2):113–121, August 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/2/113.

Hassanov:2003:MMS

[Has03a] Hikmet Gafaroglu Hassanov. A mathematical model for the separation of viscous liquid flow by magnetic field. Mathematical and Computational Applications, 8(1):13–19, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/13.

Hassanov:2003:TEI

[Has03b] Hikmet Gafaroglu Hassanov. Thermodynamic effects at interaction between laser beam and heavy-viscous liquids. Mathematical and Computational Applications, 8(1):21–26, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/21.

Hajiyev:1999:EAD

[HB99] Chingiz Hajiyev and Alper Burat. Error analysis of determining airplane location by global positioning system. Mathematical and Computational Applications, 4(2):137–144, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/137.

Hayir:2001:SSW

[HB01] Adul Hayir and Ibrahim Bakirtas. Scattering of SH waves by fixed-rigid nearly circular cylinder. Mathematical and Computational Applications, 6(1):37–45, April 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/37.

Human:2022:BBG

[HBFM22] Schalk W. Human, Andriette Bekker, Johannes T. Ferreira, and Philip Albert Mijburgh. A bivariate beta from gamma ratios for determining a potential variance change point: Inspired from a process control scenario. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/61.

Haager:1996:SNT

[HBN96] G. Haager, G. Baumann, and T. F. Nonnenmacher. Symmetries of nonlinear telegraph equations in strong fields. Mathematical and Computational Applications, 1(2):47–70, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/47.

Hu:2017:RPD

[HBS^{+}17] Min Hu, Xiangyu Bu, Xiao Sun, Zixi Yu, and Yaona Zheng. Rape plant disease recognition method of multi-feature fusion based on D–S evidence theory. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/18.

Hassan:2022:BGH

[HC22] Marwa K. H. Hassan and Christophe Chesneau. Bivariate generalized half-logistic distribution: Properties and its application in household financial affordability in KSA. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/72.

Heras-Cervantes:2020:ACF

[HCdCTAAMEJ20] Mario Heras-Cervantes, Adriana del Carmen Téllez-Anguiano, Juan Anzurez-Marín, and Elisa Espinosa-Juárez. Analysis and comparison of fuzzy models and observers for DC-DC converters applied to a distillation column heating actuator. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/55.

Hernandez-Caceres:2020:IBB

[HCGFOON20] José Luis Hernández-Caceres, René Iván González-Fernández, Marlis Ontivero-Ortega, and Guido Nolte. Introducing BisQ, a bicoherencebased nonlinear index to explore the heart rhythm. Mathematical and Computational Applications, 25(3):??, September 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/3/45.

Hedar:2020:EAE

[HDAA20] Abdel-Rahman Hedar, Wael Deabes, Majid Almaraashi, and Hesham H. Amin. Evolutionary algorithms enhanced with quadratic coding and sensing search for global optimization. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/7.

Heizmann:2023:FID

[Hei23] Nico Heizmann. On the fluctuations of internal DLA on the Sierpinski gasket graph. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/73.

Hossain:2022:NSE

[HFAAM^{+}22] Muhammad Sajjad Hossain, Muhammad Fayz-Al-Asad, Muhammad Saiful Islam Mallik, Mehmet Yavuz, Md. Abdul Alim, and Kazi Md. Khairul Basher. Numerical study of the effect of a heated cylinder on natural convection in a square cavity in the presence of a magnetic field. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/58.

Hosseini-Farid:2020:SRB

[HFATZR^{+}20] Mohammad Hosseini-Farid, MaryamSadat Amiri-Tehrani-Zadeh, Mohammadreza Ramzanpour, Mariusz Ziejewski, and Ghodrat Karami. The strain rates in the brain, brainstem, dura, and skull under dynamic loadings. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/21.

Habigt:2022:VVC

[HGS^{+}22] Moriz A. Habigt, Jonas Gesenhues, Maike Stemmler, Marc Hein, Rolf Rossaint, and Mare Mechelinck. In vivo validation of a cardiovascular simulation model in pigs. Mathematical and Computational Applications, 27(2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/2/28.

Huang:2007:ORI

[HH07] Yung-Fu Huang and Kuang-Hua Hsu. Optimal Retailer’s inventory policy under supplier credits linked to retailer payment time. Mathematical and Computational Applications, 12 (2):77–86, August 2007. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/2/77.

Hussen:2022:PMT

[HH22] Abdulmtalb Hussen and Wenjie He. Prony method for two-generator sparse expansion problem. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/4/60.

Hensawang:2021:PRP

[HIVH21] Saruda Hensawang, Sittisak Injan, Pariwate Varnakovida, and Usa Humphries. Predicting rice production in central Thailand using the WOFOST model with ENSO impact. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/72.

Huang:2010:CCI

[HJL10] Xingli Huang, Peifa Jia, and Bo Liu. Controlling chaos by an improved estimation of distribution algorithm. Mathematical and Computational Applications, 15(5):866– 871, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/866.

Hashe:2023:NAE

[HK23] Vuyo T. Hashe and Thokozani J. Kunene. Numerical analysis of the effect of the vortex finder on the Hydrocyclone’s split water ratio and separation performance. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/50.

Hong:2014:NEA

[HL14] Bao-Jian Hong and Dian-Chen Lu. New exact analytical solutions for the general KdV equation with variable coefficients. Mathematical and Computational Applications, 19(3):194– 207, December 2014. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/3/194.

Hong:2016:NAS

[HL16] Baojian Hong and Dianchen Lu. New analytic solutions for the (N + 1)-dimensional generalized Boussinesq equation. Mathematical and Computational Applications, 21(2):??, June 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/8.

Hong:2015:HAS

[HLL15] Baojian Hong, Dianchen Lu, and Yangzheng Liu. Homotopic approximate solutions for the general perturbed nonlinear Schrödinger equation. Mathematical and Computational Applications, 20(1):50–61, April 2015. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/1/50.

Hang:2017:SAB

[HLX^{+}17] Yongsheng Hang, Yue Liu, Xiaoyang Xu, Yan Chen, and Shu Mo. Sensitivity analysis based on Markovian integration by parts formula. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/4/40. See re-

traction [HLX^{+}18].

Hang:2018:RHL

[HLX^{+}18] Yongsheng Hang, Yue Liu, Xiaoyang Xu, Yan Chen, and Shu Mo. Retraction: Hang, Y.; Liu, Y.; Xu, X.; Chen, Y.; Mo, S. Sensitivity Analysis Based on Markovian Integration by Parts Formula. Math. Comput. Appl. 2017, 22,

40. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/44.  See

[HLX^{+}17].

Harley:2010:EBV

[HM10] C. Harley and E. Momoniat. Efficient boundary value problem solution for a Lane–Emden equation. Mathematical and Computational Applications, 15(4):613– 620, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/613.

Hu:2023:API

[HM23] Beichao Hu and Dwayne McDaniel. Applying physics-informed neural networks to solve Navier– Stokes equations for laminar flow around a particle. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/102.

Hosseini:2011:CAP

[HMDG11] M. M. Hosseini, Syed Tauseef Mohyud-Din, and H. Ghaneai. On the coupling of auxiliary parameter, Adomian’s polynomials and correction functional. Mathematical and Computational Applications, 16(4):959–968, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/959.

Hasan:2018:MLB

[HMKS18] Ali Hasan, Sana Moin, Ahmad Karim, and Shahaboddin Shamshirband. Machine learning-based sentiment analysis for Twitter accounts. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/11.

Hosry:2020:HFA

[HNB20] Aline Hosry, Roger Nakad, and Sachin Bhalekar. A hybrid function approach to solving a class of Fredholm and Volterra integro-differential equations. Mathematical and Computational Applications, 25(2):??, June 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/30.

Hong:2020:ABC

[Hon20]  Yi Hong.  Arbitrage

bounds on currency basket options. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/60.

Howard:2022:AAA

[How22] Roy M. Howard. Arbitrarily accurate analytical approximations for the error function. Mathematical and Computational Applications, 27(1):14:1–14:44, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/14.

Hayat:2013:SSB

[HPA13] Tasawar Hayat, Mehmet Pakdemirli, and Yiğit Aksoy. Similarity solutions for boundary layer equations of a powel-Eyring fluid. Mathematical and Computational Applications, 18 (1):62–70, April 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/1/62.

Hilth:2019:DPT

[HRM19] William Hilth, David Ryckelynck, and Claire Menet. Data pruning of tomographic data for the calibration of strain localization models. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/18.

Hervas:2020:MPA

[HSG^{+}20] Antonio Hervás, Pedro Pablo Soriano, Joan Guàrdia i Olmos, Maribel Peró, Roberto Capilla, and José Miguel Montañana. Modelling the process to access the Spanish public university system based on structural equation models. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/31.

Hamdi:2019:ALA

[HT19] Abdelouahed Hamdi and Lotfi Tadj. Augmented Lagrangian approach to the newsvendor model with component commonality. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/55.

Han:2018:CIE

[HTLY18] Zhiming Han, Mitsuharu Terashima, Bing Liu, and Hidenari Yasui. CFD investigation of the effect of the feed spacer on hydrodynamics in spiral wound membrane modules. Mathematical and Computational Applications, 23(4):??, December 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/80.

Husain:2001:GMQ

[Hus01] Shamshad Husain. On generalized mixed quasivariational inclusions. Mathematical and Computational Applications, 6(3):203–215, December 2001. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/6/3/203.

Huseyin:1999:SNC

[HWA99] Demir Hüseyin, R. W. Willianms, and F. T. Akyıldız. The singularities near the corner of a viscoelastic fluid in a 2D cavity. Mathematical and Computational Applications, 4 (1):39–44, April 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/1/39.

Herui:2015:CEA

[HXY15] Cui Herui, Peng Xu, and Zhao Yuqi. Coevolution analysis on coal-power industries cluster ecosystem based on the Lotka–Volterra model: a case study of China. Mathematical and Computational Applications, 20(2):121– 136, August 2015. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/2/121.

Ismail:2004:DAS

[IA04] I. A. Ismail and M. M.

Allan. A designed ADI software for solving Poisson’s equation. Mathematical and Computational Applications, 9(2): 157–164, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/157.

Inc:2000:NPS

[İB00] M. İnç and N. Bildik.

Non-perturbative solution of the Ginzburg– Landau equation. Mathematical and Computational Applications, 5(2): 113–117, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/113.

Inc:1998:CNO

[İBB98] Mustafa İnç, Necdet Bildik, and Hasan Bulut. A comparison of numerical ODE solvers based on Euler methods. Mathematical and Computational Applications, 3(3):

153–159, December 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/3/153.

Irshad:2021:DPL

[ICDM21] Muhammed Rasheed Irshad, Christophe Chesneau, Veena D’cruz, and Radhakumari Maya. Discrete pseudo Lindley distribution: Properties, estimation and application on INAR(1) process. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/4/76.

Iyigunler:2012:ESS

[IÇÜ12] Ismail Iyigunler, Mine Çağlar, and Gazanfer Ünal. Exact solvability of stochastic differential equations driven by finite activity Levy processes. Mathematical and Computational Applications, 17(1):68–82, April 2012. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/1/68.

Ibrahim:2014:NFD

[ID14a] Rabha W. Ibrahim and Maslina Darus. A note on fractional differential subordination based on the Srivastavaowa fractional operator. Mathematical and Computational Applications, 19(2):115–123, August 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/115.

Ibrahim:2014:NSF

[ID14b] Rabha W. Ibrahim and Maslina Darus. On a new solution of fractional differential equation using complex transform in the unit disk. Mathematical and Computational Applications, 19(2):152– 160, August 2014. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/152.

Ismail:2004:GFP

[IE04] I. A. Ismail and E. E.

Elbehady. Green’s function for parallel planes and an open rectangular channel-flow. Mathematical and Computational Applications, 9(2): 215–224, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/215.

Imrak:2010:PWR

[IE10] C. Erdem Imrak and Cengiz Erdönmez. On the problem of wire rope model generation with axial loading. Mathematical and Computational Applications, 15(2):259– 268, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/259.

Imani:2018:IEE

[IGSB18] Mahmood Hosseini Imani, Mojtaba Jabbari Ghadi, Shahaboddin Shamshirband, and Marius M. Balas. Impact evaluation of electric vehicle parking on solving securityconstrained unit commitment problem. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/13.

Ibrahim:2018:TMT

[IMKEK18] Hager A. Ibrahim, Mahmoud Riad Mahmoud, Fatma A. Khalil, and Ghada A. El-Kelany. TLmoments for Type-I censored data with an application to the Weibull distribution. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/47.

Imron:2018:MMD

[IMWY18] Chairul Imron, Lutfi Mardianto, Basuki Widodo, and Tri Yogi Yuwono. Mathematical modeling of drag coefficient reduction in circular cylinder using two passive controls at re = 1000. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/2.

Inan:2008:CIC

[İTİO08] Sait İnan, Nurettin Türkan, İlyas İnci, and Davut Olgun. Comparison of IBM-2 calculations with x(5) critical point symmetry for low-lying states in 144-154Nd. Mathematical and Computational Applications, 13(2):101– 112, August 2008. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/2/101.

Inan:2012:II

[İTM12] Sait İnan, Nureddin Türkan, and İsmail Maraş. The investigation of 130,132Te by IBM-2. Mathematical and Computational Applications, 17(1):48–55, April 2012. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/1/48.

Injan:2021:ADA

[IWH21] Sittisak Injan, Angkool Wangwongchai, and Usa Humphries. Application of data assimilation and the relationship between ENSO and precipitation. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/24.

Inan:1999:ISS

[İYK99] Sait İnan, Abdülkadir Yörük, and Atalay Küçükbursa. The IBM-1 study for some Ce isotopes. Mathematical and Computational Applications, 4(3): 205–207, December 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/205.

Iqbal:2020:ECS

[IYKA20] Naveed Iqbal, Humaira Yasmin, Bawfeh K. Kometa, and Adel A. Attiya. Effects of convection on Sisko fluid with peristalsis in an asymmetric channel. Mathematical and Computational Applications, 25 (3):??, September 2020.

CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/52.

Ismail:2004:FPR

[IZS04] I. A. Ismail, E. H.

Zahran, and M. Shehata. A formulation of Peaceman and Rachford ADI method for the threedimensional heat diffusion equation. Mathematical and Computational Applications, 9(2): 183–189, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/183.

Jabbari:2018:AEM

[Jab18] Ali Jabbari. An analytical expression for magnet shape optimization in surface-mounted permanent magnet machines. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/57.

Jamal:2021:NMB

[JAT^{+}21a] Farrukh Jamal, Ali H.

Abuzaid, Muhammad H. Tahir, Muhammad Arslan Nasir, Sadaf Khan, and Wali Khan Mashwani. New modified Burr III distribution, properties and applications.

Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/4/82.

Junior:2021:ERM

[JAT21b] Peterson Owusu Junior, Imhotep Paul Alagidede, and Aviral Kumar Tiwari. On the elicitability and risk model comparison of emerging markets equities. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/3/63.

Junqing:2015:SWS

[JB15] Meng Junqing and Nie Baisheng. Study on water seepage law of raw coal during loading process. Mathematical and Computational Applications, 20(3):217–227, December 2015. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/3/217.

Jin:2010:ETD

[JFRZ10] Wuyin Jin, Ruicheng Feng, Zhiyuan Rui, and Aihua Zhang. Effects of time delay on chaotic neuronal discharges. Mathematical and Computational Applications, 15(5):840– 845, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/840.

Jia:2013:VCA

[JH13] Jiu-Hong Jia and Tian-Qi Hang. Vibration characteristics analysis of the rub- impact rotor system with mass unbalance. Mathematical and Computational Applications, 18(2):71– 83, August 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/2/71.

Jamal:2022:GOL

[JHA^{+}22] Farrukh Jamal, Laba Handique, Abdul Hadi N. Ahmed, Sadaf Khan, Shakaiba Shafiq, and Waleed Marzouk. The generalized odd linear exponential family of distributions with applications to reliability theory. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/55.

Ji:2013:AML

[Ji13]  Li-Qun Ji.  Analysis

of a modified logistic model for describing the growth of durable customer goods in China. Mathematical and Computational Applications, 18(1):30–37, April 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/1/30.

Jiang:2017:NAM

[Jia17] Ziwu Jiang. A new approximation method with high order accuracy. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/1/11.

Joyce:2023:EGC

[JKS23] Maria Immaculate Joyce, Jagan Kandasamy, and Sivasankaran Sivanandam. Entropy generation of Cu–Al_{2}O_{3}/ water flow with convective boundary conditions through a porous stretching sheet with slip effect, Joule heating and chemical reaction. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/18.

Jantschi:2024:ACP

[JL24] Lorentz Jäntschi and Mohamed Louzazni. Accelerating convergence for the parameters of PV cell models. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/4.

Jin:2021:IMR

[JLH21] Qi-Wen Jin, Zheng Liu, and Shuan-Hai He. Investigation on the mathematical relation model of structural reliability and structural robustness. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/26.

Jahanshahloo:2012:MDE

[JLKK12] Gholamreza Jahanshahloo, Farhad Hosseinzadeh Lotfi, Masomeh Khanmohammadi, and Mansoureh Kazemimanesh. A method for discriminating efficient candidates with ranked voting data by common weights. Mathematical and Computational Applications, 17(1):1–8, April 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/1/1.

Jahanshahloo:2005:INE

[JLS^{+}05] G. R. Jahanshahloo, F. Hossienzadeh Lotfi, N. Shoja, G. Tohidi, and S. Razavyan. The identification of nondominated and efficient paths on a network. Mathematical and Computational Applications, 10(2):185– 191, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/185.

Jahanshahloo:2010:PNM

[JLS^{+}10a] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, N. Shoja, A. Gholam Abri, M. Fallah Jelodar, and Kamran Jamali Firouzabadi. Proposing a new model on data envelopment analysis by considering non discretionary factors and a review on previous models. Mathematical and Computational Applications, 15(3):344–353, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/344.

Jahanshahloo:2010:REN

[JLS^{+}10b] Golamreza Jahanshahloo, Farhad Hosseinzadeh Lotfi, Naghi Shoja, Mehdi Fallah Jelodar, and Amir Gholam Abri. Ranking extreme and non-extreme efficient decision making units in data envelopment analysis. Mathematical and Computational Applications, 15(2):299–308, August 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/299.

Jahanshahloo:2004:MSM

[JLST04] G. R. Jahanshahloo, F. Hossienzadeh Lotfi, N. Shoja, and G. Tohidi. A method for solving 0-1 multiple objective programming problem using data envelopment analysis technique. Mathematical and Computational Applications, 9(2): 147–156, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/147.

Jodlbauer:2020:PMF

[JLW20] Daniel Jodlbauer, Ulrich Langer, and Thomas Wick. Parallel matrixfree higher-order finite element solvers for phase-field fracture problems. Mathematical and Computational Applications, 25(3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/3/40.

Jahanshahloo:2004:EEI

[JLZ04] G. R. Jahanshahloo, F. Hosseinzadeh Lotfi, and M. Zohrehbandian. Estimation of efficiency and infinitesimals in data envelopment analysis. Mathematical and Computational Applications, 9(2):299–302, August 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/2/299.

Johnpillai:2003:GPF

[JM03] I. K. Johnpillai and F. M. Mahomed. Goursat problem for the factorizable hyperbolic equation in two independent variables. Mathematical and Computational Applications, 8 (1):55–62, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/55.

Jasmine:2022:PEA

[JMB22] Mansura Jasmine, Abdolmajid Mohammadian, and Hossein Bonakdari.

On the prediction of evaporation in arid climate using machine learning model. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/32.

Jazayeriy:2018:FRS

[JMS18] Hamid Jazayeriy, Saghi Mohammadi, and Shahaboddin Shamshirband. A fast recommender system for cold user using categorized items. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/1.

Jafari:2018:SFD

[JR18] Raheleh Jafari and Sina Razvarz. Solution of fuzzy differential equations using fuzzy Sumudu transforms. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/5.

Jitman:2018:AHD

[JS18] Somphong Jitman and Ekkasit Sangwisut. The average hull dimension of negacyclic codes over finite fields. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/41.

Jagan:2022:SDT

[JS22a] Kandasamy Jagan and Sivanandam Sivasankaran. Soret & Dufour and triple stratification effect on MHD flow with velocity slip towards a stretching cylinder. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/25.

Jagan:2022:TDN

[JS22b] Kandasamy Jagan and Sivanandam Sivasankaran. Three-dimensional nonlinearly thermally radiated flow of Jeffrey nanoliquid towards a stretchy surface with convective boundary and Cattaneo– Christov flux. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/98.

Juarez-Smith:2019:PBG

[JSTGV^{+}19] Perla Juárez-Smith, Leonardo Trujillo, Mario García-Valdez, Francisco Fernández de Vega, and Francisco Chávez. Pool-based genetic programming using Evospace, local search and bloat control. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/78.

Jain:2022:AAT

[JTM^{+}22] Aakanksha Jain, Iman Bahreini Toussi, Abdolmajid Mohammadian, Hossein Bonakdari, and Majid Sartaj. Applications of ANFIS-type methods in simulation of systems in marine environments. Mathematical and Computational Applications, 27(2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/29.

Joubert:2023:FIA

[JWP23a] Johannes C. Joubert, Daniel N. Wilke, and Patrick Pizette. Fourier image analysis of multiphase interfaces to quantify primary atomization. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/55.

Joubert:2023:GFD

[JWP23b] Johannes C. Joubert, Daniel N. Wilke, and Patrick Pizette. A generalized finite difference scheme for multiphase flow. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/51.

Ji:2023:FTS

[JY23] Xiaofu Ji and Xueqing Yan. Finite-time static output-feedback H_{∞} control for discrete-time singular Markov jump systems based on eventtriggered scheme. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/1.

Jin:2010:MMA

[JZL10] Fang Jin, Xin Zhang, and Peide Liu. Method for multiple attribute decision-making with continuous random variable under risk based on projection model. Mathematical and Computational Applications, 15(3):394–403, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/394.

Kahramanli:1996:AAT

[KA96a]  Ş. Ş. Kahramanlı and

N. M. Allahverdi. Algebraic approach to transformations on hypercube system. Mathematical and Computational Applications, 1(1): 50–59, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/50.

Kilic:1996:PFC

[KA96b] Arş. Gör. E. Kılıç and I. H. Altaş. Power factor correction of synchronous motor using fuzzy logic. Mathematical and Computational Applications, 1(1):66–72, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/66.

Kilic:2002:DSF

[KA02] Osman Kılıç and Alaattin Aktaş. Determination of stress functions of a curved beam subjected to an arbitrarily directed single force at the free end. Mathematical and Computational Applications, 7(2): 181–188, August 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/2/181.

Koc:2009:MGB

[KA09] Savas Koç and Zafer Aydoğmus. A MAT-LAB/GUI based fault simulation tool for power system education. Mathematical and Computational Applications, 14(3):207–217, December 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/3/207.

Kaya:2011:DQS

[KA11a] Birol Kaya and Yalcin Arisoy. Differential quadrature solution for one-dimensional aquifer flow. Mathematical and Computational Applications, 16(2):524–534, August 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/524.

Koksal:2011:SAA

[KA11b] N. Sinan Köksal and Mehmet Alkan. Stress analysis in Al based composites depending on joining quality. Mathematical and Computational Applications, 16 (1):194–202, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/194.

Kula:2011:ERC

[KA11c] Kamile Şanli Kula and Ayşen Apaydin. Estimation of the randomized complete block design parameters with fuzzy goal programming. Mathematical and Computational Applications, 16(3):721–727, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/721.

Kolkiran:2016:CKA

[KA16] Aziz Kolkiran and Girish S.

Agarwal. Correction: Kolkiran, A.; Agarwal, G. S. Amplitude Noise Reduction in a Nano-Mechanical Oscillator. Math. Comput. Appl. 2011, 16 (1), 290– 300. Mathematical and Computational Applications, 21(3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/28.  See

[Kol11].

Khan:2010:AAF

[KAAK10] Najeeb Alam Khan, Asmat Ara, Muhammad Afzal, and Azam Khan. Analytical aspect of fourth-order parabolic partial differential equations with variable coefficients. Mathematical and Computational Applications, 15(3):481– 489, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/481.

Kuukoglu:2013:MMT

[KAEÖ13] İlker Küükoğlu, Aslı Aksoy, Seval Ene, and Nursel Öztürk. A mathematical model for two dimensional loading problem in cross-docking network design. Mathematical and Computational Applications, 18(3):273– 282, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/273.

Kafini:2023:ABS

[KAGAM23] Mohammad M. Kafini, Mohammed M. Al-Gharabli, and Adel M. Al-Mahdi. Asymptotic behavior of solutions to a nonlinear swelling soil system with time delay and variable exponents. Mathematical and Computational Applications, 28 (5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/94.

Kahraman:2011:TDC

[Kah11] Bayram Kahraman. Three dimensional cutting problem: an integer programming approach. Mathematical and Computational Applications, 16 (1):105–112, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/105.

Karlik:1996:CAN

[KAK96] Bekir Karlık, Serkan Aydın, and İlker Kılıç. Comparing artificial neural networks (ANN) image compression technique with different image compression techniques. Mathematical and Computational Applications, 1(1):165–171, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/165.

Kurt:2000:CCC

[KAK00] E. Kurt, S. Acar, and R. Kasap. A comparison of chaotic circuits from a statistical approach. Mathematical and Computational Applications, 5(2): 95–103, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/95.

Khaniyev:2008:DRR

[KAKB08] Tahir Khaniyev, Rovshan Aliyev, Zafer Küçük, and Nurgul Okur Bekar. On the distributions of a renewal reward process and its additive functional. Mathematical and Computational Applications, 13(1):41–50, April 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/1/41.

Khan:2020:RPM

[KAM20] Safeer Hussain Khan, Timilehin Opeyemi Alakoya, and Oluwatosin Temitope Mewomo. Relaxed projection methods with self-adaptive step size for solving variational inequality and fixed point problems for an infinite family of multivalued relatively nonexpansive mappings in Banach spaces. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/54.

Kanoglu:2000:UAC

[Kan00] Mehmet Kanoglu. Uncertainly analysis of cryogenic turbine efficiency. Mathematical and Computational Applications, 5(3):169–177, December 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/3/169.

Kantar:2016:EVW

[Kan16] Yeliz Mert Kantar. Estimating variances in weighted least-squares estimation of distributional parameters. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/7.

Kara:2003:PSG

[Kar03a] A. H. Kara. Potential symmetry generators of some perturbed nonlinear evolution related equations. Mathematical and Computational Applications, 8 (1):89–94, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/89.

Karlik:2003:NNI

[Kar03b] Bekir Karlik. A neural network image recognition for control of manufacturing plant. Mathematical and Computational Applications, 8(2): 181–189, August 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/181.

Kara:2010:ASC

[Kar10]  A. H. Kara.  An

analysis of the symmetries and conservation laws of the class of Zakharov–Kuznetsov equations. Mathematical and Computational Applications, 15(4):658– 664, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/658.

Karaoglan:2011:INN

[Kar11] Aslan Deniz Karaoglan. An integrated neural network structure for recognizing autocorrelated and trending processes. Mathematical and Computational Applications, 16(2):514–523, August 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/514.

Karasoy:2012:BEC

[Kar12] Durdu Karasoy. Bayesian estimator of a change point in the hazard function. Mathematical and Computational Applications, 17(1):29– 38, April 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/1/29.

Kasumov:2000:ESM

[Kas00] Baba M. Kasumov. Electron subsystem microscobic modeling of multiatomic crystals: The software package “scppband”. Mathematical and Computational Applications, 5(1):41–48, April 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/1/41.

Karami:2019:MLP

[KAS19a] A. Karami, Saeid Abbasbandy, and E. Shivanian. Meshless local Petrov–Galerkin formulation of inverse Stefan problem via moving least squares approximation. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/101.

Kasumo:2019:MIU

[Kas19b] Christian Kasumo. Minimizing an Insurer’s ultimate ruin probability by reinsurance and investments. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/21.

Katebi:2017:ELF

[Kat17] Ali Katebi. Engineers in a leadership flexibility space that use data envelopment analysis. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/4/47.

Kaya:1998:PTS

[Kay98] Meryem Kaya. A presentation theorem of the spherical wave functions. Mathematical and Computational Applications, 3(3):161–167, December 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/3/161.

Karabulut:2008:LTE

[KAY08] Korhan Karabulut, Ahmet Alkan, and Ahmet S. Yilmaz. Long term energy consumption forecasting using genetic programming. Mathematical and Computational Applications, 13 (2):71–80, August 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/2/71.

Karlik:2009:CHP

[KAY09] Bekir Karlık, Alpaslan Avcı, and A. Talha Yabanıgül. Classification of helicobacter pylori according to national strains using Bayesian learning. Mathematical and Computational Applications, 14(3):241– 251, December 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/3/241.

Kaya:2015:WCR

[Kay15] Yasin Kaya. A weakly convergence result on L_{p}(x) spaces. Mathematical and Computational Applications, 20(2):106– 110, August 2015. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/2/106.

Kayri:2016:PAB

[Kay16] Murat Kayri. Predictive abilities of Bayesian regularization and Levenberg– Marquardt algorithms in artificial neural networks: a comparative empirical study on social data. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/20.

Kazaz:2004:PRU

[Kaz04] M. Kazaz. Presentations for regular unbranched cp-coverings of a Klein bottle. Mathematical and Computational Applications, 9(2): 191–199, August 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/2/191.

Kazaz:2005:SCR

[Kaz05] M. Kazaz. The sine and cosine rules for pure triangles on the dual Lorentzian unit sphere S̃_{1}^{2}. Mathematical and Computational Applications, 10(2):193–201, August 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/2/193.

Kahramanli:2003:BFS

[KB03]  Şirzat Kahramanli and

Fatih Başçiftçi. Boolean functions simplification algorithm of o(n) complexity. Mathematical and Computational Applications, 8(3):271–278, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/271.

Korukoglu:2011:IVA

[KB11] Serdar Korukoğlu and Serkan Ballı. An improved Vogel’s approximation method for the transportation problem. Mathematical and Computational Applications, 16(2):370–381, August 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/370.

Kahya:2001:FCP

[KBE01] Volkan Kahya, Ahmet Birinci, and Ragıp Erdöl. Frictionless contact problem between an elastic layer bonded to a rigid support and a rigid stamp. Mathematical and Computational Applications, 6 (1):13–22, April 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/13.

Khotanzad:2006:CMC

[KBH06] Alireza Khotanzad, Jesse W.

Bennett, and Orlando J. Hernandez. A classification methodology for color textures using multispectral random field mathematical models. Mathematical and Computational Applications, 11(2):111– 120, August 2006. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/2/111.

Kilic:1996:FDI

[KÇ96] Osman Kılıç and Ali Çalışkan. The Frenet and Darboux instantaneous rotation for curves on space-like surface. Mathematical and Computational Applications, 1(2): 77–86, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/77.

Kara:2008:HPA

[KC08] Rukiye Kara and Mehmet Can. Hopf point analysis for ratio-dependent food chain models. Mathematical and Computational Applications, 13 (1):9–22, April 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/1/9.

Kusak:2011:ADS

[KÇ11] Hatice Kuşak and Ali Çalışkan. About dual spherical wrist motion and its trajectory surface as a ruled surface. Mathematical and Computational Applications, 16 (1):309–316, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/309.

Ku-Cauich:2018:ACG

[KCMLTR18] Juan Carlos Ku-Cauich, Guillermo Morales-Luna, and Horacio Tapia-Recillas. An authentication code over Galois rings with optimal impersonation and substitution probabilities. Mathematical and Computational Applications, 23(3):??, September 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/46.

Kucuk:2003:NNU

[KD03] Ilker Kucuk and Naim Derebasi. Neural network using genetic algorithm for magnetic performance prediction of toroidal wound cores at 50 Hz. Mathematical and Computational Applications, 8(2):201– 208, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/201.

Kas:2019:ACR

[KD19] Ziyad Kas and Manohar Das. Adaptive control of resistance spot welding based on a dynamic resistance model. Mathematical and Computational Applications, 24 (4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/86.

Karlik:1998:PNR

[KE98] Bekir Karlık and Aydın Ece. Pediatric nutritional requirements determination with neural networks. Mathematical and Computational Applications, 3(3): 169–175, December 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/3/169.

Kucuk:2001:EPS

[KE01] Mümin Küçük and Ümran Esendemir. An elastoplastic stress analysis on a metal matrix composite beam of arbitrary orientation supported from two ends under a transverse uniformly distributed load. Mathematical and Computational Applications, 6(3): 155–168, December 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/3/155.

Kantar:2011:NFE

[KEA11] Erkan Kantar, R. Tuğrul Erdem, and Özgür Anıl. Nonlinear finite element analysis of impact behavior of concrete beam. Mathematical and Computational Applications, 16 (1):183–193, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/183.

Kent:1996:SBC

[Ken96] E. Fuad Kent. Satisfaction of boundary conditions on complex channel walls with quadratic minimizaion method. Mathematical and Computational Applications, 1(1): 60–65, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/60.

Kentli:2011:SFL

[Ken11] Aykut Kentli. Studies on fuzzy logic control of electrical machines in Turkish universities: an overview. Mathematical and Computational Applications, 16 (1):236–247, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/236.

Kunc:2019:CKF

[KF19a] Oliver Kunc and Felix Fritzen. Correction: Kunc, O.; Fritzen, F. Finite Strain Homogenization Using a Reduced Basis and Efficient Sampling. Math. Comput. Appl. 2019, 24, 56. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/95.  See

[KF19b].

Kunc:2019:FSH

[KF19b] Oliver Kunc and Felix Fritzen. Finite strain homogenization using a reduced basis and efficient sampling. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/56. See cor-

rection [KF19a].

Khaki:2010:ASN

[KG10] M. Khaki and D. D. Ganji. Analytical solutions of nano boundary layer flows by using He’s homotopy perturbation method. Mathematical and Computational Applications, 15(5):962– 966, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/962.

Kutuk:2013:NBT

[KG13] M. Akif Kutuk and Ibrahim Gov. Novel bracing type against seismic loading: KX type. Mathematical and Computational Applications, 18(3):438–447, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/438.

Khayretdinova:2022:ISP

[KGCFK22] Guzel Khayretdinova, Christian Gout, Théophile Chaumont-Frelet, and Sergei Kuksenko. Image segmentation with a priori conditions: Applications to medical and geophysical imaging. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/26.

Khan:2015:SFC

[KH15] Najeeb Alam Khan and Zehra Husain. Spinning flow of Casson fluid near an infinite rotating disk. Mathematical and Computational Applications, 20(3):174– 188, December 2015. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/3/174.

Khan:2022:EAH

[KH22] Sohail A. Khan and Tasawar Hayat. Entropy analysis for hydromagnetic Darcy–Forchheimer flow subject to Soret and Dufour effects. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/80.

Khaniev:1999:SRS

[Kha99] T. A. Khaniev. Some results on a stochastic process with a discrete chance interference. Mathematical and Computational Applications, 4(2):145–152, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/145.

Khalique:2003:OSI

[Kha03] C. K. Khalique. Optimal systems and invariant solutions for a class of soil water equations. Mathematical and Computational Applications, 8(2): 135–141, August 2003.

CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/135.

Ke:2006:BAQ

[KHL06] Jau-Chuan Ke, Hsin-I Huang, and Chuen-Horng Lin. A batcharrival queue with multiple servers and fuzzy parameters: Parametric programming approach. Mathematical and Computational Applications, 11(3):181–191, December 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/181.

Khan:2022:MLL

[KHR^{+}22] Sadaf Khan, Gholamhossein G. Hamedani, Hesham Mohamed Reyad, Farrukh Jamal, Shakaiba Shafiq, and Soha Othman. The minimum Lindley Lomax distribution: Properties and applications. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/16.

Kilicman:2003:CCN

[Kil03] Adem Kilicman. A comparison on the commutative neutrix products of distributions. Mathematical and Computational Applications, 8(3): 343–351, December 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/343.

Kilicman:2004:NMT

[Kil04] Adem Kiliçman. A note on Mellin transform and distributions. Mathematical and Computational Applications, 9 (1):65–72, April 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/1/65.

Kiral:2004:EWE

[Kir04] Binnur Gören Kiral. Effects of the weld electrode on the fracture behavior of welded steel beam-to-column joints. Mathematical and Computational Applications, 9(3):449–456, December 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/3/449.

Krishna:2023:AMS

[KIY^{+}23] Ram Krishna, Agbotiname Lucky Imoize, Rajveer Singh Yaduvanshi, Harendra Singh, Arun Kumar Rana, and Subhendu Kumar Pani. Analysis of multi-stacked dielectric resonator antenna with its equivalent r-l-c circuit modeling for wireless communication systems. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/4.

Kiziltug:2014:CIF

[Kız14] Sezai Kızıltuğ. On characterization of inextensible flows of curves according to type-2 Bishop frame E3. Mathematical and Computational Applications, 19(1):69– 77, April 2014. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/1/69.

Khaleghi:2012:RSS

[KJZL12] Maryamsadat Khaleghi, Gholamreza Jahanshahloo, Majid Zohrehbandian, and Farhad Hosseinzadeh Lotfi. Returns to scale and scale elasticity in two-stage Dea. Mathematical and Computational Applications, 17(3):193–202, December 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/193.

Kopac:1997:NSG

[KK97] Mehmet Kopaç and Türkan Kopaç. Numerical simulation of a gaseous fueled (methane) combustor. Mathematical and Computational Applications, 2(3):101–118, December 1997. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/2/3/101.

Kargin:2013:SRH

[KK13] Levent Kargin and Veli Kurt. Some relations on Hermite matrix polynomials. Mathematical and Computational Applications, 18(3):323– 329, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/323.

Kurt:2004:EPM

[KKA04] E. Kurt, R. Kasap, and S. Acar. Effects of periodic magnetic field to the dynamics of vibrating beam. Mathematical and Computational Applications, 9(2): 275–284, August 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/2/275.

Ke:2008:AFD

[KKC08] Jau-Chuan Ke, Ming Yang Ko, and Kuo-Ching Chiou. Analysis of 24 factorial design for a controllable M/G/1 system. Mathematical and Computational Applications, 13(3):165–174, December 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/3/165.

Konnov:2018:DMO

[KKL18] Igor Konnov, Aleksey Kashuba, and Erkki Laitinen. Dual methods for optimal allocation of telecommunication network resources with several classes of users. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/31.

Keskin:2011:ASH

[KKS11] Yıldıray Keskin, Onur Karaoğlu, and Sema Servi. The approximate solution of high-order linear fractional differential equations with variable coefficients in terms of generalized Taylor polynomials. Mathematical and Computational Applications, 16(3):617– 629, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/617.

Kula:2005:FES

[KKY05] Levent Kula, Murat Kemal Karacan, and Yusuf Yaylı. Formulas for the exponential of a semi skew-symmetric matrix of order 4. Mathematical and Computational Applications, 10 (1):99–104, April 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/99.

Klein:2018:LOD

[Kle18] Axel Klein. Logic operators on delta-sigma bit-streams. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/1/4.

Kerr:2024:ASS

[KLGP24] Gilbert Kerr, Nehemiah Lopez, and Gilberto González-Parra. Analytical solutions of systems of linear delay differential equations by the Laplace transform: Featuring limit cycles. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/11.

Kontorovich:2010:CPC

[KLRA10] V. Kontorovich, Z. Lovtchikova, and F. Ramos-Alarcon. Correlation properties of chaos: Cumulant approach. Mathematical and Computational Applications, 15(5):946– 952, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/946.

Komathiraj:2010:CCR

[KM10] K. Komathiraj and S. D.

Maharaj. A class of charged relativistic spheres. Mathematical and Computational Applications, 15(4):665– 673, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/665.

Kucukbursa:2005:ICE

[KMÇ^{+}05] Atalay Küçükbursa, Kaan Manisa, Mehmet Çiftçi, Tahsin Babacan, and İsmail Maraş. IBM-1 calculations on the eveneven 122-128Te isotopes. Mathematical and Computational Applications, 10(1):9–17, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/9.

Krishna:2022:UTD

[KMCI22] Anuresha Krishna, Radhakumari Maya, Christophe Chesneau, and Muhammed Rasheed Irshad. The unit Teissier distribution and its applications. Mathematical and Computational Applications, 27(1):??, February 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/12.

Kizilates:2013:PHM

[KN13] Gözde Kızılates and Fidan Nuriyeva. A parametric hybrid method for the traveling salesman problem. Mathematical and Computational Applications, 18(3):459– 466, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/459.

Karimi:2018:EPL

[KN18] Hossein Karimi and Parviz Nasiri. Estimation parameter of R = P (Y < X) for lengthbiased weighted Lomax distributions in the presence of outliers. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/9.

Kovac:2020:HEP

[KNHK20] Mario Kovač, Philippe Notton, Daniel Hofman, and Josip Knezović. How Europe is preparing its core solution for exascale machines and a global, sovereign, advanced computing platform. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/46.

Kune:2022:RNH

[KNRC22] Ramesh Kune, Hari Singh Naik, Borra Shashidar Reddy, and Christophe Chesneau. Role of nanoparticles and heat source/sink on MHD flow of Cu-H_{2}O nanofluid flow past a vertical plate with Soret and Dufour effects. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/102.

Kocayusufoglu:2000:CMG

[KÖ00] İsmail Kocayusufoğlu and Ertuğrul Özdamar. Connections and minimizing geodesics of taxicab geometry. Mathematical and Computational Applications, 5(3): 191–200, December 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/3/191.

Keskin:2010:RDT

[KO10] Yıldıray Keskin and Galip Oturanç. Reduced differential transform method for generalized KdV equations. Mathematical and Computational Applications, 15(3):382–393, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/382.

Kuboye:2015:DSS

[KO15] J. O. Kuboye and Z. Omar. Derivation of a six-step block method for direct solution of second order ordinary differential equations. Mathematical and Computational Applications, 20(3):151– 159, December 2015. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/3/151.

Kucuk:2000:EPS

[KÖAG00] Mümin Küçük, Reşat Özcan, Nurettin Arslan, and Mustafa Gür. An elastic-plastic stress analysis in an orthotropic composite cantilever beam loaded uniformly. Mathematical and Computational Applications, 5(2): 129–138, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/129.

Kocayusufoglu:2000:TIT

[Koc00] İsmail Kocayusufoğlu.

Trigonometry on isotaxicab geometry. Mathematical and Computational Applications, 5(3): 201–212, December 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/3/201.

Kavaka:2011:EFM

[KODM11] Aydın Kavaka, Fuat Okaya, Serdar Doğanb, and Utkan Mutmana. Earth fill modelling with finite element method. Mathematical and Computational Applications, 16(3):565–575, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/565.

Koiso:2019:UCE

[Koi19] Miyuki Koiso. Uniqueness of closed equilibrium hypersurfaces for anisotropic surface energy and application to a capillary problem. Mathematical and Computational Applications, 24 (4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/4/88.

Kolkiran:2011:ANR

[Kol11] Aziz Kolkiran. Amplitude noise reduction in a nano-mechanical oscillator. Mathematical and Computational Applications, 16(1):290– 300, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/290. See cor-

rection [KA16].

Kopac:2002:EPS

[Kop02]  Mehmet Kopaç.  Ef-

fect of plate spacing of impingement air jet on flow and heat transfer characteristics. Mathematical and Computational Applications, 7(2): 171–179, August 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/2/171.

Kirmaci:2012:FAD

[KP12] S. Kırmacı and M. Pakdemirli. Flow analysis and determination of drag forces for spikes of crops. Mathematical and Computational Applications, 17 (1):83–91, April 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/1/83.

Kilic:2004:SAL

[KPS04] M. Anaç Kiliç, M. Pakdemirli, and B. Solmaz. Similarity analysis of long wave equations with variable bottom friction. Mathematical and Computational Applications, 9 (1):117–123, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/117.

Ketsripongsa:2018:IDE

[KPSS18] Udompong Ketsripongsa, Rapeepan Pitakaso, Kanchana Sethanan, and Tassin Srivarapongse. An improved differential evolution algorithm for crop planning in the northeastern region of Thailand. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/40.

Kurpa:2024:FVA

[KPSZ24] Lidiya Kurpa, Francesco Pellicano, Tetyana Shmatko, and Antonio Zippo. Free vibration analysis of porous functionally graded material plates with variable thickness on an elastic foundation using the r-functions method. Mathematical and Computational Applications, 29(1):??, February 2024. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/10.

Kirkpatrick:2020:MCB

[KPTM20] Anna Kirkpatrick, Kalen Patton, Prasad Tetali, and Cassie Mitchell. Markov chain-based sampling for exploring RNA secondary structure under the nearest neighbor thermodynamic model and extended applications. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/67.

Khama:2023:CMG

[KR23] Mopeli Khama and Quinn Reynolds. CFD modelling of gas-solid reactions: Analysis of iron and manganese oxides reduction with hydrogen. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/43.

Kakhki:2022:UHO

[KRD^{+}22] Elyas Mirzaee Kakhki, Jalil Rezaeepazhand, Fabian Duvigneau, Lotfollah Pahlavan, Resam Makvandi, Daniel Juhre, Majid Moavenian, and Sascha Eisenträger. On the use of high-order shape functions in the SAFE method and their performance in wave propagation problems. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/63.

Klein:2018:AOD

[KS18] Axel Klein and Walter Schumacher. Algebraic operations on deltasigma bit-streams. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/49.

Kuliev:2001:EGI

[KSK01] A. A. Kuliev, C. Selam, and A. Küçükbursa. The effect of the Galileo invariance pairing on the 1- states in spherical nuclei. Mathematical and Computational Applications, 6(2):103– 111, August 2001. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/103.

Kenar:2005:FFD

[KSK05] İsmail Kenar, Cevat Selam, and Atalay Küçükbursa. 0− → 0+ first forbidden β-decay matrix elements in spherical nuclei. Mathematical and Computational Applications, 10(2):179–184, August 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/179.

Kakuli:2023:CLS

[KSM23] Molahlehi Charles Kakuli, Winter Sinkala, and Phetogo Masemola. Conservation laws and symmetry reductions of the Hunter–Saxton equation via the double reduction method. Mathematical and Computational Applications, 28 (5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/92.

Korovina:2019:PAA

[KSS19] Maria Korovina, Ilya Smirnov, and Vladimir Smirnov. On a problem arising in application of the re-quantization method to construct asymptotics of solutions to linear differential equations with holomorphic coefficients at infinity. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/16.

Kayacan:2010:SBP

[KST10] Özhan Kayacan, Salim Sahin, and Filiz Tastan. A study for boronizing process within nonextensive thermostatistics. Mathematical and Computational Applications, 15(1):14–24, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/14.

Kaewman:2018:DEA

[KSTJ18] Sasitorn Kaewman, Tassin Srivarapongse, Chalermchat Theeraviriya, and Ganokgarn Jirasirilerd. Differential evolution algorithm for multilevel assignment problem: a case study in chicken transportation. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/55.

Kaya:1998:TDS

[KT98] Ali Kaya and Mehmet Tekelioğlu. Three dimensional stress anaysis in adhesively bonded joints. Mathematical and Computational Applications, 3(2):101–111, August 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/2/101.

Kaya:1999:TDS

[KTC99] Ali Kaya, Mehmet Tekelioğlu, and Muhammet Cerit. Three dimensional stress analysis in adhesively bonded joints with rivets. Mathematical and Computational Applications, 4(3): 195–203, December 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/195.

Kula:2012:SFR

[KTD12] Kamile Şanlı Kula, Fatih Tank, and Türkan Erbay Dalkılıç. A study on fuzzy robust regression and its application to insurance. Mathematical and Computational Applications, 17(3):223– 234, December 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/3/223.

Koc:1997:SMO

[KTK97] Ramazan Koç, Hayriye Tütüncüler, and Atalay Küçükbursa. Symmetry and molecular orbitals of C60. Mathematical and Computational Applications, 2(2): 91–100, August 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/2/91.

Kacar:2011:FVA

[KTK11] Alev Kacar, H. Tugba Tan, and Metin O. Kaya. Free vibration analysis of beams on variable Winkler elastic foundation by using the differential transform method. Mathematical and Computational Applications, 16(3):773–783, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/773.

Kiral:2013:FEM

[KTS13] Binnur Gören Kıral, Mustafa Tabanoğlu, and H. Tarık Serindağ. Finite element modeling of friction stir welding in aluminum alloys joint. Mathematical and Computational Applications, 18(2):122– 131, August 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/2/122.

Kucuk:2001:TSA

[Küç01] Mümin Küçük. Thermal stress analysis in the coated upper canine. Mathematical and Computational Applications, 6(1):23–35, April 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/23.

Kekec:2013:RTM

[KUE13] Taygun Kekec, Mustafa Unel, and Hakan Erdogan. Real-time multiple object tracking using virtual shells. Mathematical and Computational Applications, 18(3):373– 382, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/373.

Kawila:2018:BFA

[KUI18] Korawut Kawila, Chaiphon Udomsetchai, and Aiyared Iampan. Bipolar fuzzy UP-algebras.

Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/69.

Kumar:2007:AMM

[Kum07] Naresh Kumar. An analytical model of mountain wave for wind with shear. Mathematical and Computational Applications, 12(1):11–19, April 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/11.

Kenar:2008:SDR

[KÜM08] İsmail Kenar, Serdar Ünlü, and Ismail Maraş. The study of dependence of the radial parts of 0− → 0+ first forbidden β-decay matrix elements on the parameters in Woods–Saxon potential. Mathematical and Computational Applications, 13(1):1–8, April 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/1/1.

Kazaz:2005:CRI

[KUÖ05] M. Kazaz, H. H. Uğurlu, and A. Özdemir. The cosine rule II for a spherical triangle on the dual unit sphere. Mathematical and Computational Applications, 10(3):313– 320, December 2005. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/3/313.

Kurt:1998:RIC

[Kur98] M. Kurt. Refractive index change produced by the Electro-optical effect in LiTaO3. Mathematical and Computational Applications, 3 (1):17–26, April 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/1/17.

Kusno:2019:CRD

[Kus19]  Kusno.  Construction

of regular developable Bézier patches. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/4.

Kumar:2023:MGD

[KVR23] B. Rushi Kumar, R. Vijayakumar, and A. Jancy Rani. A model for the generalised dispersion of synovial fluids on nutritional transport with joint impacts of electric and magnetic field.

Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/3.

Khan:2010:NVA

[KWA^{+}10] Y. Khan, Q. Wu, H. Askari, Z. Saadatnia, and M. Kalami-Yazdi. Nonlinear vibration analysis of a rigid rod on a circular surface via Hamiltonian approach. Mathematical and Computational Applications, 15(5):974– 977, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/974.

Ke:2011:NCL

[KWP11] Jau-Chuan Ke, Chia-Huang Wu, and Wen Lea Pearn. Notes on the computation of Laplacestieljes transform for lognormal and Weibull. Mathematical and Computational Applications, 16(3):641–648, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/641.

Kent:1996:RCM

[KY96] Sedef Kent and Bingül Yazgan. Reconstruction with computerized microwave diffraction tomography by using circular measurement system in the far-field region. Mathematical and Computational Applications, 1(2):71–76, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/71.

Karacan:2005:STP

[KY05] Murat Kemal Karacan and Yusuf Yaylı. Special two parameter motion. Mathematical and Computational Applications, 10(1):27–34, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/27.

Koc:2013:CSA

[KY13] Ali Aydın Koç and Özgür Yeniay. A comparative study of artificial neural networks and logistic regression for classification of marketing campaign results. Mathematical and Computational Applications, 18(3):392– 398, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/392.

Khorramabadi:2020:CEE

[KYG20] Mehdi Karami Khorramabadi, Majid Yarahmadi, and Mojtaba Ghiyasi. Cost efficiency evaluation based on a data envelopment analysis approach by considering undesirable outputs on the basis of the semidisposability assumption. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/4.

Kasap:2002:SPR

[KYK02] Emin Kasap, Salim Yüce, and Nuri Kuruoğlu. Some properties of ruled surfaces under homothety in E3. Mathematical and Computational Applications, 7(3): 235–239, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/235.

Kurt:2013:FCM

[KYS13] Ayşe Kurt, Salih Yalçınbaş, and Mehmet Sezer. Fibonacci collocation method for solving linear differential– difference equations. Mathematical and Computational Applications, 18(3):448–458, December 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/448.

Kahraman:1998:NSM

[KZAN98] R. Kahraman, H. D. Zughbi, and Y. N. Al-Nassar. A numerical simulation of melting of ice heated from above. Mathematical and Computational Applications, 3(3):127–137, December 1998. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/3/127.

Loja:2020:PNS

[LB20] Maria Amélia R. Loja and Joaquim I. Barbosa. Preface to numerical and symbolic computation: Developments and applications — 2019. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/28.

Liu:2022:MSI

[LBY^{+}22] Xiangdong Liu, Yan Bai, Cunhui Yu, Hailong Yang, Haoning Gao, Jing Wang, Qing Chang, and Xiaodong Wen. Multistrategy improved sparrow search algorithm and application. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/96.

Li:2016:CCR

[LC16] Juncheng Li and Sheng Chen. The cubic α-Catmull–Rom spline. Mathematical and Computational Applications, 21(3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/33.

Li:2018:CVS

[LC18] Yansong Li and Shougen Chen. A complex variable solution for lining stress and deformation in a non-circular deep tunnel II practical application and verification. Mathematical and Computational Applications, 23(3):??, September 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/43.

Limmun:2018:CMR

[LCB18] Wanida Limmun, Boonorm Chomtee, and John J. Borkowski. The construction of a modelrobust IV-optimal mixture designs using a genetic algorithm. Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/25.

Liao:2017:HGB

[LCL17] Xiaolian Liao, Guohua Chen, and Shangzhao Li. 3-homogeneous groups and block-transitive 7– (v, k, 3) designs. Mathematical and Computational Applications, 22 (2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/2/35.

Lim:2022:CKK

[LCW^{+}22] Yeram Lim, Tamara Chambers, Christine Walck, Safeer Siddicky, Erin Mannen, and Victor Huayamave. Challenges in kinetic-kinematic driven musculoskeletal subjectspecific infant modeling. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/36.

Lopez:2021:OPG

[LDCST21] Esmeralda López, René F. Domínguez-Cruz, and Iván Salgado-Tránsito. Optimization of power generation grids: a case of study in Eastern Mexico. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/46.

Li:2010:NPS

[LDL10] Zitian Li, Zhengde Dai, and Jun Liu. New periodic solitary-wave solutions to the (3 + 1)- dimensional Kadomtsev– Petviashvili equation. Mathematical and Computational Applications, 15(5):877–882, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/877.

Lopez-Estrada:2020:RQT

[LESEVP^{+}20] Francisco-Ronay López-Estrada, Oscar Santos-Estudillo, Guillermo Valencia-Palomo, Samuel Gómez-Peñate, and Carlos Hernández-Gutiérrez. Robust qLPV tracking fault-tolerant control of a 3 DOF mechanical Crane. Mathematical and Computational Applications, 25(3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/48.

Lopez-Estrada:2021:AMC

[LEVP21] Francisco-Ronay López-Estrada and Guillermo Valencia-Palomo. Advanced mathematics and computational applications in control systems engineering. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/20.

Lissner:2019:DDM

[LF19] Julian Lißner and Felix Fritzen. Data-driven microstructure property relations. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/57.

Lopez:2018:HDU

[LGT^{+}18] Roberto López, Luis Carlos González Gurrola, Leonardo Trujillo, Olanda Prieto, Graciela Ramírez, Antonio Posada, Perla Juárez-Smith, and Leticia Méndez. How am I driving? using genetic programming to generate scoring functions for urban driving behavior. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/19.

Li:2010:FCT

[LH10a] Zheng-Biao Li and Ji-Huan He. Fractional complex transform for fractional differential equations. Mathematical and Computational Applications, 15(5):970– 973, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/970.

Liu:2010:SCP

[LH10b] Yang-Shuai Liu and Ji-Huan He. Soft compressible porous mat for “Flying” vehicles. Mathematical and Computational Applications, 15(5):967– 969, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/967.

Li:2010:ETW

[Li10]  Zheng-Biao Li.  Ex-

act three-wave solutions for the (3+1)- dimensional Boussinesq equation. Mathematical and Computational Applications, 15(5):953– 956, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/953.

Li:2016:CCT

[Li16] Juncheng Li. A class of cubic trigonometric automatic interpolation curves and surfaces with parameters. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/18.

Liu:2011:NMM

[Liu11]  Peide Liu.  A novel

method for multiple attribute decision-making of continuous random variable under risk with attribute weight unknown. Mathematical and Computational Applications, 16(2):340– 349, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/340.

Luo:2017:DOL

[LL17] Weilin Luo and Linqiang Lan. Design optimization of the lines of the bulbous bow of a hull based on parametric modeling and computational fluid dynamics calculation. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/4.

Lin:2023:DDA

[LL23] Haojie Lin and Xuyang Lou. Data-driven active learning control for bridge cranes. Mathematical and Computational Applications, 28 (5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/101.

Lin:2024:MCR

[LL24] Yang Lin and Jin Liang.

A multi-credit-rating migration model with asymmetric migration boundaries. Mathematical and Computational Applications, 29(1):??, February 2024. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/7.

Lin:2014:TIO

[LLL14] Tsung-Jui Chiang Lin, Meng-Rong Li, and Yong Shiuan Lee. Taiex index option model by using nonlinear differential equation. Mathematical and Computational Applications, 19 (1):78–92, April 2014.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/78.

Liu:2017:AIS

[LLTC17] Zhi Liu, Chen Li, Jieqing Tan, and Xiaoyan Chen. Analysis of inflection and singular points on a parametric curve with a shape factor. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/1/9.

Lotfi:2010:RDM

[LNBM10] F. Hosseinzadeh Lotfi, N. Nematollahi, M. H. Behzadi, and M. Mirbolouki. Ranking decision making units with stochastic data by using coefficient of variation. Mathematical and Computational Applications, 15(1):148– 155, April 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/1/148.

Loja:2022:PNS

[Loj22] Maria Amélia R. Loja. Preface to numerical and symbolic computation: Developments and applications — 2021. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/107.

Lara-Ortiz:2020:HSC

[LOSCO^{+}20] Vania Lara-Ortiz, Ivan Salgado, David Cruz-Ortiz, Alejandro Guarneros, Misael Magos-Sanchez, and Isaac Chairez. Hybrid state constraint adaptive disturbance rejection controller for a mobile worm bioinspired robot. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/13.

Liang:2017:NEA

[LP17] Aihua Liang and Yu Pang.

A novel, energy-aware task duplication-based scheduling algorithm of parallel tasks on clusters. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/2.

Licciardi:2020:VTU

[LPAD20] Silvia Licciardi, Rosa Maria Pidatella, Marcello Artioli, and Giuseppe Dattoli. Voigt transform and Umbral image. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/49.

Llaguno-Roque:2023:NCN

[LRBMAM^{+}23] José-Luis Llaguno-Roque, Rocio-Erandi Barrientos-Martínez, Héctor-Gabriel Acosta-Mesa, Tania Romo-González, and Efrén Mezura-Montes. Neuroevolution of convolutional neural networks for breast cancer diagnosis using western blot strips. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/72.

Laugksch:2023:PSM

[LRL23] Kristina Laugksch, Pieter Rousseau, and Ryno Laubscher. A PINN surrogate modeling methodology for steady-state integrated thermofluid systems modeling. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/52.

Lopez:2020:IKM

[LS20] Juan Carlos Cortés López and Marc Jornet Sanz. Improving kernel methods for density estimation in random differential equations problems. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/33.

Loya:2023:BCE

[LSARZ23] José Antonio Loya, Carlos Santiuste, Josué Aranda-Ruiz, and Ramón Zaera. Buckling of cracked Euler–Bernoulli columns embedded in a Winkler elastic medium. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/87.

Lang:2020:CAL

[LSB^{+}20] Konrad Lang, Sarah Stryeck, David Bodruzic, Manfred Stepponat, Slave Trajanoski, Ursula Winkler, and Stefanie Lindstaedt. Cy-Verse Austria — a local, collaborative cyberinfrastructure. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/38.

Liu:2017:MAE

[LTYG17] Fangtao Liu, Kwok Fai Geoffrey Tso, Yongheng Yang, and Jingjing Guan. Multilevel analysis of employee satisfaction on commitment to organizational culture: Case study of Chinese stateowned enterprises. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/4/46.

Llamoza:2019:SIR

[LV19] Johan Llamoza and Desiderio A. Vasquez. Structures and instabilities in reaction fronts separating fluids of different densities. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/51.

Ligthart:2023:HDF

[LV23] Philip Frederik Ligthart and Martin Philip Venter. A hierarchical design framework for the design of soft robots. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/47.

Lurie:2019:NAN

[LVBV19] Sergey A. Lurie, Dmitriy B.

Volkov-Bogorodsky, and Valery V. Vasiliev. A new approach to nonsingular plane cracks theory in gradient elasticity. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/93.

Laubscher:2022:EPA

[LVHL22] Ryno Laubscher, Johan Van Der Merwe, Philip G. Herbst, and Jacques Liebenberg. Estimation of pulmonary arterial pressure using simulated non-invasive measurements and gradientbased optimization techniques. Mathematical and Computational Applications, 27(5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/83.

Liu:2010:VIM

[LXZ10] Hongliang Liu, Aiguo Xiao, and Yongxiang Zhao. Variational iteration method for delay differential-algebraic equations. Mathematical and Computational Applications, 15(5):834– 839, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/834.

Li:2015:DIB

[LY15] Deng-Feng Li and Jie Yang. A difference-index based ranking method of trapezoidal intuitionistic fuzzy numbers and application to multiattribute decision making. Mathematical and Computational Applications, 20(1):25–38, April 2015. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/1/25.

Limei:2015:DBM

[LYYX15] Cao Limei, Sun Yina, Shen Yanan, and Si Xinhui. Dtm-bf method for the flow and heat transfer over a nonlinearly stretching sheet with nanofluids. Mathematical and Computational Applications, 20 (1):15–24, April 2015. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/1/15.

Li:2017:RQG

[LZ17] Tao Li and Liyuan Zhang. An M/G/1 retrial G-queue with general retrial times and working breakdowns. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/15.

Li:2021:TBO

[LZ21] Ruo Li and Wei Zhong.

Towards building the OP-mapped WENO schemes: a general methodology. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/67.

Liu:2022:DPC

[LZC22] Yongli Liu, Congcong Zhao, and Hao Chao. Density peak clustering based on relative density under progressive allocation strategy. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/84.

Li:2023:ECV

[LZL^{+}23] Meicong Li, Zheng Zhang, Yangyang Li, Qiang Zhao, Mei Huang, and Xiaoping Ouyang. Effect of chrome and vanadium on the behavior of hydrogen and helium in tungsten. Mathematical and Computational Applications, 28 (4):??, August 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/81.

Meric:1997:AHP

[MAE97] Cevdet Meriç, Enver Atik, and Turgut Engez. Application of the Hall-Petch relation to microhardness measurement on AA 1030, Cu, CuSn7, CuZn30 and 6114 alloys. Mathematical and Computational Applications, 2(2):71–77, August 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/2/71.

Moitsheki:2010:SRF

[MAH10] R. J. Moitsheki, S. Abelman, and T. Hayat. Symmetry reductions of a flow with power law fluid and contaminantmodified viscosity. Mathematical and Computational Applications, 15(4):685–696, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/685.

Makaraci:1999:MOO

[Mak99] Murat Makaracı. Multiobjective optimization for design and stability of a general control system. Mathematical and Computational Applications, 4(2): 153–159, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/153.

Matebese:2010:TDF

[MAKH10] B. T. Matebese, A. R.

Adem, C. M. Khalique, and T. Hayat. Twodimensional flow in a deformable channel with porous medium and variable magnetic field. Mathematical and Computational Applications, 15(4):674–684, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/674.

Manisa:2004:VMC

[Man04] K. Manisa. Variational Monte Carlo calculations of energy per particle nuclear matter. Mathematical and Computational Applications, 9(3): 485–494, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/485.

Maneechai:2016:VCS

[Man16] Sutitar Maneechai. Variant of constants in subgradient optimization method over planar 3-index assignment problems. Mathematical and Computational Applications, 21(1):??, March 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/1/4.

Martinez:2021:SAA

[MARL21] Isaac Gibert Martínez, Frederico Afonso, Simão Rodrigues, and Fernando Lau. A sequential approach for aerodynamic shape optimization with topology optimization of airfoils. Mathematical and Computational Applications, 26(2):??, June 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/34.

Mammadzada:2011:EAI

[MAS^{+}11] Sakhavat Mammadzada, Celal Artunç, Faruk Sen, Mehmet Ali Güngör, Uğur Tekin, and Erhan Çömlekoğlu. Effect of abutment and implant shapes on stresses in dental applications using fem. Mathematical and Computational Applications, 16(2):546– 555, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/546.

Mathkour:2010:NSA

[Mat10a] Hassan I. Mathkour. A new sorting algorithm for accelerating join-based queries. Mathematical and Computational Applications, 15(2):208– 217, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/208.

Mathkour:2010:REI

[Mat10b] Hassan I. Mathkour.

RULES3-EXT improvements on RULES-3 induction algorithm. Mathematical and Computational Applications, 15(3):318–324, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/318.

Matevossian:2019:MDS

[Mat19a] Hovik Matevossian. On the mixed Dirichlet– Steklov-type and Steklovtype biharmonic problems in weighted spaces. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/25.

Matevossian:2019:MBV

[Mat19b] Hovik A. Matevossian. Mixed boundary value problems for the elasticity system in exterior domains. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/58.

Mansour:2020:CQL

[MAT20] Zeinab Mansour and Maryam Al-Towailb. The complementary q-Lidstone interpolating polynomials and applications. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/34.

Mattila:2023:SHD

[Mat23] Pertti Mattila. A survey on the Hausdorff dimension of intersections.

Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/49.

Mesecan:2011:SSP

[MBA11] İbrahim Meşecan, İhsan Ömür Bucak, and Özcan Asilkan. Searching for the shortest path through group processing for TSP. Mathematical and Computational Applications, 16 (1):53–65, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/53.

Mathew:2020:MOL

[MC20a] Jismi Mathew and Christophe Chesneau. Marshall– Olkin length-biased Maxwell distribution and its applications. Mathematical and Computational Applications, 25(4):??, December 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/65.

Mathew:2020:SNC

[MC20b] Jismi Mathew and Christophe Chesneau. Some new contributions on the Marshall–Olkin length biased Lomax distribution: Theory, modelling and data analysis. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/79.

MEO2017:ARM

[MCA17] MCA Editorial Office. Acknowledgement to reviewers of MCA in 2016. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/6.

MEO2018:ARM

[MCA18] MCA Editorial Office. Acknowledgement to reviewers of MCA in 2017. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/3.

MEO2019:ARM

[MCA19] MCA Editorial Office. Acknowledgement to reviewers of MCA in 2018. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/11.

MEO2020:ARM

[MCA20] MCA Editorial Office.

Acknowledgement to reviewers of MCA in 2019. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/6.

MEO2021:ARM

[MCA21] MCA Editorial Office.

Acknowledgment to reviewers of MCA in 2020. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/10.

MEO2022:ARM

[MCA22] MCA Editorial Office.

Acknowledgment to reviewers of MCA in 2021. Mathematical and Computational Applications, 27(1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/13.

MEO2023:ARM

[MCA23] MCA Editorial Office.

Acknowledgment to the reviewers of MCA in 2022. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/15.

Monteiro:2019:FMI

[MCB19] Patrícia Monteiro, Aldina Correia, and Vítor Braga. Factors for marketing innovation in Portuguese firms CIS 2014. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/99.

Makhanya:2023:UCF

[MCBM23] Khanyisani Mhlangano Makhanya, Simon Connell, Muaaz Bhamjee, and Neil Martinson. The use of computational fluid dynamics for assessing flow-induced acoustics to diagnose lung conditions. Mathematical and Computational Applications, 28(3):??, June 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/64.

Muse:2022:FPA

[MCNM22] Abdisalam Hassan Muse, Christophe Chesneau, Oscar Ngesa, and Samuel Mwalili. Flexible parametric accelerated hazard model: Simulation and application to censored lifetime data with crossing survival curves. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/104.

Mohyud-Din:2010:CCP

[MDNN10a] Syed Tauseef Mohyud-Din, Muhammad Aslam Noor, and Khalida Inayat Noor. Comparison and coupling of polynomials for Flierl– Petviashivili equation. Mathematical and Computational Applications, 15(2):187–198, August 2010. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/15/2/187.

Mohyud-Din:2010:MVI

[MDNN10b] Syed Tauseef Mohyud-Din, Muhammad Aslam Noor, and Khalida Inayat Noor. Modified variational iteration method for Schrodinger equations. Mathematical and Computational Applications, 15(3):309–317, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/309.

Mizrak:2012:AEM

[ME12] Mustafa Mizrak and Abdulkadir Ertaş. Application of (G′/G)-expansion method to the compound KdV–Burgers-type equations. Mathematical and Computational Applications, 17(1):18–28, April 2012. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/1/18.

Mechkour:2022:MPP

[Mec22] Houari Mechkour. Modeling of perforated piezoelectric plates. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/100.

Macias-Escobar:2021:IRS

[MECRMT^{+}21] Teodoro Macias-Escobar, Laura Cruz-Reyes, César Medina-Trejo, Claudia Gómez-Santillán, Nelson Rangel-Valdez, and Héctor Fraire-Huacuja. An interactive recommendation system for decision making based on the characterization of cognitive tasks. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/35.

Merdan:2009:HPM

[Mer09] Mehmet Merdan. Homotopy perturbation method for solving human T-cell lymphotropic virus I(HTLV-I) infection of Cd^{4+} T-cells model. Mathematical and Computational Applications, 14(2):85–96, August 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/2/85.

Mese:1996:MPE

[Mes96] Erkan Mese. Minimum permeance estimation of variable reluctance machines by using neural networks. Mathematical and Computational Applications, 1(2): 87–92, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/87.

Mortari:2021:UTF

[MF21] Daniele Mortari and Roberto Furfaro. Univariate theory of functional connections applied to component constraints. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/9.

Moghadam:2020:NMO

[MFF20] Majid Ebrahimi Moghadam, Hamid Falaghi, and Mahdi Farhadi. A novel method of optimal capacitor placement in the presence of harmonics for power distribution network using NSGA-II multi-objective genetic optimization algorithm. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/17.

Merdan:2011:SNO

[MG11a] Mehmet Merdan and Ahmet Gökdoğan. Solution of nonlinear oscillators with fractional nonlinearities by using the modified differential transformation method. Mathematical and Computational Applications, 16(3):761–772, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/761.

Misirli:2011:EFM

[MG11b] Emine Misirli and Yusuf Gurefe. Exp-function method for solving nonlinear evolution equations. Mathematical and Computational Applications, 16(1):258– 266, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/258.

Martinez-Galicia:2020:WSS

[MGGHCR^{+}20] David Martínez-Galicia, Alejandro Guerra-Hernández, Nicandro Cruz-Ramírez, Xavier Limón, and Francisco Grimaldo. Windowing as a sub-sampling method for distributed data mining. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/39.

Maras:2010:IEE

[MGT10] Ismail Maras, Ramazan Gümüs, and Nurettin Türkan. The investigation of even-even^{114−120} Xe isotopes by the framework of IBA. Mathematical and Computational Applications, 15 (1):79–88, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/79.

Mi:2021:SWV

[MHC21] Zichuan Mi, Saddam Hussain, and Christophe Chesneau. On a special weighted version of the odd Weibullgenerated class of distributions. Mathematical and Computational Applications, 26(3):??, September 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/3/62.

Mitic:2021:FAP

[Mit21a] Peter Mitic. A framework for analysis and prediction of operational risk stress. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/1/19.

Mitic:2021:ORR

[Mit21b] Peter Mitic. Operational risk reverse stress testing: Optimal solutions. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/2/38.

Moleleki:2012:SRE

[MJK12] L. D. Moleleki, A. G.

Johnpillai, and C. M. Khalique. Symmetry reductions and exact solutions of a variable coefficient (2 + 1)-Zakharov–Kuznetsov equation. Mathematical and Computational Applications, 17(2):132– 139, August 2012. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/2/132.

Mamis:1996:SRT

[MK96] M. Salih. Mamiş and Muhammet Köksal. Some renovations in transient analysis of transmission lines by state-space techniques. Mathematical and Computational Applications, 1(1):181–190, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/181.

Mamis:1997:CTL

[MK97] M. Salih Mamiş and Muhammet Köksal. Computation of transmission line transients by using fast inverse Laplace transform. Mathematical and Computational Applications, 2(2):61– 69, August 1997. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/2/61.

Muatjetjeja:2010:NPN

[MK10] Ben Muatjetjeja and Chaudry Masood Khalique. Noether, partial Noether operators and first integrals for the coupled Lane–Emden system. Mathematical and Computational Applications, 15(3):325–333, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/325.

Meric:1997:IDD

[MKK97] Cevdet Meriç, N. Sinan Köksal, and Bekir Karlık. An investigation of deep drawing of low carbon steel sheets and applications in artificial neural networks. Mathematical and Computational Applications, 2(3): 119–125, December 1997. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/2/3/119.

Manisa:2011:NER

[MKMB11] Fatma Manisa, Atalay Küçükbursa, Kaan Manisa, and Tahsin Babacan. A new expression for radial distribution function of nuclear matter. Mathematical and Computational Applications, 16(2):414– 424, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/414.

Mota:2020:PFG

[MLBR20] Ana F. Mota, Maria Amélia R. Loja, Joaquim I. Barbosa, and José A. Rodrigues. Porous functionally graded plates: an assessment of the influence of shear correction factor on static behavior. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/25.

Liu:2017:IIS

[mLkWbPxX17] Cheng ming Liu, Ze kun Wang, Hai bo Pang, and Jun xiao Xue. Image interpolation via scanning line algorithm and discontinuous B-spline. Mathematical and Computational Applications, 22(2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/2/34.

Molati:2010:GCC

[MM10a] M. Molati and F. M. Mahomed. Group classification of coupled diffusion system with applications in soil science. Mathematical and Computational Applications, 15(4):697–708, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/697.

Momoniat:2010:SRN

[MM10b] E. Momoniat and F. M. Mahomed. Symmetry reduction and numerical solution of a thirdorder ODE from thin film flow. Mathematical and Computational Applications, 15(4):709– 719, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/709.

Moremedi:2010:SDW

[MM10c] G. M. Moremedi and D. P. Mason. Streamlines and detached wakes in steady flow past a spherical liquid drop. Mathematical and Computational Applications, 15(4):543–557, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/543.

Mureithi:2010:LNS

[MM10d] E. W. Mureithi and D. P. Mason. Local nonsimilarity solutions for a forced-free boundary layer flow with viscous dissipation. Mathematical and Computational Applications, 15(4):558– 573, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/558.

Madhi:2017:ICO

[MM17] Mahdi Madhi and Norizan Mohamed. An initial condition optimization approach for improving the prediction precision of a GM(1,1) model. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/21.

Mendonca:2018:MFA

[MMF18] Gabriel A. Mendonça, Thales A. C. Maia, and Braz J. Cardoso Filho. Magnetic field analytical solution for nonhomogeneous permeability in retaining sleeve of a high-speed permanentmagnet machine. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/72.

Maposa:2023:QFB

[MMM23] Daniel Maposa, Amon Masache, and Precious Mdlongwa. A quantile functions-based investigation on the characteristics of southern African solar irradiation data. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/86.

Mkhatshwa:2019:OMD

[MMS19] Musawenkhosi Mkhatshwa, Sandile Motsa, and Precious Sibanda. Overlapping multi-domain spectral method for conjugate problems of conduction and MHD free convection flow of nanofluids over flat plates. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/75.

Mecitoglu:1996:DBS

[MÖ96] Zahit Mecitoğlu and Hasan Örenel. Dynamic behavior of a stiffened laminated conical Shell under hygrothermal loads. Mathematical and Computational Applications, 1(1):73– 84, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/73.

Manisa:2004:TCN

[MO04] K. Manisa and R. Oğul.

Transport coefficients of nuclear matter at high temperatures. Mathematical and Computational Applications, 9 (1):19–24, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/19.

Moumen:2021:WPS

[MOBA21] Abdelkader Moumen, Djamel Ouchenane, Keltoum Bouhali, and Yousif Altayeb. Well-posedness and stability results for a nonlinear damped porous–elastic system with infinite memory and distributed delay terms. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/71.

Moldenhauer:2018:IDF

[Mol18] Herbert Moldenhauer.

Integration of direction fields with standard options in finite element programs. Mathematical and Computational Applications, 23(2):??, June 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/24.

Mortari:2022:UTF

[Mor22] Daniele Mortari. Using the theory of functional connections to solve boundary value geodesic problems. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/4/64.

Moschandreou:2018:NAP

[Mos18] Terry E. Moschandreou. A new analytical procedure to solve two phase flow in tubes. Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/26.

Masroor:2017:ETM

[MPA17] Suhaib Masroor, Chen Peng, and Zain Anwar Ali. Event triggered multi-agent consensus of DC motors to regulate speed by LQR scheme. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/14.

Machado:2022:GTA

[MPA^{+}22] Pedro Machado, Sofia J.

Pinheiro, Vera Afreixo, Cristiana J. Silva, and Rui Leitão. Graph theory approach to COVID-19 transmission by municipalities and age groups. Mathematical and Computational Applications, 27(5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/86.

Mortoja:2024:RRT

[MPP^{+}24] Sk Golam Mortoja, Ayan Paul, Prabir Panja, Sabyasachi Bhattacharya, and Shyamal Kumar Mondal. Role reversals in a tri-trophic prey– predator interaction system: a model-based study using deterministic and stochastic approaches. Mathematical and Computational Applications, 29(1):??, February 2024. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/3.

Malinzi:2018:ESN

[MQ18] Joseph Malinzi and Paul Amarh Quaye. Exact solutions of nonlinear evolution models in physics and biosciences using the hyperbolic tangent method. Mathematical and Computational Applications, 23(3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/35.

Mahmoud:2005:SAC

[MR05] M. A. Mahmoud and A. A. Rasha. Stochastic approximation with compound delayed observations. Mathematical and Computational Applications, 10(2):283– 289, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/283.

Matos:2019:AEC

[MR19a] José M. A. Matos and Maria João Rodrigues. Almost exact computation of eigenvalues in approximate differential problems. Mathematical and Computational Applications, 24(4):??, December 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/96.

Mensah:2019:LRG

[MR19b] Emmanuel Kwasi Mensah and Matteo Rocca. Light robust goal programming. Mathematical and Computational Applications, 24(4):??, December 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/85.

Murillo:2023:TSL

[MRG23] Carlos Enrique Valencia Murillo, Miguel Ernesto Gutierrez Rivera, and Luis David Celaya Garcia. Thermal– structural linear static analysis of functionally graded beams using Reddy beam theory. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/84.

Majumder:2020:CNN

[MRIR20] Md. Taksir Hasan Majumder, Md. Mahabur Rahman, Anindya Iqbal, and M. Sohel Rahman. Convolutional neural network based ensemble approach for homoglyph recognition. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/71.

Molati:2005:SRE

[MS05] M. Molati and C. Wafo Soh. Similarity reduction of energy-transport models for semiconductors. Mathematical and Computational Applications, 10(2):221–230, August 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/2/221.

Melin:2023:CSM

[MSPC23] Patricia Melin, Daniela Sánchez, Martha Pulido, and Oscar Castillo. Comparative study of metaheuristic optimization of convolutional neural networks applied to face mask classification. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/6/107.

Mekahlia:2020:ROM

[MSSZ20] Abdelhak Mekahlia, Eric Semail, Franck Scuiller, and Hussein Zahr. Reducedorder model of rotor cage in multiphase induction machines: Application on the prediction of torque pulsations.

Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/11.

Mubaraki:2023:ACR

[Mub23] Ali M. Mubaraki. Asymptotic consideration of Rayleigh waves on a coated orthorhombic elastic half-space reinforced using an elastic Winkler foundation. Mathematical and Computational Applications, 28 (6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/109.

Mucuk:2000:MPU

[Muc00] Osman Mucuk. Monodromy principle and universal coverings. Mathematical and Computational Applications, 5(2): 119–127, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/119.

Mucuk:2001:TAM

[Muc01] Osman Mucuk. A topological application of the monodromy groupoid on principal bundles. Mathematical and Computational Applications, 6(3): 195–202, December 2001.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/6/3/195.

Mustafa:2016:GPN

[Mus16] Nizami Mustafa. Geometric properties of normalized Wright functions. Mathematical and Computational Applications, 21(2):??, June 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/14.

Mutlu:1998:PCU

[Mut98] Ali Mutlu. Peiffer commutators by using GAP package. Mathematical and Computational Applications, 3(1):59– 65, April 1998. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/1/59.

Mutlu:2000:FCC

[Mut00a]  Ali Mutlu.  Free 2-

crossed complexes of simplicial algebras. Mathematical and Computational Applications, 5 (1):13–22, April 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/1/13.

Mutlu:2000:JAS

[Mut00b] Ali Mutlu. Join for (augmented) simplicial group. Mathematical and Computational Applications, 5(2):105–112, August 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/105.

Mutlu:2004:FCT

[Mut04] Ali Mutlu. Freeness conditions of tensors and coproducts of groups. Mathematical and Computational Applications, 9(3):417–426, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/417.

Matos:2021:EFP

[MV21] José A. O. Matos and Paulo B. Vasconcelos. Effectiveness of floatingpoint precision on the numerical approximation by spectral methods. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/42.

Merupula:2023:PIC

[MVC23] Jahnavi Merupula, V. S.

Vaidyanathan, and Christophe Chesneau. Prediction interval for compound Conway–Maxwell–Poisson regression model with application to vehicle insurance claim data. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/39.

Meng:2016:NSS

[MWM16] Junqing Meng, Qingen Wei, and Yechao Ma. Numerical simulation study on the flow field out of a submerged abrasive water jet nozzle. Mathematical and Computational Applications, 21 (1):??, March 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/1/2.

Muchie:2022:SAE

[MWM22] Kindie Fentahun Muchie, Anthony Kibira Wanjoya, and Samuel Musili Mwalili. Small area estimation of zone-level Malnutrition among children under five in Ethiopia. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/44.

Nguyen:2020:BDS

[NAA20] Minh Nguyen, Mehmet Aktas, and Esra Akbas. Bot detection on social networks using persistent homology. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/58.

Nasiru:2022:NLD

[NAC22] Suleman Nasiru, Abdul Ghaniyyu Abubakari, and Christophe Chesneau. New lifetime distribution for modeling data on the unit interval: Properties, applications and quantile regression. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/105.

Nasiru:2023:APD

[NAC23] Suleman Nasiru, Abdul Ghaniyyu Abubakari, and Christophe Chesneau. The arctan power distribution: Properties, quantile and modal regressions with applications to biomedical data. Mathematical and Computational Applications, 28(1):??, February 2023.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/25.

Naderian:2022:IFS

[NCDM22] Hamidreza Naderian, Moe M. S. Cheung, Elena Dragomirescu, and Abdolmajid Mohammadian. Integrated finite strip computation for modelling and frequency analysis of hybrid laminated FRP structures. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/47.

Nebioglu:2014:PDB

[NÇK14] Mehmet Nebioğlu, Tarık Çakar, and Raşit Köker. A proposal for a distinction in belief functions I. Mathematical and Computational Applications, 19(1):12–20, April 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/12.

Nakamura:2019:EES

[NCM19] Gilberto M. Nakamura, Brenno Cabella, and Alexandre S. Martinez. Exact evaluation of statistical moments in superradiant emission.

Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/66.

Nebro:2022:NIR

[NGBLC22] Antonio J. Nebro, Jesús Galeano-Brajones, Francisco Luna, and Carlos A. Coello Coello. Is NSGA-II ready for largescale multi-objective optimization? Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/103.

Nazir:2018:IIC

[NISC18] Babar Nazir, Faiza Ishaq, Shahaboddin Shamshirband, and Anthony T. Chronopoulos. The impact of the implementation cost of replication in data grid job scheduling. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/28.

Narain:2010:RVP

[NK10] R. Narain and A. H.

Kara. On the redefinition of the variational and ‘partial’ variational conservation laws in a class of nonlinear PDEs with mixed derivatives. Mathematical and Computational Applications, 15(4):732–741, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/732.

Noroozi:2019:BSI

[NKI19] Ali A. Noroozi, Jaber Karimpour, and Ayaz Isazadeh. Bisimulation for secure information flow analysis of multi-threaded programs. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/64.

Narayanaswamy:2022:GMS

[NKS22a] Manoj Kumar Narayanaswamy, Jagan Kandasamy, and Sivasankaran Sivanandam. Go-MoS_{2}/water flow over a shrinking cylinder with Stefan blowing, Joule heating, and thermal radiation. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/110.

Narayanaswamy:2022:ISB

[NKS22b] Manoj Kumar Narayanaswamy, Jagan Kandasamy, and Sivasankaran Sivanandam. Impacts of Stefan blowing on hybrid nanofluid flow over a stretching cylinder with thermal radiation and Dufour and Soret effect. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/91.

Nikabadze:2019:DWV

[NLMU19] Mikhail U. Nikabadze, Sergey A. Lurie, Hovik A. Matevossian, and Armine R. Ulukhanyan. On determination of wave velocities through the eigenvalues of material objects. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/39.

Naeem:2010:AFI

[NM10] I. Naeem and F. M.

Mahomed. Approximate first integrals for a system of two coupled Van der Pol oscillators with linear diffusive coupling. Mathematical and Computational Applications, 15(4):720– 731, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/720.

Nakamura:2019:FSA

[NMCM19] Gilberto M. Nakamura, Ana Carolina P. Monteiro, George C. Cardoso, and Alexandre S. Martinez. Finite symmetries in agent-based epidemic models. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/44.

Naz:2010:CLC

[NNM10] R. Naz, I. Naeem, and F. M. Mahomed. Conservation laws and conserved quantities for laminar radial jets with swirl. Mathematical and Computational Applications, 15(4):742–761, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/4/742.

Noor:2014:DMS

[NNW14] Muhammad Aslam Noor, Khalida Inayat Noor, and Asif Waheed. Decomposition method for solving a system of thirdorder boundary value problems. Mathematical and Computational Applications, 19(3):230– 240, December 2014. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/3/230.

Nguyen:2022:SAF

[NSCTPW22] Quoc Khanh Nguyen, Stefano Serra-Capizzano, Cristina Tablino-Possio, and Eddie Wadbro. Spectral analysis of the finite element matrices approximating 3D linearly elastic structures and multigrid proposals. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/78.

Naredo:2023:CAL

[NSG^{+}23] Enrique Naredo, Candelaria Sansores, Flaviano Godinez, Francisco López, Paulo Urbano, Leonardo Trujillo, and Conor Ryan. Comprehensive analysis of learning cases in an autonomous navigation task for the evolution of general controllers. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/35.

Nagy:2021:NSE

[NSO^{+}21] Ádám Nagy, Mahmoud Saleh, Issa Omle, Humam Kareem, and Endre Kovács. New stable, explicit, shiftedhopscotch algorithms for the heat equation. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/3/61.

Nalli:2004:GRL

[NT04] Ayşe Nallı and Dursun Taşci. The GCDreciprocal LCM matrices on GCD-closed sets. Mathematical and Computational Applications, 9(1):101–106, April 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/1/101.

Nikabadze:2019:SAE

[NU19] Mikhail Nikabadze and Armine Ulukhanyan. Some applications of eigenvalue problems for tensor and tensor–block matrices for mathematical modeling of micropolar thin bodies. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/33.

Nikabadze:2019:MFL

[NUM^{+}19] Mikhail U. Nikabadze, Armine R. Ulukhanyan, Tamar Moseshvili, Ketevan Tskhakaia, Nodar Mardaleishvili, and Zurab Arkania. On the modeling of five-layer thin prismatic bodies. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/69.

Nuriyev:2003:CGO

[Nur03] Urfat G. Nuriyev. On complexity of a global optimization problem. Mathematical and Computational Applications, 8(1):27–34, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/27.

Nagarjuna:2022:NHL

[NVC22] Vasili B. V. Nagarjuna, Rudravaram Vishnu Vardhan, and Christophe Chesneau. Nadarajah– Haghighi Lomax distribution and its applications. Mathematical and Computational Applications, 27(2):??, April 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/30.

Ngin:2017:ISD

[NYH^{+}17] Kimlong Ngin, Yen Yat, Rufai O. Hafeez, Sijjad Ali Khuhro, and Mariia Sushchenko. Intelligent system for data tracking in image editing company. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/4/45.

Ozturk:2009:VDN

[OA09] Mualla Ozturk and M. Cihan Aydin. Verification of a 3-D numerical model for spillway aerator. Mathematical and Computational Applications, 14(1):21– 30, April 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/21.

Okcu:2013:EED

[OA13] Pinar Okcu and Gabil M.

Amiraliyev. Error estimates for differential difference schemes to pseudo-parabolic initialboundary value problem with delay. Mathematical and Computational Applications, 18(3):283– 292, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/283.

Ozturk:2017:TSC

[ÖB17] Taha Yasin Öztürk and Sadi Bayramov. Topology on soft continuous function spaces. Mathematical and Computational Applications, 22 (2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/2/32.

Ozturk:2005:DAS

[OC05] Ali Ugur Ozturk and Hikmet H. Catal. Dynamic analysis of semirigid frames. Mathematical and Computational Applications, 10 (1):1–8, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/1.

Ozkurt:2009:ATD

[OC09] Celil Ozkurt and Fatih Camci. Automatic traffic density estimation and vehicle classification for traffic surveillance systems using neural networks. Mathematical and Computational Applications, 14(3):187–196, December 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/3/187.

Oz:2006:PVC

[ÖD06] H. R. Öz and M. T. Daş. In-plane vibrations of circular curved beams with a transverse open crack. Mathematical and Computational Applications, 11(1):1–10, April 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/1/1.

Ouamara:2019:PME

[OD19] Daoud Ouamara and Frédéric Dubas. Permanentmagnet eddy-current losses: a global revision of calculation and analysis. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/3/67.

Ozturk:2003:SSI

[ÖE03] Savaş Öztürk and E. Murat Esin. Simulation of swarm intelligence and possible applications in engineering. Mathematical and Computational Applications, 8(3):361– 368, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/361.

Ouakka:2019:TMU

[OF19] Slimane Ouakka and Nicholas Fantuzzi. Trustworthiness in modeling unreinforced and reinforced t-joints with finite elements. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/27.

Oliveira:2021:AEA

[OFCP21] Pedro N. Oliveira, Elza M. M. Fonseca, Raul D. S. G. Campilho, and Paulo A. G. Piloto. Analytical equations applied to the study of steel profiles under fire according to different nominal temperature-time curves. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/48.

Ozer:2003:AGM

[ÖG03] Halil Özer and Durmuş Günay. Aspects of the Gurson model and its applications. Mathematical and Computational Applications, 8(3): 327–334, December 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/3/327.

Orcen:2013:EPS

[ÖG13] Gurbet Örçen and Mustafa Gür. Elasto-plastic stress analysis in laminated thermoplastic composite plates with an elliptic hole. Mathematical and Computational Applications, 18(1):38– 49, April 2013. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/1/38.

Ozdemir:2005:HSC

[ÖK05] A. Özdemir and M. Kazaz. Hyperbolic sine and cosine rules for geodesic triangles on the hyperbolic unit sphere H_{0}^{2}. Mathematical and Computational Applications, 10(2):203–209, August 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/2/203.

Ozbay:2003:WBD

[ÖKK03] Yüksel Özbay, Bekir Karlik, and A. Resit Kavsaoglu. A Windowsbased digital filter design. Mathematical and Computational Applications, 8(3):287–294, December 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/287.

Okumus:2011:EPS

[Oku11a] Fuat Okumuş. Elastoplastic stress analysis in silicon carbide fiber reinforced magnesium metal matrix composite with a square hole. Mathematical and Computational Applications, 16 (1):216–225, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/216.

Okumus:2011:ECS

[Oku11b] Fuat Okumuş. Experimental and comparative study of different fibers on the strength and failure of quasi-isotropic plastic composite with a hole. Mathematical and Computational Applications, 16(2):404– 413, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/404.

Okumus:2011:ICL

[Oku11c] Fuat Okumuş. Investigation of coating layer to reduce thermal stresses in steel fiber reinforced aluminum metal matrix composite. Mathematical and Computational Applications, 16(1):226– 235, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/226.

Okumus:2011:IGM

[Oku11d] Fuat Okumuş. Investigation of geometric and material effects on critical load of thermoplasticmatrix composite plates with a square hole under in-plane loadings. Mathematical and Computational Applications, 16(3):576–587, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/576.

Oumellal:2021:CSC

[OL21] Fatima Oumellal and Abdellah Lamnii. Curve and surface construction using Hermite trigonometric interpolant. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/11.

Olejnik:2020:CMI

[Ole20] Paweł Olejnik. Computational methods in interdisciplinary applications of nonlinear dynamics. Mathematical and Computational Applications, 25(4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/62.

Ota:2020:EPM

[OM20] Yasushi Ota and Naoki Mizutani. Estimating parameters in mathematical model for societal booms through Bayesian inference approach. Mathematical and Computational Applications, 25(3):??, September 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/42.

Oke:2018:OCA

[OMX18] Segun Isaac Oke, Maba Boniface Matadi, and Sibusiso Southwell Xulu. Optimal control analysis of a mathematical model for breast cancer. Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/21.

Oz:2005:NFB

[ÖÖ05] H. R. Öz and E. Özkaya. Natural frequencies of beam-mass systems in transverse motion for different end conditions. Mathematical and Computational Applications, 10(3):369–376, December 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/3/369.

Olugbade:2022:RAI

[OOI^{+}22] Samuel Olugbade, Stephen Ojo, Agbotiname Lucky Imoize, Joseph Isabona, and Mathew O. Alaba. A review of artificial intelligence and machine learning for incident detectors in road transport systems. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/5/77.

Ozyigit:2004:LFP

[ÖÖT04] H. A. Özyiğit, H. R. Öz, and M. Tekelioğlu. Linear forced in-plane and out-of plane vibrations of frames having a curved member. Mathematical and Computational Applications, 9(3): 371–380, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/371.

Orhan:2011:AES

[OPUÖ11] Zeynep Orhan, İlknur Pehlivan, Volkan Uslan, and Pınar Önder. Automated extraction of semantic word relations in Turkish lexicon. Mathematical and Computational Applications, 16 (1):13–21, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/13.

Olivier:2019:MTT

[ORC19] Clément Olivier, David Ryckelynck, and Julien Cortial. Multiple tensor train approximation of parametric constitutive equations in elastoviscoplasticity. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/17.

Ozturk:2005:ISR

[ÖS05] Ali Uğur Öztürk and Mutlu Seçer. An investigation for semi-rigid frames by different connection models. Mathematical and Computational Applications, 10 (1):35–44, April 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/35.

Oguntala:2018:AAA

[OSAAA18] George Oguntala, Gbeminiyi Sobamowo, Yinusa Ahmed, and Raed Abd-Alhameed. Application of approximate analytical technique using the homotopy perturbation method to study the inclination effect on the thermal behavior of porous fin heat sink. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/62.

Ozalp:2002:HIN

[ÖT02] Nuri Özalp and Fatma Taşdelen. The Harnack inequalities for nonnegative solutions of an elliptic and ultra-hyperbolic equation. Mathematical and Computational Applications, 7(3):229–234, December 2002. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/229.

Okuducu:1999:SHN

[OTŞ99]  Şeref Okuducu, Güneş

Tanır, and Başar Şarer. The sensitivity of halo nuclei to the nuclear radius. Mathematical and Computational Applications, 4(2):107–111, August 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/107.

Ouimet:2023:MNM

[Oui23] Frédéric Ouimet. Moments of the negative multinomial distribution. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/85.

Overmars:2018:EGR

[OV18] Anthony Overmars and Sitalakshmi Venkatraman. An efficient Golden ratio method for secure cryptographic applications. Mathematical and Computational Applications, 23(4):??, December 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/58.

Overmars:2019:FFS

[OV19] Anthony Overmars and Sitalakshmi Venkatraman. A fast factorisation of semi-primes using sum of squares. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/62.

Overmars:2020:MAR

[OV20] Anthony Overmars and Sitalakshmi Venkatraman. Mathematical attack of RSA by extending the sum of squares of primes to factorize a semi-prime. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/63.

Opoku:2023:SAC

[OWO^{+}23] Michael O. Opoku, Eric N. Wiah, Eric Okyere, Albert L. Sackitey, Emmanuel K. Essel, and Stephen E. Moore. Stability analysis of Caputo fractional order viral dynamics of hepatitis B cellular infection. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/24.

Ozyilmaz:2001:IIT

[ÖY01] Emin Özyılmaz and Yusuf Yaylı. On the integral invariants of a timelike ruled surface. Mathematical and Computational Applications, 6(2): 137–145, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/137.

Okur:2011:TSA

[OYCS11] Ahmet Okur, Recep Yigit, Erdal Celik, and Onur Sayman. Thermal stress analysis in ZrO2 insulation coatings on Cr-Ni substrates during cooling process. Mathematical and Computational Applications, 16(3):598–604, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/598.

Oz:2000:CNF

[Öz00a]  H. R. Öz.  Calcula-

tion of the natural frequencies of a beam–mass system using finite element method. Mathematical and Computational Applications, 5 (2):67–76, August 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/2/67.

Oz:2000:NTV

[Öz00b] H. R. Öz. On the nonlinear transverse vibrations and stability of an axially accelerating beam. Mathematical and Computational Applications, 5(3): 157–167, December 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/3/157.

Ozbay:2002:EPS

[Özb02] Mahmut Özbay. Elasticplastic stress analysis and residual stresses in metal matrix laminated plates under inplane loading. Mathematical and Computational Applications, 7 (1):23–32, April 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/1/23.

Ozcan:2000:EPS

[Özc00] R. Özcan. Elastic-plastic stress analysis in thermoplastic composite cantilever loaded by a single force. Mathematical and Computational Applications, 5(2):77– 89, August 2000. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/2/77.

Ozdamar:1996:BHB

[Özd96a] Ing. Aydoğan Özdamar.

Buckling of heavy bars. Mathematical and Computational Applications, 1(2):93–99, December 1996. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/1/2/93.

Ozdemir:1996:PGM

[Özd96b] Ahmet Ş. Özdemir. On the partial generalization of the measure of transcendence of some formal Laurent series. Mathematical and Computational Applications, 1(2): 100–110, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/100.

Ozdamar:1997:DEH

[Özd97] Aydoğan Özdamar. The development of Euler hyperbola in heavy bars. Mathematical and Computational Applications, 2(3):127–130, December 1997. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/2/3/127.

Ozer:2000:SBP

[Öze00]  Teoman Özer.  Solu-

tion of Boussinesq problem using lie symmetries. Mathematical and Computational Applications, 5(1):1–11, April 2000. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/1/1.

Ozer:2002:ATC

[Öze02] Mahmut Özer. Analysis of time-course of the recovery from inactivation of ionic currents in cerebellar punkinje cell. Mathematical and Computational Applications, 7(1):53–63, April 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/1/53.

Ozer:2003:PPA

[Oze03] Mahmut Ozer. Phaseplane analysis for a simplified model of Purkinje cell dendrite. Mathematical and Computational Applications, 8 (1):71–78, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/71.

Ozel:2010:NDM

[Öze10] Mustafa Özel. A new decomposition method for solving system of nonlinear equations. Mathematical and Computational Applications, 15 (1):89–95, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/89.

Ozgan:2010:MSA

[Özg10]  Ercan Özgan.  Mod-

elling the stability of asphalt concrete with fuzzy logic and statistical methods for various freezing and thawing cycles. Mathematical and Computational Applications, 15(2):176– 186, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/176.

Ozkaya:2001:LTV

[Özk01] E. Özkaya. Linear transverse vibrations of a simply supported beam carrying concentrated masses. Mathematical and Computational Applications, 6(2):147–152, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/147.

Ozyilmaz:1998:TCF

[Özy98]  E. Özyılmaz.  The

tangent conoids family which depends on the ruled surface. Mathematical and Computational Applications, 3 (1):49–57, April 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/1/49.

Ozyigit:2009:LVF

[Özy09] H. Alper Özyiğit. Linear vibrations of frames carrying a concentrated mass. Mathematical and Computational Applications, 14(3):197–206, December 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/3/197.

Ozyilmaz:2011:CDG

[Özy11] Emin Özyılmaz. Classical differential geometry of curves according to type-2 Bishop trihedra. Mathematical and Computational Applications, 16(4):858– 867, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/4/858.

Ozyilmaz:2014:DTR

[Özy14] Emin Özyilmaz. The Darboux trihedrons of regular curves on a regular time-like surface. Mathematical and Computational Applications, 19(2):172–183, August 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/172.

Pakdemirli:2011:NPI

[PAB11] Mehmet Pakdemirli, Yiğit Aksoy, and Hakan Boyacı. A new perturbationiteration approach for first order differential equations. Mathematical and Computational Applications, 16(4):890– 899, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/890.

Pakdemirli:2002:GCS

[Pak02] Mehmet Pakdemirli. Group classification and some similarity solutions for a nonlinear filtration equation. Mathematical and Computational Applications, 7(1):1–4, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/1.

Pakdemirli:2016:NPA

[Pak16] Mehmet Pakdemirli. A new perturbation approach to optimal polynomial regression. Mathematical and Computational Applications, 21 (1):??, March 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/1/1.

Pyy:2018:SOH

[PALL18] Johanna Pyy, Anssi Ahtikoski, Alexander Lapin, and Erkki Laitinen. Solution of optimal harvesting problem by finite difference approximations of size-structured population model. Mathematical and Computational Applications, 23 (2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/2/22.

Pal:2023:ICI

[PAR23] Soumyadip Pal, Fahad Al Basir, and Santanu Ray. Impact of cooperation and intra-specific competition of prey on the stability of prey– predator models with refuge. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/88.

Pakdemirli:1997:GAC

[PB97] M. Pakdemirli and H. Boyacı.

A generalized approach to coupled nonlinear vibrations of continuous systems. Mathematical and Computational Applications, 2(3):141–154, December 1997. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/3/141.

Pakdemirli:1999:CDV

[PB99] Mehmet Pakdemirli and Hakan Boyacı. A comparison of different versions of the method of multiple scales for an arbitrary model of odd nonlinearities. Mathematical and Computational Applications, 4(3): 273–282, December 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/273.

Pakdemirli:2001:VSB

[PB01] M. Pakdemirli and H. Boyacı.

Vibrations of a stretched beam with non-ideal boundary conditions. Mathematical and Computational Applications, 6(3):217–220, December 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/3/217.

Pakdemirli:2003:VSS

[PB03] M. Pakdemirli and H. Boyaci. Vibrations of a simply supported beam with a non-ideal support at an intermediate point. Mathematical and Computational Applications, 8(2):159– 164, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/159.

Pakdemirli:1997:CSO

[PBY97] Mehmet Pakdemirli, Hakan Boyacı, and Mehmet Yılmaz. Continuous systems with odd nonlinearities: a general solution procedure. Mathematical and Computational Applications, 2 (2):85–90, August 1997. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/2/2/85.

Pakdemirli:2007:PDC

[PBY07] M. Pakdemirli, H. Boyacı, and H. A. Yurtsever. Perturbative derivation and comparisons of rootfinding algorithms with fourth order derivatives. Mathematical and Computational Applications, 12(2):117–124, August 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/2/117.

Pakdemirli:2008:RFA

[PBY08] M. Pakdemirli, H. Boyacı, and H. A. Yurtsever. A root-finding algorithm with fifth order derivatives. Mathematical and Computational Applications, 13(2):123–128, August 2008. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/13/2/123.

Promma:2023:FFM

[PC23] Nat Promma and Nawinda Chutsagulprom. Forecasting financial and macroeconomic variables using an adaptive parameter VAR-KF model. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/19.

Pasinli:2011:MCM

[PCCT11] Ahmet Pasinli, Osman Culha, Erdal Celik, and Mustafa Toparli. A model for the calculation of mechanical properties of hydroxyapatite coatings on Ti6Al4V substrate using finite element methods. Mathematical and Computational Applications, 16 (1):136–147, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/136.

Peitz:2018:SRT

[PD18] Sebastian Peitz and Michael Dellnitz. A survey of recent trends in multiobjective optimal control — surrogate models, feedback control and objective reduction. Mathematical and Computational Applications, 23(2):??, June 2018. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/30.

Piegari:2019:FCC

[PD19] Estefanía Piegari and Silvina Ponce Dawson. Functional ca^{2+} channels between channel clusters are necessary for the propagation of IP_{3} r-mediated ca^{2+} waves. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/61.

Pekiner:1996:UGM

[Pek96] Fahri Okan Pekiner. Using the growth models for optimization of energy distributions systems. Mathematical and Computational Applications, 1(2):169–173, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/169.

Pracek:2010:MMY

[PF10a] Stanislav Praček and Sluga Franci. Matematical model for yarn unwinding from packages. Mathematical and Computational Applications, 15(5):853–858, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/853.

Pracek:2010:NSY

[PF10b] Stanislav Praček and Sluga Franci. Numerical simulations of yarn unwinding from packages. Mathematical and Computational Applications, 15(5):846–852, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/846.

Pereira:2019:MMS

[PFR19] José Carlos Garcia Pereira, Jorge Cruz Fernandes, and Luís Guerra Rosa. Mathematical models for simulation and optimization of high-flux solar furnaces. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/65.

Pirim:2011:PEC

[PGB^{+}11] Harun Pirim, Dilip Gautam, Tanmay Bhowmik, Andy D. Perkins, Burak Ekşioglu, and Ahmet Alkan. Performance of an ensemble clustering algorithm on biological data sets. Mathematical and Computational Applications, 16 (1):87–96, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/87.

Perestrelo:2021:MFF

[PGRL21] Sara Perestrelo, Maria C. Grácio, Nuno A. Ribeiro, and Luís M. Lopes. Modelling forest fires using complex networks. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/4/68.

Pinault:2022:MCW

[Pin22] Jean-Louis Pinault. Morlet cross-wavelet analysis of climatic state variables expressed as a function of latitude, longitude, and time: New light on extreme events. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/50.

Pakdemirli:2009:NPA

[PKB09] M. Pakdemirli, M. M. F.

Karahan, and H. Boyacı. A new perturbation algorithm with better convergence properties: Multiple scales Lindstedt Poincaré method. Mathematical and Computational Applications, 14 (1):31–44, April 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/31.

Pakdemirli:2011:FVS

[PKB11] M. Pakdemirli, M. M. F.

Karahan, and H. Boyacı. Forced vibrations of strongly nonlinear systems with multiple scales Lindstedt Poincaré method. Mathematical and Computational Applications, 16(4):879–889, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/879.

Pinter:2022:FCS

[PKC22] János D. Pintér, Frank J. Kampas, and Ignacio Castillo. Finding the conjectured sequence of largest small n-polygons by numerical optimization. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/42.

Peker:2011:DTM

[PKO11] Haldun Alpaslan Peker, Onur Karaoğlu, and Galip Oturanç. The differential transformation method and Padé approximant for a form of Blasius equation. Mathematical and Computational Applications, 16(2):507–513, August 2011. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/16/2/507.

Pulido-Luna:2020:DNN

[PLLRCC20] Jesus R. Pulido-Luna, Jorge A. López-Rentería, and Nohe R. Cazarez-Castro. Design of a nonhomogeneous nonlinear synchronizer and its implementation in reconfigurable hardware. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/51.

Phillipson:2021:AIP

[PLW21] Frank Phillipson, Peter Langenkamp, and Reinder Wolthuis. Alternative initial probability tables for elicitation of Bayesian belief networks. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/54.

Pal:2022:CNG

[PM22] Swadesh Pal and Roderick Melnik. Coupled neural–glial dynamics and the role of astrocytes in Alzheimer’s disease. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/33.

Polucci:2020:NNI

[PMF20] Daniele Polucci, Michele Marchetti, and Simone Fiori. A novel nonisotonic statistical bivariate regression method — application to stratigraphic data modeling and interpolation. Mathematical and Computational Applications, 25 (1):??, March 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/1/15.

Phaneendra:2014:SFD

[PMR14] K. Phaneendra, K. Madhulatha, and Y. N. Reddy. Special finite difference method for singular perturbation problems with oneend boundary layer. Mathematical and Computational Applications, 19(3):208–217, December 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/3/208.

Pooe:2003:ISC

[PMS03] C. A. Pooe, F. M. Mahomed, and C. Wafo Soh. Invariant solutions and conservation laws of the Black–Scholes equation. Mathematical and Computational Applications, 8(1):63–70, April 2003. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/1/63.

Pakdemirli:1998:ABL

[PÖ98] M. Pakdemirli and E. Özkaya. Approximate boundary layer solution of a moving beam problem. Mathematical and Computational Applications, 3(2): 93–100, August 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/2/93.

Popa:1998:SFP

[Pop98]  Valeriu Popa.  Some

fixed point theorems in Hilbert spaces. Mathematical and Computational Applications, 3 (1):11–16, April 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/1/11.

Petitgirard:2020:SST

[PPB^{+}20] Julien Petitgirard, Tony Piguet, Philippe Baucour, Didier Chamagne, Eric Fouillien, and Jean-Christophe Delmare. Steady state and 2D thermal equivalence circuit for winding heads — a new modelling approach. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/70.

Pittella:2022:MCM

[PPRF22] Erika Pittella, Emanuele Piuzzi, Pietro Russo, and Francesco Fabbrocino. Microwave characterization and modelling of PA6/GNPs composites. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/41.

Plata:2020:CSH

[PPRVC20] Corina Plata, Pablo J. Prieto, Ramon Ramirez-Villalobos, and Luis N. Coria. Chaos synchronization for hyperchaotic Lorenz-type system via fuzzy-based sliding-mode observer. Mathematical and Computational Applications, 25(1):??, March 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/1/16.

Pijls:2023:CMM

[PQ23] Henk Pijls and Le Phuong Quan. A computational method with Maple for finding the maximum curvature of a Bézierspline curve. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/56.

Perez-Rodriguez:2021:HED

[PR21] Ricardo Pérez-Rodríguez.

A hybrid estimation of distribution algorithm for the quay Crane scheduling problem. Mathematical and Computational Applications, 26(3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/64.

Perez-Rodriguez:2023:RHE

[PRFH23] Ricardo Pérez-Rodríguez and Sergio Frausto-Hernández. A radial hybrid estimation of distribution algorithm for the truck and trailer routing problem. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/27.

Paez-Rueda:2023:EPM

[PRFP^{+}23] Carlos-Iván Páez-Rueda, Arturo Fajardo, Manuel Pérez, German Yamhure, and Gabriel Perilla. Exploring the potential of mixed Fourier series in signal processing applications using one-dimensional smooth closed-form functions with compact support: a comprehensive tutorial. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/93.

Patel:2019:TAA

[PRT^{+}19] Reena Patel, Guillermo Riveros, David Thompson, Edward Perkins, Jan Jeffery Hoover, John Peters, and Antoinette Tordesillas. A transdisciplinary approach for analyzing stress flow patterns in biostructures. Mathematical and Computational Applications, 24(2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/47.

Pakdemirli:2005:ASH

[PŞ05] M. Pakdemirli and A. Z. Şahin. Approximate symmetries of hyperbolic heat conduction equation with temperature dependent thermal properties. Mathematical and Computational Applications, 10 (1):139–145, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/139.

Pakdemirli:2006:SSF

[PŞ06] M. Pakdemirli and A. Z.

Şahin. A similarity solution of fin equation with variable thermal conductivity and heat transfer coefficient. Mathematical and Computational Applications, 11 (1):25–30, April 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/1/25.

Pakdemirli:2015:SQN

[PS15] Mehmet Pakdemirli and Gözde Sarı. Solution of quadratic nonlinear problems with multiple scales Lindstedt– Poincaré method. Mathematical and Computational Applications, 20(2):137–150, August 2015. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/20/2/137.

Prawoto:2019:ACS

[PS19] Yunan Prawoto and Agus Suhartono. The Archimedes’ constant, π seen by mechanical engineers. Mathematical and Computational Applications, 24(3):??, September 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/72.

Pelzer:2020:FCP

[PS20] Anouk F. G. Pelzer and Alef E. Sterk. Finite cascades of pitchfork bifurcations and multistability in generalized lorenz-96 models. Mathematical and Computational Applications, 25 (4):??, December 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/4/78.

Patel:2022:STF

[PS22] Himanshukumar R. Patel and Vipul A. Shah. Shadowed type-2 fuzzy sets in dynamic parameter adaption in cuckoo search and flower pollination algorithms for optimal design of fuzzy fault-tolerant controllers. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/89.

Pollicott:2023:SFD

[PS23] Mark Pollicott and Julia Slipantschuk. Sierpiński fractals and the dimension of their Laplacian spectrum. Mathematical and Computational Applications, 28(3):??, June 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/70.

Pedrosa:2022:LMP

[PSF22] Matheus V. A. Pedrosa, Tristan Schneider, and Kathrin Flaßkamp. Learning motion primitives automata for autonomous driving applications. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/54.

Pranowo:2024:LBM

[PSW24] Pranowo, Djoko Budiyanto Setyohadi, and Agung Tri Wijayanta. The lattice Boltzmann method using parallel computation: a great potential solution for various complicated acoustic problems. Mathematical and Computational Applications, 29(1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/12.

Parovik:2021:SAN

[PT21] Roman Parovik and Dmitriy Tverdyi. Some aspects of numerical analysis for a model nonlinear fractional variable order equation. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/55.

Peng:2018:FBT

[PTLZ18] Kaijun Peng, Jieqing Tan, Zhiming Li, and Li Zhang. Fractal behavior of a ternary 4point rational interpolation subdivision scheme. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/65.

Pury:2021:TCC

[Pur21] Pedro A. Pury. Time to critical condition in emergency services. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/70.

Perez-Villafuerte:2021:EPD

[PVCRRV^{+}21] Mercedes Perez-Villafuerte, Laura Cruz-Reyes, Nelson Rangel-Valdez, Claudia Gomez-Santillan, and Héctor Fraire-Huacuja. Effect of the profile of the decision maker in the search for solutions in the decision-making process. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/28.

Pakdemirli:2000:ASW

[PY00] M. Pakdemirli and M. Yürüsoy.

On approximate symmetries of a wave equation with quadratic nonlinearity. Mathematical and Computational Applications, 5(3):179–184, December 2000. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/3/179.

Pannerselvam:2022:SSB

[PYR22]  Kiran  Pannerselvam,

Deepanshu Yadav, and Palaniappan Ramu. Scarce sample-based reliability estimation and optimization using importance sampling. Mathematical and Computational Applications, 27(6):??, December 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/99.

Qadir:2024:CCB

[Qad24] Asghar Qadir. Complex connections between symmetry and singularity analysis. Mathematical and Computational Applications, 29(1):??, February 2024. ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/15.

Quiroz-Castellanos:2023:NEO

[QCdlFL^{+}23] Marcela Quiroz-Castellanos, Luis Gerardo de la Fraga, Adriana Lara, Leonardo Trujillo, and Oliver Schütze. Numerical and evolutionary optimization 2021. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/71.

Qi:2016:SPF

[QM16] Feng Qi and Mansour Mahmoud. Some properties of a function originating from geometric probability for pairs of hyperplanes intersecting with a convex body. Mathematical and Computational Applications, 21(3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/27.

Quan:2018:ACA

[QN18] Le Phuong Quan and Thái Anh Nhan. Applying computer algebra systems in approximating the trigonometric functions. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/37.

Quiroz:2022:NEO

[QRdlFS22] Marcela Quiroz, Juan Gabriel Ruiz, Luis Gerardo de la Fraga, and Oliver Schütze. Numerical and evolutionary optimization 2020. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/70.

Shi:2013:NES

[qS13] Ye qiong Shi. New exact solutions of the (2 + 1)- dimensional Ginzburg– Landau equation. Mathematical and Computational Applications, 18(2):103–111, August 2013. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/18/2/103.

Qi:2019:UGC

[QSC19] Feng Qi, Xiao-Ting Shi, and Pietro Cerone. A unified generalization of the Catalan, Fuss, and Fuss — Catalan numbers. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/49.

Quan:2018:CMM

[Qua18] Le Phuong Quan. A computational method with MAPLE for a piecewise polynomial approximation to the trigonometric functions. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/63.

Rouhparvar:2010:EUS

[RAA10] H. Rouhparvar, S. Abbasbandy, and T. Allahviranloo. Existence and uniqueness of solution of an uncertain characteristic Cauchy reaction-diffusion equation by Adomian decomposition method. Mathematical and Computational Applications, 15(3):404–419, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/404.

Rajagopal:2010:NCM

[Raj10] K. R. Rajagopal. On a new class of models in elasticity. Mathematical and Computational Applications, 15(4):506– 528, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/4/506.

Rabah:2022:SCN

[RAK22] Fatima Rabah, Marwan Abukhaled, and Suheil A. Khuri. Solution of a complex nonlinear fractional biochemical reaction model. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/45.

Rawashdeh:2014:ASC

[Raw14] Mahmoud Rawashdeh.

Approximate solutions for coupled systems of nonlinear PDEs using the reduced differential transform method. Mathematical and Computational Applications, 19(2):161–171, August 2014. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/19/2/161.

Rasulov:1999:EAE

[RC99] M. Rasulov and E. Coşkun. An effective algorithm for Euler system in a class of discontinuous functions. Mathematical and Computational Applications, 4(2):185– 188, August 1999. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/2/185.

Ren:2016:CRA

[RC16] Yaming Ren and Zhongxian Chen. On the O(1/n) convergence rate of the auxiliary problem principle for separable convex programming and its application to the power systems multi-area economic dispatch problem. Mathematical and Computational Applications, 21 (3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/3/35.

Ramirez-Cardenas:2020:SST

[RCTR20] Oscar-David Ramírez-Cárdenas and Felipe Trujillo-Romero. Sensorless speed tracking of a brushless DC motor using a neural network. Mathematical and Computational Applications, 25(3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/57.

Rostamiyan:2010:AAI

[RFD^{+}10] Yasser Rostamiyan, A. Fereidoon, M. R. Davoudabadi, H. Yaghoobi, and D. D. Ganji. Analytical approach to investigation of deflection of circular plate under uniform load by homotopy perturbation method. Mathematical and Computational Applications, 15(5):816– 821, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/816.

Ramos-Figueroa:2023:ESG

[RFQCMMCR23] Octavio Ramos-Figueroa, Marcela Quiroz-Castellanos, Efrén Mezura-Montes, and Nicandro Cruz-Ramírez. An experimental study of grouping mutation operators for the unrelated parallelmachine scheduling problem. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/6.

Rozhkov:2023:CMT

[RG23] Aleksandr N. Rozhkov and Vera V. Galishnikova. Construction and modification of topological tables for digital models of linear complexes. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/37.

Razvarz:2017:IAM

[RJ17] Sina Razvarz and Raheleh Jafari. ICA and ANN modeling for photocatalytic removal of pollution in wastewater. Mathematical and Computational Applications, 22(3):??, September 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/3/38.

Ramchandraiah:2022:DDC

[RKR^{+}22a] Chirnam Ramchandraiah, Naikoti Kishan, Gundlapally Shiva Kumar Reddy, Kiran Kumar Paidipati, and Christophe Chesneau. Doublediffusive convection in bidispersive porous medium with Coriolis effect. Mathematical and Computational Applications, 27(4):??, August 2022.

CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/56.

Reddy:2022:DDC

[RKR^{+}22b] Gundlapally Shiva Kumar Reddy, Nilam Venkata Koteswararao, Ragoju Ravi, Kiran Kumar Paidipati, and Christophe Chesneau. Dissolutiondriven convection in a porous medium due to vertical axis of rotation and magnetic field. Mathematical and Computational Applications, 27(3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/53.

Reddy:2023:ICM

[RKS23] Vinodh Srinivasa Reddy, Jagan Kandasamy, and Sivasankaran Sivanandam. Impacts of Casson model on hybrid nanofluid flow over a moving thin needle with Dufour and Soret and thermal radiation effects. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/2.

Reynolds:2023:CMM

[RKX23] Quinn G. Reynolds, Thokozile P. Kekana, and Buhle S. Xakalashe. A computational magnetohydrodynamic modelling study on plasma arc behaviour in gasification applications. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/60.

Rui:2017:OST

[RLY^{+}17] Xiaodong Rui, Yue Liu, Aijun Yang, Hongqiang Yang, and Chengcui Zhang. Optimization of setting take-profit levels for derivative trading. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/1.

Resca:2020:GGP

[RM20] Lorenzo G. Resca and Nicholas A. Mecholsky. Geometry and geodesy on the primary visual cortex as a surface of revolution. Mathematical and Computational Applications, 25(4):??, December 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/64.

Rodriguez-Mata:2020:FHG

[RMBTGH^{+}20] Abraham Efraim Rodriguez-Mata, Yaneth Bustos-Terrones, Victor Gonzalez-Huitrón, Pablo Antonio Lopéz-Peréz, Omar Hernández-González, and Leonel Ernesto Amabilis-Sosa. A fractional highgain nonlinear observer design — application for rivers environmental monitoring model. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/44.

Robitzsch:2022:FPG

[Rob22] Alexander Robitzsch.

Four-parameter guessing model and related item response models. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/95.

Rodrigues:2020:IAF

[Rod20] José A. Rodrigues. Isogeometric analysis for fluid shear stress in cancer cells. Mathematical and Computational Applications, 25(2):??, June 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/19.

Roussos:2021:VBL

[Rou21] Evangelos Roussos. Variational Bayesian learning of SMoG s: Modelling and their application to synthetic aperture radar. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/2/45.

Raza:2016:NSS

[RRO16a] Jawad Raza, Azizah Mohd Rohni, and Zurni Omar. A note on some solutions of copper-water (cu-water) nanofluids in a channel with slowly expanding or contracting walls with heat transfer. Mathematical and Computational Applications, 21(2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/24.

Raza:2016:NIC

[RRO16b] Jawad Raza, Azizah Mohd Rohni, and Zurni Omar. Numerical investigation of copper-water (cuwater) nanofluid with different shapes of nanoparticles in a channel with stretching wall: Slip effects. Mathematical and Computational Applications, 21(4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/43.

Rehman:2022:FMV

[RRR^{+}22] Aziz Ur Rehman, Muhammad Bilal Riaz, Wajeeha Rehman, Jan Awrejcewicz, and Dumitru Baleanu. Fractional modeling of viscous fluid over a moveable inclined plate subject to exponential heating with singular and non-singular kernels. Mathematical and Computational Applications, 27(1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/8.

Reynolds:2021:DII

[RS21] Robert Reynolds and Allan Stauffer. Definite integral involving rational functions of powers and exponentials expressed in terms of the Lerch function. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/58.

Rozza:2023:FPC

[RSF23] Gianluigi Rozza, Oliver Schütze, and Nicholas Fantuzzi. Feature paper collection of mathematical and computational applications — 2022. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/16.

Rashidi:2021:SAS

[RSS^{+}21] Mohammad Mehdi Rashidi, Mikhail A. Sheremet, Maryam Sadri, Satyaranjan Mishra, Pradyumna Kumar Pattnaik, Faranak Rabiei, Saeid Abbasbandy, Hussein Sahihi, and Esmaeel Erfani. Semi-analytical solution of two-dimensional viscous flow through expanding/contracting gaps with permeable walls. Mathematical and Computational Applications, 26(2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/2/41.

Ramatlo:2023:DTH

[RWL23] Dineo A. Ramatlo, Daniel N. Wilke, and Philip W. Loveday. Digital twin hybrid modeling for enhancing guided wave ultrasound inspection signals in welded rails. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/58.

Rugamba:2018:GFL

[RZ18] Jean Rugamba and Yanni Zeng. Green’s function of the linearized logarithmic Keller–Segel–Fisher/ KPP system. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/56.

Siddiqi:1997:GCP

[SA97] A. H. Siddiqi and R. Ahmad. On generalized complementarity problem. Mathematical and Computational Applications, 2(1):31–43, April 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/1/31.

Salti:2016:HTF

[SA16] Mustafa Salti and Oktay Aydogdu. Holographic tachyon in fractal geometry. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/21.

Shahzadi:2019:HBP

[SA19] Gulfam Shahzadi and Muhammad Akram. Hypergraphs based on Pythagorean fuzzy soft model. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/4/100.

Sagiroglu:1998:ANN

[Sağ98]  Şeref Sağıroğlu.  Ar-

tificial neural networks in robotic applications. Mathematical and Computational Applications, 3(2):67–81, August 1998. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/3/2/67.

Siddiqi:1998:PAG

[SAH98] A. H. Siddiqi, Rais Ahmad, and S. Husain. A perturbated algorithm for generalized nonlinear quasi-variational inclusions. Mathematical and Computational Applications, 3(3):177–184, December 1998. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/3/177.

Sabbah:2004:IPC

[SAI04a] A. S. Sabbah, M. A.

Abdou, and A. S. Ismail. An infinite plate with a curvilinear hole and flowing heat. Mathematical and Computational Applications, 9(2): 321–332, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/321.

Salahuddin:2004:ETG

[SAI04b] Hanif Salahuddin, Rais Ahmad, and S. S. Irfan. An existence theorem for general mixed multivalued mildly nonlinear variational inequalities without convexity. Mathematical and Computational Applications, 9(2): 201–203, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/201.

Sajid:2018:BCR

[Saj18] Mohammad Sajid. Bifurcation and chaos in real dynamics of a twoparameter family arising from generating function of generalized Apostoltype polynomials. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/7.

Salahuddin:2001:MIT

[SAK01] Hanif Salahuddin, R. Ahmad, and M. F. Khan. Mann and Ishikawa type perturbed iterative algorithms for generalized nonlinear variational inclusions. Mathematical and Computational Applications, 6(1):47– 52, April 2001. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/47.

Salahuddin:2002:GNM

[Sal02a] Hanif Salahuddin. General nonlinear mixed variational-like inequality problem involving relaxed Lipschitz operators. Mathematical and Computational Applications, 7(2):151–156, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/151.

Salahuddin:2002:GPA

[Sal02b] Hanif Salahuddin. On a general projection algorithm for variational inequalities involving relaxed Lipschitz and relaxed monotone mappings. Mathematical and Computational Applications, 7(2):133–137, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/133.

Salahli:2003:MAM

[Sal03] M. Ali Salahli. Multilevel architecture for multidimensional data base. Mathematical and Computational Applications, 8(3):263–269, December 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/263.

Salahuddin:2004:PMQ

[Sal04] Hanif Salahuddin. Projection methods for quasi-variational inequalities. Mathematical and Computational Applications, 9(2):125–131, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/125.

Salahuddin:2006:SAG

[Sal06a] Hanif Salahuddin. Sensitivity analysis of general mixed multivalued mildly nonlinear variational inequalities. Mathematical and Computational Applications, 11(2):121–125, August 2006. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/11/2/121.

Salem:2006:NSI

[Sal06b]  S. A. Salem.  On

the numerical solution of the incompressible Navier–Stokes equations in primitive variables using grid generation techniques. Mathematical and Computational Applications, 11(2):127– 136, August 2006. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/2/127.

Salahli:2009:AMS

[Sal09] Mehmet Ali Salahli. An approach for measuring semantic relatedness between words via related terms. Mathematical and Computational Applications, 14 (1):55–63, April 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/1/55.

Samedov:1999:ASF

[Sam99] Rafig Samedov. An approach to the support of the fault-tolerance of the double redundant computer control systems. Mathematical and Computational Applications, 4(2):175– 184, August 1999. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/2/175.

SoaresDosSantos:2022:ASM

[SAM^{+}22] Lindomar Soares Dos Santos, José Renato Alcarás, Lucas Murilo Da Costa, Mateus Mendonça Ramos Simões, and Alexandre Souto Martinez. Analytical solutions of microplastic particles dispersion using a Lotka– Volterra predator–prey model with time-varying intraspecies coefficients. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/66.

Saruhan:2009:DOO

[Sar09] Hamit Saruhan. Designing optimum oil thickness in artificial human knee joint by simulated annealing. Mathematical and Computational Applications, 14(2):109– 117, August 2009. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/2/109.

Sari:2011:DQS

[Sar11] Murat Sari. Differential quadrature solutions of the generalized Burgers– Fisher equation with a strong stability preserving high-order time integration. Mathematical and Computational Applications, 16(2):477– 486, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/477.

Sato:2023:HEC

[Sat23]  Junya Sato.  Hand–

eye calibration using a tablet computer. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/22.

Savas:1999:SCS

[Sav99] Ekrem Savaş. On statistically convergent sequences of fuzzy numbers. Mathematical and Computational Applications, 4(3):283–287, December 1999. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/4/3/283.

Savas:2000:SDS

[Sav00] Elcrem Savas. Spaces of double sequences of fuzzy numbers. Mathematical and Computational Applications, 5 (1):35–40, April 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/1/35.

Savas:2010:SDL

[Sav10] Ekrem Savas. On some double lacunary sequence spaces of fuzzy numbers. Mathematical and Computational Applications, 15(3):439–448, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/439.

Savas:2012:SAC

[Sav12]  Ekrem Savaş.  On

strongly almost convergent sequence spaces of fuzzy numbers. Mathematical and Computational Applications, 17 (2):92–99, August 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/2/92.

Savas:2013:SCS

[Sav13]  Ekrem Savaş.  Some

i-convergent sequence spaces of fuzzy numbers defined by infinite matrix. Mathematical and Computational Applications, 18 (2):84–93, August 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/2/84.

Savas:2014:LSC

[Sav14] Ekrem Savaş. Lacunary statistical convergence of multiple sequences in intuitionistic fuzzy normed spaces. Mathematical and Computational Applications, 19(3):241– 254, December 2014. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/3/241.

Savas:2016:GDS

[Sav16] Ekrem Savaş. On generalized double statistical convergence of order α in intuitionistic fuzzy normed spaces. Mathematical and Computational Applications, 21 (3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/3/36.

Sarwar:2018:DMA

[SAZ18] Musavarah Sarwar, Muhammad Akram, and Fariha Zafar. Decision making approach based on competition graphs and extended TOPSIS method under bipolar fuzzy environment. Mathematical and Computational Applications, 23(4):??, December 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/68.

Solak:2002:SBM

[SB02] Süleyman Solak and Durmuş Bozkurt. Some bounds on ℓ_{p} matrix and ℓ_{p} operator norms of almost circulant, Cauchy– Toeplitz and Cauchy– Hankel matrices. Mathematical and Computational Applications, 7(3): 211–218, December 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/211.

Selam:2004:ILF

[SBA^{+}04] C. Selam, T. Babacan, H. A. Aygör, H. Bircan, A. Küçübursa, and I. Maraş. The investigation of the log(ft) values for the allowed Gamow– Teller transitions of some deformed nuclei. Mathematical and Computational Applications, 9 (1):79–90, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/79.

Singh:2023:HOM

[SBB23] Gurjeet Singh, Sonia Bhalla, and Ramandeep Behl. Higher-order multiplicative derivative iterative scheme to solve the nonlinear problems. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/23.

Smedberg:2022:KDM

[SBDN^{+}22] Henrik Smedberg, Carlos Alberto Barrera-Diaz, Amir Nourmohammadi, Sunith Bandaru, and Amos H. C. Ng. Knowledge-driven multiobjective optimization for reconfigurable manufacturing systems. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/106.

Sonmez:1996:CAC

[SBF96] A. İhsan Sönmez, Adil Baykasoğlu, and İ. Hüseyin Filiz. Computer aided constrained optimisation of cutting conditions in drilling operations on a CNC lathe by using geometric programming. Mathematical and Computational Applications, 1(1):97–104, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/97.

Seflek:2010:DES

[SÇ10] Ali Yavuz Seflek and Kazım Çarman. A design of an expert system for selecting pumps used in agricultural irrigation. Mathematical and Computational Applications, 15(1):108– 116, April 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/108.

Sanderson:2019:PSM

[SC19] Conrad Sanderson and Ryan Curtin. Practical sparse matrices in C++ with hybrid storage and template-based expression optimisation. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/70.

Sen:2012:PBP

[SÇGA12] Baha Sen, Abdullah Çavusoğlu, Haldun Göktas, and Nesrin Aydin. Population based procedural artificial city generation using beta distribution. Mathematical and Computational Applications, 17(1):9–17, April 2012.

CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/1/9.

Shi:2013:APS

[SD13] Peilin Shi and Lingzhen Dong. Anti-periodic solutions for neural networks with delays and impulses. Mathematical and Computational Applications, 18(1):50–61, April 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/1/50.

Sahin:2017:CSN

[ŞE17] Mehmet Şahin and Rızvan Erol. A comparative study of neural networks and ANFIS for forecasting attendance rate of soccer games. Mathematical and Computational Applications, 22 (4):??, December 2017. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/4/43.

Senturk:2005:EIR

[Şen05]  Ş. Şentürk. Enlarged

ionization region and wiley-McLaren type geometry. Mathematical and Computational Applications, 10(1):133– 138, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/133.

Soureshjani:2024:TDM

[SEZ24] Mohammad Khodabakhshi Soureshjani, Hermann J. Eberl, and Richard G. Zytner. Three-dimensional model for bioventing: Mathematical solution, calibration and validation. Mathematical and Computational Applications, 29(1):??, February 2024. ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/16.

Sepahvandzadeh:2018:NSS

[SGA18] Amaneh Sepahvandzadeh, Bahman Ghazanfari, and Nader Asadian. Numerical solution of stochastic generalized fractional diffusion equation by finite difference method. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/53.

Sanchez-Garcia:2023:MLC

[SGBR23] Marta M. Sánchez-García, Gonzalo Barderas, and Pilar Romero. Modelization of low-cost maneuvers for an areostationary preliminary mission design. Mathematical and Computational Applications, 28 (6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/105.

Singla:2023:CFD

[SGC23] Rohan Singla, Shubham Gupta, and Arnab Chanda. A computational fluid dynamicsbased model for assessing rupture risk in cerebral arteries with varying aneurysm sizes. Mathematical and Computational Applications, 28 (4):??, August 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/90.

Salvador-Gonzalez:2023:ESS

[SGCRRL^{+}23] María Concepción Salvador-González, Juana Canul-Reich, Rafael Rivera-López, Efrén Mezura-Montes, and Erick de la Cruz-Hernandez. Evolutionary selection of a set of association rules considering biological constraints describing the prevalent elements in bacterial vaginosis. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/75.

Sagiroglu:1998:NCB

[SGE98]  Şeref Sağıroğlu, Kerim

Güney, and Mehmet Erler. Neural computing of the bandwith of resonant rectangular microstrip antennas. Mathematical and Computational Applications, 3 (1):37–47, April 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/1/37.

Sekmen:2006:PPS

[SGEP06] Yakup Sekmen, Mustafa Gölcü, Perihan Erduranlı, and Yaşar Pancar. Prediction of performance and smoke emission using artificial neural network in a Diesel engine. Mathematical and Computational Applications, 11(3):205– 214, December 2006. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/205.

Sari:2010:HOF

[SGZ10] Murat Sari, Gürhan Gürarslan, and Asuman Zeytinoğlu. High-order finite difference schemes for solving the advectiondiffusion equation. Mathematical and Computational Applications, 15(3):449–460, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/449. See er-

ratum [SGZ11].

Sari:2011:EH

[SGZ11] Murat Sari, Gürhan Gürarslan, and Asuman Zeytinoğlu. Erratum to: “High-Order Finite Difference Schemes for Solving the Advection-Diffusion Equation, Mathematical and Computational Applications, Vol. 15, No. 3, 449–460, 2010.”. Mathematical and Computational Applications, 16(4):979–??, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/979.  See

[SGZ10].

Singh:2018:NHW

[SH18] Narinder Singh and Hanaa Hachimi. A new hybrid whale optimizer algorithm with mean strategy of grey Wolf optimizer for global optimization. Mathematical and Computational Applications, 23 (1):??, March 2018. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/14.

Shamilov:1999:TGB

[Sha99] A. Kh. Shamilov. On the theory of generalized ballistical problem for nonlinear abstract differential equations. Mathematical and Computational Applications, 4 (1):69–74, April 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/1/69.

Shahid:2019:ASA

[Sha19] Nazish Shahid. An analytical study of adversely affecting radiation and temperature parameters on a magnetohydrodynamic elastoviscous fluid. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/31.

Sultan:2005:CAM

[SHE05] A. M. Sultan, N. A.

Hassan, and N. M. Elhamy. Computational analysis of a multiserver bulk arrival with two modes server breakdown. Mathematical and Computational Applications, 10(2):249–259, August 2005. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/2/249.

Sanchez-Hernandez:2021:PPM

[SHFSGB^{+}21] Juan P. Sánchez-Hernández, Juan Frausto-Solís, Juan J. González-Barbosa, Diego A. Soto-Monterrubio, Fanny G. Maldonado-Nava, and Guadalupe Castilla-Valdez. A peptides prediction methodology for tertiary structure based on simulated annealing. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/2/39.

Shinde:2006:PCS

[Shi06] Shilpa Shinde. Propagation of cylindrical shock waves in a non-uniform rotating stellar atmosphere under the action of monochromatic radiation and gravitation. Mathematical and Computational Applications, 11(2):95–102, August 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/2/95.

Shokrollahi:2022:EWP

[Sho22] Foad Shokrollahi. Equity warrants pricing formula for uncertain financial market. Mathematical and Computational Applications, 27 (2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/18.

Shrivastava:1998:ECC

[Shr98] H. S. P. Shrivastava.

Evaluation of certain class of Eulerian integrals of the multivariable h-function. Mathematical and Computational Applications, 3(2): 113–125, August 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/2/113.

Shrivastava:2001:SGF

[Shr01] HSP Shrivastava. Some generating function relations of multiindex Hermite polynomials. Mathematical and Computational Applications, 6(3): 189–194, December 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/3/189.

Shrivastava:2002:MMH

[Shr02] HSP Shrivastava. Multiindex multivariable Hermite polynomials. Mathematical and Computational Applications, 7(2): 139–149, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/139.

Shrivastava:2004:SCP

[Shr04] HSP Shrivastava. On sister Celine’s polynomials of several variables. Mathematical and Computational Applications, 9(2):309–320, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/309.

Shrivastava:2005:ECC

[Shr05] HSP Shrivastava. Evaluation of certain class of Eulerian integrals of multivariable sister Celine’s polynomials. Mathematical and Computational Applications, 10(2):239– 248, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/239.

Salahuddin:2004:PMG

[SI04] Hanif Salahuddin and S. S. Irfan. Proximal methods for generalized nonlinear quasivariational inclusions. Mathematical and Computational Applications, 9(2):165–171, August 2004. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/9/2/165.

Simsek:2013:FFD

[ŞİH13] Barış Şimşek, Y. Tansel İç, and Emir H.Şimşek. A full factorial design based desirability function approach for optimization of properties of C 40/50 concrete class. Mathematical and Computational Applications, 18(3):330– 339, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/330.

Sinir:2010:BCS

[Sin10] B. Gültekin Sinir. Bifurcation and chaos of slightly curved pipes. Mathematical and Computational Applications, 15(3):490–502, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/490.

Sivakumar:2019:SNH

[SJ19] Parimala Sivakumar and Jayakumar Jayaraman. Some new higher order weighted Newton methods for solving nonlinear equation with applications. Mathematical and Computational Applications, 24(2):??, June 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/2/59.

Sivasankaran:2022:NSM

[SJ22] Sivanandam Sivasankaran and Kandasamy Janagi. Numerical study on mixed convection flow and energy transfer in an inclined channel cavity: Effect of baffle size. Mathematical and Computational Applications, 27(1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/1/9.

Solis:2020:SSF

[SJKCD20] Francisco Solis, Silvia Jerez, Roberto Ku-Carrillo, and Sandra Delgadillo. Structural stability of a family of exponential polynomial maps. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/20.

Saglam:2003:CTC

[SK03] Haci Sağlam and Hülya Kaçar. Cutting tool condition monitoring using surface texture via neural network. Mathematical and Computational Applications, 8(2): 235–243, August 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/235.

Sultan:2005:NMG

[SK05] Ahmed M. Sultan and Hala Mahmoud Khaleel. A new modified goodness of fit tests for type 2 censored sample from normal population. Mathematical and Computational Applications, 10(2):165– 177, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/165.

Sadat:2018:ESD

[SK18] Rahma Sadat and Magda Kassem. Explicit solutions for the (2 + 1)-dimensional Jaulent– Miodek equation using the integrating factors method in an unbounded domain. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/1/15.

Subasi:2005:NNC

[SKAK05] Abdulhamit Subasi, M. Kemal Kiymik, Ahmet Alkan, and Etem Koklukaya. Neural network classification of EEG signals by using AR with MLE preprocessing for epileptic seizure detection. Mathematical and Computational Applications, 10(1):57– 70, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/57.

Sharma:2022:ETS

[SKB22] Himani Sharma, Munish Kansal, and Ramandeep Behl. An efficient two-step iterative family adaptive with memory for solving nonlinear equations and their applications. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/97.

Sharma:2023:EOD

[SKB23] Himani Sharma, Munish Kansal, and Ramandeep Behl. An efficient optimal derivativefree fourth-order method and its memory variant for non-linear models and their dynamics. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/48.

Sherman:2022:CSC

[SKGP22] Michelle Sherman, Gilbert Kerr, and Gilberto González-Parra. Comparison of symbolic computations for solving linear delay differential equations using the Laplace transform method. Mathematical and Computational Applications, 27 (5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/81.

Sriboonchandr:2019:IDE

[SKK19] Prasert Sriboonchandr, Nuchsara Kriengkorakot, and Preecha Kriengkorakot. Improved differential evolution algorithm for flexible job shop scheduling problems. Mathematical and Computational Applications, 24(3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/3/80.

Sresracoo:2018:USA

[SKKC18] Poontana Sresracoo, Nuchsara Kriengkorakot, Preecha Kriengkorakot, and Krit Chantarasamai. Ushaped assembly line balancing by using differential evolution algorithm. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/79.

Shafiq:2022:BND

[SKM^{+}22] Shakaiba Shafiq, Sadaf Khan, Waleed Marzouk, Jiju Gillariose, and Farrukh Jamal. The binomial–natural discrete Lindley distribution: Properties and application to count data. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/4/62.

Smiechowicz:2019:LEE

[ŚLO19] Wojciech Śmiechowicz, Théo Loup, and Paweł Olejnik. Lyapunov exponents of early stage dynamics of parametric mutations of a rigid pendulum with harmonic excitation. Mathematical and Computational Applications, 24(4):??, December 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/4/90.

Soh:1999:PGC

[SM99] C. Wafo Soh and F. M. Mahhomed. Preliminary group classification of y′′ = f(x)y^{2}. Mathematical and Computational Applications, 4(3): 289–295, December 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/289.

Sytnyk:2021:MMN

[SM21] Dmytro Sytnyk and Roderick Melnik. Mathematical models with nonlocal initial conditions: an exemplification from quantum mechanics. Mathematical and Computational Applications, 26 (4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/73.

Soykasap:1996:DRC

[SMB96] Ömer Soykasap, Zahit Mecitoğlu, and Oğuz Borat. Dynamic response of composite cylindrical shells to shock loading. Mathematical and Computational Applications, 1(1):85–96, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/85.

Sarisahin:2014:PN

[SN14] Tugba Sarisahin and Ayse Nalli. On the pentanacci numbers. Mathematical and Computational Applications, 19(3):255–262, December 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/3/255.

Senturk:2011:MIB

[ŞÖ11] Utku Şentürk and Aydoğan Özdamar. Modelling the interaction between water waves and the oscillating water column wave energy device. Mathematical and Computational Applications, 16(3):630–640, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/630.

Seker:2011:DRC

[ŞÖK11] Soner Şeker, Ali Uğur Öztürk, and Celal Kozanoğlu. Determination of reducing coefficient values of semi-rigid frames using artificial neural network. Mathematical and Computational Applications, 16(3):649–658, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/649.

Solis:2019:EEP

[Sol19] Francisco Solis. Evolution of an exponential polynomial family of discrete dynamical systems. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/13.

Solomyak:2023:NTM

[Sol23] Boris Solomyak. Notes on the transversality method for iterated function systems — a survey. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/3/65.

Selvi:2013:PAD

[SÖS13] Eşref Selvi, Merve Özdemir, and M. Alper Selver. Performance analysis of distance transform based inter-slice similarity information on segmentation of medical image series. Mathematical and Computational Applications, 18(3):511– 520, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/511.

Signes-Pont:2021:EGM

[SPCPMM21] Maria Teresa Signes-Pont, José Juan Cortés-Plana, and Higinio Mora-Mora. An epidemic grid model to address the spread of Covid-19: a comparison between Italy, Germany and France. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/14.

Santos:2018:FOH

[SPD18] Isabelle Santos, Stéphane Puechmorel, and Guillaume Dufour. First order Hadamard variation of the harmonic navigation function on a sphere world. Mathematical and Computational Applications, 23 (3):??, September 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/3/48.

Safdar:2011:ICL

[SQA11] M. Safdar, Asghar Qadir, and S. Ali. Inequivalence of classes of linearizable systems of second order ordinary differential equations obtained by real and complex symmetry analysis. Mathematical and Computational Applications, 16(4):923– 934, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/923.

Santos-Ruiz:2020:SOE

[SRLEPVP20] Ildeberto Santos-Ruiz, Francisco-Ronay López-Estrada, Vicenç Puig, and Guillermo Valencia-Palomo. Simultaneous optimal estimation of roughness and minor loss coefficients in a pipeline. Mathematical and Computational Applications, 25(3):??, September 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/56.

Soyhan:1996:CMD

[SS96] Hakan Serhad Soyhan and Cern Soruşbay. Conbustion modelling in dedicated naturel gas engines. Mathematical and Computational Applications, 1(1):119–125, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/119.

Savas:2003:SSS

[SS03] Ekrem Savaş and Rahmet Savaş. On some sequence spaces and lacunary σ-statistical convergence. Mathematical and Computational Applications, 8(2):165– 172, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/165.

Sen:2006:EPT

[SS06] Faruk Sen and Metin Sayer. Elasto-plastic thermal stress analysis in a thermoplastic composite disc under uniform temperature using FEM. Mathematical and Computational Applications, 11(1):31–39, April 2006. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/11/1/31.

Sekmen:2007:MMS

[SS07] Perihan Sekmen and Yakup Sekmen. Mathematical modeling of a SI engine cycle with actual air-fuel cycle analyses. Mathematical and Computational Applications, 12(3):161–171, December 2007. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/3/161.

Sadek:2022:SNM

[SSA22a] Lakhlifa Sadek, El Mostafa Sadek, and Hamad Talibi Alaoui. On some numerical methods for solving large differential nonsymmetric Stein matrix equations. Mathematical and Computational Applications, 27 (4):??, August 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/4/69.

Stipsitz:2022:ASS

[SSA22b] Monika Stipsitz and Hèlios Sanchis-Alepuz. Approximating the steadystate temperature of 3D electronic systems with convolutional neural networks. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/7.

Sinisterra-Sierra:2023:CDA

[SSGCPR23] Santiago Sinisterra-Sierra, Salvador Godoy-Calderón, and Miriam Pescador-Rojas. COVID-19 data analysis with a multiobjective evolutionary algorithm for causal association rule mining. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/12.

Salah:2023:CMH

[SSV23] Faisal Salah, Abdelmgid O. M. Sidahmed, and K. K. Viswanathan. Chemical MHD hiemenz flow over a nonlinear stretching sheet and Brownian motion effects of nanoparticles through a porous medium with radiation effect. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/1/21.

Sahin:1996:ETD

[ŞT96] Salim Şahin and Mustafa Toparlı. Examination of temperature distribution as a result of hardbrazing of two different materials by finite elements method. Mathematical and Computational Applications, 1 (1):172–180, June 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/172.

Solmaz:2011:ENS

[ST11] Murat Yavuz Solmaz and Aydın Turgut. An experimental and numerical study on the effects of taper angles and overlap length on the failure and stress distribution of adhesively-bonded single-lap joints. Mathematical and Computational Applications, 16 (1):159–170, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/159.

Seminara:2018:IPG

[ST18] Silvia Alejandra Seminara and María Inés Troparevsky. Iterated Petrov–Galerkin method with regular pairs for solving Fredholm integral equations of the second kind. Mathematical and Computational Applications, 23 (4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/23/4/73.

Sotiropoulos:2022:IBC

[ST22] Dimitris G. Sotiropoulos and Konstantinos Tserpes. Interval-based computation of the uncertainty in the mechanical properties and the failure analysis of unidirectional composite materials. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/38.

Sozinando:2023:ESC

[STA23a] Desejo Filipeson Sozinando, Bernard Xavier Tchomeni, and Alfayo Anyika Alugongo. Experimental study of coupled torsional and lateral vibration of vertical rotorto-stator contact in an inviscid fluid. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/2/44.

Stark:2023:NAC

[Sta23b] Sebastian Stark. Numerical aspects of a continuum sintering model formulated in the standard dissipative framework. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/69.

Solak:2004:UBS

[STB04] Süleyman Solak, Ramazan Türkmen, and Durmuş Bozkurt. Upper bounds for the spectral and l_{p} norms of Cauchy– Toeplitz and Cauchy– Hankel matrices. Mathematical and Computational Applications, 9 (1):41–47, April 2004.

CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/41.

Sen:2007:ETS

[STK07] Faruk Sen, Mustafa Toparlı, and N. Sinan Köksal. The evaluation of thermal stresses in thermoplastic laminated composites with different square holes. Mathematical and Computational Applications, 12 (1):59–68, April 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/59.

Sahin:2013:PIV

[ŞÜ13] Caner Şahin and Mustafa Ünel. Performance improvement in vslam using stabilized feature points. Mathematical and Computational Applications, 18(3):361–372, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/361.

Sinir:2013:ESB

[SUE13] B. Gültekin Sınır, F. Esra Uz, and Sefa Ergun. Exact solution and buckling configuration of nanotubes containing internal flowing fluid. Mathematical and Computational Applications, 18(3):539–547, December 2013. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/539.

Sultan:2007:MCB

[Sul07] Ahmed M. M. Sultan. Multi-channel Bi-level heterogeneous servers bulk arrival queueing system with Erlangian service time. Mathematical and Computational Applications, 12(2):97– 105, August 2007. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/2/97.

Sultan:2008:DDP

[Sul08] Ahmed M. M. Sultan. A data driven parameter estimation for the three-parameter Weibull population from censored samples. Mathematical and Computational Applications, 13(2):129– 136, August 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/2/129.

Song:2010:BSD

[SW10] Ruili Song and Bo Wang. Backward stochastic differential equation on hedging American contingent claims. Mathematical and Computational Applications, 15(5):895–900, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/895.

Sinha:2022:MEH

[SW22] Ankur Sinha and Jyrki Wallenius. MCDM, EMO and hybrid approaches: Tutorial and review. Mathematical and Computational Applications, 27(6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/112.

Sarmadi:2022:SNN

[SWA^{+}22] Sorena Sarmadi, James J.

Winkle, Razan N. Alnahhas, Matthew R. Bennett, Krešimir Josić, Andreas Mang, and Robert Azencott. Stochastic neural networks for automatic cell tracking in microscopy image sequences of bacterial colonies. Mathematical and Computational Applications, 27(2):??, April 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/22.

Song:2022:TDA

[SWJS22] Jia Song, Fujie Wang, Kemin Jia, and Haohao Shen. A timedomain artificial boundary for near-field wave problem of fluid saturated porous media. Mathematical and Computational Applications, 27(4):??, August 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/4/71.

Sekmen:2011:AEE

[SY11] Perihan Sekmen and Zeki Yılbaşı. Application of energy and exergy analyses to a CI engine using biodiesel fuel. Mathematical and Computational Applications, 16(4):797– 808, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/4/797.

Soureshjani:2023:DRB

[SZ23] Mohammad Khodabakhshi Soureshjani and Richard G. Zytner. Developing a robust bioventing model. Mathematical and Computational Applications, 28(3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/76.

Shi:2023:SHD

[SZTY23] Xiaowen Shi, Xiangyu Zhang, Renwu Tang, and Juan Yang. Solve highdimensional reflected partial differential equations by neural network method. Mathematical and Computational Applications, 28 (4):??, August 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/79.

Tekin:1996:FEA

[TA96] Muhammed D. Tekin and M. Fatih Altan. The finite element analysis of buckling of laminated rectangular reinforced concrete plates with circular hole. Mathematical and Computational Applications, 1(2): 164–168, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/164.

Turetken:1999:PWD

[TA99] Bahattin Türetken and Ali Alkumru. Plane wave diffraction by an open parallel plate waveguide with dielectric loading. Mathematical and Computational Applications, 4(3):241–249, December 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/241.

Torre:2019:IMP

[TABT19] María Susana Torre, Anais Acquaviva, Jean-Marc Boyer, and Jorge Tredicce. Influence of a modulated parameter on hantavirus infection. Mathematical and Computational Applications, 24(3):??, September 2019. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/3/68.

Tantawy:2008:IMS

[Tan08] Said Tantawy. An iterative method for solving linear fraction programming (LFP) problem with sensitivity analysis. Mathematical and Computational Applications, 13(3):147–151, December 2008. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/3/147.

Tariq:2023:FHH

[TAS^{+}23] Muhammad Tariq, Hijaz Ahmad, Asif Ali Shaikh, Sotiris K. Ntouyas, Evren Hınçal, and Sania Qureshi. Fractional Hermite–Hadamardtype inequalities for differentiable preinvex mappings and applications to modified Bessel and q-digamma functions. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/108.

Taylan:2001:BSS

[Tay01] Metin Taylan. Behavior of ships in shallow and restricted Waters. Mathematical and Computational Applications, 6(1):1–11, April 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/1/1.

Taylan:2004:EFS

[Tay04] Metin Taylan. Effect of forward speed on ship rolling and stability. Mathematical and Computational Applications, 9(2):133–145, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/133.

Turkmen:2004:NNG

[TB04] Ramazan Türkmen and Durmuş Bozkurt. A note on the norms of the GCD matrix. Mathematical and Computational Applications, 9(2): 303–308, August 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/2/303.

Togun:2016:NVN

[TB16] Necla Togun and Süleyman Murat Bağdatlı. Nonlinear vibration of a nanobeam on a Pasternak elastic foundation based on nonlocal Euler–Bernoulli beam theory. Mathematical and Computational Applications, 21 (1):??, March 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/1/3.

Turuelo:2021:SAE

[TB21] Constantino Grau Turuelo and Cornelia Breitkopf. Simple algebraic expressions for the prediction and control of high-temperature annealed structures by linear perturbation analysis. Mathematical and Computational Applications, 26(2):??, June 2021. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/26/2/43.

Tarhan:2004:MTP

[TÇ04] Sefa Tarhan and Kazim Çarman. Modeling the torque and power requirements of traction tires of horticultural tractors using dimensional analysis. Mathematical and Computational Applications, 9(3):427–434, December 2004. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/427.

Taskin:2006:MMV

[TÇ06] Mustafa Taşkın and Ugur Çalıgülü. Modelling of microhardness values by means of artificial neural networks of Al/ Sicp metal matrix composite material couples processed with diffusion method. Mathematical and Computational Applications, 11(3):163– 172, December 2006. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/163.

Torkul:1996:CSE

[TCA96] O. Torkul, I. H. Cedimoğlu, and M. S. Aksoy. A comparative study of existing manufacturing, planning and control approaches. Mathematical and Computational Applications, 1(2): 117–126, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/117.

Toma:2020:MCC

[TCALK20] Milan Toma, Rosalyn Chan-Akeley, Christopher Lipari, and Sheng-Han Kuo. Mechanism of coup and contrecoup injuries induced by a knock-out punch. Mathematical and Computational Applications, 25 (2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/22.

Ting:2011:OLS

[TCC11] Chia-Kuan Ting, Yuan-Shyi Peter Chiu, and Chu-Chai Henry Chan. Optimal lot sizing with scrap and random breakdown occurring in backorder replenishing period. Mathematical and Computational Applications, 16(2):329–339, August 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/329.

Tadj:2023:EMB

[TCC^{+}23] Mohammed Tadj, Lakhdar Chaib, Abdelghani Choucha, Al-Motasem Aldaoudeyeh, Ahmed Fathy, Hegazy Rezk, Mohamed Louzazni, and Attia El-Fergany. Enhanced MPPT-based fractionalorder PID for PV systems using Aquila optimizer. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/99.

Tomy:2021:STE

[TCM21] Lishamol Tomy, Christophe Chesneau, and Amritha K. Madhav. Statistical techniques for environmental sciences: a review. Mathematical and Computational Applications, 26(4):??, December 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/74.

Tutar:2006:MEI

[TCY06] M. Tutar, I. Celik, and I. Yavuz. Modelling of effect of inflow turbulence on large eddy simulation of bluff body flows. Mathematical and Computational Applications, 11(3):225–234, December 2006. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/11/3/225.

Tanisli:2012:BDS

[TD12] Murat Tanişli and Süleyman Demir. Biquaternionic description of the Schrödinger equation. Mathematical and Computational Applications, 17(2):176– 181, August 2012. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/17/2/176.

Taskin:2008:ANN

[TDC08] Mustafa Taskin, Halil Dikbas, and Ugur Caligulu. Artificial neural network (ANN) approach to prediction of diffusion bonding behavior (shear strength) of Niti alloys manufactured by powder metalurgy method. Mathematical and Computational Applications, 13(3):183– 191, December 2008. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/13/3/183.

Tokatli:2011:EHA

[TDCT11] Gorkem Tokatli, Pinar Dundar, Moharram Challenger, and Tufan Turaci. An extended heuristic algorithm to settle reacting objects on a planar surface. Mathematical and Computational Applications, 16 (1):279–289, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/279.

Tekin:2011:SRB

[TDT^{+}11] M. Tekin, A. Demir, T. Turalı, H. Nohutçu, and M. Bağcı. Strengthening of RC beams with prefabricated RC rectangular cross-sectional plates. Mathematical and Computational Applications, 16(4):868–878, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/868.

Taskesen:2005:CAN

[TE05] Ahmet Taskesen and Yucel Ercan. Computer aided nonlinear analysis of machine tool vibrations and a developed computer software. Mathematical and Computational Applications, 10(3):377–385, December 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/377.

Topac:2011:RSC

[TEK11] M. Murat Topaç, H. Eren Enginar, and N. Sefa Kuralay. Reduction of stress concentration at the corner bends of the anti-roll bar by using parametric optimisation. Mathematical and Computational Applications, 16 (1):148–158, April 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/148.

Tomy:2021:SML

[TGC21] Lishamol Tomy, Veena G, and Christophe Chesneau. The sine modified Lindley distribution. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/26/4/81.

Tomy:2022:ASM

[TGC22] Lishamol Tomy, Veena G, and Christophe Chesneau. Applications of the sine modified Lindley distribution to biomedical data. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/3/43.

Torpi:1996:NNA

[TGG96] Hamid Torpi, Filiz Güneş, and Fikret Gürgen. Neural network approach for the characterisation of the active microwave devices. Mathematical and Computational Applications, 1(1):113– 118, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/113.

Tanrioven:1996:ARP

[TGK96] Muğdeşem Tanrıöven, Erdin Gökalp, and Celal Kocatepe. Active and reactive power adjustment of a synchronous generator connected to infinite bus system by fuzzy logic controller. Mathematical and Computational Applications, 1 (1):105–112, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/105.

Tezcan:1998:AUF

[TGS98] Tekin Tezcan, Ümit Gökkuş, and Gültekin Sınır. Analysis of unsteady flow in complex pipe system by the method of characteristics. Mathematical and Computational Applications, 3(1):27– 36, April 1998. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/1/27.

Torres:2018:SSO

[TJCGAPA18] Lizeth Torres, Javier Jiménez-Cabas, José Francisco Gómez-Aguilar, and Pablo Pérez-Alcazar. A simple spectral observer. Mathematical and Computational Applications, 23(2):??, June 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/2/23.

Telli:2004:MMB

[TK04] Sevda Telli and Osman Kopmaz. On the mathematical modeling of beams rotating about a fixed axis. Mathematical and Computational Applications, 9(3): 333–347, December 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/3/333.

Trejo:2019:MMM

[TKCC19] Imelda Trejo, Hristo Kojouharov, and Benito Chen-Charpentier. Modeling the macrophagemediated inflammation involved in the bone fracture healing process. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/1/12.

Taddele:2020:HSR

[TKGS20] Guash Haile Taddele, Poom Kumam, Anteneh Getachew Gebrie, and Kanokwan Sitthithakerngkiet. Halfspace relaxation projection method for solving multiple-set split feasibility problem. Mathematical and Computational Applications, 25 (3):??, September 2020. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/47.

Tamilzharasan:2022:MMC

[TKK^{+}22] B. M. Tamilzharasan, S. Karthikeyan, Mohammed K. A. Kaabar, Mehmet Yavuz, and Fatma Özköse. Magneto mixed convection of Williamson nanofluid flow through a double stratified porous medium in attendance of activation energy. Mathematical and Computational Applications, 27 (3):??, June 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/46.

Turkan:2010:BGI

[TM10] Nureddin Turkan and Ismail Maras. E(5) behaviour of the Ge isotopes. Mathematical and Computational Applications, 15(3):428–438, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/428.

Turkan:2011:SRI

[TM11] Nureddin Turkan and Ismail Maras. Search on results of IBM for region between 120 ≤ A ≤ 150: 120–128 Te and 122–134 Xe nuclei. Mathematical and Computational Applications, 16(2):467– 476, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/467.

Turkmen:1996:NSR

[TMB96] Halit S. Türkmen, Zahit Mecitoğlu, and Oğuz Borat. Nonlinear structural response of laminated composite panels subjected to blast loadings. Mathematical and Computational Applications, 1(1):126–133, June 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/126.

Toussi:2021:PMP

[TMK21] Iman Bahreini Toussi, Abdolmajid Mohammadian, and Reza Kianoush. Prediction of maximum pressure at the roofs of rectangular water tanks subjected to harmonic base excitation using the multigene genetic programming method. Mathematical and Computational Applications, 26 (1):??, March 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/6.

Tanbay:2013:NSM

[TO13] Tayfun Tanbay and Bilge Ozgener. Numerical solution of the multigroup neutron diffusion equation by the meshless RBF collocation method. Mathematical and Computational Applications, 18(3):399–407, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/399.

Taisho:2022:FSC

[TOM^{+}22] Ayumu Taisho, Keiko Ono, Erina Makihara, Naoya Ikushima, and Sohei Yamakawa. Furniture style compatibility estimation by multi-branch deep Siamese network. Mathematical and Computational Applications, 27(5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/76.

Tchoumi:2022:III

[TRD^{+}22] Stéphane Yanick Tchoumi, Herieth Rwezaura, Mamadou Lamine Diagne, Gilberto González-Parra, and Jean Tchuenche. Impact of infective immigrants on COVID-19 dynamics. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/11.

Tsao:2009:PMM

[Tsa09] Yu-Chung Tsao. Profit maximization multi-item inventory models considering trade credit and sales learning curve. Mathematical and Computational Applications, 14(1):45–53, April 2009. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/14/1/45.

Toparli:2011:NIR

[TSC11] Mustafa Toparli, Onur Sayman, and Erdal Celik. Numerical investigation of residual stresses in YBCO/CeO2/YSZ/ CeO2/Ni architextured materials for coated conductors. Mathematical and Computational Applications, 16(1):125– 135, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/1/125.

Trachoo:2017:LTH

[TSS17]  Kamonchat  Trachoo,

Wannika Sawangtong, and Panumart Sawangtong. Laplace transform homotopy perturbation method for the two dimensional Black– Scholes model with European call option. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/23.

Tantay:2011:CBC

[TT11] Bahadır Tantay and Ferhat Taş. The curvature of a Bézier control polyline. Mathematical and Computational Applications, 16(2):350– 358, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/350.

Turkan:1997:IIS

[TU97] N. Turkan and İ. Uluer.

An IBM(2) investigation of some properties of 100Ru nuclei. Mathematical and Computational Applications, 2 (2):79–84, August 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/2/79.

Tashi:2019:GMP

[TUK19] Tashi Tashi, AMM Sharif Ullah, and Akihiko Kubo. Geometric modeling and 3D printing using recursively generated point cloud. Mathematical and Computational Applications, 24 (3):??, September 2019. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/24/3/83.

Turaci:2016:ALD

[Tur16] Tufan Turaci. The average lower 2-domination number of wheels related graphs and an algorithm. Mathematical and Computational Applications, 21(3):??, September 2016. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/3/29.

Teng:2018:HBR

[TW18] Zhongming Teng and Xuansheng Wang. Heavy ball restarted CMRH methods for linear systems. Mathematical and Computational Applications, 23(1):??, March 2018. CODEN ???? ISSN 2297-8747. URL

                                        https://www.mdpi.com/
                                        2297-8747/23/1/10.

Tsao:2010:EJR

[TWL10] Yu-Chung Tsao, Ming-Yu Wang, and Pei-Ling Lee. Effects of joint replenishment policy on company cost under permissible delay in payments. Mathematical and Computational Applications, 15(2):248–258, August 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/2/248.

Tan:2017:FPS

[TWS17] Jieqing Tan, Bo Wang, and Jun Shi. A fivepoint subdivision scheme with two parameters and a four-point shapepreserving scheme. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/22.

Turan:2011:OOC

[TY11] Mustafa Erkan Turan and Mehmet Ali Yurdusev. Optimization of open canal cross sections by differential evolution algorithm. Mathematical and Computational Applications, 16 (1):77–86, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/77.

Tastekin:2015:FCM

[TY15] Dilek Taştekin and Salih Yalçınbaş. Fermat collocation method for nonlinear system of first order boundary value problems. Mathematical and Computational Applications, 20(2):94– 105, August 2015. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/2/94.

Tastekin:2013:TCM

[TYS13] Dilek Taştekin, Salih Yalçınbaş, and Mehmet Sezer. Taylor collocation method for solving a class of the first order nonlinear differential equations. Mathematical and Computational Applications, 18(3):383– 391, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/18/3/383.

Tarlakovskii:2019:EOE

[TZ19] Dmitry Tarlakovskii and Andrei Zemskov. An elastodiffusive orthotropic Euler–Bernoulli beam considering diffusion flux relaxation. Mathematical and Computational Applications, 24(1):??, March 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/23.

Usman:2021:NSA

[UAI21] Muhammad Usman, Shaaban Abdallah, and Mudassar Imran. Nonlinear and stability analysis of a ship with general roll-damping using an asymptotic perturbation method. Mathematical and Computational Applications, 26 (2):??, June 2021. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/2/33.

Uddin:2019:ANB

[UAT19] Marjan Uddin, Hazrat Ali, and Muhammad Taufiq. On the approximation of a nonlinear biological population model using localized radial basis function method. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/54.

Uslan:2010:MIS

[UB10] Volkan Uslan and Ihsan Ömür Bucak. Microarray image segmentation using clustering methods. Mathematical and Computational Applications, 15(2):240– 247, August 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/2/240.

Ubeyli:2004:BMP

[Übe04] Mustafa Übeyli. 233U breeding in a modified Prometheus-l IFE reactor. Mathematical and Computational Applications, 9(3):463–473, December 2004. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/9/3/463.

Ugurlu:1996:SMD

[UÇ96] H. Hüseyin Uğurlu and Ali Çalışkan. The study mapping for directed space-like and time-like in Minkowski 3-space R13. Mathematical and Computational Applications, 1(2):142–148, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/142.

Ugur:1999:BEF

[UC99] Ahmet Ugur and Michael Conrad. Building evolution friendliness into cellular automation dynamics: The cytomatrix neuron model. Mathematical and Computational Applications, 4(3): 305–314, December 1999. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/305.

Urkmez:2010:DSP

[UC10] Abdullah Urkmez and Nurettin Cetinkaya. Determining spot price and economic dispatch in deregulated power systems. Mathematical and Computational Applications, 15(1):25–33, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/25.

Ustundag:2011:RSN

[ÜC11] Dursun Üstündağ and Mehmet Cevri. Recovering sinusoids from noisy data using Bayesian inference with simulated annealing. Mathematical and Computational Applications, 16(2):382– 391, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/382.

Uzun:2019:NFF

[UC19] Büşra Uzun and Ömer Civalek. Nonlocal FEM formulation for vibration analysis of nanowires on elastic matrix with different materials. Mathematical and Computational Applications, 24 (2):??, June 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/2/38.

Unlu:2004:DFC

[ÜDMA04] Bekir Sadik Ünlü, Hülya (Kaçar) Durmuş, Cevdet Meriç, and Enver Atik. Determination of friction coefficient at journal bearings by experimental and by means of artificial neural networks method. Mathematical and Computational Applications, 9(3): 399–408, December 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/3/399.

Ugurlu:2000:CHS

[UG00] H. H. Uğurlu and H. Gündoğan. The cosine hyperbolic and sine hyperbolic rules for dual hyperbolic spherical trigonometry. Mathematical and Computational Applications, 5(3):185–190, December 2000. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/5/3/185.

Ugurlu:2013:NHA

[Ugu13] Onur Ugurlu. A new heuristic algorithm to solve the maximum independent set problem. Mathematical and Computational Applications, 18(3):495–501, December 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/495.

Ubeyli:2003:AFA

[ÜIG03] Elif Derya Übeyli, Hakan Işık, and İnan Güler. Application of FFT and Arma spectral analysis to arterial Doppler signals. Mathematical and Computational Applications, 8(3):311–318, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/311.

Ucar:1996:CEM

[UK96a] F. Nazan Uçar and Mustafa Karaman. A cost-effective method for ultrasound volumetric imaging. Mathematical and Computational Applications, 1(2):127–132, December 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/127.

Ugurlu:1996:FDI

[UK96b] H. Hüseyin Uğurlu and Hüseyin Kocayiğit. The Frenet and Darboux instantaneous rotation vectors of curves on timelike surface. Mathematical and Computational Applications, 1(2): 133–141, December 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/2/133.

Ugur:2009:GAB

[UKÇ^{+}09] Aybars Uğur, Serdar Korukoğlu, Ali Çalıskan, Muhammed Cinsdikici, and Ali Alp. Genetic algorithm based solution for TSP on a sphere. Mathematical and Computational Applications, 14(3):219–228, December 2009. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/14/3/219.

Ugurlu:2005:FTH

[UKÖ05] H. H. Uğurlu, M. Kazaz, and A. Özdemir. Fundamental theorems for the hyperbolic geodesic triangles. Mathematical and Computational Applications, 10(2):231– 238, August 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/2/231.

Uzunoglu:1996:ANN

[UKY96] Mehmet Uzunoğlu, Celal Kocatepe, and Recep Yumurtaca. An artificial neural network based preestimation fitler for bad data defection, identification and elimination in state estimation. Mathematical and Computational Applications, 1(1):159–164, June 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/159.

Ullah:2017:SRM

[Ull17] A. M. M. Sharif Ullah.

Surface roughness modeling using q-sequence. Mathematical and Computational Applications, 22(2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/2/33.

Umul:2011:DWW

[Umu11] Yusuf Ziya Umul. Diffraction of waves by a wedge with phase conjugate faces. Mathematical and Computational Applications, 16(3):748–752, December 2011. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/748.

Unal:1996:DMM

[Üna96]  Adem Ünal.  Deter-

mination of mathematical model of electric power system using linear graph. Mathematical and Computational Applications, 1(1):134– 139, June 1996. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/134.

Unal:2003:AFI

[Üna03]  Gazanfer Ünal.  Ap-

proximate first integrals of a Hamiltonian system with two-degrees of freedom. Mathematical and Computational Applications, 8(2):151– 158, August 2003. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/8/2/151.

Unver:2007:IMT

[Unv07] Ihsan Unver. On the inventory model with two delaying barriers. Mathematical and Computational Applications, 12(3):125–134, December 2007. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/12/3/125.

Ustundag:2003:RIT

[ÜÖ03] Dursun Üstündağ and Ersen Özbaşaran. Recovering images from traveltime data. Mathematical and Computational Applications, 8 (1):95–102, April 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/1/95.

Uulu:2014:EMC

[UÖKE14] Shahriar Shamil Uulu, Halil Rıdvan Öz, Ferhat Karaca, and Cahit A. Evrensel. Estimation of mucus clearance in pulmonary airways by means of a regression model. Mathematical and Computational Applications, 19(2):144–151, August 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/2/144.

Ugurlu:1996:RBD

[UT96] H. Hüseyin Uğurlu and Ali Topal. Relation between Darboux instantaneous rotation vectors of curves on time-like surface. Mathematical and Computational Applications, 1(2):149–157, December 1996. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/149.

Unal:2005:WMS

[ÜU05] Füsun Ünal and Aynur Uysal. Weyl manifolds with semi-symmetric connection. Mathematical and Computational Applications, 10(3):351– 358, December 2005. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/3/351.

Umul:2003:AEE

[UY03] Y. Zyra Umul and U. Yalçin. Asymptotic evaluation of the edge diffraction in cylindric paraboloidal reflector antennas. Mathematical and Computational Applications, 8(2): 143–150, August 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/2/143.

Usal:2011:CEA

[UY11] Melek Usal and Lokman Yünlü. On constitutive equations for anisotropic nonlinearly piezoelectric materials. Mathematical and Computational Applications, 16(4):839– 848, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/4/839.

UlHaq:2022:SCF

[UYAA22] Ihtisham Ul Haq, Mehmet Yavuz, Nigar Ali, and Ali Akgül. A SARS-CoV-2 fractional-order mathematical model via the modified Euler method. Mathematical and Computational Applications, 27(5):??, October 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/5/82.

Vanani:2012:NSF

[VA12] Solat Karimi Vanani and Azim Aminataei. On the numerical solution of fractional partial differential equations. Mathematical and Computational Applications, 17(2):140–151, August 2012. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/17/2/140.

Venter:2023:IEL

[VC23a] Martin Philip Venter and Naudé Thomas Conradie. Intermediate encoding layers for the generative design of 2D soft robot actuators: a comparison of CPPN’s, l-systems and random generation. Mathematical and Computational Applications, 28(3):??, June 2023. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/68.

Vijay:2023:CPR

[VC23b] Vijay and Arya Kumar Bedabrata Chand. Convexity-preserving rational cubic zipper fractal interpolation curves and surfaces. Mathematical and Computational Applications, 28 (3):??, June 2023. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/28/3/74.

Vahaplar:2011:EDL

[VÇÖ11] Alper Vahaplar, C. Cengiz Çelikoğlu, and Murat Özgören. Entropy in dichotic listening EEEG recordings. Mathematical and Computational Applications, 16(1):43– 52, April 2011. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/43.

Vargas-Hakim:2020:EMO

[VHMMG20] Gustavo-Adolfo Vargas-Hákim, Efrén Mezura-Montes, and Edgar Galván. Evolutionary multi-objective energy production optimization: an empirical comparison. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/32.

Venter:2023:GDS

[VJ23] Martin Philip Venter and Izak Johannes Joubert. Generative design of soft robot actuators using ESP. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/53.

Vargas-Martinez:2023:PAM

[VMRVF^{+}23] Manuel Vargas-Martínez, Nelson Rangel-Valdez, Eduardo Fernández, Claudia Gómez-Santillán, and María Lucila Morales-Rodríguez. Performance analysis of multiobjective simulated annealing based on decomposition. Mathematical and Computational Applications, 28 (2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/38.

Valcke:2022:BRL

[VPJ22] Sophie Valcke, Andrea Piacentini, and Gabriel Jonville. Benchmarking regridding libraries used in Earth system modelling. Mathematical and Computational Applications, 27(2):??, April 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/2/31.

Vahaj:2019:ETD

[VRMJG19] AmirAbbas Vahaj, Akbar Rahideh, Hossein Moayed-Jahromi, and AliReza Ghaffari. Exact two-dimensional analytical calculations for magnetic field, electromagnetic torque, UMF, back-EMF, and inductance of outer rotor surface inset permanent magnet machines. Mathematical and Computational Applications, 24 (1):??, March 2019. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/1/24.

Vazquez-Santiago:2023:VMM

[VSAMMM23]  Diana-Itzel  Vázquez-

Santiago, Héctor-Gabriel Acosta-Mesa, and Efrén Mezura-Montes. Vehicle make and model recognition as an openset recognition problem and new class discovery. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/80.

Visser:2023:SGB

[VVN23] Carl-Hein Visser, Gerhard Venter, and Melody Neaves. A strain-gauge-based method for the compensation of out-of-plane motions in 2D digital image correlation. Mathematical and Computational Applications, 28(2):??, April 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/2/40.

VanTonder:2023:NMI

[VVV23] John Dean Van Tonder, Martin Philip Venter, and Gerhard Venter. A new method for improving inverse finite element method material characterization for the Mooney–Rivlin material model through constrained optimization. Mathematical and Computational Applications, 28(4):??, August 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/4/78.

Wu:2001:WNN

[WA01] Mingyang Wu and Hojjat Adeli. Waveletneural network model for automatic traffic incident detection. Mathematical and Computational Applications, 6 (2):85–96, August 2001. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/6/2/85.

Wang:2013:NMS

[Wan13] Minghui Wang. A new method for solving matrix equation AXB + CXT D = E. Mathematical and Computational Applications, 18 (1):12–18, April 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/1/12.

Wang:2023:OBS

[WCH23] Wanlin Wang, Jinxiong Chen, and Zhenkun Huang. Observer-based state estimation for recurrent neural networks: an output-predicting and LPV-based approach. Mathematical and Computational Applications, 28(6):??, December 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/104.

Wang:2023:HNM

[WED^{+}23] Hao Wang, Michael Emmerich, André Deutz, Víctor Adrián Sosa Hernández, and Oliver Schütze. The hypervolume Newton method for constrained multi-objective optimization problems. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/10.

Willems:2021:FEA

[WFV21] Robin Willems, Léo A. J. Friedrich, and Clemens V. Verhoosel. Finite element analysis of laminar heat transfer within an axial-flux permanent magnet machine. Mathematical and Computational Applications, 26(1):??, March 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/1/23.

Wu:2010:SSS

[WH10] Ling-Ye Wu and Ji-Huan He. Study on the stability of Steiner tree structure of explosion-proof textiles. Mathematical and Computational Applications, 15(5):936– 939, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/936.

Wang:2023:KTB

[WJ23] Xilu Wang and Yaochu Jin. Knowledge transfer based on particle filters for multi-objective optimization. Mathematical and Computational Applications, 28 (1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/14.

Wang:2010:APS

[WLW^{+}10] Yuzhi Wang, Bo Li, Renqing Wang, Jing Su, and Xiaoxia Rong. Application of the phase space reconstruction in ecology. Mathematical and Computational Applications, 15(5):872–876, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/872.

Walker:2021:ASR

[WN21] Anthony S. Walker and Kyle E. Niemeyer. Applying the swept rule for solving two-dimensional partial differential equations on heterogeneous architectures. Mathematical and Computational Applications, 26 (3):??, September 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/3/52.

Wen:2015:MHM

[WOL15] Xingzi Wen, Junlin Ouyang, and Youjin Liu. A method of hybrid multiple attributes group decision making with risk considering decision-makers’ confidence. Mathematical and Computational Applications, 20(1):62–75, April 2015. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/1/62.

Wang:2017:DCV

[WSLZ17] Jun Wang, Dagang Sun, Shizhong Liu, and Xin Zhang. Damping characteristics of viscoelastic damping structure under coupled condition. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/27.

Wu:2015:AMM

[WT15] Wenqing Wu and Yinghui Tang. Analysis of an M^{(λ1,λ2)} /M/1/WV queue with controlled vacation interruption and variable arrival rate. Mathematical and Computational Applications, 20(1):39–49, April 2015. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/1/39.

Wu:2010:VAS

[Wu10] Yue Wu. Variational approach to solitary solutions using Jacobi-elliptic functions. Mathematical and Computational Applications, 15(5):910– 923, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/910.

Wu:2017:AHF

[Wu17] Chunqing Wu. Analysis of a hand-foot-mouth disease model with standard incidence rate and estimation for basic reproduction number. Mathematical and Computational Applications, 22 (2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/22/2/29.

Wen:2020:VFB

[WW20] Conghua Wen and Junwei Wei. Volatility forecasting based on cyclical two-component model: Evidence from Chinese futures markets and sector stocks. Mathematical and Computational Applications, 25(3):??, September 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/3/59.

Wang:2017:ROT

[WZ17] Jiangtao Wang and Jingai Zhang. Robust object tracking in infrared video via adaptive weighted patches. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/3.

Wang:2023:EIF

[WZ23] Yanan Wang and Shuying Zhai. Explicit integrating factor Runge– Kutta method for the extended Fisher–Kolmogorov equation. Mathematical and Computational Applications, 28(6):??, December 2023. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/6/110.

Wang:2020:EGA

[WZT20] Yue Wang, Jinchuan Zhou, and Jingyong Tang. Existence of generalized augmented Lagrange multipliers for constrained optimization problems. Mathematical and Computational Applications, 25(2):??, June 2020. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/25/2/24.

Xiao:2010:HFO

[Xia10] Jian-Hua Xiao. Hard fling objector contact with surface of fluid as a ricochet. Mathematical and Computational Applications, 15(5):940– 945, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/940.

Xing:2022:MOO

[XS22] Siyuan Xing and Jian-Qiao Sun. Multiobjective optimization of an elastic rod with viscous termination. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/94.

Xu:2010:AHA

[Xu10] Lan Xu. Application of Hamiltonian approach to an oscillation of a mass attached to a stretched elastic wire. Mathematical and Computational Applications, 15(5):901– 906, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/901.

Xiao:2019:PPA

[XWCF19] Xiaodong Xiao, Ting Wu, Yuanfang Chen, and Xingyue Fan. Privacypreserved approximate classification based on homomorphic encryption. Mathematical and Computational Applications, 24(4):??, December 2019. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/24/4/92.

Yalcinbas:2011:HSS

[YA11] Salih Yalçınbaş and Müge Aynigül. Hermite series solutions of linear Fredholm integral equations. Mathematical and Computational Applications, 16(2):497– 506, August 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/2/497.

Yuzukirmizi:2014:FDS

[YA14] Mustafa Yuzukirmizi and Hakan Arslan. Fault diagnosis of shaft- ball bearing system using one-way analysis of variance. Mathematical and Computational Applications, 19(1):37–49, April 2014. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/19/1/37.

Yaagoub:2023:GSM

[YA23] Zakaria Yaagoub and Karam Allali. Global stability of multi-strain SEIR epidemic model with vaccination strategy. Mathematical and Computational Applications, 28(1):??, February 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/1/9.

Yalcinbas:2010:LSS

[YAA10] Salih Yalçınbas, Müge Aynigül, and Tuğçe Akkaya. Legendre series solutions of Fredholm integral equations. Mathematical and Computational Applications, 15(3):371–381, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/3/371.

Yigit:1996:FEM

[YAAK96] A. S. Yigit, M. D. Al-Ansary, and M. Khalid. Finite element modeling of drillstrings. Mathematical and Computational Applications, 1(2): 158–163, December 1996. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/2/158.

Yagiz:2002:RCS

[Yag02]  Nurkan Yagiz.  Ro-

bust control of a spatial robot using sliding modes. Mathematical and Computational Applications, 7(3):219–228, December 2002. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/3/219.

Yalcinbas:2013:ASL

[Yal13] Salih Yalçınbaş. Approximate solutions of linear Fredholm integral equations system with variable coefficients. Mathematical and Computational Applications, 18 (1):19–29, April 2013. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/1/19.

Yan:2016:ABC

[Yan16] Lanlan Yan. Adjustable Bézier curves with simple geometric continuity conditions. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/44.

Yao:2015:PAP

[Yao15] Zhijian Yao. Positive almost periodic solution for impulsive Nicholson’s blowflies model with linear harvesting term on the bounded domain. Mathematical and Computational Applications, 20(1):1–14, April 2015. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/1/1.

Yavicoli:2022:SNT

[Yav22] Alexia Yavicoli. A survey on newhouse thickness, fractal intersections and patterns. Mathematical and Computational Applications, 27 (6):??, December 2022. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/27/6/111.

Yayli:2000:MFV

[Yay00] Yusuf Yayli. On the motion of the Frenet vectors and spacelike ruled surfaces in the Minkowski 3-space. Mathematical and Computational Applications, 5(1):49– 55, April 2000. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/5/1/49.

Yazici:2016:DPD

[Yaz16] Hakan Yazici. Design of a parameter-dependent optimal vibration control of a non-linear vehicle suspension system. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/13.

Yilmaz:1997:INR

[YB97]  Ş. Yılmaz and S. Bozdemir.

Investigation of NMR relaxation data using dipole correlation function. Mathematical and Computational Applications, 2(3):131–139, December 1997. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/2/3/131.

Yaldiz:2003:DMD

[YB03] Ercan Yaldiz and Mehmet Bayrak. A different method determining dielectric constant of soil and its FDTD simulation. Mathematical and Computational Applications, 8(3):303–310, December 2003. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/303.

Yang:2017:NSN

[YBTH17] Touna Yang, Nguyen Thanh Binh, Tran Manh Thang, and Duong Thi Hoa. A new smoothing nonlinear penalty function for constrained optimization. Mathematical and Computational Applications, 22(2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/2/31.

Yang:2021:LAC

[YBYL21] Xue Yang, Hong Bian, Haizheng Yu, and Dandan Liu. The local antimagic chromatic numbers of some join graphs. Mathematical and Computational Applications, 26(4):??, December 2021. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/26/4/80.

Yigit:2011:ARM

[YD11] Hüseyin Yiğit and Fatih Dikmen. Analytical regularization method for rigorous diffraction analysis of knife edges for TM waves. Mathematical and Computational Applications, 16(3):738– 747, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/738.

Yurdem:2011:DSD

[YDD11] Huseyin Yurdem, Vedat Demir, and Adnan Degirmencioglu. Development of a software to determine the emitter characteristics and the optimum length of new designed drip irrigation laterals. Mathematical and Computational Applications, 16(3):728– 737, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/728.

Yagiz:2003:ECM

[YE03] Nurkan Yagiz and Cuneyt Ertal. Evaluation of control methods on a structural system. Mathematical and Computational Applications, 8(3): 369–376, December 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/369.

Yeniay:2005:PFM

[Yen05] Özgür Yeniay. Penalty function methods for constrained optimization with genetic algorithms. Mathematical and Computational Applications, 10(1):45–56, April 2005. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/10/1/45.

Yeniay:2014:CSA

[Yen14] Özgür Yeniay. Comparative study of algorithms for response surface optimization. Mathematical and Computational Applications, 19 (1):93–104, April 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/1/93.

Yeni:2002:ISS

[YG02] E. Çınar Yeni and Binnur Gören. Investigation of the stresses and strains of steel column-beam connections subjected to seismic loading. Mathematical and Computational Applications, 7(2): 123–132, August 2002. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/7/2/123.

Yaman:2011:CDD

[YG11] Metin Yaman and Şevket Gür. Continuous dependence for the damped nonlinear hyperbolic equation. Mathematical and Computational Applications, 16(2):437–442, August 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/2/437.

Yalcinbas:2015:BCM

[YG15] Salih Yalçınbaş and Huriye Gürler. Bernstein collocation method for solving the first order nonlinear differential equations with the mixed non-linear conditions. Mathematical and Computational Applications, 20(3):160–173, December 2015. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/20/3/160.

Yuzbasi:2013:MLP

[YGS13]  Şuayip Yüzbaşı, Em-

rah Gök, and Mehmet Sezer. Müntz-Legendre polynomial solutions of linear delay Fredholm integro-differential equations and residual correction. Mathematical and Computational Applications, 18(3):476– 485, December 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/3/476.

Yaliniz:2005:NGH

[YH05] Ayse Funda Yalınız and H. Hilmi Hacısalihoğlu. Null generalized helices in L3 and L4, 3 and 4dimensional Lorentzian space. Mathematical and Computational Applications, 10(1):105– 111, April 2005. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/10/1/105.

Yildiz:2011:IFE

[Yil11] Hatice Duran Yildiz. Infrared free electron laser, resonator parameters optimization with GENE-SIS, OPC and GLAD codes. Mathematical and Computational Applications, 16(3):659–668, December 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/3/659.

Yavuz:1996:FVR

[YK96a] A. Kadir Yavuz and Temel Kotil. Free vibrations of rotating adhesively bonded orthotropic plates. Mathematical and Computational Applications, 1 (1):140–146, June 1996. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/1/1/140.

Yigit:1996:VCM

[YK96b] Ahmet S. Yiğit and Mohammad Khalid. Vibration control methods by mode localization: a comparative study. Mathematical and Computational Applications, 1(1):147–153, June 1996.

CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/1/1/147.

Yimer:2020:PGM

[YKG20] Seifu Endris Yimer, Poom Kumam, and Anteneh Getachew Gebrie. Proximal gradient method for solving bilevel optimization problems. Mathematical and Computational Applications, 25(4):??, December 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/4/66.

Yang:2016:FGP

[YL16] Lian Yang and Zhangping Lu. Fuzzy grey prediction-based particle filter for object tracking. Mathematical and Computational Applications, 21(3):??, September 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/3/37.

Yang:2013:MST

[YLX^{+}13] Shu-Hong Yang, Chui-Gui Li, Dong-Xue Xia, Meng Wang, and Xiao-Rong Wang. Modeling and simulation of a traffic flow model considering the influence of traffic interruption. Mathematical and Computational Applications, 18(2):112– 121, August 2013. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/18/2/112.

Yalcinbas:2010:ASH

[YÖS10] Salih Yalçınbas, Nesrin Özsoy, and Mehmet Sezer. Approximate solution of higher order linear differential equations by means of a new rational Chebyshev collocation method. Mathematical and Computational Applications, 15 (1):45–56, April 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/1/45.

Yin:2022:EOP

[YQG22] Shengwen Yin, Haogang Qin, and Qiang Gao. An efficient orthogonal polynomial method for auxetic structure analysis with epistemic uncertainties. Mathematical and Computational Applications, 27(3):??, June 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/3/49.

Yu:2016:VGM

[YRB16] Chunxiao Yu, Cuihuan Ren, and Xueting Bai. VRP-GMRES(m) iteration algorithm for fast multipole boundary element method. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/49.

Yildinm:2002:ANN

[YS02]  Ş. Yıldınm and Ş. Sağıroğlu.

Artificial neural networks in robot control systems: a survey paper. Mathematical and Computational Applications, 7(2): 103–112, August 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/2/103.

Yu:2017:OHO

[YS17] Xing Yu and Hongguo Sun. Optimal hedging with options and futures against price risk and background risk. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/5.

Yazar:2007:CBI

[YU07] Harun Resit Yazar and İhsan Uluer. A correspondence between IBAi and IBA-II model and electromagnetic transitions of some erbium isotopes. Mathematical and Computational Applications, 12(2):69– 76, August 2007. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/2/69.

Yuksel:2003:RAC

[YUKA03] Ramazan Yüksel, Mehmet Uzunoğlu, Celal Kocatepe, and Mehmet Argin. The resonance analysis caused by harmonics in power systems including thyristor controlled reactor. Mathematical and Computational Applications, 8(3): 335–341, December 2003. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/8/3/335.

Yurdusev:1999:AEC

[Yur99] Mehmet Ali Yurdusev. Assessing enviromental consequences of water projects based on a pseudo expert system approach. Mathematical and Computational Applications, 4(3):251–259, December 1999. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/4/3/251.

Yurusoy:2002:PDS

[Yür02]  M. Yürüsoy.  Pres-

sure distribution in a slider bearing lubricated with second and third grade fluids. Mathematical and Computational Applications, 7 (1):15–22, April 2002. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/7/1/15.

Yurusoy:2004:FTG

[Yür04] M. Yürüsoy. Flow of a third grade fluid between concentric circular cylinders. Mathematical and Computational Applications, 9 (1):11–17, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/11.

Yao:2024:NMB

[YW24] Mengli Yao and Zhifeng Weng. A numerical method based on operator splitting collocation scheme for nonlinear Schrödinger equation. Mathematical and Computational Applications, 29(1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/29/1/6.

Yenicerioglu:2007:SDI

[YY07] Ali Fuat Yeniçerioğlu and Salih Yalçınbaş. On the stability of delay integrodifferential equations. Mathematical and Computational Applications, 12(1):51–58, April 2007. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/12/1/51.

Yilmaz:2011:UDE

[YY11] Barış Yılmaz and Mehmet Ali Yurdusev. Use of data envelopment analysis as a multi criteria decision tool – a case of irrigation management. Mathematical and Computational Applications, 16(3):669– 679, December 2011. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/16/3/669.

Yin:2016:POH

[YYWX16] Wenjun Yin, Yinwei Yang, Zhanyu Wang, and Jing Xu. Parameters optimization of a hydraulic buffer system for belt arrestor in downward belt conveyors. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/42.

Yao:2016:RSI

[YZBZ16] Xunxiang Yao, Yunfeng Zhang, Fangxun Bao, and Caiming Zhang. Rational spline image upscaling with constraint parameters. Mathematical and Computational Applications, 21(4):??, December 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/48.

Yu:2017:QIO

[YZHY17] Rengui Yu, Chungang Zhu, Xianmin Hou, and Li Yin. Quasiinterpolation operators for bivariate quintic spline spaces and their applications. Mathematical and Computational Applications, 22 (1):??, March 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/10.

Zor:1998:SAA

[ZA98] Mehmet Zor and Sami Aksoy. Stres analysis around dental implant in human mandible. Mathematical and Computational Applications, 3 (2):83–91, August 1998. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/3/2/83.

Zana:2018:TAP

[ZBBZ18] Roland Reginald Zana, Bálint Bodor, László Bencsik, and Ambrus Zelei. A tutorial for the analysis of the piecewisesmooth dynamics of a constrained multibody model of vertical hopping. Mathematical and Computational Applications, 23(4):??, December 2018. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/23/4/74.

Zhou:2015:SBA

[ZCC15] Liangqiang Zhou, Fangqi Chen, and Yushu Chen. Stability and bifurcation analysis of a pipe conveying pulsating fluid with combination parametric and internal resonances. Mathematical and Computational Applications, 20(3):200– 216, December 2015. CO-DEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/20/3/200.

Zarate:2022:ADE

[ZdlÁGLTT22] Josué Enríquez Zárate, María de los Ángeles Gómez López, Javier Alberto Carmona Troyo, and Leonardo Trujillo. Analysis and detection of erosion in wind turbine blades. Mathematical and Computational Applications, 27 (1):??, February 2022. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/1/5.

Zahir:2016:RSE

[ZE16] Jihad Zahir and Abderrahim El Qadi. A recommendation system for execution plans using machine learning. Mathematical and Computational Applications, 21 (2):??, June 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/2/23.

Zhang:2017:GMB

[ZG17] Li-Tao Zhang and Tong-Xiang Gu. Global modulus-based synchronous multisplitting multi-parameters TOR methods for linear complementarity problems. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/
2297-8747/22/1/20.

Zengin:2010:IGR

[ZGV10] Zeynep Zengin, Fikret Gürgen, and Füsun Varol. Intrauterine growth restriction (IUGR) risk decision based on support vector machines. Mathematical and Computational Applications, 15(3):472–480, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/3/472.

Zhou:2010:VAC

[ZH10] Lin-Hong Zhou and J. H. He. The variational approach coupled with an ancient Chinese mathematical method to the relativistic oscillator. Mathematical and Computational Applications, 15(5):930–935, December 2010. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/15/5/930.

Zhang:2014:NTP

[Zha14] Kun Zhang. A network traffic prediction model based on quantum inspired Pso and wavelet neural network. Mathematical and Computational Applications, 19(3):218–229, December 2014. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/19/3/218.

Zhang:2010:CMT

[ZHN10] Kai Zhang, Lizhong Huang, and Deming Nie.

Cylindrical microparticle transport and deposition from electrokinetic microflow in a 90 degree bend. Mathematical and Computational Applications, 15(5):822– 827, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/822.

Zhu:2010:EEM

[Zhu10] Shundong Zhu. The extended G′/G-expansion method and travelling wave solutions of nonlinear evolution equations. Mathematical and Computational Applications, 15(5):924–929, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/924.

Zhuang:2017:SDD

[Zhu17] Kejun Zhuang. Spatiotemporal dynamics of a delayed and diffusive viral infection model with logistic growth. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/7.

Zeng:2010:MAB

[ZJ10] Tao Zeng and Lili Jiang.

Mechanics analysis of 3D braided composites based on the helix geometry model. Mathematical and Computational Applications, 15(5):883– 888, December 2010. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/883.

Zhang:2020:PPA

[ZLLX20] Weiming Zhang, Dapan Li, Xuyang Lou, and Dezhi Xu. Prescribed performance adaptive backstepping control for winding segmented permanent magnet linear synchronous motor. Mathematical and Computational Applications, 25(2):??, June 2020. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/25/2/18.

Zhang:2017:SRA

[ZLZ^{+}17a] Hong-Bo Zhang, Qing Lei, Bi-Neng Zhong, Ji-Xiang Du, and Duan-Sheng Chen. A sparse representation algorithm for effective photograph retrieval. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/8.

Zhang:2017:MSP

[ZLZ17b] Yong Ou Zhang, Xu Li, and Tao Zhang. Modeling sound propagation using the corrective smoothed particle method with an acoustic boundary treatment technique. Mathematical and Computational Applications, 22(1):??, March 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/1/26.

Zaman:2023:LSC

[ZMA23] Fiazuddin D. Zaman, Fazal M. Mahomed, and Faiza Arif. Lie symmetry classification, optimal system, and conservation laws of damped Klein– Gordon equation with power law non-linearity. Mathematical and Computational Applications, 28(5):??, October 2023. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/28/5/96.

Zdanowski:2024:STR

[ZMCL24] Francisco Zdanowski, Isabel Malico, Paulo Canhoto, and Rui Pedro Lima. Simulation of a thermal recuperative incinerator of VOCs with a special focus on the heat exchanger. Mathematical and Computational Applications, 29 (1):??, February 2024. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/29/1/1.

Zarei:2011:SFD

[ZP11] Jafar Zarei and Javad Poshtan. Sensor fault detection and diagnosis of a process using unknown input observer. Mathematical and Computational Applications, 16(1):31–42, April 2011. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/16/1/31.

Zhang:2016:PAP

[ZPL16] Tianwei Zhang, Liyan Pang, and Yongzhi Liao. Positive almost periodic solutions for a delayed predator–prey model with Hassell– Varley type functional response. Mathematical and Computational Applications, 21(2):??, June 2016. CODEN ???? ISSN 2297-8747. URL https:

                                        //www.mdpi.com/2297-
                                        8747/21/2/10.

Zhao:2022:DCN

[ZPW^{+}22] Xiaoqing Zhao, Lei Pang, Lianming Wang, Sen Men, and Lei Yan. Deep convolutional neural network for detection and prediction of waxy corn seed viability using hyperspectral reflectance imaging. Mathematical and Computational Applications, 27(6):??, December 2022. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/27/6/109.

Zughbi:2004:NSF

[ZS04] Habib D. Zughbi and Saiffudin Sheikh. Numerical simulations of flow distribution in unpacked and partially packed vessels. Mathematical and Computational Applications, 9 (1):25–39, April 2004. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/9/1/25.

Zhao:2016:RNI

[ZS16] Jia Zhao and Gang Sun.

Reliable network interdiction models with multiple unit costs. Mathematical and Computational Applications, 21 (4):??, December 2016. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/50.

Zhou:2016:MOO

[ZWZ16] Junchao Zhou, Chun Wang, and Junjun Zhu. Multi-objective optimization of a spring diaphragm clutch on an automobile based on the non-dominated sorting genetic algorithm (NSGA-II). Mathematical and Computational Applications, 21(4):??, December 2016. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/21/4/47.

Zhao:2017:FCF

[ZWZ17] Wenling Zhao, Ruyu Wang, and Hongxiang Zhang. Finite convergence for feasible solution sequence of variational inequality problems. Mathematical and Computational Applications, 22 (2):??, June 2017. CO-DEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/2/36.

Zhong:2010:FAR

[ZZ10] Ting Zhong and Jingjing Zhang. Frequencyamplitude relationship for nonlinear oscillator with discontinuity. Mathematical and Computational Applications, 15(5):907–909, December 2010. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/15/5/907.

Zhao:2017:NMO

[ZZ17] Hongyan Zhao and Lian Zhou. A novel method of offset approximation along the normal direction with high precision. Mathematical and Computational Applications, 22(3):??, September 2017. CODEN ???? ISSN 2297-8747. URL https:

//www.mdpi.com/2297-
8747/22/3/39.